### v1

# A Calculus of Overlays

Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, Overlay-Calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. It trivially embeds the $\lambda$-calculus, although the entire semantics builds on only naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance simply does not arise. This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence. Overlay-Calculus is distilled from the Overlay language, a practical implementation in which we observed further emergent phenomena: the Expression Problem dissolves, programs are CPS-agnostic, records natively encode random-access memory, and self-reference resolves to multiple targets. These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
declarative programming, inheritance, denotational semantics, self-referential records, expression problem, $\lambda$-calculus, Böhm trees, configuration languages
This work was conducted independently prior to the author’s employment at Figure AI.
$c$s: Theory of computation Denotational semanticsccs: Theory of computation Object oriented constructsccs: Theory of computation Functional constructsccs: Theory of computation Program schemesccs: Software and its engineering Formal language definitionsccs: Software and its engineering Extensible languagesccs: Software and its engineering Object oriented languages
1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or overlay. Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves remarkable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself—Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory explains why this mechanism is so powerful.
The $\lambda$-calculus serves as the foundational computational model for functional programming. No analogous calculus exists for declarative programming. This gap matters because the two paradigms differ in fundamental ways. Configuration languages are inherently declarative: their values are immutable and they have no first-class functions. At first glance, Turing completeness appears incompatible with these constraints. The three classical models of computation each violate at least one of them:
Model Turing complete Immutable No $\lambda$
Turing Machine $\checkmark$ $\times$ $\checkmark$
$\lambda$-calculus $\checkmark$ $\checkmark$ $\times$
RAM Machine $\checkmark$ $\times$ $\checkmark$
The Turing Machine and RAM Machine require mutable state; the $\lambda$-calculus requires first-class functions. A computational model that is Turing complete, immutable, and function-free would sit in the empty cell of this table. The NixOS module system already suggests that such a model exists: its inheritance-based composition over recursive records, without explicit functions, is expressive enough to configure entire operating systems.
Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures—initial algebras in the sense of universal algebra. We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures—in the spirit of F-coalgebras (Rutten, 2000)—where semantics is determined by the observer and a finite prefix suffices for any finite observation. This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006)—over 100,000 packages—behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set. Guided by this observation and by the NixOS module system’s recursive merging mechanism, we built the Overlay language (Yang, 2025b)—a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
In developing real programs in the Overlay language—Church-encoded data types (Appendix B), an Expression Problem solution (Section 5), trie-based random-access memory (Appendix A), and continuation patterns (Appendix G)—we discovered that several long-standing problems in programming language theory do not arise. The linearization problem that has plagued mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) does not exist—not because it is resolved, but because the three primitives make it structurally impossible (Section 3). Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system). Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction. These are not properties we engineered; they are phenomena we observed in working programs.
This paper distills these observations into Overlay-Calculus, a minimal computational model for declarative programming. It turns out that function application has a straightforward translation into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument. Since the $\lambda$-calculus embeds into Overlay-Calculus, Overlay-Calculus is Turing complete. The translation reveals an asymmetry: while embedding the $\lambda$-calculus into Overlay-Calculus is trivial, the reverse is non-trivial because Overlay-Calculus provides primitive operations—self-referential records, open recursion, symmetric inheritance—that have no direct counterparts in the $\lambda$-calculus. Moreover, Overlay-Calculus naturally encodes random-access memory: records are tries, and inheritance is trie union, yielding an immutable analogue of the RAM machine (Appendix A). This asymmetry suggests that Overlay-Calculus is not merely a syntactic variant of the $\lambda$-calculus but captures different computational patterns.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, Overlay-Calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
All constructions in this paper have been implemented and tested in the Overlay language, available as open source (Yang, 2025b), with a test suite covering every example. All tests are mechanically verified against expected output snapshots.
Contributions
  • We present Overlay-Calculus, a minimal computational model for declarative programming that contains only overlay constructs (record, definition, inheritance) and no functions or scalar types.
  • Inheritance operates at the tree level: a record is a set of elements—definitions and inheritance sources—so inheriting from two overlays that share a label recursively merges their subtrees (union mount semantics), rather than overriding at the method level as in mixin and trait calculi. Since records are sets, inheritance is inherently commutative, idempotent, and associative. The linearization problem inherent in prior mixin and trait calculi simply does not arise.
  • The semantics is purely observational: there are no reduction rules, no evaluation order, and no notion of a result that a term reduces to. The observer drives computation by querying paths in a lazily constructed tree.
  • Self-referential records require no fixed-point combinator. The $\lambda$-calculus translation maps each application to a inheritance; the resulting inheritance tree is the call stack. The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree—a first-order set-theoretic query, not a fixed-point construction.
  • Self-reference resolves to multiple objects, not one. Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala, Appendix C; NixOS module system, Appendix D). Since inheritance is idempotent, all objects contribute equally—no disambiguation is needed.
  • The $\lambda$-calculus (in A-normal form) translates into Overlay-Calculus in five rules, so Overlay-Calculus is Turing complete. Composing this translation with the semantic functions induces a semantics on the $\lambda$-calculus that requires no function spaces, no substitution, and no reduction rules—only set membership in powerset lattices (Section 4). The translation is adequate and fully abstract with respect to Böhm tree equivalence (Section 4.2): two terms have the same overlay-convergence behavior in all contexts if and only if they have the same Böhm tree. This yields a fully abstract semantics of the lazy $\lambda$-calculus that lives entirely in naive set theory—powerset lattices and the Knaster–Tarski fixed-point theorem—without domain theory, inverse limits, or Scott continuity. Standard data types (booleans, natural numbers) can be Church-encoded without scalars (Appendix B).
  • We observe several emergent properties in practice: the Expression Problem (Wadler, 1998) does not arise (Section 5), programs are CPS-agnostic (Appendix G), and records natively encode random-access memory (Appendix A).
Section 5 demonstrates these properties through a case study of the Expression Problem.
2. Syntax
Let $e$ denote an expression. Let $\ell$ denote a label (property name). Let $k$ denote a non-negative integer (path length).
$e\quad::=\quad$ $\{c_{1},\;\ldots,\;c_{m}\}$ $\text{(}m\geq 0\text{)}$
$\mid\quad$ $c$ $\text{(sugar for }\{c\}\text{)}$
$c\quad::=\quad$ $\ell\mapsto e$ (definition)
$\mid\quad$ $[\ell_{\mathrm{up}}.\mathbf{this}.\mid\uparrow^{\!n}.]\;\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ $\text{(inheritance, }k\geq 0\text{ with qualifier,}$
$k\geq 1\text{ without)}$
An expression is a set of elements enclosed in braces. Each element is either a definition $\ell\mapsto e$ or an inheritance source. Because expressions are sets, element order is irrelevant and duplicates have no effect—inheritance is inherently commutative, idempotent, and associative.
The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding. There is no variable binding in Overlay-Calculus. The Overlay language also provides scalar types through a foreign-function interface (FFI); Overlay-Calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI—scalars simply did not arise.
Inheritance
An inheritance is written as a reference that navigates the scope hierarchy upward and then projects properties downward. The upward navigation admits two equivalent notations:
  • Named form: $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, analogous to Java’s Outer.this.field. Each $\{\ldots\}$ record has a label in its enclosing scope; $\ell_{\mathrm{up}}$ names the target enclosing scope.
  • Indexed form: $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n$ is a de Bruijn index (de Bruijn, 1972) counting the number of enclosing scope levels to traverse.
Both forms retrieve the target scope—the fully inherited record after all inheritance has been applied—and then project properties $\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ from it. The named form desugars to the indexed form during parsing (Section 3).
A new scope level is created at each $\{\ldots\}$ (record literal). The scope contains all properties of that overlay, including those inherited via inheritance sources. For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single overlay whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$. Inheritance sources do not create additional scope levels.
  • $\mathrm{Foo}.\mathbf{this}$ (equivalently, $\uparrow^{\!n}$ where $\mathrm{Foo}$ is $n$ scope levels above) retrieves the enclosing record named $\mathrm{Foo}$.
  • $\mathrm{Foo}.\mathbf{this}.\ell$ (equivalently, $\uparrow^{\!n}.\ell$) projects property $\ell$ from $\mathrm{Foo}$.
  • When $k=0$, the reference inherits the entire enclosing scope, which may produce an infinitely deep tree.
When the qualifier is unambiguous, it may be omitted: the shorthand $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ ($k\geq 1$) resolves $\ell_{\mathrm{down},1}$ against the innermost enclosing scope whose $\mathtt{defines}$ contains it, and desugars to the corresponding indexed form. Since this uses $\mathtt{defines}$, the shorthand can only access properties defined in record literals, not those inherited through inheritance; one of the two qualified forms is required for inherited properties or to bypass variable shadowing.
3. Overlay Trees
Given an AST, what are the properties visible at each position in the tree? This section answers the question by defining a collection of mutually recursive functions over paths.
[…]
Most of these functions have direct analogues in object-oriented languages: $\mathtt{Path}$ identifies a position in the tree, $\mathtt{init}$ and $\mathtt{last}$ navigate paths, $\mathtt{defines}$ extracts the locally defined members from the AST, $\mathtt{inherits}$ extracts the inheritance declarations from the AST, $\mathtt{properties}$ computes all visible members at a path including both locally defined and inherited ones, $\mathtt{supers}$ computes the transitive inheritance closure, and $\mathtt{bases}$ collects the direct base classes. One function has no object-oriented analogue: $\mathtt{overlays}$ collects the identity-sharing paths that arise when inheritance introduces multiple definitions of the same label at the same scope level. Two functions, the reference resolution function $\mathtt{resolve}$ and the qualified this resolution function $\mathtt{this}$, have object-oriented analogues whose semantics differ significantly; they are discussed in detail below.
All functions are pure and may be cached (memoised) without changing semantics.
Path
A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ell_{2},\ldots,\ell_{n})$ that identifies a position in the tree. [^1] Since paths are referentially transparent, an implementation may intern paths so that structural equality reduces to pointer equality. We write $p$ for a path. The root path is the empty sequence $()$. Given a path $p$ and a label $\ell$, $p\mathbin{\triangleright}\ell$ is the sequence $p$ extended with $\ell$. For a non-root path, the parent path $\mathtt{init}(p)$ is $p$ with its last element removed, and the final label $\mathtt{last}(p)$ is the last element of $p$. For convenience, we sometimes write $\ell$ for $\mathtt{last}(p)$ when the path is clear from context. With paths in hand, we can state what the AST provides at each path.
AST
An expression $e$ (Section 2) is parsed into an AST. The AST provides two primitive functions at each path $p$:
  • $\mathtt{defines}(p)$: the set of labels that have subtrees at $p$.
  • $\mathtt{inherits}(p)$: the set of reference pairs $(n,\;\ell_{\mathrm{down},*})$, where $n$ is the de Bruijn index (de Bruijn, 1972) (zero-based) and $\ell_{\mathrm{down},*}$ is the list of downward projections.
During parsing, all three syntactic forms are resolved to de Bruijn index pairs $(n,\;\ell_{\mathrm{down},*})$. The indexed form $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ already carries the de Bruijn index $n$ directly; no resolution is needed. A named reference $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the last occurrence of $\ell_{\mathrm{up}}$ among the labels of $p$. Let $p_{\mathrm{def}}$ be the prefix of $p$ up to and including that occurrence. The de Bruijn index is $n=|p|-|p_{\mathrm{def}}|$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. A lexical reference $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the nearest prefix $p^{\prime}$ of $p$ such that $\ell_{\mathrm{down},1}\in\mathtt{defines}(p^{\prime})$. The de Bruijn index is $n=|p|-|p^{\prime}|$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. All three forms produce the same representation; the semantic functions below operate only on de Bruijn index pairs.
Parsing populates $\mathtt{defines}$ and $\mathtt{inherits}$ from each element of a record. A definition $\ell\mapsto e$ contributes $\ell$ to $\mathtt{defines}(p)$. A reference contributes $(n,\;\ell_{\mathrm{down},*})$ to $\mathtt{inherits}(p)$. Since records are sets, the order of elements is irrelevant and duplicate elements have no effect. Both $\mathtt{defines}$ and $\mathtt{inherits}$ are pure data; they are not functions of runtime state. Given these two primitives, we can now answer the question posed at the beginning of this section.
Properties
The properties of a path are its own properties together with those inherited from all supers:
(1) $\mathtt{properties}(p)=\bigl\{\;\ell\;\big|\;(\_,\;p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\;\bigr\}$
The remainder of this section defines $\mathtt{supers}$, the transitive inheritance closure, and its dependencies.
Supers
Intuitively, $\mathtt{supers}(p)$ collects every path that $p$ inherits from: the identity-sharing paths $\mathtt{overlays}$ of $p$ itself, plus the $\mathtt{overlays}$ of each direct base $\mathtt{bases}$ of $p$, and so on transitively. Each result is paired with the inheritance-site context $\mathtt{init}(p_{\mathrm{base}})$ through which it is reached. This provenance is needed by qualified this resolution, i.e. the $\mathtt{this}$ function defined below, to map a definition-site overlay back to the inheritance-site paths that incorporate it:
(2) $\mathtt{supers}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{base}}),\;p_{\mathrm{overlay}})\;\big|\;p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{overlay}}\in\mathtt{overlays}(p_{\mathrm{base}})\;\bigr\}$
The $\mathtt{supers}$ formula depends on two functions: the identity-sharing paths $\mathtt{overlays}$ and the one-hop reference targets $\mathtt{bases}$. We define $\mathtt{overlays}$ first.
Overlays
Intuitively, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overlays}(p)$ collects all such paths that share the same identity as $p$. Concretely, the overlays of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
(3) $\mathtt{overlays}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &(\_,\;p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
It remains to define $\mathtt{bases}$, the other dependency of $\mathtt{supers}$.
Bases
Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references—analogous to the direct base classes in object-oriented languages. Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overlays}$ one step:
[…]
The $\mathtt{bases}$ formula calls the reference resolution function $\mathtt{resolve}$, which we define next.
Reference resolution
Intuitively, $\mathtt{resolve}$ turns a syntactic reference into the set of paths it points to in the fully inherited tree. It takes an inheritance-site path $p_{\mathrm{site}}$, a definition-site path $p_{\mathrm{def}}$, a de Bruijn index $n$, and downward projections $\ell_{\mathrm{down},*}$. Resolution proceeds in two phases: first, the qualified this resolution function $\mathtt{this}$ performs $n$ upward steps starting from the enclosing scope $\mathtt{init}(p_{\mathrm{def}})$, mapping the definition-site path to inheritance-site paths; then the downward projections $\ell_{\mathrm{down},*}$ are appended:
(5) $\mathtt{resolve}(p_{\mathrm{site}},\;p_{\mathrm{def}},\;n,\;\ell_{\mathrm{down},*})=\\ \bigl\{\;p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\ell_{\mathrm{down},k}\;\big|\;p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\;\mathtt{init}(p_{\mathrm{def}}),\;n)\;\bigr\}$
When multiple routes exist (due to multi-path inheritance), they may yield different target paths; all are collected. We now define $\mathtt{this}$.
Qualified this resolution
Intuitively, $\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” The design of $\mathtt{this}$ reconciles two historically opposing approaches to scope resolution. Early Lisp implementations used dynamic scope (McCarthy, 1978), where variable references are resolved against the run-time call stack; this is flexible but unpredictable. Scheme (Sussman and Steele Jr., 1975) corrected this with lexical scope, where a de Bruijn index resolves in a single step by indexing into the statically determined environment; this is predictable but rigid. Scala (Odersky et al., 2004) and the NixOS module system (NixOS Contributors, 2024) follow the lexical approach: each this reference is bound to a single statically known path. When inheritance introduces multiple inheritance routes to the same scope, Scala rejects the ambiguity at compile time (Appendix C) and the NixOS module system rejects the duplicate option declaration (Appendix D)—neither can express that all routes are equally valid.
Overlay-Calculus combines the strengths of both approaches. The reference itself is lexically determined at the definition site as a de Bruijn index, so what $\mathtt{this}$ looks up is predictable. But because inheritance interleaves scopes from different overlays, the result depends on the inherited tree structure: $\mathtt{this}$ walks upward one scope level at a time, consulting $\mathtt{supers}$ at each level, and naturally tracks a set of inheritance-site paths rather than a single one. All routes contribute equally to the inherited result, and since inheritance is idempotent, duplicated contributions are harmless. The only observable is property presence, making multi-path resolution well-defined without disambiguation.
Concretely, each step finds, among the supers of every path in the frontier $S$, those whose overlay component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier. After $n$ steps the frontier contains the answer:
(6) $\small\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a non-negative integer, the recursion terminates.
Alternative approaches considered
Our first three implementations of the Overlay language used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple inheritance routes to the same enclosing scope.
In retrospect, the difficulty was inevitable. A closure captures a single environment, so a fixed-point combinator solves for a single self-reference; when multiple routes exist, it attempts to unify them into one value. A stack-based environment is a linear chain—each scope level has one parent—so looking up a de Bruijn index returns a single value. Both data structures embed a single-valued assumption that is incompatible with the multi-path situation that arises when independent overlays inherit from the same enclosing scope (Appendix C). Equation (6) emerged from recognizing this: it operates on paths as first-order data and tracks a frontier set $S$ rather than a single current scope.
This completes the chain of definitions needed to compute $\mathtt{properties}(p)$.
Observational semantics without reduction
The definitions above constitute the complete semantics of Overlay-Calculus. There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments. This is a deliberate design choice, not an omission. In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur. In Overlay-Calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything—it is the value. The functions $\mathtt{properties}$, $\mathtt{supers}$, etc. determine what is observable at each path, and that is all. Computation arises not from rewriting terms but from querying an inductively deeper path in a lazily constructed tree: the observer drives the computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand. Formally, the mutually recursive equations (1)–(6) define set membership inductively: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these equations. The corresponding operator on the product lattice of path sets is monotone, and its least fixed point exists by the Knaster–Tarski theorem (Tarski, 1955). The detailed well-definedness argument—identifying the lattice, verifying positivity of each equation, and handling the transitive closure $\mathtt{bases}^{*}$—is given in Appendix E. When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point—analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness. This observation model is inspired by F-coalgebras (Rutten, 2000), where identity is determined by observable behavior rather than by internal structure.
4. Translation from $\lambda$-Calculus
The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to Overlay-Calculus. In ANF, every intermediate result is bound to a name and each expression contains at most one application:
[…]
Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results, and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus. The ANF structure matches Overlay-Calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
Let $\mathcal{T}$ denote the translation function. Each $\lambda$-abstraction introduces one scope level in the translated overlay tree.
$x$ ($\lambda$-bound, index $n$)   $\longrightarrow$ $\uparrow^{\!n}.\mathrm{argument}$
$x$ (let-bound)   $\longrightarrow$ $x.\mathrm{result}$
$\lambda x.\,M$   $\longrightarrow$ $\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$   $\longrightarrow$ $\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$V_{1}\;V_{2}$ (tail call)   $\longrightarrow$ $\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
A $\lambda$-bound variable $x$ with de Bruijn index $n$ (the number of enclosing $\lambda$-abstractions between the reference and its binder) becomes $\uparrow^{\!n}.\mathrm{argument}$. The indexed reference $\uparrow^{\!n}$ traverses $n$ scope levels upward to reach the abstraction that binds $x$, giving access to the $\mathrm{argument}$ slot after all inheritance. A let-bound variable $x$ becomes the lexical reference $x.\mathrm{result}$, projecting the application result from the sibling property $x$ in the enclosing record.
An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$; we call this the abstraction shape. In the $\mathbf{let}$-binding rule, the name $x$ binds to the inherited record $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$, and $\mathcal{T}(M)$ may reference $x.\mathrm{result}$ to obtain the value of the application. In the tail-call rule, the fresh label $\mathrm{tail}$ serves the same purpose, and $\mathrm{result}$ projects the answer. In both cases the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to Overlay-Calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is Overlay-Calculus; Appendix A further shows that Overlay-Calculus primitives directly encode a random-access machine.
An elementary semantics of the $\lambda$-calculus
Composing the translation $\mathcal{T}$ with the semantic functions of Section 3 induces a semantics on the $\lambda$-calculus itself. Given a $\lambda$-term $M$, one translates it to $\mathcal{T}(M)$ and then determines convergence behavior by querying $\ell\in\mathtt{properties}(p)$ via equations (1)–(6), following the $\mathrm{result}$ chain until the abstraction shape appears or the fixed point yields the empty set.
This induced semantics is more elementary than traditional approaches in the mathematical machinery it requires:
  • No function spaces. Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices. Here, the semantic domain is a product of powerset lattices—no function spaces, no continuity conditions, no inverse limits. The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix E), which requires only set-theoretic monotonicity—strictly weaker than Scott continuity.
  • No substitution. Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi. Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree—no substitution occurs at any point.
  • No reduction rules. There are no $\beta$-reduction steps, no evaluation contexts, and no choice between call-by-name and call-by-value (Plotkin, 1975). The semantics is observational: one queries path membership, and the mutually recursive equations unfold on demand.
That a semantics of the $\lambda$-calculus can be given without function spaces is perhaps unexpected: function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function. Section 4.2 shows that the convergence behavior of the translated overlay tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate (a term converges iff it has a head normal form) and fully abstract (two terms are overlay-contextually equivalent iff they have the same Böhm tree).
Combining these two facts—elementary machinery and full abstraction—yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics that lives entirely in naive set theory. The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem. No continuous lattices, no inverse limits, no Scott topology, no game-theoretic machinery. Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Hyland and Ong’s game semantics (Hyland and Ong, 2000) achieves full abstraction but at the cost of even greater mathematical complexity. The overlay tree semantics achieves full abstraction with strictly simpler foundations than either.
The reverse direction is non-trivial, as we discuss next.
4.1. Expressive Asymmetry
The translation above consists of five rules; the full computational power of the $\lambda$-calculus embeds into Overlay-Calculus with no additional machinery.
The reverse direction—encoding Overlay-Calculus in pure $\lambda$-calculus—is non-trivial. Defining the meaning of Overlay-Calculus requires machinery for lazy allocation, recursive data structures, and set-valued resolution that have no direct counterparts in the $\lambda$-calculus. This sophistication is not accidental—it reflects three features:
  1. (1)
    Self-referential records with lazy evaluation. An overlay’s properties can reference other properties and enclosing scopes through qualified this without explicit binding. In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
  2. (2)
    Open recursion with symmetric inheritance. Inheritance merges definitions from independent sources, all sharing the same inherited overlay tree. Encoding this in the $\lambda$-calculus requires maintaining an extensible dictionary of methods that can be merged from multiple directions. Cook’s denotational semantics of inheritance (Cook, 1989) devoted an entire dissertation to formalizing this pattern in a functional setting.
  3. (3)
    Commutative, idempotent merge. Inheritance in Overlay-Calculus is symmetric: a record’s elements form a set, so reordering them produces the same observable result. In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery.
Overlay-Calculus is not the $\lambda$-calculus in disguise: it provides primitive operations for computational patterns that are derivable but complex in the $\lambda$-calculus.
The reverse asymmetry also exists: nested expressions such as f(gx)f\;(g\;x) are primitive in the $\lambda$-calculus but require manual A-normal form conversion in Overlay-Calculus—the intermediate result gxg\;x must be given an explicit name. However, ANF conversion is a trivial, mechanical transformation (Flanagan et al., 1993); a surface language built on Overlay-Calculus can support nested expressions as syntactic sugar without extending the calculus itself.
4.2. Böhm Tree Correspondence
The translation $\mathcal{T}$ induces a semantics on the $\lambda$-calculus via the overlay tree equations (1)–(6). We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
An overlay tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties—the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$. We write $T{\Uparrow}$ if no such depth exists.
Definition 4.1 (Overlay-convergence).
An overlay tree $T$ converges, written $T{\Downarrow}$, if there exists $n\geq 0$ such that:
$\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{properties}\!\bigl(\,\underbrace{\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
in the least fixed point of Section 3.
Theorem 4.2 (Adequacy).
A closed ANF $\lambda$-term $M$ has a head normal form if and only if $\mathcal{T}(M){\Downarrow}$.
Theorem 4.3 (Full Abstraction).
For closed $\lambda$-terms $M$ and $N$, define $\mathcal{T}(M)\approx\mathcal{T}(N)$ iff for every closing context $C[\cdot]$, $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$. Then:
$\mathcal{T}(M)\approx\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
Together, Adequacy and Full Abstraction establish that the overlay tree semantics is a computation mechanism for Böhm trees (Barendregt, 1984)—the canonical fully abstract model of the lazy $\lambda$-calculus—using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory. This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract—it identifies terms that observational equivalence distinguishes (Plotkin, 1975). The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix F.
5. Case Study: The Expression Problem
The Expression Problem (Wadler, 1998) asks how to extend a data type with both new cases and new operations, without modifying existing code. Existing solutions require object algebras (d. S. Oliveira and Cook, 2012; Wang and d. S. Oliveira, 2016), finally tagless interpreters (Carette et al., 2009), or type-class machinery. In a typed setting the problem additionally requires static type safety; Overlay-Calculus is untyped, so we focus on the extensibility aspect. Suppose a standard library provides a $\mathrm{binnat}$ module for binary natural arithmetic (see Appendix B). We define an expression language with an evaluation operation:
$\{\quad\mathrm{expression}$ $\mapsto\{\mathrm{Constant}\mapsto\{\mathrm{magnitude}\mapsto\{\}\},\;\mathrm{Addition}\mapsto\{\mathrm{left}\mapsto\{\},\;\mathrm{right}\mapsto\{\}\}\},$
$\mathrm{evaluation}$ $\mapsto\{\mathrm{Constant}\mapsto\{\mathrm{expression}.\mathrm{Constant},\;\mathrm{outcome}\mapsto\mathrm{Constant}.\mathbf{this}.\mathrm{magnitude}\},$
$\qquad\mathrm{Addition}\mapsto\{\mathrm{expression}.\mathrm{Addition},\;$
$\qquad\qquad\mathrm{left\_outcome}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{left}.\mathrm{outcome},$
$\qquad\qquad\mathrm{right\_outcome}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{right}.\mathrm{outcome},$
$\qquad\qquad\mathrm{sum}\mapsto\{\mathrm{binnat}.\mathrm{Addition},\;\mathrm{augend}\mapsto\mathrm{sum}.\mathbf{this}.\mathrm{left\_outcome},\;\mathrm{addend}\mapsto\mathrm{sum}.\mathbf{this}.\mathrm{right\_outcome}\},$
$\qquad\qquad\mathrm{outcome}\mapsto\mathrm{sum}.\mathrm{outcome}\}\},$
$\mathrm{binnat}$ $\mapsto\{\mathrm{Addition}\mapsto\{\mathrm{augend}\mapsto\{\},\;$
$\qquad\qquad\mathrm{addend}\mapsto\{\},\;\mathrm{outcome}\mapsto\{\}\}\}\quad\}$
A new operation (e.g., display) is added by defining a new module that inherits from the same $\mathrm{expression}$ schemas:
$\mathrm{display}\mapsto\{\mathrm{Constant}$ $\mapsto\{\mathrm{expression}.\mathrm{Constant},\;\mathrm{representation}\mapsto\mathrm{Constant}.\mathbf{this}.\mathrm{magnitude}\},$
$\mathrm{Addition}$ $\mapsto\{\mathrm{expression}.\mathrm{Addition},\;$
$\qquad\mathrm{representation}\mapsto\{$
$\qquad\qquad\mathrm{left}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{left}.\mathrm{representation},\;$
$\qquad\qquad\mathrm{right}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{right}.\mathrm{representation}\}\}\}$
A new case (e.g., negation) is added with its operation handlers and an empty API for the arithmetic primitive:
$\mathrm{negation}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{operand}\mapsto\{\}\}\},$
$\mathrm{negation\_evaluation}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{negation}.\mathrm{Negation},\;$
$\qquad\mathrm{operand\_outcome}\mapsto\mathrm{Negation}.\mathbf{this}.\mathrm{operand}.\mathrm{outcome},$
$\qquad\mathrm{negated}\mapsto\{\mathrm{binnat}.\mathrm{Negation},\;\mathrm{operand}\mapsto\mathrm{operand\_outcome}\},$
$\qquad\mathrm{outcome}\mapsto\mathrm{negated}.\mathrm{outcome}\}\},$
$\mathrm{binnat}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{operand}\mapsto\{\},\;\mathrm{outcome}\mapsto\{\}\}\}$
All modules compose freely; none requires modification of existing definitions. A complete executable version of this example—including both the evaluation and display operations, a negation extension, and their free inheritance—is available in the implementation (Yang, 2025b), along with test cases verifying evaluation and display for nested expressions mixing old and new cases.
Note that $\mathrm{magnitude}\mapsto\{\}$ in $\mathrm{Constant}$ resembles a type annotation but is purely structural in Overlay-Calculus: the evaluator works as long as the required properties appear somewhere in the inheritance chain, regardless of any schema declarations.
6. Discussion
6.1. Relationship to $\lambda$-Calculus
Three facts establish the relationship between Overlay-Calculus and the $\lambda$-calculus.
More elementary and stronger than Scott domains.
The semantics induced by the translation (Section 4) requires only powerset lattices and the Knaster–Tarski fixed-point theorem—no function spaces, no inverse limits, no Scott continuity. Yet this semantics is fully abstract (Theorem 4.3), whereas Scott’s $D_{\infty}$ model (Scott, 1976; Wadsworth, 1976) is adequate but not fully abstract—it identifies terms that observational equivalence distinguishes (Plotkin, 1975).
$\lambda$-Calculus embeds into Overlay-Calculus.
The five-rule translation of Section 4 maps every ANF $\lambda$-term to an overlay tree. The translation uses only single-target qualified this—the full multi-target semantics is not needed to embed $\lambda$-calculus.
Multi-target qualified this is the natural primitive.
If the semantic functions of Section 3 were restricted to single-target resolution (as in DOT’s single-valued self variable (Amin et al., 2016)), the formulas would need to reject various structurally legitimate cases—adding complexity, not removing it. Multi-target qualified this is not an artificial extension; it is what falls out when no restrictions are imposed. The $\lambda$-calculus translation uses only single-target qualified this, suggesting that $\lambda$-calculus itself represents an artificially restricted fragment of the natural model. Appendix H presents a detailed analysis of the asymmetry in problem difficulty between the two systems.
6.2. Observed Phenomena
Several properties emerge in practice without having been designed:
  • The Expression Problem does not arise (Section 5, Appendix H).
  • Programs are CPS-agnostic: empty record slots serve as dependency injection points, making the same program text CPS or direct-style depending on assembly (Appendix G).
  • Records are tries, and inheritance is trie union—an immutable analogue of random-access memory (Appendix A).
  • Unlike SKI combinators (Schönfinkel, 1924; Curry and Feys, 1958), which are isomorphic to $\lambda$-calculus in problem difficulty, Overlay-Calculus exhibits genuine asymmetry (Section 4.1).
6.3. Properties of Inheritance
Because records are sets, inheritance is inherently commutative and idempotent, making Overlay-Calculus transparent to linearization order. Deduplication of inherited overlays is purely an optimization.
6.4. Relationship to Practical Systems
Several classes of declarative systems that have emerged independently in practice turn out to be instances of Overlay-Calculus.
Configuration and object-oriented languages
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024)—whose recursive attribute merging motivated this work—can be understood as an implementation of Overlay-Calculus with additional features (type checking, error reporting, FFI to the Nix language). Detailed comparisons with CUE, Jsonnet, Dhall, with mixin and trait calculi, and with the NixOS module system’s lazy evaluation limitations under multiple inheritance routes are given in Section 7.
Effect systems
Algebraic effects with handlers (Plotkin and Power, 2003; Pretnar, 2015) and monad transformers structure computational effects through composition. The CPS-agnostic property (Appendix G) shows how Overlay-Calculus naturally supports effect handler composition: empty API slots can be filled with effect handlers (CPS interpretation) or direct implementations (direct-style interpretation). The same pattern appears in algebraic effect systems, without explicit monad machinery.
Modular software architectures
Plugin systems, component frameworks, and dependency injection frameworks enable software extensibility through declarative inheritance. The symmetric, associative nature of inheritance (Section 3) explains why such systems work: components can be inherited in any order without changing behavior. These systems implement Overlay-Calculus patterns ad hoc.
Union file systems
Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view. Their semantics resembles Overlay-Calculus—later layers override earlier layers, files from all layers remain accessible—but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers. We implemented Overlay-Calculus directly as a union file system (Yang, 2025a), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives. In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance—a file-system-as-compiler in which the directory tree is both the source program and the compilation target.
7. Related Work
Denotational semantics of inheritance
Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records and inheritance as composition of generators (functions from self to complete object) and wrappers (functions that modify generators), solved by a fixed-point construction. Cook’s key insight—that inheritance is a general mechanism applicable to any form of recursive definition, not only object-oriented methods—is one of the starting points of the present work. The model has four primitives (records, functions, generators, wrappers) and requires a fixed-point combinator; composition is asymmetric (a wrapper modifies a generator, not vice versa). Whether the same insight can be captured with fewer primitives and symmetric inheritance was a natural question.
Mixin-based inheritance
Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses—functions from a superclass parameter to a subclass—unifying the inheritance mechanisms of Smalltalk, Beta, and CLOS under a single model. Because mixin application is function composition, it is neither commutative nor idempotent: applying the same mixin twice may produce different results, and applying two mixins in different orders may yield different linearizations. The C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order; whether linearization is inherent to inheritance or an artifact of non-commutativity was worth examining. Mixin composition operates at the method level: applying a mixin replaces or extends individual methods in a flat namespace. There is no recursive merging of nested structure—composing two mixins that define the same method name results in an override, not a union mount. Finally, mixins presuppose the $\lambda$-calculus as their computational foundation: method bodies are functions, and mixin application is function composition. Mixins organize and compose functions but cannot replace them—without the underlying $\lambda$-calculus, there is nothing to compose. The mixin mechanism itself is not Turing complete.
Traits
Schärli et al. (Schärli et al., 2003) introduced traits as composable units of behavior. Trait composition is symmetric and commutative: the sum of two traits is order-independent, which is a significant improvement over mixins. However, when two traits provide methods with the same name, a conflict arises and must be explicitly resolved by the programmer through exclusion or aliasing. Ducasse et al. (Ducasse et al., 2006) extended the model with a flattening property: the semantics of a class is independent of whether its methods originate from traits or are defined directly. Like mixins, trait composition is flat: it operates on individual method names, not on nested structures, so same-name definitions are conflicts rather than opportunities for deeper merging. And like mixins, traits presuppose the $\lambda$-calculus: trait methods are functions, and the computational power resides in those functions, not in the trait mechanism itself.
Family polymorphism and virtual classes
Ernst (Ernst, 2001) introduced family polymorphism, where families of mutually dependent classes can be refined together in subclasses. Ernst, Ostermann, and Cook (Ernst et al., 2006) formalized this in the virtual class calculus, where classes are class-valued attributes of objects, accessed via path expressions of the form this.out.C. The path this.out navigates to an enclosing scope, analogous to qualified this ($\ell_{\mathrm{up}}.\mathbf{this}$) in this paper. A notable restriction is that each such path resolves to a single class in a single enclosing object. When multiple inheritance routes lead to the same scope—a common situation under inheritance—single-valued resolution must choose one route and discard the others, losing information. Whether qualified this can resolve to a set of paths, letting all routes contribute equally, had not been explored.
Object calculi
Abadi and Cardelli (Abadi and Cardelli, 1996) developed a comprehensive theory of objects, treating objects—rather than classes—as the primitive notion. Their calculi include method update and self-referential objects, but methods are fundamentally functions (taking self as a parameter), and object extension is asymmetric (method override replaces the previous definition). Boudol (Boudol, 2004) studied the recursive record semantics of objects in a call-by-value setting, showing that self-referential records require an unsafe fixed-point operator. Whether self-reference can be handled without a fixed-point operator in the object language—for instance, through coalgebraic observation rather than reduction—was a question these calculi left open.
Record calculi
Harper and Pierce (Harper and Pierce, 1991) presented a record calculus based on symmetric concatenation with a type system that tracks the absence of fields. When two records define the same label, concatenation is undefined (a type error); the calculus avoids same-label composition rather than giving it meaning. Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) gave a type system for records and variants in ML. These calculi operate on flat records (no nesting) and treat records as data (initial algebras): a record is a finite map from labels to values, fully determined at construction time. Nested records, self-reference, and lazy observation—the properties needed for records to serve as a computational substrate—are outside their scope.
DOT calculus
Amin et al. (Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala’s type system. DOT paths navigate through objects to reach nested type members, providing a rich dependent type structure. DOT’s self variable in the object constructor $\nu(x:T)\,d$ binds $x$ to a single object; when multiple inheritance routes lead to the same enclosing scope, the resulting intersection type is unrealizable and Scala rejects it at compile time (Appendix C). DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure overlay inheritance (see Section 8).
Configuration languages
CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification (&) is commutative, associative, and idempotent. CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification. CUE’s lattice includes scalar types with a conflict semantics (unifying incompatible scalars yields $\bot$); whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question. Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative (the right-hand side wins on scalar conflicts), and deep merging requires explicit opt-in (+: syntax) on a per-field basis—so union mount semantics is available but not the default. Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
Module systems and union mount semantics
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override—the same union mount semantics that distinguishes tree-level inheritance from mixin and trait composition. When two modules define the same nested path, their subtrees are merged rather than overridden. The NixOS module system got nearly all of the high-level design right: union mount inheritance, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top. Its practical success is remarkable—the same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity. The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance. When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix D). The module system cannot express that both routes contribute equally to the same option—the multi-path situation that recursive merging naturally gives rise to—which motivated the coalgebraic semantics of Section 3.
SKI combinator calculus
Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus can be reformulated without bound variables using three combinators (S, K, I). SKI is function-free at the object level and Turing complete. However, bracket abstraction translates between SKI and the $\lambda$-calculus preserving computational structure, so the two systems are isomorphic in terms of problem difficulty (Section 4): what is hard in one remains hard in the other. Whether a function-free calculus can exhibit a genuine shift in problem difficulty—making some problems structurally easier than in the $\lambda$-calculus, rather than merely encoding them differently—is the question that motivates the comparison in Section 4.
8. Future Work
The Overlay language (Yang, 2025b) is an executable implementation of Overlay-Calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the inheritance, and a foreign-function interface (FFI) that introduces scalar values from the host language. The future work described below concerns the distance between this implementation and a fully practical language—spanning both theoretical foundations and library-level encodings.
Type system
The Overlay language’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result. Formalizing the soundness of these checks—proving that well-typed programs do not produce dangling references at runtime—is one direction of future work. Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how Overlay-Calculus relates to the $\lambda$-calculus: a Typed Overlay-Calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with overlay inheritance. A formal type system would be amenable to mechanization in proof assistants (Coq, Agda) for verifying type safety, progress, and preservation properties.
Beyond soundness of the existing checks, additional type system features would strengthen the language. A key example is totality checking: verifying that a inheritance provides implementations for all required slots, not merely that references resolve. In untyped Overlay-Calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary—it defines a structural slot but imposes no constraint on what must fill it. Totality checking would enforce such declarations statically, rejecting inheritances that leave required slots unfilled.
Standard library
The remaining items concern the standard library: they require no extensions to the calculus or type system, but involve nontrivial encodings within the existing framework.
Overlay-Calculus is open by default: inheritance can freely merge any two overlays, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$). This is not a defect—it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 4.1). However, many operations (e.g., equality testing) assume that values are linear—inhabiting exactly one constructor. The observer pattern (Appendix B) does not enforce this: applied to a multi-constructor value, all callbacks fire and their results are inherited.
Schema validation can be implemented at the library level using existing primitives. An observer can test whether a particular constructor is present, returning a Boolean; a second Boolean dispatch then branches on the result. By chaining such tests, a factory can validate that a value matches exactly one constructor and carries the required fields. This does not restrict inheritance itself—it remains unconstrained—but provides a way to detect invariant violations before they propagate.
The same technique—observer returning Boolean, followed by Boolean dispatch—yields closed pattern matching. The observer pattern of Appendix B is inherently open: new constructor cases can be added through inheritance. Closed matching, where exactly one branch is taken, can be encoded as a visitor chain: a linked list of if-then-else nodes, each testing one constructor and falling through on mismatch. This is analogous to GHC.Generics’ $\mathrm{(:+:)}$ sum representation (Magalhães et al., 2010), where each link corresponds to one summand. An important consequence is that such chains do not commute—they have a fixed order—and therefore cannot be extended through open inheritance, which is the correct semantics for closed dispatch.
The trie operations in Appendix A are hardcoded for binary natural number keys. Since every overlay is a trie (Appendix A) and insertion is simply inheritance, only operations that cannot be expressed as inheritance need explicit implementation: deletion, dynamic lookup, and prefix lookup. Generalizing to arbitrary key types requires each key factory to provide two interfaces—SelectTrieChild (select a subtrie by constructor) and ReplaceTrieChild (replace a subtrie by constructor while preserving the rest)—corresponding to the getter and setter of a lens (Foster et al., 2007) focused on the constructor-determined child. Prefix lookup is structurally analogous to equality testing: both synchronously traverse a key (a single-path trie) alongside another structure, dispatching at each level. The difference is the return type—a Boolean for equality, a subtrie for prefix lookup.
Appendix A Trie Operations
[…]
$\mathrm{Zero}$ $\;\longrightarrow\;\mathrm{value}$
$\mathrm{Odd}(n)$ $\;\longrightarrow\;\mathrm{at\_odd},\;\text{then path for }n$
$\mathrm{Even}(n)$ $\;\longrightarrow\;\mathrm{at\_even},\;\text{then path for }n$
Write (insertion)
Inserting an entry with value $v$ at key $k$ is inheriting a singleton trie whose nesting mirrors the bit structure of $k$. For example, inserting at key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$:
$\mathrm{trie}^{\prime}=\{\mathrm{trie},\;\mathrm{at\_odd}\mapsto\{\mathrm{at\_even}\mapsto\{\mathrm{value}\mapsto v\}\}\}$
Each inheritance creates a new immutable trie; the original is unchanged.
Read (static lookup)
When the key is known statically, retrieving the entry at key $k$ is reference-path navigation: the reference follows the same label sequence that was used for insertion. For the same key $5$:
$\mathrm{enclosing\_scope}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}.\mathrm{at\_even}.\mathrm{value}$
where $\mathrm{enclosing\_scope}$ is the qualified-this scope name containing the trie definition. The reference mechanism traverses the nested record structure—which is the trie—returning the value at the final node.
Read (dynamic lookup)
When the key is a BinNat value constructed at runtime via $\mathrm{Zero}$, $\mathrm{Odd}$, and $\mathrm{Even}$, the lookup path must be computed dynamically. This is achieved by folding the BinNat key into a Selector—an overlay that, given a trie, navigates to the appropriate node and returns the stored entry. A Selector declares the trie shape and a result slot:
$\mathrm{Selector}\mapsto\{\mathrm{trie}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\},\;\mathrm{result}\mapsto\{\}\}$
Three observer callbacks build Selectors bottom-up. These callbacks ($\mathrm{on\_zero}$, $\mathrm{on\_odd}$, $\mathrm{on\_even}$) are interpretation points in the sense of Appendix G: they can be filled externally by an orchestrator (CPS interpretation) or inherited locally (direct interpretation), demonstrating the CPS-agnostic property. The $\mathrm{on\_zero}$ callback returns a Selector that reads $\mathrm{value}$ from the trie:
$\mathrm{lookup\_on\_zero}\mapsto\{\mathrm{Selector},\;\mathrm{result}\mapsto\mathrm{lookup\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{value}\}$
The $\mathrm{on\_odd}$ callback takes a Selector (via the $\mathrm{argument}/\mathrm{result}$ encoding of Section 4) and returns a new Selector that navigates to $\mathrm{at\_odd}$, then delegates to the argument Selector:
$\mathrm{lookup\_on\_odd}$ $\mapsto\{\mathrm{argument}\mapsto\mathrm{Selector},\;\mathrm{result}\mapsto\{\mathrm{Selector},\;$
$\qquad\mathrm{applied}\mapsto\{\mathrm{argument},\;\mathrm{trie}\mapsto\mathrm{lookup\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\}\}$
The $\mathrm{on\_even}$ callback is symmetric, navigating to $\mathrm{at\_even}$ instead.
The $\mathrm{Lookup}$ combinator folds a BinNat key with these callbacks to obtain a Selector, then applies it to the trie:
$\mathrm{Lookup}$ $\mapsto\{$
$\quad\mathrm{key}\mapsto\mathrm{trie\_module}.\mathbf{this}.\mathrm{BinNat},$
$\quad\mathrm{trie\_input}\mapsto\{\},$
$\quad\mathrm{applied\_key}\mapsto\{\mathrm{key},\;$
$\qquad\quad\mathrm{on\_zero}\mapsto\mathrm{lookup\_on\_zero},\;$
$\qquad\quad\mathrm{on\_odd}\mapsto\mathrm{lookup\_on\_odd},\;$
$\qquad\quad\mathrm{on\_even}\mapsto\mathrm{lookup\_on\_even}\},$
$\quad\mathrm{applied\_selector}\mapsto\{\mathrm{applied\_key}.\mathrm{result},\;$
$\qquad\quad\mathrm{trie}\mapsto\mathrm{trie\_input}\},$
$\quad\mathrm{result}\mapsto\mathrm{applied\_selector}.\mathrm{result}\}$
For example, looking up key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$ folds into a Selector that navigates $\mathrm{at\_odd}.\mathrm{at\_even}.\mathrm{value}$—the same path as the static reference, but computed from the key value. The entire construction uses only records (with inheritance sources) and references; no additional primitives are needed.
Delete
Since inheritance is additive—it can only merge properties, never remove them—deletion cannot be expressed as a single inheritance. Instead, deletion reconstructs the trie, copying every subtrie except at the targeted key.
We fold the BinNat key into a Rebuilder—an overlay that, given an input trie, produces a new trie as its result. A Rebuilder declares the trie shape on both its input and output:
$\mathrm{Rebuilder}\mapsto\{$ $\mathrm{trie}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\},\;$
$\mathrm{result}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\}\}$
Three observer callbacks build Rebuilders bottom-up. The $\mathrm{on\_zero}$ callback returns a Rebuilder that copies the subtries but omits the value:
$\mathrm{delete\_on\_zero}\mapsto\{$ $\mathrm{Rebuilder},\;\mathrm{result}\mapsto\{$
$\quad\mathrm{at\_odd}\mapsto\mathrm{delete\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd},\;$
$\quad\mathrm{at\_even}\mapsto\mathrm{delete\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_even}\}\}\}$
The $\mathrm{result}.\mathrm{value}$ slot inherits only the empty declaration from $\mathrm{Rebuilder}$, effectively deleting the entry.
The $\mathrm{on\_odd}$ callback takes a Rebuilder (the recursive result for the subtrie) and returns a new Rebuilder that preserves $\mathrm{value}$ and $\mathrm{at\_even}$, but replaces $\mathrm{at\_odd}$ with the recursively rebuilt subtrie:
$\mathrm{delete\_on\_odd}$ $\mapsto\{\mathrm{argument}\mapsto\mathrm{Rebuilder},\;$
$\quad\mathrm{result}\mapsto\{\mathrm{Rebuilder},\;$
$\qquad\mathrm{applied}\mapsto\{\mathrm{argument},\;$
$\qquad\quad\mathrm{trie}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}\},$
$\qquad\mathrm{result}\mapsto\{$
$\qquad\quad\mathrm{value}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{value},\;$
$\qquad\quad\mathrm{at\_odd}\mapsto\mathrm{applied}.\mathrm{result},\;$
$\qquad\quad\mathrm{at\_even}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_even}\}\}\}$
The $\mathrm{on\_even}$ callback is symmetric.
The $\mathrm{Delete}$ combinator folds a BinNat key with these callbacks to obtain a Rebuilder, then applies it to the trie:
$\mathrm{Delete}$ $\mapsto\{$
$\quad\mathrm{key}\mapsto\mathrm{trie\_module}.\mathbf{this}.\mathrm{BinNat},$
$\quad\mathrm{trie\_input}\mapsto\{\},$
$\quad\mathrm{applied\_key}\mapsto\{\mathrm{key},\;$
$\qquad\quad\mathrm{on\_zero}\mapsto\mathrm{delete\_on\_zero},\;$
$\qquad\quad\mathrm{on\_odd}\mapsto\mathrm{delete\_on\_odd},\;$
$\qquad\quad\mathrm{on\_even}\mapsto\mathrm{delete\_on\_even}\},$
$\quad\mathrm{applied\_rebuilder}\mapsto\{\mathrm{applied\_key}.\mathrm{result},\;$
$\qquad\quad\mathrm{trie}\mapsto\mathrm{trie\_input}\},$
$\quad\mathrm{result}\mapsto\mathrm{applied\_rebuilder}.\mathrm{result}\}$
Since inheritance is idempotent, the empty scaffolding left by the Rebuilder schema merges harmlessly with any subsequent insertions.
Update
Updating the entry at key $k$ to a new value $v$ combines deletion and insertion. Because $\mathrm{Delete}$ is a combinator whose output lives in the result property, one must first project out the rebuilt trie before inheriting the new singleton. In ANF style:
$\mathrm{deleted}$ $\mapsto\{\mathrm{Delete},\;\mathrm{key}\mapsto k,\;\mathrm{trie\_input}\mapsto\mathrm{trie}\}$
$\mathrm{result}$ $\mapsto\{\mathrm{deleted}.\mathrm{result},\;\mathrm{at\_odd}\mapsto\{\mathrm{at\_even}\mapsto\{\mathrm{value}\mapsto v\}\}\}$
where the second line uses key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$ as a concrete example. The projection $\mathrm{result}$ is essential: $\mathrm{Delete}$ carries internal scaffolding ($\mathrm{key}$, $\mathrm{trie\_input}$, $\mathrm{applied\_key}$, etc.), and only its $\mathrm{result}$ property holds the rebuilt trie. This intermediate projection prevents Update from being a single symmetric inheritance.
Appendix B Church Encoding of Standard Types
This appendix demonstrates that standard data types can be encoded in Overlay-Calculus without scalar types, using Church encoding. These encodings serve as a proof of concept that a standard library can be built entirely within Overlay-Calculus. Each encoding is presented as a module overlay. In practical implementations, more efficient representations (e.g., binary naturals) or FFI to host language primitives would typically be used. The implementation (Yang, 2025b) includes executable implementations of these encodings with test cases for $\mathrm{Not}(\mathrm{True})$, $\mathrm{And}(\mathrm{True},\mathrm{True})$, $\mathrm{Or}(\mathrm{False},\mathrm{True})$, and natural number addition, all verified against expected output snapshots.
B.1. Boolean Module
The boolean module contains a schema ($\mathrm{Boolean}$), constructors ($\mathrm{True}$, $\mathrm{False}$), and operations ($\mathrm{Not}$, $\mathrm{And}$, $\mathrm{Or}$). All definitions are siblings within a single module overlay.
$\{\quad\mathrm{boolean}\mapsto\{\quad\mathrm{Boolean}$ $\mapsto\{\mathrm{on\_true}\mapsto\{\},\;\mathrm{on\_false}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},$
$\mathrm{True}$ $\mapsto\{\mathrm{Boolean},\;\mathrm{result}\mapsto\mathrm{True}.\mathbf{this}.\mathrm{on\_true}\},$
$\mathrm{False}$ $\mapsto\{\mathrm{Boolean},\;\mathrm{result}\mapsto\mathrm{False}.\mathbf{this}.\mathrm{on\_false}\},$
$\mathrm{Not}$ $\mapsto\{\mathrm{Boolean},$
$\qquad\mathrm{operand}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{applied\_operand}\mapsto\{\mathrm{operand},$
$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{Not}.\mathbf{this}.\mathrm{on\_false},$
$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{Not}.\mathbf{this}.\mathrm{on\_true}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_operand}.\mathrm{result}\},$
$\mathrm{And}$ $\mapsto\{\mathrm{Boolean},$
$\qquad\mathrm{left}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{right}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{applied\_left}\mapsto\{\mathrm{left},$
$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{right},$
$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{False}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_left}.\mathrm{result}\},$
$\mathrm{Or}$ $\mapsto\{\mathrm{Boolean},$
$\qquad\mathrm{left}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{right}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{applied\_left}\mapsto\{\mathrm{left},$
$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{True},$
$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{right}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_left}.\mathrm{result}\}\quad\}\quad\}$
Reference explanation
Within the $\mathrm{boolean}$ module, own properties such as $\mathrm{Boolean}$, $\mathrm{True}$, $\mathrm{False}$, $\mathrm{Not}$, $\mathrm{And}$, and $\mathrm{Or}$ are accessed via lexical references (e.g., the inheritance source $\mathrm{Boolean}$ in each definition). Qualified this is required in two situations:
  • Inherited properties. Inside True’s definition, $\mathrm{True}.\mathbf{this}.\mathrm{on\_true}$ requires qualified this because $\mathrm{on\_true}$ is inherited from Boolean via inheritance, not defined as an own property. Similarly, $\mathrm{Not}.\mathbf{this}.\mathrm{on\_false}$ and $\mathrm{Not}.\mathbf{this}.\mathrm{on\_true}$ access inherited callbacks.
  • Variable shadowing. Inside Not’s $\mathrm{applied\_operand}$, the labels $\mathrm{on\_true}$ and $\mathrm{on\_false}$ are own properties of $\mathrm{applied\_operand}$ (with swapped values), so a lexical reference $\mathrm{on\_false}$ would resolve to the local definition rather than Not’s inherited callback. Qualified this $\mathrm{Not}.\mathbf{this}.\mathrm{on\_false}$ bypasses the shadowing.
Worked example
Consider the expression $\{\mathrm{boolean}.\mathrm{Not},\;\mathrm{operand}\mapsto\mathrm{boolean}.\mathrm{True},\;\mathrm{on\_true}\mapsto A,\;\mathrm{on\_false}\mapsto B\}$ at the root scope level, where $A$ and $B$ are arbitrary expressions. The result is an object whose $\mathrm{result}$ label contains the same observable tree as $B$, because $\mathrm{Not}$ swaps $\mathrm{on\_true}$ and $\mathrm{on\_false}$, and $\mathrm{True}$ selects $\mathrm{on\_true}$, which after swapping becomes $B$.
B.2. Natural Number Module
The natural number module contains a schema ($\mathrm{Nat}$), constructors ($\mathrm{Zero}$, $\mathrm{Succ}$), and an operation ($\mathrm{Add}$).
$\{\quad\mathrm{natural}\mapsto\{\quad\mathrm{Nat}$ $\mapsto\{\mathrm{successor}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},\;\mathrm{zero}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},$
$\mathrm{Zero}$ $\mapsto\{\mathrm{Nat},\;\mathrm{result}\mapsto\mathrm{Zero}.\mathbf{this}.\mathrm{zero}\},$
$\mathrm{Succ}$ $\mapsto\{\mathrm{Nat},$
$\qquad\mathrm{predecessor}\mapsto\mathrm{Nat},$
$\qquad\mathrm{applied\_predecessor}\mapsto\{\mathrm{predecessor},$
$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Succ}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{zero}\mapsto\mathrm{Succ}.\mathbf{this}.\mathrm{zero}\},$
$\qquad\mathrm{applied\_successor}\mapsto\{\mathrm{Succ}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{argument}\mapsto\mathrm{applied\_predecessor}.\mathrm{result}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_successor}.\mathrm{result}\},$
$\mathrm{Add}$ $\mapsto\{\mathrm{Nat},$
$\qquad\mathrm{augend}\mapsto\mathrm{Nat},$
$\qquad\mathrm{addend}\mapsto\mathrm{Nat},$
$\qquad\mathrm{applied\_addend}\mapsto\{\mathrm{addend},$
$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{zero}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{zero}\},$
$\qquad\mathrm{applied\_augend}\mapsto\{\mathrm{augend},$
$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{zero}\mapsto\mathrm{applied\_addend}.\mathrm{result}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_augend}.\mathrm{result}\}\quad\}\quad\}$
Concrete example
At the root scope level,
$\mathrm{one}\mapsto\{\mathrm{natural}.\mathrm{Succ},\;\mathrm{predecessor}\mapsto\mathrm{natural}.\mathrm{Zero}\}.$
Applying this with a Peano-style successor and zero: $\{\mathrm{one},\;\mathrm{successor}\mapsto S,\;\mathrm{zero}\mapsto Z\}$ produces an object whose $\mathrm{result}$ label contains the same observable tree as $\{S,\;\mathrm{argument}\mapsto Z.\mathrm{result}\}.\mathrm{result}$, i.e., successor applied once to zero. Here $S$ and $Z$ are arbitrary expressions representing a successor operation and a zero value respectively.
Appendix C Multi-Path Self-Reference in Scala
This appendix demonstrates that Scala 3 rejects the multi-path self-reference pattern that Overlay-Calculus handles naturally. Consider two objects that independently extend an outer class, each providing its own copy of an inner trait:
class MyOuter:
  trait MyInner:
    def outer = MyOuter.this
object Object1 extends MyOuter
object Object2 extends MyOuter
object MyObjectA extends Object1.MyInner
                     with Object2.MyInner
Scala 3 rejects MyObjectA with the error:
trait MyInner is extended twice
object MyObjectA cannot be instantiated since
  it has conflicting base types
  Object1.MyInner and Object2.MyInner
The rejection is not a surface-level restriction but a consequence of DOT’s type-theoretic foundations (Amin et al., 2016). In DOT, an object is constructed as $\nu(x:T)\,d$, where the self variable $x$ binds to a single object. Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer self-reference to a different object. Their intersection requires outer to simultaneously return Object1 and Object2—but DOT’s self variable is single-valued, making this intersection unrealizable.
The equivalent Overlay-Calculus definition is:
$\{\quad\mathrm{MyOuter}$ $\mapsto\{\mathrm{MyInner}\mapsto\{\mathrm{outer}\mapsto\mathrm{MyOuter}.\mathbf{this}\}\},$
$\mathrm{Object1}$ $\mapsto\{\mathrm{MyOuter}\},$
$\mathrm{Object2}$ $\mapsto\{\mathrm{MyOuter}\},$
$\mathrm{MyObjectA}$ $\mapsto\{\mathrm{Object1}.\mathrm{MyInner},\;\mathrm{Object2}.\mathrm{MyInner}\}\quad\}$
This is well-defined in Overlay-Calculus. Inside $\mathrm{MyInner}$, the reference $\mathrm{MyOuter}.\mathbf{this}$ has de Bruijn index $n=1$ (one scope level up from $\mathrm{MyInner}$ to $\mathrm{MyOuter}$). When $\mathrm{MyObjectA}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{MyObjectA})$ for overlay paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths—one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$—and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{MyObjectA}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Since inheritance is idempotent, the two routes contribute the same properties and no disambiguation is needed.
Appendix D Multi-Path Self-Reference in the NixOS Module System
This appendix demonstrates that the NixOS module system rejects the multi-path self-reference pattern with a static error. The translation uses the module system’s own abstractions: deferredModule for traits (unevaluated module values imported into other fixpoints) and submoduleWith for objects (evaluated in their own fixpoints).
let
  lib = (import <nixpkgs> { }).lib;
  result = lib.evalModules {
    modules = [
      (toplevel@{ config, ... }: {
        # class MyOuter { trait MyInner {
        #   def outer = MyOuter.this } }
        options.MyOuter = lib.mkOption {
          default = { };
          type = lib.types.deferredModuleWith {
            staticModules = [
              (MyOuter: {
                options.MyInner = lib.mkOption {
                  default = { };
                  type = lib.types.deferredModuleWith {
                    staticModules = [
                      (MyInner: {
                        options.outer = lib.mkOption {
                          default = { };
                          type =
                            lib.types.deferredModuleWith {
                              staticModules =
                                [ toplevel.config.MyOuter ];
        # object Object1 extends MyOuter
        options.Object1 = lib.mkOption {
          default = { };
          type = lib.types.submoduleWith {
            modules = [ toplevel.config.MyOuter ];
        # object Object2 extends MyOuter
        options.Object2 = lib.mkOption {
          default = { };
          type = lib.types.submoduleWith {
            modules = [ toplevel.config.MyOuter ];
        # object MyObjectA extends
        #   Object1.MyInner with Object2.MyInner
        options.MyObjectA = lib.mkOption {
          default = { };
          type = lib.types.submoduleWith {
            modules = [
              toplevel.config.Object1.MyInner
              toplevel.config.Object2.MyInner
in
  builtins.attrNames result.config.MyObjectA.outer
The NixOS module system rejects this with:
error: The option ‘MyObjectA.outer’
  in ‘<unknown-file>’
  is already declared
  in ‘<unknown-file>’.
The rejection occurs because Object1.MyInner and Object2.MyInner each carry the full staticModules of the MyInner deferred module, including the declaration options.outer. When MyObjectA imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error—it cannot express that both declarations originate from the same trait and should be unified.
The rejection is analogous to Scala’s “conflicting base types” error (Appendix C): both systems assume that each option (or type member) has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently. In Overlay-Calculus, the same pattern is well-defined because $\mathtt{overlays}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
Appendix E Well-Definedness of the Semantic Functions
This appendix proves that the mutually recursive equations (1)–(6) have a unique least fixed point, and therefore the semantics of Section 3 is well-defined.
E.1. The Product Lattice
Fix an AST with its primitive functions $\mathtt{defines}$ and $\mathtt{inherits}$. Let $\mathcal{L}$ denote the set of all labels and $\mathcal{P}$ the set of all paths (finite sequences of labels). Define the following domains, one for each semantic function:
$D_{\mathtt{properties}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{L})$ (sets of labels)
$D_{\mathtt{supers}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P}\times\mathcal{P})$ (sets of site–overlay pairs)
$D_{\mathtt{overlays}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{bases}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{resolve}}$ $=(\mathcal{P}\times\mathcal{P}\times\mathbb{N}\times\mathcal{L}^{*})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{this}}$ $=(\mathcal{P}(\mathcal{P})\times\mathcal{P}\times\mathbb{N})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
Each $D_{i}$ is a complete lattice under the pointwise subset ordering: $f\sqsubseteq g$ iff $f(x)\subseteq g(x)$ for all $x$. The product $D=D_{\mathtt{properties}}\times D_{\mathtt{supers}}\times D_{\mathtt{overlays}}\times D_{\mathtt{bases}}\times D_{\mathtt{resolve}}\times D_{\mathtt{this}}$ is a complete lattice under the componentwise ordering, with bottom element $\bot=(\lambda x.\,\varnothing,\ldots,\lambda x.\,\varnothing)$.
E.2. The Operator
Equations (1)–(6) define a function $F:D\to D$ that takes a tuple of candidate interpretations $(\mathtt{properties},\mathtt{supers},\mathtt{overlays},\mathtt{bases},\mathtt{resolve},\mathtt{this})$ and produces a new tuple by evaluating the right-hand side of each equation. The transitive closure $\mathtt{bases}^{*}$ appearing in equation (2) is derived from $\mathtt{bases}$: for any fixed $\mathtt{bases}$, the set $\mathtt{bases}^{*}(p)$ is defined as the least fixed point of the monotone operator $X\mapsto\{p\}\cup\bigcup_{q\in X}\mathtt{bases}(q)$ on $\mathcal{P}(\mathcal{P})$. If $\mathtt{bases}\sqsubseteq\mathtt{bases}^{\prime}$ (pointwise), then $\mathtt{bases}^{*}\sqsubseteq\mathtt{bases}^{\prime*}$ (pointwise), since enlarging the one-step relation can only enlarge the transitive closure.
The $\mathtt{this}$ function involves a recursion on the de Bruijn index $n$. Since $n$ decreases by one at each step and is a non-negative integer, this recursion terminates for any fixed interpretation of $\mathtt{supers}$. The result of $\mathtt{this}(S,p_{\mathrm{def}},n)$ for all valid $(S,p_{\mathrm{def}},n)$ is therefore well-defined as a function of $\mathtt{supers}$, and it remains to show that it is monotone in $\mathtt{supers}$.
E.3. Monotonicity
Proposition E.1.
The operator $F$ is monotone: if $\vec{d}\sqsubseteq\vec{d}^{\prime}$ in $D$, then $F(\vec{d})\sqsubseteq F(\vec{d}^{\prime})$.
Proof.
It suffices to show that each component of $F$ is monotone in every argument, i.e., that enlarging any of the input interpretations (in the $\subseteq$ ordering) can only enlarge the output. We verify this by inspecting the right-hand side of each equation and confirming that every occurrence of a mutually defined function appears in a positive position—as the domain of an existential quantifier in a set comprehension (Aczel, 1977).
Equation (1): $\mathtt{properties}$
$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\,\bigr\}$. The function $\mathtt{supers}$ appears as the set being iterated over ($\in\mathtt{supers}(p)$). The predicate $\ell\in\mathtt{defines}(\cdot)$ involves only the AST primitive $\mathtt{defines}$, not any mutually defined function. Enlarging $\mathtt{supers}(p)$ can only add pairs, hence can only add labels to the result. Positive in $\mathtt{supers}$.
Equation (2): $\mathtt{supers}$
$\mathtt{supers}(p)=\bigl\{\,(\mathtt{init}(p_{\mathrm{base}}),p_{\mathrm{overlay}})\bigm|p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{overlay}}\in\mathtt{overlays}(p_{\mathrm{base}})\,\bigr\}$. Both $\mathtt{bases}^{*}$ and $\mathtt{overlays}$ appear as sets being iterated over. As noted above, $\mathtt{bases}^{*}$ is monotone in $\mathtt{bases}$. Enlarging either $\mathtt{bases}$ or $\mathtt{overlays}$ can only add elements to the result. Positive in $\mathtt{bases}$ and $\mathtt{overlays}$.
Equation (3): $\mathtt{overlays}$
For $p\neq()$:
$\mathtt{overlays}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|(\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $\mathtt{last}(p)\in\mathtt{defines}(\cdot)$ involves only the AST primitive. Enlarging $\mathtt{supers}$ can only add branches, hence can only add paths to the result. Positive in $\mathtt{supers}$.
Equation (4): $\mathtt{bases}$
$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{overlay}}\in\mathtt{overlays}(p),\;(n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{overlay}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
Both $\mathtt{overlays}$ and $\mathtt{resolve}$ appear as sets being iterated over. The function $\mathtt{inherits}$ is an AST primitive. Enlarging either $\mathtt{overlays}$ or $\mathtt{resolve}$ can only add targets to the result. Positive in $\mathtt{overlays}$ and $\mathtt{resolve}$.
Equation (5): $\mathtt{resolve}$
$\mathtt{resolve}(\ldots)=\bigl\{\,p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\bigm|p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\mathtt{init}(p_{\mathrm{def}}),n)\,\bigr\}.$
The function $\mathtt{this}$ appears as the set being iterated over. Enlarging $\mathtt{this}$ can only add paths to the result. Positive in $\mathtt{this}$.
Equation (6): $\mathtt{this}$
For $n>0$: $\mathtt{this}(S,p_{\mathrm{def}},n)=\mathtt{this}(S^{\prime},\mathtt{init}(p_{\mathrm{def}}),n-1)$ where $S^{\prime}=\bigl\{\,p_{\mathrm{site}}\bigm|p_{\mathrm{current}}\in S,\;(p_{\mathrm{site}},p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\,\bigr\}$. The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $p_{\mathrm{overlay}}=p_{\mathrm{def}}$ is an equality test on path data, not involving any mutually defined function. Enlarging $\mathtt{supers}$ can only enlarge $S^{\prime}$, and by induction on $n$ (the base case $n=0$ returns $S$ unchanged), the recursive call is monotone in $S^{\prime}$. Positive in $\mathtt{supers}$.
In summary, every mutually defined function appears on the right-hand side of every equation only inside the pattern $x\in f(\ldots)$ within a set comprehension—i.e., as the domain of an existential quantifier. No complement, set difference, or negated membership test appears in any equation. Therefore $F$ is a monotone operator on the complete lattice $D$. ∎
E.4. Existence of the Least Fixed Point
Theorem E.2.
The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overlays}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined as the least fixed point of $F$.
Proof.
By Proposition E.1, $F$ is monotone on the complete lattice $D$. By the Knaster–Tarski theorem (Tarski, 1955), $F$ has a least fixed point $\vec{d}^{*}=\bigsqcup_{n\geq 0}F^{n}(\bot)$, where $\bot$ assigns the empty set to every input. The components of $\vec{d}^{*}$ are the semantic functions of Section 3.
Concretely, the least fixed point is the union of the finite approximation chain: $\bot\sqsubseteq F(\bot)\sqsubseteq F^{2}(\bot)\sqsubseteq\cdots$. A membership $\ell\in\mathtt{properties}(p)$ holds in $\vec{d}^{*}$ if and only if it appears at some finite stage $F^{k}(\bot)$, which corresponds to a finite derivation tree using the equations. When no finite derivation exists, the membership is absent from the least fixed point and the query is undefined—the analogue of divergence. ∎
Appendix F Böhm Tree Correspondence: Proofs
This appendix contains the full proofs of Adequacy (Theorem 4.2) and Full Abstraction (Theorem 4.3) stated in Section 4.2.
F.1. Böhm Trees
We briefly recall the definition of Böhm trees (Barendregt, 1984). A head reduction $M\to_{h}M^{\prime}$ contracts the outermost $\beta$-redex only: if $M$ has the form $(\lambda x.\,e)\;v\;M_{1}\cdots M_{k}$, then $M\to_{h}e[v/x]\;M_{1}\cdots M_{k}$. A term $M$ is in head normal form (HNF) if it has no head redex, i.e., $M=\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ where $y$ is a variable.
The Böhm tree $\mathrm{BT}(M)$ of a $\lambda$-term $M$ is an infinite labeled tree defined by:
$\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\text{if }M\to^{*}_{h}\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{cases}$
Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Barendregt, 1984; Hyland and Ong, 2000).
F.2. Path Encoding
We define a correspondence between positions in the Böhm tree and paths in the overlay tree. When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting overlay tree has a specific shape:
  • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$ (the abstraction shape).
  • A $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ translates to a record with own properties $\{x,\mathrm{result}\}$, where $x$ holds the encapsulated application $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$.
  • A tail call $V_{1}\;V_{2}$ translates to a record with own properties $\{\_\mathrm{tail},\mathrm{result}\}$, where $\mathrm{tail}$ holds the encapsulated application and $\mathrm{result}$ projects $\mathrm{tail}.\mathrm{result}$.
  • A variable $x$ with de Bruijn index $n$ translates to an indexed reference $\uparrow^{\!n}.\mathrm{argument}$.
Each Böhm tree position is a sequence of navigation steps from the root. In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body (peeling off one abstraction layer) and entering the $i$-th argument $M_{i}$ of the head variable. In the overlay tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
Rather than comparing absolute paths across different overlay trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection. Recall the definition of overlay-convergence (Definition 4.1). Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation. The abstraction shape at the end signals that a weak head normal form has been reached.
F.3. Single-Path Lemma
The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path overlay inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
Lemma F.1 (Single Path).
For any closed ANF $\lambda$-term $M$, at every invocation of $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)$ during evaluation of $\mathcal{T}(M)$, the frontier set $S$ contains exactly one path.
Proof.
By induction on the structure of the ANF term.
Case $\mathcal{T}(\lambda x.\,M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
This is a record literal with $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$ and no inheritance sources, so $\mathtt{inherits}=\varnothing$. At the root path $p$, we have $\mathtt{overlays}(p)=\{p\}$ (a single path) since there are no inheritance sources to introduce additional branches. Inside $\mathcal{T}(M)$, any reference to $x$ is $\uparrow^{\!n}.\mathrm{argument}$ where $n$ is the de Bruijn index pointing to this scope level. Since $\mathcal{T}(\lambda x.\,M)$ introduces exactly one scope level, there is exactly one overlay at that level, hence $\mathtt{this}$ finds exactly one matching pair $(p_{\mathrm{site}},\;p_{\mathrm{overlay}})$ in $\mathtt{supers}$.
Case $\mathcal{T}(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level ($\mathtt{inherits}=\varnothing$), so $\mathtt{overlays}(\text{root})=\{\text{root}\}$. Inside the $x$ subtree, the inheritance $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$ has exactly one inheritance source ($\mathcal{T}(V_{1})$), creating exactly one $\mathtt{bases}$ entry. By induction on $V_{1}$, the single-path property holds inside $\mathcal{T}(V_{1})$. Inside $\mathcal{T}(M)$, the induction hypothesis applies directly.
Case $\mathcal{T}(V_{1}\;V_{2})=\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
Identical to the $\mathbf{let}$-binding case: the outer record has own properties $\{\_\mathrm{tail},\mathrm{result}\}$ with no inheritance sources at the outer level, and the $\mathrm{tail}$ subtree contains exactly one inheritance source.
Case $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$
This is a reference, not a record. The $\mathtt{this}$ resolution walks up $n$ scope levels. By the inductive hypothesis on the enclosing translations, each step of $\mathtt{this}$ encounters exactly one matching pair in $\mathtt{supers}$.
The key structural invariant is that the translation $\mathcal{T}$ never produces two independent inheritance sources at the same scope level that both define the same label. In every case, application inheritances are encapsulated inside a named property ($x$ or $\mathrm{tail}$), and each such encapsulated inheritance has exactly one inheritance source. This prevents diamond inheritance and ensures single-path resolution throughout. ∎
F.4. Substitution Lemma
The core mechanism of the translation is that inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ plays the role of substitution. We make this precise.
Lemma F.2 (Substitution).
Let $M$ be an ANF term in which $x$ may occur free, and let $V$ be a closed value. Define the inherited tree $C=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. Then for every path $p$ under $C\mathbin{\triangleright}\mathrm{result}$ and every label $\ell$:
$\ell\in\mathtt{properties}(p)\text{ in }C\mathbin{\triangleright}\mathrm{result}\quad\Longleftrightarrow\quad\ell\in\mathtt{properties}(p)\text{ in }\mathcal{T}(M[V/x])$
where $M[V/x]$ is the usual capture-avoiding substitution.
Proof.
By structural induction on $M$.
Base case: $M=x$
Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference. In the inherited tree $C$, $\mathtt{this}$ resolves the reference by walking up $n$ scope levels to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ overlays the $\mathrm{argument}$ slot with $\mathcal{T}(V)$. Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma F.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree. Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
Base case: $M=y$ where $y\neq x$
Then $\uparrow^{\!m}.\mathrm{argument}$ (where $m$ is the de Bruijn index of $y$, $m\neq n$) walks up to a different scope level, so the inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ at the binding scope of $x$ has no effect. Since $M[V/x]=y$, the properties are identical.
Inductive case: $M=\lambda y.\,M^{\prime}$
Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$. The subtree at $C\mathbin{\triangleright}\mathrm{result}$ is another abstraction-shaped record. References to $x$ inside $\mathcal{T}(M^{\prime})$ resolve through $\mathtt{this}$ in exactly the same way (with one additional scope level from the inner $\lambda$), and by induction on $M^{\prime}$, the properties under $C\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M^{\prime}[V/x])$. Since $(\lambda y.\,M^{\prime})[V/x]=\lambda y.\,(M^{\prime}[V/x])$ (assuming $y$ is fresh), the result follows.
Inductive case: $M=\mathbf{let}\;z=V_{1}\;V_{2}$ $\mathbf{in}\;M^{\prime}$
Then $\mathcal{T}(M)=\{z\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$. The substitution distributes: $M[V/x]=\mathbf{let}\;z=V_{1}[V/x]\;V_{2}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$. By induction on $V_{1}$, $V_{2}$, and $M^{\prime}$, the properties under each subtree ($z$ and $\mathrm{result}$) coincide, since references to $x$ inside each subtree resolve to $\mathcal{T}(V)$ via the same $\mathtt{this}$ mechanism.
Inductive case: $M=V_{1}\;V_{2}$ (tail call)
Analogous to the $\mathbf{let}$-binding case with $\mathrm{tail}$ in place of $z$. ∎
F.5. Convergence Preservation
The convergence criterion (Definition 4.1) follows the $\mathrm{result}$ chain from the root. Each step in this chain corresponds to one head reduction step in the $\lambda$-calculus. We make this precise.
Lemma F.3 (Result-step).
Let $(\lambda x.\,M)\;V$ be a tail call (a $\beta$-redex). Then for every path $p$ and label $\ell$:
$\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)\;\Longleftrightarrow\;\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}(M[V/x])$
That is, following one $\mathrm{result}$ projection in the tail call’s overlay tree yields the same properties as the root of the reduct’s overlay tree.
Proof.
The translation gives:
$\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)=\{\_\mathrm{tail}\mapsto\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
The $\mathrm{result}$ label at the root is defined as $\mathrm{tail}.\mathrm{result}$, so $\text{root}\mathbin{\triangleright}\mathrm{result}$ resolves to the $\mathrm{result}$ path inside the inheritance $\mathrm{tail}=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. This inheritance is exactly the inherited tree $C$ of Lemma F.2, and $\mathrm{tail}\mathbin{\triangleright}\mathrm{result}$ is $C\mathbin{\triangleright}\mathrm{result}$. By Lemma F.2, the properties under $C\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M[V/x])$.
Let-binding variant. An analogous result holds for $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$ where $V_{1}=\lambda y.\,M^{\prime\prime}$. The translation gives:
$\mathcal{T}\bigl(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\bigr)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$
Since $\mathrm{result}\mapsto\mathcal{T}(M^{\prime})$ is a property definition, the subtree at $\text{root}\mathbin{\triangleright}\mathrm{result}$ is $\mathcal{T}(M^{\prime})$. Inside $\mathcal{T}(M^{\prime})$, each reference to $x$ is $x.\mathrm{result}$, which resolves to the $\mathrm{result}$ path inside the $x$ subtree $\{\mathcal{T}(\lambda y.\,M^{\prime\prime}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$. This is the inherited tree $C$ of Lemma F.2, so $x.\mathrm{result}$ has the same properties as $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. In the $\lambda$-calculus, $M=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$, and the head reduct is $N=M^{\prime}[(V_{1}\;V_{2})/x]$. In $\mathcal{T}(N)$, each occurrence of $x$ is replaced by $(V_{1}\;V_{2})$, whose result, by the tail-call variant above, also yields $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. Therefore the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$ for all $p$. ∎
Theorem F.4 (Convergence preservation).
If $M\to_{h}N$ (head reduction), then $\mathcal{T}(M){\Downarrow}$ if and only if $\mathcal{T}(N){\Downarrow}$.
Proof.
A head reduction step in ANF takes one of two forms: a tail call $(\lambda x.\,M^{\prime})\;V\to_{h}M^{\prime}[V/x]$, or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[(V_{1}\;V_{2})/x]$ (since $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$ in the $\lambda$-calculus). In both cases, Lemma F.3 (tail-call and let-binding variants respectively) shows that the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n}$ in $\mathcal{T}(M)$ coincide with the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n-1}$ in $\mathcal{T}(N)$ for all $n\geq 1$, where $N$ is the head reduct. Hence the abstraction shape appears at depth $n$ in the pre-reduct if and only if it appears at depth n1n-1 in the post-reduct. ∎
F.6. Adequacy
A closed ANF term at the top level is either an abstraction (a value) or a computation (a $\mathbf{let}$-binding or tail call). An abstraction translates to a record whose root immediately has the abstraction shape; a computation’s root has the form $\{x,\mathrm{result}\}$ or $\{\_\mathrm{tail},\mathrm{result}\}$, and one must follow the $\mathrm{result}$ chain to find the eventual value.
Proof of Theorem 4.2.
($\Rightarrow$) Suppose $M\to^{*}_{h}\lambda x.\,M^{\prime}$ in $k$ head reduction steps. We show $\mathcal{T}(M)$ converges at depth $\leq k$ by induction on $k$.
Base case ($k=0$): $M$ is already an abstraction $\lambda x.\,M^{\prime}$. Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$, whose root has $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$. These labels appear at stage $F^{1}(\bot)$ of the approximation chain, so $\mathcal{T}(M)$ converges at depth $n=0$.
Inductive step ($k\geq 1$): $M$ is not an abstraction, so $M$ is either a tail call or a $\mathbf{let}$-binding, both of which are $\beta$-redexes. Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$. By Lemma F.3 (tail-call or let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$. By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
($\Leftarrow$) Suppose $\mathcal{T}(M){\Downarrow}$ at depth $n$. We show $M$ has a head normal form by induction on $n$.
Base case ($n=0$): the root of $\mathcal{T}(M)$ has abstraction shape $\{\mathrm{argument},\mathrm{result}\}$. Only the abstraction rule $\mathcal{T}(\lambda x.\,M^{\prime})$ produces a root with own property $\mathrm{argument}$. (For tail calls and $\mathbf{let}$-bindings, the root has $\mathtt{defines}=\{\_\mathrm{tail},\mathrm{result}\}$ or $\{x,\mathrm{result}\}$, and $\mathrm{argument}$ cannot appear via $\mathtt{supers}$ because $\mathtt{inherits}=\varnothing$ at the root, giving $\mathtt{bases}(\text{root})=\varnothing$.) Therefore $M$ is an abstraction, hence already in head normal form.
Inductive step ($n\geq 1$): the abstraction shape appears at depth $n$ but not at depth 0. By the base case argument, $M$ is not an abstraction, so $M$ is either a tail call $V_{1}\;V_{2}$ or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$. In both cases, $M$ is a $\beta$-redex (since $V_{1}=\lambda x.\,M^{\prime}$ or $V_{1}=\lambda y.\,M^{\prime\prime}$ respectively, the only closed ANF value). By Lemma F.3 (tail-call or let-binding variant respectively), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct. So $\mathcal{T}(N)$ converges at depth n1n-1. By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$. ∎
F.7. Full Abstraction
Adequacy relates a single term’s head normal form to its overlay-convergence. Full abstraction lifts this to an equivalence between terms: the translation $\mathcal{T}$ preserves and reflects observational equivalence.
Definition F.5 (Overlay-contextual equivalence).
For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
$\mathcal{T}(C[M]){\Downarrow}\;\Longleftrightarrow\;\mathcal{T}(C[N]){\Downarrow}$
This definition observes only convergence (a single bit), but quantifying over all contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior.
Proof of Theorem 4.3.
The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the overlay tree $\mathcal{T}(C[M])$ is determined by the overlay tree $\mathcal{T}(M)$ and the translation of the context. This means $\mathcal{T}$ preserves the context structure needed for the argument.
($\Leftarrow$) Suppose $\mathrm{BT}(M)=\mathrm{BT}(N)$. Böhm tree equivalence is a congruence (Barendregt, 1984), so for every context $C[\cdot]$, $\mathrm{BT}(C[M])=\mathrm{BT}(C[N])$. In particular, $C[M]$ has a head normal form iff $C[N]$ does. By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$. Hence $\mathcal{T}(M)\approx\mathcal{T}(N)$.
($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$. Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Hyland and Ong, 2000): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa). By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$. Hence $\mathcal{T}(M)\not\approx\mathcal{T}(N)$. ∎
The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Hyland and Ong, 2000). The overlay tree semantics inherits full abstraction from the $\lambda$-calculus without needing to define or compare absolute paths inside overlay trees.
Appendix G CPS-Agnostic Programs
In the $\lambda$-calculus, a program is either in continuation-passing style (CPS) or direct style. Converting between these representations requires explicit transformation—the CPS conversion (Plotkin, 1975; Sabry and Felleisen, 1993). In Overlay-Calculus, the distinction does not exist: the same program text is CPS or direct-style depending on how dependencies are assembled.
Consider a generic computation step:
$\mathrm{Step}\mapsto\{\mathrm{input}\mapsto\{\},\;\mathrm{process}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\}$
The empty $\mathrm{process}$ slot is an interpretation point. If filled externally by an orchestrator, it acts as a continuation parameter (CPS interpretation). If inherited locally with an implementation, it acts as a direct dependency (direct-style interpretation). The $\mathrm{Step}$ code remains unchanged.
CPS interpretation via external orchestration
When $\mathrm{process}$ is injected from outside, the computation passes its intermediate result to an external handler:
$\{\mathrm{Step}.\mathbf{this},\;$
$\quad\mathrm{input}\mapsto\{\mathrm{data}\mapsto\{\}\},\;$
$\quad\mathrm{process}\mapsto\{$
$\qquad\mathrm{result}\mapsto\{\mathrm{handled}\mapsto\{\}\}$
$\quad\}$
The $\mathrm{process}$ slot acts as a continuation: it receives the result and determines what happens next. This is CPS-style programming.
Direct-style interpretation via local inheritance
Alternatively, $\mathrm{process}$ can be filled by local inheritance:
$\{\mathrm{Step}.\mathbf{this},\;$
$\quad\mathrm{input}\mapsto\{\mathrm{data}\mapsto\{\}\},\;$
$\quad\mathrm{process.result}\mapsto\{\mathrm{computed}\mapsto\{\}\}$
Here $\mathrm{process}$ is implemented directly within the inheritance. This is direct-style programming. The $\mathrm{Step}$ code is identical in both cases—only the assembly strategy differs.
Representation independence
This property is a form of representation independence (Reynolds, 1983) applied to control flow: client code using $\mathrm{Step}$ need not know whether $\mathrm{process}$ will be filled continuation-style or directly. Library authors write code once; users choose the control flow style at assembly time.
The $\lambda$-calculus lacks this flexibility. A function written in direct style must be explicitly transformed (CPS conversion) before it can be used continuation-style. Overlay-Calculus programs are inherently in A-normal form (Flanagan et al., 1993), where every intermediate result has a name.
However, Overlay-Calculus ANF differs critically from ANF in functional programming. In functional language compilers, ANF conversion requires specifying a target monad—either a concrete monad or an abstract one (Moggi, 1991). Compiler-internal ANF transformations (the ANF form used in intermediate representations) typically target the identity monad, which cannot express CPS transformations. This is analogous to how imperative compilers use SSA (Static Single Assignment) form, which similarly lacks control flow flexibility. Overlay-Calculus ANF, by contrast, has built-in dependency injection: empty API slots can be filled by external orchestrators. This capability enables the same ANF program to serve both CPS and direct-style interpretations without monad specification or transformation.
Multiple symmetric continuations
When interpreted continuation-style, Overlay-Calculus naturally supports multiple exit points:
$\mathrm{Choice}\mapsto\{\mathrm{on\_success}$ $\mapsto\{\},\;$
$\mathrm{on\_failure}$ $\mapsto\{\},\;$
$\mathrm{on\_timeout}$ $\mapsto\{\},\;$
$\mathrm{result}$ $\mapsto\{\}\}$
All continuation slots are symmetric elements of the same record. Traditional CPS in the $\lambda$-calculus distinguishes a single continuation parameter; additional continuations must be encoded (e.g., as pairs or disjoint sums). Overlay-Calculus treats all continuation slots uniformly, eliminating the need for specialized encodings.
Abstract and concrete are syntactically identical
From a category-theoretic perspective, abstract algebra and concrete implementation have the same syntactic structure in Overlay-Calculus:
Abstract: $\mathrm{API}\mapsto\{\}$
Concrete: $\mathrm{API}\mapsto\{\text{implementation}\}$
This is the abstract/concrete correspondence: the same functor applies to both. In traditional languages, abstract interfaces (e.g., Java interfaces, Haskell type classes) have different syntax from concrete implementations. In Overlay-Calculus, they are syntactically identical—only the interpretation differs. An empty API is simultaneously an abstract requirement (when viewed as a dependency) and a concrete extension point (when viewed as an injection site).
The Selector and Rebuilder combinators (Appendix A) demonstrate this flexibility: their callback slots ($\mathrm{on\_zero}$, $\mathrm{on\_odd}$, $\mathrm{on\_even}$) can be filled externally (CPS interpretation) or inherited locally (direct interpretation).
Practical implications
A library component with empty API slots can be used in multiple contexts without modification:
  • In synchronous code, fill slots with direct implementations
  • In asynchronous code, fill slots with continuation handlers
  • In effect systems, compose with algebraic effect handlers (Plotkin and Power, 2003)
  • In test code, fill slots with mock implementations
No dual APIs are needed. The same component serves all use cases. This eliminates a common source of code duplication in traditional languages, where async and sync versions of the same functionality must be maintained separately.
The implementation (Yang, 2025b) includes executable examples of all three patterns described above: escape continuations analogous to call/cc, success/failure continuations analogous to exception handling, and protected computations with multiple symmetric exit points. Each example includes test cases verifying that the correct continuation path is taken.
CPS-agnosticism is not a designed feature of Overlay-Calculus; it is an emergent property of records and inheritance. The calculus defines no notion of continuation—the three syntax constructs and six semantic functions of Section 3 are the entire language.
Appendix H Asymmetry in Problem Difficulty
A skeptical reader might ask whether Overlay-Calculus is simply the $\lambda$-calculus with different syntax. The test of whether two computational models are “the same” is whether they make the same problems easy and the same problems hard. They do not:
System Expression Problem Random-access memory
$\lambda$-calculus Object algebras (d. S. Oliveira and Cook, 2012), Explicit trie encoding
finally tagless (Carette et al., 2009)
SKI combinators Same difficulty as $\lambda$ Same difficulty as $\lambda$
Overlay-Calculus Direct inheritance Records are tries
Overlay-Calculus requires no additional machinery for either problem class; the $\lambda$-calculus requires elaborate encodings for both.
The semantic domain of Overlay-Calculus (Section 3) is entirely first-order: paths are label sequences, overlays are the record trees produced by evaluation, AST nodes are pure data, and the resolution functions ($\mathtt{overlays}$, $\mathtt{supers}$, $\mathtt{resolve}$, etc.) are ordinary set-valued functions over these structures. No function types appear in the object language or its semantics. The crucial question is not what notation is used, but what computational patterns are primitive in the language itself.
In Overlay-Calculus, the primitive operations are record construction, qualified-this reference, and inheritance. Function application is derived from these primitives (Section 4), not primitive itself. The asymmetry of Section 4.1 makes this concrete: embedding the $\lambda$-calculus into Overlay-Calculus requires five rules; the reverse requires sophisticated encoding.
Why tree-level inheritance dissolves the Expression Problem.
Each operation is an independent overlay that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited overlays. Crucially, this applies to return types: if operation $A$ returns values from factory $F$, and an independently defined overlay $B$ adds a property to $F$, then inheriting from $A$ and $B$ causes the values returned by $A$ to carry that property automatically—without modifying either $A$ or $B$. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
Multi-path qualified this (equation 6) is essential to this dissolution—not merely to solving the Expression Problem (for which many solutions exist (d. S. Oliveira and Cook, 2012; Carette et al., 2009)), but to making it not arise in the first place. In the case study of Section 5, $\mathrm{evaluation}.\mathrm{Constant}$ and $\mathrm{display}.\mathrm{Constant}$ both inherit $\mathrm{expression}.\mathrm{Constant}$. When the two operations are inherited, the resulting $\mathrm{Constant}$ inherits the $\mathrm{expression}$ schema through two independent routes. The reference $\mathrm{Constant}.\mathbf{this}.\mathrm{magnitude}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties. In Scala, the same pattern—two independently defined inner classes extending the same outer class’s trait—triggers a “conflicting base types” rejection (Appendix C), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously. The Expression Problem can still be solved in Scala using object algebras or similar patterns, but it cannot be made to not arise: free inheritance of independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.

### v2

@@
 Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, overlay-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming.
-It trivially embeds the $\lambda$-calculus, although the entire semantics builds on only naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance simply does not arise.
-This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence.
-Overlay-Calculus is distilled from the Overlay language, a practical implementation in which we observed further emergent phenomena: the Expression Problem dissolves, programs are CPS-agnostic, records natively encode random-access memory, and self-reference resolves to multiple targets.
+It trivially embeds the $\lambda$-calculus, although the entire semantics rests solely on naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance does not arise.
+This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence (Barendregt, 1984).
+Overlay-calculus is distilled from the Overlay language, a practical implementation in which we observed further emergent phenomena: the Expression Problem dissolves, programs are function color blind (Nystrom, 2015), ordinary arithmetic yields the relational semantics of logic programming, and self-reference resolves to multiple targets, making overlay-calculus strictly more expressive than the $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991).
 These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
@@
 Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or overlay.
-Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves remarkable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself—Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively.
-Yet no computational theory explains why this mechanism is so powerful.
+Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively.
+Yet no computational theory explains why this mechanism is so expressive.
 The $\lambda$-calculus serves as the foundational computational model for functional programming.
@@
 The NixOS module system already suggests that such a model exists: its inheritance-based composition over recursive records, without explicit functions, is expressive enough to configure entire operating systems.
-Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures—initial algebras in the sense of universal algebra.
-We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures—in the spirit of F-coalgebras (Rutten, 2000)—where semantics is determined by the observer and a finite prefix suffices for any finite observation.
-This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006)—over 100,000 packages—behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set.
-Guided by this observation and by the NixOS module system’s recursive merging mechanism, we built the Overlay language (Yang, 2025b)—a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
-In developing real programs in the Overlay language—Church-encoded data types (Appendix B), an Expression Problem solution (Section 5), trie-based random-access memory (Appendix A), and continuation patterns (Appendix G)—we discovered that several long-standing problems in programming language theory do not arise.
-The linearization problem that has plagued mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) does not exist—not because it is resolved, but because the three primitives make it structurally impossible (Section 3).
+Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures, that is, initial algebras in the sense of universal algebra.
+We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures, in the spirit of F-coalgebras (Rutten, 2000), where semantics is determined by the observer and a finite prefix suffices for any finite observation.
+This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set.
+Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives.
+Through iterative refinement, beginning with an early version implemented in Nix, the system evolved into the Overlay language (Yang, 2025b), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
+In developing real programs in the Overlay language, including boolean logic and natural number arithmetic (Section 5) and a web application (Section 5), we found that the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) is structurally absent: because records are sets, inheritance is inherently commutative, idempotent, and associative (Section 3).
 Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system).
 Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction.
-These are not properties we engineered; they are phenomena we observed in working programs.
+We also found that boolean logic and arithmetic, implemented as separate overlay files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism; that Church-encoded values automatically form trie structures; and that operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$, which is the relational semantics of logic programming (Section 5); and that the same web application overlay file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015), equivalently CPS-agnostic behavior (Section 5).
 This paper distills these observations into overlay-calculus, a minimal computational model for declarative programming.
-It turns out that function application has a straightforward translation into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument.
+Function application translates directly into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument.
 Since the $\lambda$-calculus embeds into overlay-calculus, overlay-calculus is Turing complete.
-The translation reveals an asymmetry: while embedding the $\lambda$-calculus into Overlay-Calculus is trivial, the reverse is non-trivial because Overlay-Calculus provides primitive operations—self-referential records, open recursion, symmetric inheritance—that have no direct counterparts in the $\lambda$-calculus.
-Moreover, Overlay-Calculus naturally encodes random-access memory: records are tries, and inheritance is trie union, yielding an immutable analogue of the RAM machine (Appendix A).
-This asymmetry suggests that Overlay-Calculus is not merely a syntactic variant of the $\lambda$-calculus but captures different computational patterns.
+The translation reveals an asymmetry: embedding the $\lambda$-calculus into overlay-calculus requires five local, syntax-directed rules (Section 4), but the reverse is impossible—overlay enables open extension that changes operational equivalence on $\lambda$-terms, establishing a strict separation in Felleisen’s sense of expressiveness (Felleisen, 1991) (Section 4.2).
 Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, overlay-calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
 All constructions in this paper have been implemented and tested in the Overlay language, available as open source (Yang, 2025b), with a test suite covering every example.
+[^1] The Overlay language is statically typed and therefore stricter than the untyped overlay-calculus.
+All examples in this paper are ported from the Overlay language and conform to both semantics.
+Elements added to satisfy the type checker are harmless for the algorithms presented here and may aid readability.
 All tests are mechanically verified against expected output snapshots.
@@
   • We present overlay-calculus, a minimal computational model for declarative programming that contains only overlay constructs (record, definition, inheritance) and no functions or scalar types.
-  • Inheritance operates at the tree level: a record is a set of elements—definitions and inheritance sources—so inheriting from two overlays that share a label recursively merges their subtrees (union mount semantics), rather than overriding at the method level as in mixin and trait calculi.
+  • Inheritance operates at the tree level: a record is a set of elements, both definitions and inheritance sources, so inheriting from two overlays that share a label recursively merges their subtrees (deep merge), rather than overriding at the method level as in mixin and trait calculi.
 Since records are sets, inheritance is inherently commutative, idempotent, and associative.
-The linearization problem inherent in prior mixin and trait calculi simply does not arise.
+The linearization problem inherent in prior mixin and trait calculi does not arise.
   • The semantics is purely observational: there are no reduction rules, no evaluation order, and no notion of a result that a term reduces to.
@@
 The $\lambda$-calculus translation maps each application to a inheritance; the resulting inheritance tree is the call stack.
-The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree—a first-order set-theoretic query, not a fixed-point construction.
+The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree: a first-order set-theoretic query, not a fixed-point construction.
   • Self-reference resolves to multiple objects, not one.
-Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala, Appendix C; NixOS module system, Appendix D).
-Since inheritance is idempotent, all objects contribute equally—no disambiguation is needed.
-  • The $\lambda$-calculus (in A-normal form) translates into Overlay-Calculus in five rules, so Overlay-Calculus is Turing complete.
-Composing this translation with the semantic functions induces a semantics on the $\lambda$-calculus that requires no function spaces, no substitution, and no reduction rules—only set membership in powerset lattices (Section 4).
-The translation is adequate and fully abstract with respect to Böhm tree equivalence (Section 4.2): two terms have the same overlay-convergence behavior in all contexts if and only if they have the same Böhm tree.
-This yields a fully abstract semantics of the lazy $\lambda$-calculus that lives entirely in naive set theory—powerset lattices and the Knaster–Tarski fixed-point theorem—without domain theory, inverse limits, or Scott continuity.
-Standard data types (booleans, natural numbers) can be Church-encoded without scalars (Appendix B).
-  • We observe several emergent properties in practice: the Expression Problem (Wadler, 1998) does not arise (Section 5), programs are CPS-agnostic (Appendix G), and records natively encode random-access memory (Appendix A).
-Section 5 demonstrates these properties through a case study of the Expression Problem.
+Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala and the NixOS module system, Appendix B).
+Since inheritance is idempotent, all objects contribute equally; no disambiguation is needed.
+  • The $\lambda$-calculus (in A-normal form) translates into overlay-calculus in five rules (Section 4).
+The translation is fully abstract with respect to Böhm tree equivalence (Section 4.1), yielding a semantics of the lazy $\lambda$-calculus that requires no domain theory.
+  • We observe further properties: extensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section 5).
+  • Open extension via overlay changes the operational equivalence of $\lambda$-terms, establishing via Felleisen’s Theorem 3.14 (Felleisen, 1991) that the $\lambda$-calculus cannot macro-express overlay (Section 4.2).
+Section 5 demonstrates these properties through a case study of Nat arithmetic.
 2. Syntax
@@
 Each element is either a definition $\ell\mapsto e$ or an inheritance source.
-Because expressions are sets, element order is irrelevant and duplicates have no effect—inheritance is inherently commutative, idempotent, and associative.
+Because expressions are sets, element order is irrelevant and duplicates have no effect; inheritance is inherently commutative, idempotent, and associative.
 The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding.
@@
 The Overlay language also provides scalar types through a foreign-function interface (FFI); overlay-calculus is what remains after removing the FFI.
-None of the examples in this paper uses the FFI—scalars simply did not arise.
+None of the examples in this paper uses the FFI; scalars did not arise.
 Inheritance
@@
 Each $\{\ldots\}$ record has a label in its enclosing scope; $\ell_{\mathrm{up}}$ names the target enclosing scope.
-  • Indexed form: $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n$ is a de Bruijn index (de Bruijn, 1972) counting the number of enclosing scope levels to traverse.
-Both forms retrieve the target scope—the fully inherited record after all inheritance has been applied—and then project properties $\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ from it.
+  • Indexed form: $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n\geq 0$ is a de Bruijn index (de Bruijn, 1972): $n=0$ refers to the reference’s enclosing scope, and each increment of $n$ moves one scope level further out.
+Both forms retrieve the target scope, that is, the fully inherited record after all inheritance has been applied, and then project properties $\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ from it.
 The named form desugars to the indexed form during parsing (Section 3).
-A new scope level is created at each $\{\ldots\}$ (record literal).
+A new scope level is created at each $\{\ldots\}$ record literal.
 The scope contains all properties of that overlay, including those inherited via inheritance sources.
@@
 Inheritance sources do not create additional scope levels.
-  • $\mathrm{Foo}.\mathbf{this}$ (equivalently, $\uparrow^{\!n}$ where $\mathrm{Foo}$ is $n$ scope levels above) retrieves the enclosing record named $\mathrm{Foo}$.
+  • $\mathrm{Foo}.\mathbf{this}$ (equivalently, $\uparrow^{\!n}$ where $\mathrm{Foo}$ is $n$ steps above the enclosing scope) retrieves the record named $\mathrm{Foo}$.
   • $\mathrm{Foo}.\mathbf{this}.\ell$ (equivalently, $\uparrow^{\!n}.\ell$) projects property $\ell$ from $\mathrm{Foo}$.
@@
 Given an AST, what are the properties visible at each position in the tree? This section answers the question by defining a collection of mutually recursive functions over paths.
-[…]
+Unlike the $\lambda$-calculus, which requires reduction rules ($\beta$-reduction) to specify how terms compute, overlay-calculus has no reduction steps.
+There is no rewriting of terms, no evaluation order, and no notion of a “result” that a term reduces to.
+Instead, the semantics is purely observational: one queries whether a label $\ell$ is a property at a path $p$ in the inheritance tree.
+The mutually recursive functions defined below, namely $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overlays}$, $\mathtt{bases}$, $\mathtt{resolve}$, and $\mathtt{this}$, constitute the entire semantics.
+This is the observational perspective introduced in Section 1: like an F-coalgebra (Rutten, 2000), an overlay tree’s behavior is determined by observation (property presence), not by reduction.
 Most of these functions have direct analogues in object-oriented languages: $\mathtt{Path}$ identifies a position in the tree, $\mathtt{init}$ and $\mathtt{last}$ navigate paths, $\mathtt{defines}$ extracts the locally defined members from the AST, $\mathtt{inherits}$ extracts the inheritance declarations from the AST, $\mathtt{properties}$ computes all visible members at a path including both locally defined and inherited ones, $\mathtt{supers}$ computes the transitive inheritance closure, and $\mathtt{bases}$ collects the direct base classes.
@@
   • $\mathtt{defines}(p)$: the set of labels that have subtrees at $p$.
-  • $\mathtt{inherits}(p)$: the set of reference pairs $(n,\;\ell_{\mathrm{down},*})$, where $n$ is the de Bruijn index (de Bruijn, 1972) (zero-based) and $\ell_{\mathrm{down},*}$ is the list of downward projections.
+  • $\mathtt{inherits}(p)$: the set of reference pairs $(n,\;\ell_{\mathrm{down},*})$, where $n$ is the de Bruijn index and $\ell_{\mathrm{down},*}$ is the list of downward projections.
 During parsing, all three syntactic forms are resolved to de Bruijn index pairs $(n,\;\ell_{\mathrm{down},*})$.
@@
 A named reference $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the last occurrence of $\ell_{\mathrm{up}}$ among the labels of $p$.
-Let $p_{\mathrm{def}}$ be the prefix of $p$ up to and including that occurrence.
-The de Bruijn index is $n=|p|-|p_{\mathrm{def}}|$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$.
+Let $p_{\mathrm{target}}$ be the prefix of $p$ up to and including that occurrence.
+The de Bruijn index is $n=|p|-|p_{\mathrm{target}}|-1$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$.
 A lexical reference $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the nearest prefix $p^{\prime}$ of $p$ such that $\ell_{\mathrm{down},1}\in\mathtt{defines}(p^{\prime})$.
-The de Bruijn index is $n=|p|-|p^{\prime}|$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$.
+The de Bruijn index is $n=|p|-|p^{\prime}|-1$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$.
 All three forms produce the same representation; the semantic functions below operate only on de Bruijn index pairs.
@@
 (2) $\mathtt{supers}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{base}}),\;p_{\mathrm{overlay}})\;\big|\;p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{overlay}}\in\mathtt{overlays}(p_{\mathrm{base}})\;\bigr\}$
-The $\mathtt{supers}$ formula depends on two functions: the identity-sharing paths $\mathtt{overlays}$ and the one-hop reference targets $\mathtt{bases}$. We define $\mathtt{overlays}$ first.
+Here $\mathtt{bases}^{*}$ denotes the reflexive-transitive closure of $\mathtt{bases}$.
+The $\mathtt{supers}$ formula depends on two functions: the identity-sharing paths $\mathtt{overlays}$ and the one-hop reference targets $\mathtt{bases}$.
+We define $\mathtt{overlays}$ first.
 Overlays
-Intuitively, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overlays}(p)$ collects all such paths that share the same identity as $p$.
+Intuitively, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overlays}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge.
 Concretely, the overlays of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
@@
 Bases
-Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references—analogous to the direct base classes in object-oriented languages.
+Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references, analogous to the direct base classes in object-oriented languages.
 Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overlays}$ one step:
@@
 It takes an inheritance-site path $p_{\mathrm{site}}$, a definition-site path $p_{\mathrm{def}}$, a de Bruijn index $n$, and downward projections $\ell_{\mathrm{down},*}$.
-Resolution proceeds in two phases: first, the qualified this resolution function $\mathtt{this}$ performs $n$ upward steps starting from the enclosing scope $\mathtt{init}(p_{\mathrm{def}})$, mapping the definition-site path to inheritance-site paths; then the downward projections $\ell_{\mathrm{down},*}$ are appended:
+Resolution proceeds in two phases: $\mathtt{this}$ performs $n$ upward steps starting from the enclosing scope $\mathtt{init}(p_{\mathrm{def}})$, mapping the definition-site path to inheritance-site paths; then the downward projections $\ell_{\mathrm{down},*}$ are appended:
 (5) $\mathtt{resolve}(p_{\mathrm{site}},\;p_{\mathrm{def}},\;n,\;\ell_{\mathrm{down},*})=\\ \bigl\{\;p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\ell_{\mathrm{down},k}\;\big|\;p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\;\mathtt{init}(p_{\mathrm{def}}),\;n)\;\bigr\}$
-When multiple routes exist (due to multi-path inheritance), they may yield different target paths; all are collected. We now define $\mathtt{this}$.
+When multiple routes exist due to multi-path inheritance, they may yield different target paths; all are collected. We now define $\mathtt{this}$.
 Qualified this resolution
-Intuitively, $\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” The design of $\mathtt{this}$ reconciles two historically opposing approaches to scope resolution.
-Early Lisp implementations used dynamic scope (McCarthy, 1978), where variable references are resolved against the run-time call stack; this is flexible but unpredictable.
-Scheme (Sussman and Steele Jr., 1975) corrected this with lexical scope, where a de Bruijn index resolves in a single step by indexing into the statically determined environment; this is predictable but rigid.
-Scala (Odersky et al., 2004) and the NixOS module system (NixOS Contributors, 2024) follow the lexical approach: each this reference is bound to a single statically known path.
-When inheritance introduces multiple inheritance routes to the same scope, Scala rejects the ambiguity at compile time (Appendix C) and the NixOS module system rejects the duplicate option declaration (Appendix D)—neither can express that all routes are equally valid.
-Overlay-Calculus combines the strengths of both approaches.
-The reference itself is lexically determined at the definition site as a de Bruijn index, so what $\mathtt{this}$ looks up is predictable.
-But because inheritance interleaves scopes from different overlays, the result depends on the inherited tree structure: $\mathtt{this}$ walks upward one scope level at a time, consulting $\mathtt{supers}$ at each level, and naturally tracks a set of inheritance-site paths rather than a single one.
-All routes contribute equally to the inherited result, and since inheritance is idempotent, duplicated contributions are harmless.
-The only observable is property presence, making multi-path resolution well-defined without disambiguation.
-Concretely, each step finds, among the supers of every path in the frontier $S$, those whose overlay component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier.
+$\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Each step finds, among the supers of every path in the frontier $S$, those whose overlay component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier.
 After $n$ steps the frontier contains the answer:
-(6) $\small\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
+(6) $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
 At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a non-negative integer, the recursion terminates.
-Alternative approaches considered
-Our first three implementations of the Overlay language used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972).
-All three produced subtle bugs when inheritance introduced multiple inheritance routes to the same enclosing scope.
-In retrospect, the difficulty was inevitable.
-A closure captures a single environment, so a fixed-point combinator solves for a single self-reference; when multiple routes exist, it attempts to unify them into one value.
-A stack-based environment is a linear chain—each scope level has one parent—so looking up a de Bruijn index returns a single value.
-Both data structures embed a single-valued assumption that is incompatible with the multi-path situation that arises when independent overlays inherit from the same enclosing scope (Appendix C).
-Equation (6) emerged from recognizing this: it operates on paths as first-order data and tracks a frontier set $S$ rather than a single current scope.
 This completes the chain of definitions needed to compute $\mathtt{properties}(p)$.
@@
 There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments.
-This is a deliberate design choice, not an omission.
 In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur.
-In Overlay-Calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything—it is the value.
+In overlay-calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything: it is the value.
 The functions $\mathtt{properties}$, $\mathtt{supers}$, etc.
@@
 The corresponding operator on the product lattice of path sets is monotone, and its least fixed point exists by the Knaster–Tarski theorem (Tarski, 1955).
-The detailed well-definedness argument—identifying the lattice, verifying positivity of each equation, and handling the transitive closure $\mathtt{bases}^{*}$—is given in Appendix E.
-When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point—analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness.
+The detailed well-definedness argument, which identifies the lattice, verifies positivity of each equation, and handles the transitive closure $\mathtt{bases}^{*}$, is given in Appendix C.
+When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point, analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness.
 This observation model is inspired by F-coalgebras (Rutten, 2000), where identity is determined by observable behavior rather than by internal structure.
-4. Translation from $\lambda$-Calculus
+4. Embedding the $\lambda$-Calculus
 The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to overlay-calculus.
@@
 $V_{1}\;V_{2}$ (tail call)   $\longrightarrow$ $\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
-A $\lambda$-bound variable $x$ with de Bruijn index $n$ (the number of enclosing $\lambda$-abstractions between the reference and its binder) becomes $\uparrow^{\!n}.\mathrm{argument}$.
-The indexed reference $\uparrow^{\!n}$ traverses $n$ scope levels upward to reach the abstraction that binds $x$, giving access to the $\mathrm{argument}$ slot after all inheritance.
-A let-bound variable $x$ becomes the lexical reference $x.\mathrm{result}$, projecting the application result from the sibling property $x$ in the enclosing record.
+[…]
 An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$; we call this the abstraction shape.
@@
 The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to overlay-calculus: every ANF term has an image, and the construction is compositional.
-Since the $\lambda$-calculus is Turing complete, so is Overlay-Calculus; Appendix A further shows that Overlay-Calculus primitives directly encode a random-access machine.
+Since the $\lambda$-calculus is Turing complete, so is overlay-calculus.
 An elementary semantics of the $\lambda$-calculus
@@
 Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices.
-Here, the semantic domain is a product of powerset lattices—no function spaces, no continuity conditions, no inverse limits.
-The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix E), which requires only set-theoretic monotonicity—strictly weaker than Scott continuity.
+Here, the semantic domain is a product of powerset lattices, with no function spaces, no continuity conditions, and no inverse limits.
+The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix C), which requires only set-theoretic monotonicity, which is strictly weaker than Scott continuity.
   • No substitution.
 Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi.
-Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree—no substitution occurs at any point.
+Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree; no substitution occurs at any point.
   • No reduction rules.
@@
 The semantics is observational: one queries path membership, and the mutually recursive equations unfold on demand.
-That a semantics of the $\lambda$-calculus can be given without function spaces is perhaps unexpected: function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function.
-Section 4.2 shows that the convergence behavior of the translated overlay tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate (a term converges iff it has a head normal form) and fully abstract (two terms are overlay-contextually equivalent iff they have the same Böhm tree).
-Combining these two facts—elementary machinery and full abstraction—yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics that lives entirely in naive set theory.
+The key step is that function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function; no function spaces are required.
+Section 4.1 shows that the convergence behavior of the translated overlay tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate, meaning a term converges iff it has a head normal form, and fully abstract, meaning two terms are overlay-contextually equivalent iff they have the same Böhm tree.
+Combining these two facts, elementary machinery and full abstraction, yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics that lives entirely in naive set theory.
 The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem.
 No continuous lattices, no inverse limits, no Scott topology, no game-theoretic machinery.
-Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Hyland and Ong’s game semantics (Hyland and Ong, 2000) achieves full abstraction but at the cost of even greater mathematical complexity.
+Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Abramsky and McCusker’s game semantics (Abramsky and McCusker, 1995) achieves full abstraction but at the cost of even greater mathematical complexity.
 The overlay tree semantics achieves full abstraction with strictly simpler foundations than either.
-The reverse direction is non-trivial, as we discuss next.
-4.1. Expressive Asymmetry
-The translation above consists of five rules; the full computational power of the $\lambda$-calculus embeds into Overlay-Calculus with no additional machinery.
-The reverse direction—encoding Overlay-Calculus in pure $\lambda$-calculus—is non-trivial.
-Defining the meaning of Overlay-Calculus requires machinery for lazy allocation, recursive data structures, and set-valued resolution that have no direct counterparts in the $\lambda$-calculus.
-This sophistication is not accidental—it reflects three features:
-  1. (1)
-    Self-referential records with lazy evaluation.
-An overlay’s properties can reference other properties and enclosing scopes through qualified this without explicit binding.
-In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
-  2. (2)
-    Open recursion with symmetric inheritance.
-Inheritance merges definitions from independent sources, all sharing the same inherited overlay tree.
-Encoding this in the $\lambda$-calculus requires maintaining an extensible dictionary of methods that can be merged from multiple directions.
-Cook’s denotational semantics of inheritance (Cook, 1989) devoted an entire dissertation to formalizing this pattern in a functional setting.
-  3. (3)
-    Commutative, idempotent merge.
-Inheritance in Overlay-Calculus is symmetric: a record’s elements form a set, so reordering them produces the same observable result.
-In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery.
-Overlay-Calculus is not the $\lambda$-calculus in disguise: it provides primitive operations for computational patterns that are derivable but complex in the $\lambda$-calculus.
-The reverse asymmetry also exists: nested expressions such as f(gx)f\;(g\;x) are primitive in the $\lambda$-calculus but require manual A-normal form conversion in Overlay-Calculus—the intermediate result gxg\;x must be given an explicit name.
-However, ANF conversion is a trivial, mechanical transformation (Flanagan et al., 1993); a surface language built on Overlay-Calculus can support nested expressions as syntactic sugar without extending the calculus itself.
 4.1. Böhm Tree Correspondence
@@
 We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
-An overlay tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties—the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$.
+An overlay tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$.
 We write $T{\Uparrow}$ if no such depth exists.
@@
 $\mathcal{T}(M)\approx\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
-Together, Adequacy and Full Abstraction establish that the overlay tree semantics is a computation mechanism for Böhm trees (Barendregt, 1984)—the canonical fully abstract model of the lazy $\lambda$-calculus—using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory.
-This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract—it identifies terms that observational equivalence distinguishes (Plotkin, 1975).
-The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix F.
-5. Case Study: The Expression Problem
-The Expression Problem (Wadler, 1998) asks how to extend a data type with both new cases and new operations, without modifying existing code.
-Existing solutions require object algebras (d.
+Together, Adequacy and Full Abstraction establish that the overlay tree semantics is a computation mechanism for Böhm trees (Barendregt, 1984), the canonical fully abstract model of the lazy $\lambda$-calculus, using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory.
+This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract, as it identifies terms that observational equivalence distinguishes (Plotkin, 1975).
+The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix D.
+4.2. Expressive Asymmetry
+Two $\lambda$-terms that are Böhm-tree equivalent—and therefore indistinguishable in the $\lambda$-calculus—can be separated in overlay-calculus by overlaying new observable projections onto an existing definition via open extension.
+By Felleisen’s expressiveness criterion (Felleisen, 1991), overlay-calculus is strictly more expressive than the $\lambda$-calculus: the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
+The formal definitions, constructions, and proofs are given in Appendix A; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
+The underlying mechanism is open extension: overlay definitions compose via open recursion, so existing definitions gain new observable properties without modification.
+In the $\lambda$-calculus, a function’s result is closed—there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
+5. Case Study: Nat Arithmetic
+We implemented boolean logic and natural number arithmetic, both unary and binary, as overlay files (Yang, 2025b).
+This section traces the unary Nat case in detail.
+The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names are constructors and operations; lowerCamelCase names are fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action.
+Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
+NatData
+Following the Church-encoding convention (Yang, 2025b), we define a factory containing the abstract product type and its two constructors.
+A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field.
+The file name $\mathrm{NatData}$ is itself the top-level scope:
+$\mathrm{NatData}\mapsto\{\;$ $\mathrm{NatFactory}\mapsto\{\mathrm{Product}\mapsto\{\},$
+$\quad\mathrm{Zero}\mapsto\{\mathrm{Product}\},$
+$\quad\mathrm{Successor}\mapsto\{\mathrm{Product},\;\mathrm{predecessor}\mapsto\{\mathrm{Product}\}\}\},$
+$\mathrm{Nat}\mapsto\{\mathrm{NatFactory},\;\mathrm{NatFactory}.\mathrm{Product}\}\}$
+NatPlus
+Addition is a separate overlay file that inherits $\mathrm{NatData}$ and extends each constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field.
+The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
+$\mathrm{NatPlus}\mapsto\{\;$ $\mathrm{NatData},$
+$\mathrm{NatFactory}\mapsto\{$
+$\quad\mathrm{Product}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{sum}\mapsto\{\mathrm{Product}\}\}\},$
+$\quad\mathrm{Zero}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{addend}\mapsto\{\mathrm{Product}\},\;\mathrm{sum}\mapsto\{\mathrm{addend}\}\}\},$
+$\quad\mathrm{Successor}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{addend}\mapsto\{\mathrm{Product}\},$
+$\qquad\mathrm{\_increasedAddend}\mapsto\{\mathrm{Successor},\;\mathrm{predecessor}\mapsto\{\mathrm{addend}\}\},$
+$\qquad\mathrm{\_recursiveAddition}\mapsto\{\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}.\mathrm{Plus},$
+$\qquad\quad\mathrm{addend}\mapsto\{\mathrm{\_increasedAddend}\}\},$
+$\qquad\mathrm{sum}\mapsto\{\mathrm{\_recursiveAddition}.\mathrm{sum}\}\}\}\}\}$
+NatEquality
+To test that $2+3=5$, we needed equality on Nat values.
+Equality requires case analysis on the constructor; we use the Visitor pattern.
+$\mathrm{NatVisitor}$ adds a $\mathrm{Visitor}$ operation to each constructor, with a $\mathrm{Visit}$ field that delegates to the appropriate $\mathrm{VisitZero}$ or $\mathrm{VisitSuccessor}$ branch:
+$\mathrm{NatVisitor}\mapsto\{\;$ $\mathrm{NatData},$
+$\mathrm{NatFactory}\mapsto\{$
+$\quad\mathrm{Product}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{Visit}\mapsto\{\}\}\},$
+$\quad\mathrm{Zero}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitZero}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitZero}\}\}\},$
+$\quad\mathrm{Successor}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitSuccessor}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitSuccessor}\}\}\}\}\}$
+$\mathrm{BooleanData}$ defines the output sort:
+$\mathrm{BooleanData}\mapsto\{\;$ $\mathrm{BooleanFactory}\mapsto\{\mathrm{Product}\mapsto\{\},\;\mathrm{True}\mapsto\{\mathrm{Product}\},\;\mathrm{False}\mapsto\{\mathrm{Product}\}\},$
+$\mathrm{Boolean}\mapsto\{\mathrm{BooleanFactory},\;\mathrm{BooleanFactory}.\mathrm{Product}\}\}$
+$\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each constructor.
+$\mathrm{Zero}$ is equal only to another $\mathrm{Zero}$; $\mathrm{Successor}$ is equal only to another $\mathrm{Successor}$ with an equal predecessor.
+Case analysis on $\mathrm{other}$ is done via its $\mathrm{Visitor}$:
+$\mathrm{NatEquality}\mapsto\{\;$ $\mathrm{NatVisitor},\;\mathrm{BooleanData},$
+$\mathrm{NatFactory}\mapsto\{$
+$\qquad\mathrm{Product}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product}\},$
+$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
+$\qquad\mathrm{Zero}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product}\},$
+$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
+$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{True}\}\},$
+$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
+$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
+$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\},$
+$\qquad\mathrm{Successor}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product},\;\mathrm{predecessor}\mapsto\{\mathrm{Product}\}\},$
+$\hskip 18.49988pt\mathrm{RecursiveEquality}\mapsto\{\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}.\mathrm{Equal},$
+$\hskip 18.49988pt\;\mathrm{other}\mapsto\{\mathrm{Equal}.\mathbf{this}.\mathrm{other}.\mathrm{predecessor}\}\},$
+$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
+$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
+$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{RecursiveEquality}.\mathrm{equal}\}\},$
+$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
+$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\}\}\}$
+Composing $\mathrm{NatPlus}$ with $\mathrm{NatEquality}$ by inheriting both, with no modifications to either file, gives the resulting Nat values both $\mathrm{Plus}$ and $\mathrm{Equal}$ automatically.
+The concrete numerals live in a shared file that inherits only $\mathrm{NatData}$:
+$\mathrm{NatConstants}\mapsto\{\;$ $\mathrm{NatData},$
+$\mathrm{One}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
+$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Zero}\}\},$
+$\mathrm{Two}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
+$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{One}\}\},$
+$\mathrm{Three}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
+$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Two}\}\},$
+$\mathrm{Four}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
+$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Three}\}\},$
+$\mathrm{Five}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
+$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Four}\}\}\}$
+The arithmetic test inherits $\mathrm{NatConstants}$ together with the operations it needs:
+$\mathrm{Test}\mapsto\{\;$ $\mathrm{NatConstants},\;\mathrm{NatPlus},\;\mathrm{NatEquality},$
+$\mathrm{Addition}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Two}.\mathrm{Plus},\;\mathrm{addend}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Three}\}\},$
+$\mathrm{Test2plus3}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Five}.\mathrm{Equal},\;\mathrm{other}\mapsto\{\mathrm{Addition}.\mathrm{sum}\}\}\}$
+Querying $\mathrm{Test}.\mathrm{Test2plus3}.\mathrm{equal}$ via the host-language API returns $\mathrm{True}$.
+No existing file was modified.
+$\mathrm{NatEquality}$, a new operation, and $\mathrm{Boolean}$, a new data type, each required only a new file: extensibility in both dimensions of the Expression Problem (Wadler, 1998).
+Tree-level inheritance and return types
+Each operation is an independent overlay that extends a factory’s subtree.
+Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited overlays.
+Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory.
+Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either file.
+In object algebras (d.
 S.
-Oliveira and Cook, 2012; Wang and d.
-S.
-Oliveira, 2016), finally tagless interpreters (Carette et al., 2009), or type-class machinery.
-In a typed setting the problem additionally requires static type safety; Overlay-Calculus is untyped, so we focus on the extensibility aspect.
-Suppose a standard library provides a $\mathrm{binnat}$ module for binary natural arithmetic (see Appendix B).
-We define an expression language with an evaluation operation:
-$\{\quad\mathrm{expression}$ $\mapsto\{\mathrm{Constant}\mapsto\{\mathrm{magnitude}\mapsto\{\}\},\;\mathrm{Addition}\mapsto\{\mathrm{left}\mapsto\{\},\;\mathrm{right}\mapsto\{\}\}\},$
-$\mathrm{evaluation}$ $\mapsto\{\mathrm{Constant}\mapsto\{\mathrm{expression}.\mathrm{Constant},\;\mathrm{outcome}\mapsto\mathrm{Constant}.\mathbf{this}.\mathrm{magnitude}\},$
-$\qquad\mathrm{Addition}\mapsto\{\mathrm{expression}.\mathrm{Addition},\;$
-$\qquad\qquad\mathrm{left\_outcome}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{left}.\mathrm{outcome},$
-$\qquad\qquad\mathrm{right\_outcome}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{right}.\mathrm{outcome},$
-$\qquad\qquad\mathrm{sum}\mapsto\{\mathrm{binnat}.\mathrm{Addition},\;\mathrm{augend}\mapsto\mathrm{sum}.\mathbf{this}.\mathrm{left\_outcome},\;\mathrm{addend}\mapsto\mathrm{sum}.\mathbf{this}.\mathrm{right\_outcome}\},$
-$\qquad\qquad\mathrm{outcome}\mapsto\mathrm{sum}.\mathrm{outcome}\}\},$
-$\mathrm{binnat}$ $\mapsto\{\mathrm{Addition}\mapsto\{\mathrm{augend}\mapsto\{\},\;$
-$\qquad\qquad\mathrm{addend}\mapsto\{\},\;\mathrm{outcome}\mapsto\{\}\}\}\quad\}$
-A new operation (e.g., display) is added by defining a new module that inherits from the same $\mathrm{expression}$ schemas:
-$\mathrm{display}\mapsto\{\mathrm{Constant}$ $\mapsto\{\mathrm{expression}.\mathrm{Constant},\;\mathrm{representation}\mapsto\mathrm{Constant}.\mathbf{this}.\mathrm{magnitude}\},$
-$\mathrm{Addition}$ $\mapsto\{\mathrm{expression}.\mathrm{Addition},\;$
-$\qquad\mathrm{representation}\mapsto\{$
-$\qquad\qquad\mathrm{left}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{left}.\mathrm{representation},\;$
-$\qquad\qquad\mathrm{right}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{right}.\mathrm{representation}\}\}\}$
-A new case (e.g., negation) is added with its operation handlers and an empty API for the arithmetic primitive:
-$\mathrm{negation}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{operand}\mapsto\{\}\}\},$
-$\mathrm{negation\_evaluation}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{negation}.\mathrm{Negation},\;$
-$\qquad\mathrm{operand\_outcome}\mapsto\mathrm{Negation}.\mathbf{this}.\mathrm{operand}.\mathrm{outcome},$
-$\qquad\mathrm{negated}\mapsto\{\mathrm{binnat}.\mathrm{Negation},\;\mathrm{operand}\mapsto\mathrm{operand\_outcome}\},$
-$\qquad\mathrm{outcome}\mapsto\mathrm{negated}.\mathrm{outcome}\}\},$
-$\mathrm{binnat}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{operand}\mapsto\{\},\;\mathrm{outcome}\mapsto\{\}\}\}$
-All modules compose freely; none requires modification of existing definitions.
-A complete executable version of this example—including both the evaluation and display operations, a negation extension, and their free inheritance—is available in the implementation (Yang, 2025b), along with test cases verifying evaluation and display for nested expressions mixing old and new cases.
-Note that $\mathrm{magnitude}\mapsto\{\}$ in $\mathrm{Constant}$ resembles a type annotation but is purely structural in Overlay-Calculus: the evaluator works as long as the required properties appear somewhere in the inheritance chain, regardless of any schema declarations.
+Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
+Multi-path qualified this (equation 6) is essential.
+In the test file above, $\mathrm{NatPlus}.\mathrm{Successor}$ and $\mathrm{NatEquality}.\mathrm{Successor}$ both inherit $\mathrm{NatData}.\mathrm{Successor}$.
+When the two operations are composed, the resulting $\mathrm{Successor}$ inherits the $\mathrm{NatData}$ schema through two independent routes.
+The reference $\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties.
+In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix B), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously.
+The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.
+Church-encoded Nats are tries
+A Church-encoded Nat is a record whose structure mirrors the constructor path used to build it: $\mathrm{Zero}$ is a flat record; $\mathrm{S}(\mathrm{Zero})$ is a record with a $\mathrm{predecessor}$ field pointing to a $\mathrm{Zero}$ record; and so on.
+This is exactly the shape of a trie.
+A scope that simultaneously inherits two Nat values is a trie union.
+The following test applies $\mathrm{Plus}$ to a trie union of $\{1,2\}$ with addend a trie union of $\{3,4\}$, then compares the result to itself with $\mathrm{Equal}$:
+$\mathrm{CartesianTest}\mapsto\{\;$ $\mathrm{NatConstants},\;\mathrm{NatPlus},\;\mathrm{NatEquality},$
+$\mathrm{OneOrTwo}\mapsto\{\mathrm{CartesianTest}.\mathbf{this}.\mathrm{One},\;\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Two}\},$
+$\mathrm{ThreeOrFour}\mapsto\{\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Three},\;\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Four}\},$
+$\mathrm{Result}\mapsto\{\mathrm{OneOrTwo}.\mathrm{Plus},\;\mathrm{addend}\mapsto\{\mathrm{ThreeOrFour}\}\},$
+$\mathrm{Check}\mapsto\{\mathrm{Result}.\mathrm{sum}.\mathrm{Equal},\;\mathrm{other}\mapsto\{\mathrm{Result}.\mathrm{sum}\}\}\}$
+$\mathrm{Result}.\mathrm{sum}$ evaluates to the trie $\{4,5,6\}$, the set of all pairwise sums.
+$\mathrm{Equal}$ then applies to every pair drawn from $\{4,5,6\}\times\{4,5,6\}$; querying $\mathrm{CartesianTest}.\mathrm{Check}.\mathrm{equal}$ via the host-language API returns $\{\mathrm{True},\mathrm{False}\}$.
+The Cartesian product arises directly from the semantic equations.
+$\mathrm{OneOrTwo}$ inherits both $\mathrm{One}$ and $\mathrm{Two}$, so $\mathtt{bases}(\mathrm{OneOrTwo})$ contains two paths.
+When $\mathrm{Result}$ inherits $\mathrm{OneOrTwo}.\mathrm{Plus}$, the $\mathtt{bases}$ of $\mathrm{Result}$ include the $\mathrm{Plus}$ of both $\mathrm{One}$ and $\mathrm{Two}$.
+Each $\mathrm{Plus}$ resolves its $\mathrm{addend}$ via $\mathtt{this}$, reaching $\mathrm{ThreeOrFour}$, which is itself a trie union of $\mathrm{Three}$ and $\mathrm{Four}$: the $\mathtt{supers}$ of the $\mathrm{addend}$ path contain both values.
+Since $\mathrm{Successor}.\mathrm{Plus}$ recursively delegates to $\mathrm{predecessor}.\mathrm{Plus}$ with an incremented addend, and $\mathrm{predecessor}$ also resolves through the trie union, the recursion branches at every trie node.
+The result $\mathrm{sum}$ therefore collects all paths reachable through any combination of the two operands, which is the Cartesian product.
+The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: $\mathtt{overlays}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.
 6. Discussion
-6.1. Relationship to $\lambda$-Calculus
-Three facts establish the relationship between Overlay-Calculus and the $\lambda$-calculus.
-More elementary and stronger than Scott domains.
-The semantics induced by the translation (Section 4) requires only powerset lattices and the Knaster–Tarski fixed-point theorem—no function spaces, no inverse limits, no Scott continuity.
-Yet this semantics is fully abstract (Theorem 4.3), whereas Scott’s $D_{\infty}$ model (Scott, 1976; Wadsworth, 1976) is adequate but not fully abstract—it identifies terms that observational equivalence distinguishes (Plotkin, 1975).
-$\lambda$-Calculus embeds into Overlay-Calculus.
-The five-rule translation of Section 4 maps every ANF $\lambda$-term to an overlay tree.
-The translation uses only single-target qualified this—the full multi-target semantics is not needed to embed $\lambda$-calculus.
-Multi-target qualified this is the natural primitive.
-If the semantic functions of Section 3 were restricted to single-target resolution (as in DOT’s single-valued self variable (Amin et al., 2016)), the formulas would need to reject various structurally legitimate cases—adding complexity, not removing it.
-Multi-target qualified this is not an artificial extension; it is what falls out when no restrictions are imposed.
-The $\lambda$-calculus translation uses only single-target qualified this, suggesting that $\lambda$-calculus itself represents an artificially restricted fragment of the natural model.
-Appendix H presents a detailed analysis of the asymmetry in problem difficulty between the two systems.
-6.2. Observed Phenomena
-Several properties emerge in practice without having been designed:
-  • The Expression Problem does not arise (Section 5, Appendix H).
-  • Programs are CPS-agnostic: empty record slots serve as dependency injection points, making the same program text CPS or direct-style depending on assembly (Appendix G).
-  • Records are tries, and inheritance is trie union—an immutable analogue of random-access memory (Appendix A).
-  • Unlike SKI combinators (Schönfinkel, 1924; Curry and Feys, 1958), which are isomorphic to $\lambda$-calculus in problem difficulty, Overlay-Calculus exhibits genuine asymmetry (Section 4.1).
-6.3. Properties of Inheritance
-Because records are sets, inheritance is inherently commutative and idempotent, making Overlay-Calculus transparent to linearization order.
-Deduplication of inherited overlays is purely an optimization.
-6.4. Relationship to Practical Systems
-Several classes of declarative systems that have emerged independently in practice turn out to be instances of Overlay-Calculus.
-Configuration and object-oriented languages
-The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024)—whose recursive attribute merging motivated this work—can be understood as an implementation of Overlay-Calculus with additional features (type checking, error reporting, FFI to the Nix language).
-Detailed comparisons with CUE, Jsonnet, Dhall, with mixin and trait calculi, and with the NixOS module system’s lazy evaluation limitations under multiple inheritance routes are given in Section 7.
-Effect systems
-Algebraic effects with handlers (Plotkin and Power, 2003; Pretnar, 2015) and monad transformers structure computational effects through composition.
-The CPS-agnostic property (Appendix G) shows how Overlay-Calculus naturally supports effect handler composition: empty API slots can be filled with effect handlers (CPS interpretation) or direct implementations (direct-style interpretation).
-The same pattern appears in algebraic effect systems, without explicit monad machinery.
+6.1. Emergent Phenomena
+Several features of overlay-calculus are primitive yet have no direct counterpart in the $\lambda$-calculus. Encoding any one of them requires pervasive, non-local changes to program structure.
+  1. (1)
+    Self-referential records with lazy evaluation.
+An overlay’s properties can reference other properties and enclosing scopes through qualified this without explicit binding.
+In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (Boudol, 2004) (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
+  2. (2)
+    Open recursion with symmetric inheritance.
+Inheritance merges definitions from independent sources, all sharing the same inherited overlay tree.
+Encoding this in the $\lambda$-calculus requires maintaining an extensible dictionary of methods that can be merged from multiple directions.
+Cook’s denotational semantics of inheritance (Cook, 1989) devoted an entire dissertation to formalizing this pattern in a functional setting.
+  3. (3)
+    Commutative, idempotent merge.
+Inheritance in overlay-calculus is symmetric: a record’s elements form a set, so reordering them produces the same observable result.
+In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery (Harper and Pierce, 1991; Cardelli, 1994).
+  4. (4)
+    Relational semantics.
+Every overlay-calculus operation is defined by recursion on overlay trees, not on individual values.
+When an overlay tree is viewed as a trie, the correspondence with relational algebra becomes precise: a path from the root to a leaf encodes a single value (e.g., $\mathrm{Successor}\to\mathrm{predecessor}\to\mathrm{Zero}$ encodes the natural number $1$); a trie union encodes a set of such values, i.e., a unary relation.
+An operation that takes two trie-union operands traverses every combination of their paths through $\mathtt{bases}$ (equation 4), which is a cross-join; the transitive closure $\mathtt{bases}^{*}$ in $\mathtt{supers}$ (equation 2) computes the union of all reachable paths; and $\mathtt{properties}$ (equation 1) projects the observable labels.
+The carrier is a trie rather than a flat relation, and observation is coinductive—the observer drives computation by querying progressively deeper paths—rather than Datalog’s bottom-up materialization, but the relational structure of the computation is the same.
+In the Case Study (Section 5), the $\mathrm{Plus}$ and $\mathrm{Equal}$ operations are defined on single Nat values.
+When applied to set-valued overlays $\{1,2\}$ and $\{3,4\}$, the same definitions, without any modification, produce the set of all pairwise sums $\{4,5,6\}$.
+This is relational semantics in action: each operation distributes over every combination of its inputs.
+The behavior is native: the semantic equations of Section 3 resolve every property to a set of paths, so single-valued and set-valued computation are the same mechanism.
+No code change is needed to move from one to the other.
+  5. (5)
+    Function color blindness.
+We implemented a web application as overlay files (Yang, 2025b), initially targeting Python’s synchronous sqlite3 and http.server modules.
+The business logic overlay file contained no async-specific code.
+Later, we replaced the FFI layer with one built on aiosqlite and starlette; the business logic overlay file did not change.
+This is an instance of function color blindness (Nystrom, 2015): async and sync functions are distinct types in the host language, yet the overlay file is agnostic to the distinction because its empty property slots are filled at assembly time, with coroutine handlers under asyncio, or with direct implementations under synchronous execution.
+More generally, empty slots can be filled with effect handlers (Plotkin and Power, 2003; Pretnar, 2015) under a CPS interpretation, or with direct implementations under a direct-style interpretation, without explicit monad machinery.
+Each of these features is primitive in overlay-calculus.
+In the $\lambda$-calculus, encoding them requires pervasive changes to program structure: every scope must thread a mutable store for lazy allocation, every method invocation must consult a dynamically constructed dictionary for open recursion, and every composition must be wrapped in order-insensitive infrastructure for commutativity.
+These are global, cascading transformations, not local rewrites of individual terms.
+6.2. Classification of Semantic Variants
+The goals of declarative programming rule out several properties of the model: separate primitives for methods, classes, types, and objects; mutable reference cells; domain-theoretic approximation; and asymmetric override.
+Ruling these out leaves a minimal syntax of three constructs: records, definitions, and inheritance.
+Two degrees of freedom remain.
+These dimensions are independent, and each combination yields a coherent semantics.
+They arise in any system whose definitions form a tree of named scopes, where each node is a path and references are path-dependent, even in systems ruled out above, such as Scala, which has separate primitives for methods, classes, objects, and types.
+Existing designs occupy several of the resulting cells:
+No deep merge Deep merge
+Early binding CUE (van Lohuizen, 2019)
+Dynamic scope union file systems, Jsonnet (Cunningham, 2014)
+Lexical scope Scala (Odersky et al., 2004) NixOS modules (NixOS Contributors, 2024)
+Multi-target $\mathtt{this}$ overlay-calculus
+Early binding
+With early binding, references resolve at definition time to a fixed path, independently of how the record is later inherited.
+CUE (van Lohuizen, 2019) exemplifies this: field references are statically resolved, and a field in a nested record cannot refer back to properties of its enclosing context contributed by later composition.
+Early binding combined with deep merge is coherent and practically useful, but cross-overlay cooperation requires explicit wiring rather than inheritance: a definition cannot observe properties contributed by overlays composed after it.
+Dynamic scope
+With dynamic scope, references resolve against the context at the point of use rather than the definition site.
+Early Lisp implementations used dynamic scope (McCarthy, 1978), where variable references are resolved against the run-time call stack.
+Union file systems exhibit the same property at the file-system level: a symbolic link’s relative path resolves against the mount context at dereference time, not against the layer that defined the link.
+Jsonnet (Cunningham, 2014) follows this approach: self always refers to the final merged object, not the object at the definition site.
+When the Scala multi-path example (Appendix B) is translated to Jsonnet, it does not error, but outer resolves to the merged MyObjectA rather than to Object1 or Object2 separately, silently discarding the distinction between the two inheritance routes.
+Dynamic scope is flexible but loses the ability to track which inheritance route a reference came from.
+Deep merge
+Without the second case of equation (3), $\mathtt{overlays}(p)$ would always return $\{p\}$, a single element.
+When two independently authored overlays both define a label with the same name at the same scope level, only the definition at the literal path $p$ would contribute to $\mathtt{supers}$ and hence to $\mathtt{properties}$; the other definitions would be silently invisible.
+This is the situation in most composition systems: pytest permits at most one fixture per name per scope; trait systems (Schärli et al., 2003) flag same-name methods as conflicts.
+Under any of these rules, two independently authored overlays that both extend an entity with the same name, which is the situation that arises naturally when solving the Expression Problem (Wadler, 1998), cannot be composed.
+The non-trivial case of equation (3) removes this limitation: it collects all paths that share the same label and merges their definitions rather than shadowing each other.
+This is the record-level counterpart of method override in object-oriented languages, and the mechanism by which the calculus admits extensibility in both dimensions of the Expression Problem (Wadler, 1998) (Section 5).
+Multi-target qualified this
+The qualified this reference $\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Scheme (Sussman and Steele Jr., 1975) introduced lexical scope, where a de Bruijn index resolves in a single step by indexing into the statically determined environment; this is predictable but rigid.
+Scala (Odersky et al., 2004) and the NixOS module system (NixOS Contributors, 2024) follow the lexical approach: each this reference is bound to a single statically known path.
+When inheritance introduces multiple inheritance routes to the same scope, Scala rejects the ambiguity at compile time (Appendix B) and the NixOS module system rejects the duplicate option declaration (Appendix B); neither can express that all routes are equally valid.
+Overlay-calculus combines the predictability of lexical scope with the flexibility of dynamic scope.
+The reference itself is lexically determined at the definition site as a de Bruijn index, so what $\mathtt{this}$ looks up is predictable.
+But because inheritance interleaves scopes from different overlays, the result depends on the inherited tree structure: $\mathtt{this}$ walks upward one scope level at a time, consulting $\mathtt{supers}$ at each level, and naturally tracks a set of inheritance-site paths rather than a single one.
+All routes contribute equally to the inherited result, and since inheritance is idempotent, duplicated contributions are harmless.
+The only observable is property presence, making multi-path resolution well-defined without disambiguation.
+Our first three implementations of the Overlay language used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972).
+All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope.
+The difficulty was structural, not incidental: a closure captures a single environment, so a fixed-point combinator solves for a single self-reference; a stack-based environment is a linear chain where each scope level has one parent.
+Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix B).
+Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope.
+If the equations were restricted to single-target resolution, as in DOT’s single-valued self variable (Amin et al., 2016), they would need to reject structurally legitimate cases, adding complexity rather than removing it.
+Consequences of the alternative choices
+Choosing differently in any dimension requires introducing machinery that is absent from the equations of Section 3.
+With early binding, the equations must freeze reference targets at definition time, so a definition cannot observe properties contributed by later-composed overlays.
+With dynamic scope, the equations must resolve references against a runtime context rather than the definition-site structure, losing the ability to distinguish inheritance routes.
+Restricting $\mathtt{overlays}(p)$ to $\{p\}$ means same-label definitions from different inheritance sources no longer share identity, requiring either a method resolution order (MRO) (Barrett et al., 1996) to linearize them or a shadowing rule to select one and discard the rest, which is why systems without deep merge rarely support path-dependent multiple inheritance in practice.
+Restricting $\mathtt{this}$ to a single path requires a mechanism to select among inheritance routes (Scala rejects the ambiguity at compile time; the NixOS module system raises a duplicate-declaration error), and loses contributions from all non-selected routes.
+These mechanisms, including early-binding freezing, runtime context threading, MRO, shadowing, route selection, and error reporting, are more artificial than their absence.
+The equations of Section 3 contain none of them.
+7. Related Work
+Minimal Turing-complete models
+Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus reduces to three combinators $S$, $K$, II (II is redundant: $I=SKK$).
+The bracket abstraction that compiles $\lambda$-terms to SKI is mechanical but incurs exponential blow-up in term size; no new computational patterns emerge beyond those already present in the $\lambda$-calculus.
+Dolan (Dolan, 2013) showed that the x86 mov instruction alone is Turing complete by exploiting addressing modes as implicit arithmetic on mutable memory; the resulting programs are correct but provide no abstraction mechanisms beyond those of the RAM Machine.
+The Turing Machine itself, while universal, lacks random access: reading a cell at distance $d$ requires $d$ sequential tape movements.
+Each of these models achieves Turing completeness by removing structure from an existing model: fewer abstractions, worse ergonomics, no new expressiveness.
+None of these reductions exhibits emergent phenomena absent from the model it reduces.
+Denotational semantics of inheritance
+Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records and inheritance as composition of generators, which are functions from self to complete object, and wrappers, which are functions that modify generators, solved by a fixed-point construction.
+Cook’s key insight, that inheritance is a general mechanism applicable to any form of recursive definition and not only object-oriented methods, is one of the starting points of the present work.
+The model has four primitives—records, functions, generators, and wrappers—and requires a fixed-point combinator; composition is asymmetric, as a wrapper modifies a generator but not vice versa.
+Mixin-based inheritance
+Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses, that is, functions from a superclass parameter to a subclass, unifying the inheritance mechanisms of Smalltalk, Beta, and CLOS under a single model.
+Because mixin application is function composition, it is neither commutative nor idempotent: applying the same mixin twice may produce different results, and applying two mixins in different orders may yield different linearizations.
+The C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order; whether linearization is inherent to inheritance or an artifact of non-commutativity had not been established.
+Mixin composition operates at the method level: applying a mixin replaces or extends individual methods in a flat namespace.
+There is no recursive merging of nested structure: composing two mixins that define the same method name results in an override, not deep merge.
+Finally, mixins presuppose the $\lambda$-calculus as their computational foundation: method bodies are functions, and mixin application is function composition.
+Mixins organize and compose functions but cannot replace them; without the underlying $\lambda$-calculus, there is nothing to compose.
+The mixin mechanism itself is not Turing complete.
+Traits
+Schärli et al.
+(Schärli et al., 2003) introduced traits as composable units of behavior.
+Trait composition is symmetric and commutative: the sum of two traits is order-independent, which is a departure from mixin ordering.
+However, when two traits provide methods with the same name, a conflict arises and must be explicitly resolved by the programmer through exclusion or aliasing.
+Ducasse et al.
+(Ducasse et al., 2006) extended the model with a flattening property: the semantics of a class is independent of whether its methods originate from traits or are defined directly.
+Like mixins, trait composition is flat: it operates on individual method names, not on nested structures, so same-name definitions are conflicts rather than recursive merges.
+And like mixins, traits presuppose the $\lambda$-calculus: trait methods are functions, and the computational power resides in those functions, not in the trait mechanism itself.
+Family polymorphism and virtual classes
+Ernst (Ernst, 2001) introduced family polymorphism, where families of mutually dependent classes can be refined together in subclasses.
+Ernst, Ostermann, and Cook (Ernst et al., 2006) formalized this in the virtual class calculus, where classes are class-valued attributes of objects, accessed via path expressions of the form this.out.C.
+The path this.out navigates to an enclosing scope, analogous to qualified this ($\ell_{\mathrm{up}}.\mathbf{this}$) in this paper.
+A notable restriction is that each such path resolves to a single class in a single enclosing object.
+When multiple inheritance routes lead to the same scope, a common situation under inheritance, single-valued resolution must choose one route and discard the others, losing information.
+The framework provides no mechanism for set-valued resolution.
+Object calculi
+Abadi and Cardelli (Abadi and Cardelli, 1996) developed a comprehensive theory of objects, treating objects rather than classes as the primitive notion.
+Their calculi include method update and self-referential objects, but methods are fundamentally functions that take self as a parameter, and object extension is asymmetric since method override replaces the previous definition.
+Boudol (Boudol, 2004) studied the recursive record semantics of objects in a call-by-value setting, showing that self-referential records require an unsafe fixed-point operator.
+Whether self-reference can be handled without a fixed-point operator in the object language, for instance through coalgebraic observation rather than reduction, was not explored in these calculi.
+Record calculi
+Harper and Pierce (Harper and Pierce, 1991) presented a record calculus based on symmetric concatenation with a type system that tracks the absence of fields.
+When two records define the same label, concatenation is undefined and raises a type error; the calculus leaves same-label composition undefined.
+Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) gave a type system for records and variants in ML.
+These calculi operate on flat records with no nesting and treat records as data, that is, initial algebras: a record is a finite map from labels to values, fully determined at construction time.
+Nested records, self-reference, and lazy observation, which are the properties needed for records to serve as a computational substrate, are outside their scope.
+DOT calculus
+Amin et al.
+(Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala’s type system.
+DOT paths navigate through objects to reach nested type members, providing a rich dependent type structure.
+DOT’s self variable in the object constructor $\nu(x:T)\,d$ binds $x$ to a single object; when multiple inheritance routes lead to the same enclosing scope, the resulting intersection type is unrealizable and Scala rejects it at compile time (Appendix B).
+DOT stratifies its syntax into terms and types, with types further divided into abstract type members bounded by $L..U$ and concrete definitions; each stratum requires its own subtyping rules.
+Whether these three roles could be unified under a single construct remains an open question in the DOT line of work.
+DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure overlay inheritance (see Section 8).
+Configuration languages
+CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification (&) is commutative, associative, and idempotent.
+CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification.
+CUE’s lattice includes scalar types with a conflict semantics where unifying incompatible scalars yields $\bot$; whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question.
+Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative since the right-hand side wins on scalar conflicts, and deep merging requires explicit opt-in via the +: syntax) on a per-field basis, so deep merge is available but not the default.
+Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
+Module systems and deep merge
+The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same overlay semantics that distinguishes tree-level inheritance from mixin and trait composition.
+When two modules define the same nested path, their subtrees are merged rather than overridden.
+The NixOS module system incorporates deep merge, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top.
+The same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity.
+The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance.
+When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix B).
+The module system cannot express that both routes contribute equally to the same option, that is, the multi-path situation that recursive merging naturally gives rise to, which motivated the coalgebraic semantics of Section 3.
 Modular software architectures
@@
 Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view.
-Their semantics resembles Overlay-Calculus—later layers override earlier layers, files from all layers remain accessible—but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers.
+Their semantics resembles overlay-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers.
 We implemented overlay-calculus directly as a union file system (Yang, 2025a), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives.
-In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance—a file-system-as-compiler in which the directory tree is both the source program and the compilation target.
-7. Related Work
-Denotational semantics of inheritance
-Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records and inheritance as composition of generators (functions from self to complete object) and wrappers (functions that modify generators), solved by a fixed-point construction.
-Cook’s key insight—that inheritance is a general mechanism applicable to any form of recursive definition, not only object-oriented methods—is one of the starting points of the present work.
-The model has four primitives (records, functions, generators, wrappers) and requires a fixed-point combinator; composition is asymmetric (a wrapper modifies a generator, not vice versa).
-Whether the same insight can be captured with fewer primitives and symmetric inheritance was a natural question.
-Mixin-based inheritance
-Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses—functions from a superclass parameter to a subclass—unifying the inheritance mechanisms of Smalltalk, Beta, and CLOS under a single model.
-Because mixin application is function composition, it is neither commutative nor idempotent: applying the same mixin twice may produce different results, and applying two mixins in different orders may yield different linearizations.
-The C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order; whether linearization is inherent to inheritance or an artifact of non-commutativity was worth examining.
-Mixin composition operates at the method level: applying a mixin replaces or extends individual methods in a flat namespace.
-There is no recursive merging of nested structure—composing two mixins that define the same method name results in an override, not a union mount.
-Finally, mixins presuppose the $\lambda$-calculus as their computational foundation: method bodies are functions, and mixin application is function composition.
-Mixins organize and compose functions but cannot replace them—without the underlying $\lambda$-calculus, there is nothing to compose.
-The mixin mechanism itself is not Turing complete.
-Traits
-Schärli et al.
-(Schärli et al., 2003) introduced traits as composable units of behavior.
-Trait composition is symmetric and commutative: the sum of two traits is order-independent, which is a significant improvement over mixins.
-However, when two traits provide methods with the same name, a conflict arises and must be explicitly resolved by the programmer through exclusion or aliasing.
-Ducasse et al.
-(Ducasse et al., 2006) extended the model with a flattening property: the semantics of a class is independent of whether its methods originate from traits or are defined directly.
-Like mixins, trait composition is flat: it operates on individual method names, not on nested structures, so same-name definitions are conflicts rather than opportunities for deeper merging.
-And like mixins, traits presuppose the $\lambda$-calculus: trait methods are functions, and the computational power resides in those functions, not in the trait mechanism itself.
-Family polymorphism and virtual classes
-Ernst (Ernst, 2001) introduced family polymorphism, where families of mutually dependent classes can be refined together in subclasses.
-Ernst, Ostermann, and Cook (Ernst et al., 2006) formalized this in the virtual class calculus, where classes are class-valued attributes of objects, accessed via path expressions of the form this.out.C.
-The path this.out navigates to an enclosing scope, analogous to qualified this ($\ell_{\mathrm{up}}.\mathbf{this}$) in this paper.
-A notable restriction is that each such path resolves to a single class in a single enclosing object.
-When multiple inheritance routes lead to the same scope—a common situation under inheritance—single-valued resolution must choose one route and discard the others, losing information.
-Whether qualified this can resolve to a set of paths, letting all routes contribute equally, had not been explored.
-Object calculi
-Abadi and Cardelli (Abadi and Cardelli, 1996) developed a comprehensive theory of objects, treating objects—rather than classes—as the primitive notion.
-Their calculi include method update and self-referential objects, but methods are fundamentally functions (taking self as a parameter), and object extension is asymmetric (method override replaces the previous definition).
-Boudol (Boudol, 2004) studied the recursive record semantics of objects in a call-by-value setting, showing that self-referential records require an unsafe fixed-point operator.
-Whether self-reference can be handled without a fixed-point operator in the object language—for instance, through coalgebraic observation rather than reduction—was a question these calculi left open.
-Record calculi
-Harper and Pierce (Harper and Pierce, 1991) presented a record calculus based on symmetric concatenation with a type system that tracks the absence of fields.
-When two records define the same label, concatenation is undefined (a type error); the calculus avoids same-label composition rather than giving it meaning.
-Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) gave a type system for records and variants in ML.
-These calculi operate on flat records (no nesting) and treat records as data (initial algebras): a record is a finite map from labels to values, fully determined at construction time.
-Nested records, self-reference, and lazy observation—the properties needed for records to serve as a computational substrate—are outside their scope.
-DOT calculus
-Amin et al.
-(Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala’s type system.
-DOT paths navigate through objects to reach nested type members, providing a rich dependent type structure.
-DOT’s self variable in the object constructor $\nu(x:T)\,d$ binds $x$ to a single object; when multiple inheritance routes lead to the same enclosing scope, the resulting intersection type is unrealizable and Scala rejects it at compile time (Appendix C).
-DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure overlay inheritance (see Section 8).
-Configuration languages
-CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification (&) is commutative, associative, and idempotent.
-CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification.
-CUE’s lattice includes scalar types with a conflict semantics (unifying incompatible scalars yields $\bot$); whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question.
-Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative (the right-hand side wins on scalar conflicts), and deep merging requires explicit opt-in (+: syntax) on a per-field basis—so union mount semantics is available but not the default.
-Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
-Module systems and union mount semantics
-The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override—the same union mount semantics that distinguishes tree-level inheritance from mixin and trait composition.
-When two modules define the same nested path, their subtrees are merged rather than overridden.
-The NixOS module system got nearly all of the high-level design right: union mount inheritance, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top.
-Its practical success is remarkable—the same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity.
-The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance.
-When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix D).
-The module system cannot express that both routes contribute equally to the same option—the multi-path situation that recursive merging naturally gives rise to—which motivated the coalgebraic semantics of Section 3.
-SKI combinator calculus
-Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus can be reformulated without bound variables using three combinators (S, K, I).
-SKI is function-free at the object level and Turing complete.
-However, bracket abstraction translates between SKI and the $\lambda$-calculus preserving computational structure, so the two systems are isomorphic in terms of problem difficulty (Section 4): what is hard in one remains hard in the other.
-Whether a function-free calculus can exhibit a genuine shift in problem difficulty—making some problems structurally easier than in the $\lambda$-calculus, rather than merely encoding them differently—is the question that motivates the comparison in Section 4.
+In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance: a file-system-as-compiler in which the directory tree is both the source program and the compilation target.
 8. Future Work
 The Overlay language (Yang, 2025b) is an executable implementation of overlay-calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the inheritance, and a foreign-function interface (FFI) that introduces scalar values from the host language.
-The future work described below concerns the distance between this implementation and a fully practical language—spanning both theoretical foundations and library-level encodings.
+The future work described below concerns the distance between this implementation and a fully practical language, spanning both theoretical foundations and library-level encodings.
 Type system
 The Overlay language’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result.
-Formalizing the soundness of these checks—proving that well-typed programs do not produce dangling references at runtime—is one direction of future work.
+Formalizing the soundness of these checks, that is, proving that well-typed programs do not produce dangling references at runtime, is one direction of future work.
 Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how overlay-calculus relates to the $\lambda$-calculus: a Typed overlay-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with overlay inheritance.
@@
 A key example is totality checking: verifying that a inheritance provides implementations for all required slots, not merely that references resolve.
-In untyped Overlay-Calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary—it defines a structural slot but imposes no constraint on what must fill it.
+In untyped overlay-calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary: it defines a structural slot but imposes no constraint on what must fill it.
 Totality checking would enforce such declarations statically, rejecting inheritances that leave required slots unfilled.
@@
 Overlay-calculus is open by default: inheritance can freely merge any two overlays, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$).
-This is not a defect—it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 4.1).
-However, many operations (e.g., equality testing) assume that values are linear—inhabiting exactly one constructor.
-The observer pattern (Appendix B) does not enforce this: applied to a multi-constructor value, all callbacks fire and their results are inherited.
+This is not a defect; it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 5).
+However, many operations (e.g., equality testing) assume that values are linear, inhabiting exactly one constructor.
+The observer pattern (Section 5) does not enforce this: applied to a multi-constructor value, all callbacks fire and their results are inherited.
 Schema validation can be implemented at the library level using existing primitives.
@@
 By chaining such tests, a factory can validate that a value matches exactly one constructor and carries the required fields.
-This does not restrict inheritance itself—it remains unconstrained—but provides a way to detect invariant violations before they propagate.
-The same technique—observer returning Boolean, followed by Boolean dispatch—yields closed pattern matching.
-The observer pattern of Appendix B is inherently open: new constructor cases can be added through inheritance.
+This does not restrict inheritance itself, which remains unconstrained, but provides a way to detect invariant violations before they propagate.
+The same technique, an observer returning Boolean followed by Boolean dispatch, yields closed pattern matching.
+The observer pattern of Section 5 is inherently open: new constructor cases can be added through inheritance.
 Closed matching, where exactly one branch is taken, can be encoded as a visitor chain: a linked list of if-then-else nodes, each testing one constructor and falling through on mismatch.
 This is analogous to GHC.Generics’ $\mathrm{(:+:)}$ sum representation (Magalhães et al., 2010), where each link corresponds to one summand.
-An important consequence is that such chains do not commute—they have a fixed order—and therefore cannot be extended through open inheritance, which is the correct semantics for closed dispatch.
-The trie operations in Appendix A are hardcoded for binary natural number keys.
-Since every overlay is a trie (Appendix A) and insertion is simply inheritance, only operations that cannot be expressed as inheritance need explicit implementation: deletion, dynamic lookup, and prefix lookup.
-Generalizing to arbitrary key types requires each key factory to provide two interfaces—SelectTrieChild (select a subtrie by constructor) and ReplaceTrieChild (replace a subtrie by constructor while preserving the rest)—corresponding to the getter and setter of a lens (Foster et al., 2007) focused on the constructor-determined child.
-Prefix lookup is structurally analogous to equality testing: both synchronously traverse a key (a single-path trie) alongside another structure, dispatching at each level.
-The difference is the return type—a Boolean for equality, a subtrie for prefix lookup.
-Appendix A Trie Operations
+An important consequence is that such chains do not commute, as they have a fixed order, and therefore cannot be extended through open inheritance, which is the correct semantics for closed dispatch.
+9. Conclusion
+The goals of declarative programming forced the semantics of overlay-calculus upon us.
+The emergent phenomena exhibit an expressive asymmetry between overlay-calculus and the $\lambda$-calculus.
+Inheritance is not a mechanism for organizing functions; functions are syntactic sugar for inheritance.
+Appendix A Expressive Asymmetry: Proofs
+Definition A.1 (Macro-expressibility, after Felleisen (Felleisen, 1991)).
+A construct $C$ of language $\mathscr{L}_{1}$ is macro-expressible in language $\mathscr{L}_{0}$ if there exists a translation $\mathcal{E}$ such that for every program $P$ containing occurrences of $C$, the translation $\mathcal{E}(P)$ is obtained by replacing each occurrence of $C$ with an $\mathscr{L}_{0}$ expression that depends only on $C$’s subexpressions, leaving all other constructs of $P$ unchanged.
+Theorem A.2 (Forward macro-expressibility).
+The translation $\mathcal{T}$ (Section 4) is a macro-expressible embedding of the $\lambda$-calculus into overlay-calculus.
+Proof.
+Each of the five rules of $\mathcal{T}$ replaces one $\lambda$-calculus construct with an overlay-calculus expression that depends only on the construct’s immediate subexpressions.
+No rule inspects the surrounding context.
+∎
+Theorem A.3 (Non-expressibility of overlay).
+There is no macro-expressible translation from overlay-calculus to the $\lambda$-calculus.
+Proof.
+Let $\mathscr{L}_{1}$ be the overlay-calculus and let $\mathscr{L}_{0}$ be the sublanguage consisting of overlay trees in the image of the translation $\mathcal{T}$ (Section 4), so that $\mathscr{L}_{0}\subset\mathscr{L}_{1}$.
+Define the operational equivalence $\cong_{0}$ on $\mathscr{L}_{0}$ via Böhm trees: $\mathcal{T}(M)\cong_{0}\mathcal{T}(N)$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$.
+Define $\cong_{1}$ on $\mathscr{L}_{1}$ as overlay-contextual equivalence: $S\cong_{1}S^{\prime}$ iff for every overlay context $C[\cdot]$, $C[S]{\Downarrow}\Leftrightarrow C[S^{\prime}]{\Downarrow}$.
+By the Böhm Tree Correspondence (Theorems 4.2 and 4.3), $\mathscr{L}_{1}$ conservatively extends $\mathscr{L}_{0}$: for closed $\lambda$-terms $M$, $\mathcal{T}(M){\Downarrow}$ if and only if $M$ has a head normal form.
+Define Church booleans via the translation $\mathcal{T}$ (Section 4): $\mathrm{true}=\lambda t.\,\lambda f.\,t$ and $\mathrm{false}=\lambda t.\,\lambda f.\,f$. In de Bruijn notation:
+$\mathrm{true}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}$
+$\mathrm{false}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!0}.\mathrm{argument}\}\}$
+Define Church boolean equality $\mathrm{eq}=\lambda a.\,\lambda b.\,(a\;b)\;(b\;\mathrm{false}\;\mathrm{true})$. In ANF:
+$\mathrm{eq}=\lambda a.\,\lambda b.\,\mathbf{let}\;x_{1}=b\;\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;x_{2}=x_{1}\;\mathrm{true}\;\mathbf{in}\;\mathbf{let}\;x_{3}=a\;b\;\mathbf{in}\;x_{3}\;x_{2}$
+Applying $\mathcal{T}$ (Section 4):
+$\mathrm{eq}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},$
+$\quad x_{1}\mapsto\{\uparrow^{\!0}.\mathrm{argument},\;\mathrm{argument}\mapsto\mathrm{false}\},$
+$\quad x_{2}\mapsto\{x_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
+$\quad x_{3}\mapsto\{\uparrow^{\!1}.\mathrm{argument},\;\mathrm{argument}\mapsto\uparrow^{\!0}.\mathrm{argument}\},$
+$\quad\_\mathrm{tail}\mapsto\{x_{3}.\mathrm{result},\;\mathrm{argument}\mapsto x_{2}.\mathrm{result}\},$
+$\quad\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}\}$
+Here $\uparrow^{\!1}.\mathrm{argument}$ in $x_{3}$ reaches the outer $\lambda a$ scope (one $\lambda$ up from $\lambda b$), and $\uparrow^{\!0}.\mathrm{argument}$ reaches $\lambda b$ itself.
+In overlay-calculus, applying a function record to an argument is: inherit the function, override its $\mathrm{argument}$ slot, and project $\mathrm{result}$.
+Define overlay trees $e_{1}$ and $e_{2}$ that apply $\mathrm{eq}$ to its two arguments via this mechanism:
+$e_{1}\;=\;\{$ $\mathrm{eqFalse}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{false}\},$
+$\mathrm{applied}\mapsto\{\mathrm{eqFalse}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{false}\},$
+$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
+$e_{2}\;=\;\{$ $\mathrm{eqTrue}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{true}\},$
+$\mathrm{applied}\mapsto\{\mathrm{eqTrue}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
+$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
+That is, $e_{1}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{false}\;\mathrm{false}))$ and $e_{2}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{true}\;\mathrm{true}))$, where $\mathrm{ANF}$ converts a $\lambda$-term to A-normal form by naming every intermediate application (Section 4).
+Since both $\mathrm{ANF}$ and $\mathcal{T}$ are macro-expressible, their composition is too, so $e_{1},e_{2}\in\mathscr{L}_{0}$.
+Both $e_{1}.\mathrm{result}$ and $e_{2}.\mathrm{result}$ evaluate to Church $\mathrm{true}$.
+By Full Abstraction (Theorem 4.3), $\mathrm{BT}(e_{1})=\mathrm{BT}(e_{2})$, so $e_{1}\cong_{0}e_{2}$.
 […]
-$\mathrm{Zero}$ $\;\longrightarrow\;\mathrm{value}$
-$\mathrm{Odd}(n)$ $\;\longrightarrow\;\mathrm{at\_odd},\;\text{then path for }n$
-$\mathrm{Even}(n)$ $\;\longrightarrow\;\mathrm{at\_even},\;\text{then path for }n$
-Write (insertion)
-Inserting an entry with value $v$ at key $k$ is inheriting a singleton trie whose nesting mirrors the bit structure of $k$.
-For example, inserting at key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$:
-$\mathrm{trie}^{\prime}=\{\mathrm{trie},\;\mathrm{at\_odd}\mapsto\{\mathrm{at\_even}\mapsto\{\mathrm{value}\mapsto v\}\}\}$
-Each inheritance creates a new immutable trie; the original is unchanged.
-Read (static lookup)
-When the key is known statically, retrieving the entry at key $k$ is reference-path navigation: the reference follows the same label sequence that was used for insertion. For the same key $5$:
-$\mathrm{enclosing\_scope}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}.\mathrm{at\_even}.\mathrm{value}$
-where $\mathrm{enclosing\_scope}$ is the qualified-this scope name containing the trie definition.
-The reference mechanism traverses the nested record structure—which is the trie—returning the value at the final node.
-Read (dynamic lookup)
-When the key is a BinNat value constructed at runtime via $\mathrm{Zero}$, $\mathrm{Odd}$, and $\mathrm{Even}$, the lookup path must be computed dynamically.
-This is achieved by folding the BinNat key into a Selector—an overlay that, given a trie, navigates to the appropriate node and returns the stored entry.
-A Selector declares the trie shape and a result slot:
-$\mathrm{Selector}\mapsto\{\mathrm{trie}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\},\;\mathrm{result}\mapsto\{\}\}$
-Three observer callbacks build Selectors bottom-up.
-These callbacks ($\mathrm{on\_zero}$, $\mathrm{on\_odd}$, $\mathrm{on\_even}$) are interpretation points in the sense of Appendix G: they can be filled externally by an orchestrator (CPS interpretation) or inherited locally (direct interpretation), demonstrating the CPS-agnostic property.
-The $\mathrm{on\_zero}$ callback returns a Selector that reads $\mathrm{value}$ from the trie:
-$\mathrm{lookup\_on\_zero}\mapsto\{\mathrm{Selector},\;\mathrm{result}\mapsto\mathrm{lookup\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{value}\}$
-The $\mathrm{on\_odd}$ callback takes a Selector (via the $\mathrm{argument}/\mathrm{result}$ encoding of Section 4) and returns a new Selector that navigates to $\mathrm{at\_odd}$, then delegates to the argument Selector:
-$\mathrm{lookup\_on\_odd}$ $\mapsto\{\mathrm{argument}\mapsto\mathrm{Selector},\;\mathrm{result}\mapsto\{\mathrm{Selector},\;$
-$\qquad\mathrm{applied}\mapsto\{\mathrm{argument},\;\mathrm{trie}\mapsto\mathrm{lookup\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}\},$
-$\qquad\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\}\}$
-The $\mathrm{on\_even}$ callback is symmetric, navigating to $\mathrm{at\_even}$ instead.
-The $\mathrm{Lookup}$ combinator folds a BinNat key with these callbacks to obtain a Selector, then applies it to the trie:
-$\mathrm{Lookup}$ $\mapsto\{$
-$\quad\mathrm{key}\mapsto\mathrm{trie\_module}.\mathbf{this}.\mathrm{BinNat},$
-$\quad\mathrm{trie\_input}\mapsto\{\},$
-$\quad\mathrm{applied\_key}\mapsto\{\mathrm{key},\;$
-$\qquad\quad\mathrm{on\_zero}\mapsto\mathrm{lookup\_on\_zero},\;$
-$\qquad\quad\mathrm{on\_odd}\mapsto\mathrm{lookup\_on\_odd},\;$
-$\qquad\quad\mathrm{on\_even}\mapsto\mathrm{lookup\_on\_even}\},$
-$\quad\mathrm{applied\_selector}\mapsto\{\mathrm{applied\_key}.\mathrm{result},\;$
-$\qquad\quad\mathrm{trie}\mapsto\mathrm{trie\_input}\},$
-$\quad\mathrm{result}\mapsto\mathrm{applied\_selector}.\mathrm{result}\}$
-For example, looking up key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$ folds into a Selector that navigates $\mathrm{at\_odd}.\mathrm{at\_even}.\mathrm{value}$—the same path as the static reference, but computed from the key value.
-The entire construction uses only records (with inheritance sources) and references; no additional primitives are needed.
-Delete
-Since inheritance is additive—it can only merge properties, never remove them—deletion cannot be expressed as a single inheritance.
-Instead, deletion reconstructs the trie, copying every subtrie except at the targeted key.
-We fold the BinNat key into a Rebuilder—an overlay that, given an input trie, produces a new trie as its result. A Rebuilder declares the trie shape on both its input and output:
-$\mathrm{Rebuilder}\mapsto\{$ $\mathrm{trie}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\},\;$
-$\mathrm{result}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\}\}$
-Three observer callbacks build Rebuilders bottom-up. The $\mathrm{on\_zero}$ callback returns a Rebuilder that copies the subtries but omits the value:
-$\mathrm{delete\_on\_zero}\mapsto\{$ $\mathrm{Rebuilder},\;\mathrm{result}\mapsto\{$
-$\quad\mathrm{at\_odd}\mapsto\mathrm{delete\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd},\;$
-$\quad\mathrm{at\_even}\mapsto\mathrm{delete\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_even}\}\}\}$
-The $\mathrm{result}.\mathrm{value}$ slot inherits only the empty declaration from $\mathrm{Rebuilder}$, effectively deleting the entry.
-The $\mathrm{on\_odd}$ callback takes a Rebuilder (the recursive result for the subtrie) and returns a new Rebuilder that preserves $\mathrm{value}$ and $\mathrm{at\_even}$, but replaces $\mathrm{at\_odd}$ with the recursively rebuilt subtrie:
-$\mathrm{delete\_on\_odd}$ $\mapsto\{\mathrm{argument}\mapsto\mathrm{Rebuilder},\;$
-$\quad\mathrm{result}\mapsto\{\mathrm{Rebuilder},\;$
-$\qquad\mathrm{applied}\mapsto\{\mathrm{argument},\;$
-$\qquad\quad\mathrm{trie}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}\},$
-$\qquad\mathrm{result}\mapsto\{$
-$\qquad\quad\mathrm{value}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{value},\;$
-$\qquad\quad\mathrm{at\_odd}\mapsto\mathrm{applied}.\mathrm{result},\;$
-$\qquad\quad\mathrm{at\_even}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_even}\}\}\}$
-The $\mathrm{on\_even}$ callback is symmetric.
-The $\mathrm{Delete}$ combinator folds a BinNat key with these callbacks to obtain a Rebuilder, then applies it to the trie:
-$\mathrm{Delete}$ $\mapsto\{$
-$\quad\mathrm{key}\mapsto\mathrm{trie\_module}.\mathbf{this}.\mathrm{BinNat},$
-$\quad\mathrm{trie\_input}\mapsto\{\},$
-$\quad\mathrm{applied\_key}\mapsto\{\mathrm{key},\;$
-$\qquad\quad\mathrm{on\_zero}\mapsto\mathrm{delete\_on\_zero},\;$
-$\qquad\quad\mathrm{on\_odd}\mapsto\mathrm{delete\_on\_odd},\;$
-$\qquad\quad\mathrm{on\_even}\mapsto\mathrm{delete\_on\_even}\},$
-$\quad\mathrm{applied\_rebuilder}\mapsto\{\mathrm{applied\_key}.\mathrm{result},\;$
-$\qquad\quad\mathrm{trie}\mapsto\mathrm{trie\_input}\},$
-$\quad\mathrm{result}\mapsto\mathrm{applied\_rebuilder}.\mathrm{result}\}$
-Since inheritance is idempotent, the empty scaffolding left by the Rebuilder schema merges harmlessly with any subsequent insertions.
-Update
-Updating the entry at key $k$ to a new value $v$ combines deletion and insertion.
-Because $\mathrm{Delete}$ is a combinator whose output lives in the result property, one must first project out the rebuilt trie before inheriting the new singleton.
-In ANF style:
-$\mathrm{deleted}$ $\mapsto\{\mathrm{Delete},\;\mathrm{key}\mapsto k,\;\mathrm{trie\_input}\mapsto\mathrm{trie}\}$
-$\mathrm{result}$ $\mapsto\{\mathrm{deleted}.\mathrm{result},\;\mathrm{at\_odd}\mapsto\{\mathrm{at\_even}\mapsto\{\mathrm{value}\mapsto v\}\}\}$
-where the second line uses key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$ as a concrete example.
-The projection $\mathrm{result}$ is essential: $\mathrm{Delete}$ carries internal scaffolding ($\mathrm{key}$, $\mathrm{trie\_input}$, $\mathrm{applied\_key}$, etc.), and only its $\mathrm{result}$ property holds the rebuilt trie.
-This intermediate projection prevents Update from being a single symmetric inheritance.
-Appendix B Church Encoding of Standard Types
-This appendix demonstrates that standard data types can be encoded in Overlay-Calculus without scalar types, using Church encoding.
-These encodings serve as a proof of concept that a standard library can be built entirely within Overlay-Calculus.
-Each encoding is presented as a module overlay.
-In practical implementations, more efficient representations (e.g., binary naturals) or FFI to host language primitives would typically be used.
-The implementation (Yang, 2025b) includes executable implementations of these encodings with test cases for $\mathrm{Not}(\mathrm{True})$, $\mathrm{And}(\mathrm{True},\mathrm{True})$, $\mathrm{Or}(\mathrm{False},\mathrm{True})$, and natural number addition, all verified against expected output snapshots.
-B.1. Boolean Module
-The boolean module contains a schema ($\mathrm{Boolean}$), constructors ($\mathrm{True}$, $\mathrm{False}$), and operations ($\mathrm{Not}$, $\mathrm{And}$, $\mathrm{Or}$).
-All definitions are siblings within a single module overlay.
-$\{\quad\mathrm{boolean}\mapsto\{\quad\mathrm{Boolean}$ $\mapsto\{\mathrm{on\_true}\mapsto\{\},\;\mathrm{on\_false}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},$
-$\mathrm{True}$ $\mapsto\{\mathrm{Boolean},\;\mathrm{result}\mapsto\mathrm{True}.\mathbf{this}.\mathrm{on\_true}\},$
-$\mathrm{False}$ $\mapsto\{\mathrm{Boolean},\;\mathrm{result}\mapsto\mathrm{False}.\mathbf{this}.\mathrm{on\_false}\},$
-$\mathrm{Not}$ $\mapsto\{\mathrm{Boolean},$
-$\qquad\mathrm{operand}\mapsto\mathrm{Boolean},$
-$\qquad\mathrm{applied\_operand}\mapsto\{\mathrm{operand},$
-$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{Not}.\mathbf{this}.\mathrm{on\_false},$
-$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{Not}.\mathbf{this}.\mathrm{on\_true}\},$
-$\qquad\mathrm{result}\mapsto\mathrm{applied\_operand}.\mathrm{result}\},$
-$\mathrm{And}$ $\mapsto\{\mathrm{Boolean},$
-$\qquad\mathrm{left}\mapsto\mathrm{Boolean},$
-$\qquad\mathrm{right}\mapsto\mathrm{Boolean},$
-$\qquad\mathrm{applied\_left}\mapsto\{\mathrm{left},$
-$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{right},$
-$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{False}\},$
-$\qquad\mathrm{result}\mapsto\mathrm{applied\_left}.\mathrm{result}\},$
-$\mathrm{Or}$ $\mapsto\{\mathrm{Boolean},$
-$\qquad\mathrm{left}\mapsto\mathrm{Boolean},$
-$\qquad\mathrm{right}\mapsto\mathrm{Boolean},$
-$\qquad\mathrm{applied\_left}\mapsto\{\mathrm{left},$
-$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{True},$
-$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{right}\},$
-$\qquad\mathrm{result}\mapsto\mathrm{applied\_left}.\mathrm{result}\}\quad\}\quad\}$
-Reference explanation
-Within the $\mathrm{boolean}$ module, own properties such as $\mathrm{Boolean}$, $\mathrm{True}$, $\mathrm{False}$, $\mathrm{Not}$, $\mathrm{And}$, and $\mathrm{Or}$ are accessed via lexical references (e.g., the inheritance source $\mathrm{Boolean}$ in each definition).
-Qualified this is required in two situations:
-  • Inherited properties.
-Inside True’s definition, $\mathrm{True}.\mathbf{this}.\mathrm{on\_true}$ requires qualified this because $\mathrm{on\_true}$ is inherited from Boolean via inheritance, not defined as an own property.
-Similarly, $\mathrm{Not}.\mathbf{this}.\mathrm{on\_false}$ and $\mathrm{Not}.\mathbf{this}.\mathrm{on\_true}$ access inherited callbacks.
-  • Variable shadowing.
-Inside Not’s $\mathrm{applied\_operand}$, the labels $\mathrm{on\_true}$ and $\mathrm{on\_false}$ are own properties of $\mathrm{applied\_operand}$ (with swapped values), so a lexical reference $\mathrm{on\_false}$ would resolve to the local definition rather than Not’s inherited callback.
-Qualified this $\mathrm{Not}.\mathbf{this}.\mathrm{on\_false}$ bypasses the shadowing.
-Worked example
-Consider the expression $\{\mathrm{boolean}.\mathrm{Not},\;\mathrm{operand}\mapsto\mathrm{boolean}.\mathrm{True},\;\mathrm{on\_true}\mapsto A,\;\mathrm{on\_false}\mapsto B\}$ at the root scope level, where $A$ and $B$ are arbitrary expressions.
-The result is an object whose $\mathrm{result}$ label contains the same observable tree as $B$, because $\mathrm{Not}$ swaps $\mathrm{on\_true}$ and $\mathrm{on\_false}$, and $\mathrm{True}$ selects $\mathrm{on\_true}$, which after swapping becomes $B$.
-B.2. Natural Number Module
-The natural number module contains a schema ($\mathrm{Nat}$), constructors ($\mathrm{Zero}$, $\mathrm{Succ}$), and an operation ($\mathrm{Add}$).
-$\{\quad\mathrm{natural}\mapsto\{\quad\mathrm{Nat}$ $\mapsto\{\mathrm{successor}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},\;\mathrm{zero}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},$
-$\mathrm{Zero}$ $\mapsto\{\mathrm{Nat},\;\mathrm{result}\mapsto\mathrm{Zero}.\mathbf{this}.\mathrm{zero}\},$
-$\mathrm{Succ}$ $\mapsto\{\mathrm{Nat},$
-$\qquad\mathrm{predecessor}\mapsto\mathrm{Nat},$
-$\qquad\mathrm{applied\_predecessor}\mapsto\{\mathrm{predecessor},$
-$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Succ}.\mathbf{this}.\mathrm{successor},$
-$\qquad\qquad\mathrm{zero}\mapsto\mathrm{Succ}.\mathbf{this}.\mathrm{zero}\},$
-$\qquad\mathrm{applied\_successor}\mapsto\{\mathrm{Succ}.\mathbf{this}.\mathrm{successor},$
-$\qquad\qquad\mathrm{argument}\mapsto\mathrm{applied\_predecessor}.\mathrm{result}\},$
-$\qquad\mathrm{result}\mapsto\mathrm{applied\_successor}.\mathrm{result}\},$
-$\mathrm{Add}$ $\mapsto\{\mathrm{Nat},$
-$\qquad\mathrm{augend}\mapsto\mathrm{Nat},$
-$\qquad\mathrm{addend}\mapsto\mathrm{Nat},$
-$\qquad\mathrm{applied\_addend}\mapsto\{\mathrm{addend},$
-$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{successor},$
-$\qquad\qquad\mathrm{zero}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{zero}\},$
-$\qquad\mathrm{applied\_augend}\mapsto\{\mathrm{augend},$
-$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{successor},$
-$\qquad\qquad\mathrm{zero}\mapsto\mathrm{applied\_addend}.\mathrm{result}\},$
-$\qquad\mathrm{result}\mapsto\mathrm{applied\_augend}.\mathrm{result}\}\quad\}\quad\}$
-Concrete example
-At the root scope level,
-$\mathrm{one}\mapsto\{\mathrm{natural}.\mathrm{Succ},\;\mathrm{predecessor}\mapsto\mathrm{natural}.\mathrm{Zero}\}.$
-Applying this with a Peano-style successor and zero: $\{\mathrm{one},\;\mathrm{successor}\mapsto S,\;\mathrm{zero}\mapsto Z\}$ produces an object whose $\mathrm{result}$ label contains the same observable tree as $\{S,\;\mathrm{argument}\mapsto Z.\mathrm{result}\}.\mathrm{result}$, i.e., successor applied once to zero.
-Here $S$ and $Z$ are arbitrary expressions representing a successor operation and a zero value respectively.
-Appendix C Multi-Path Self-Reference in Scala
-This appendix demonstrates that Scala 3 rejects the multi-path self-reference pattern that Overlay-Calculus handles naturally.
+$C={}$ $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument},\;\mathrm{secondOperand}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}\}\}$
+The de Bruijn indices count scope levels from $\mathrm{repr}$: $\uparrow^{\!0}$ is $\mathrm{repr}$ itself, $\uparrow^{\!1}$ is the $\lambda b$ scope, and $\uparrow^{\!2}$ is the $\lambda a$ scope.
+Thus $\mathrm{firstOperand}$ projects $\lambda a$’s argument and $\mathrm{secondOperand}$ projects $\lambda b$’s argument.
+In the composite scope $\{S,\,C\}$, the definition of $\mathrm{eq}$ gains $\mathrm{repr}$ as a new sibling of $x_{1},x_{2},x_{3}$ inside the $\lambda b$ scope.
+By open recursion, the de Bruijn references $\uparrow^{\!2}.\mathrm{argument}$ and $\uparrow^{\!1}.\mathrm{argument}$ resolve through the same scope chain as the existing let-bindings.
+Since $e_{1}$ and $e_{2}$ inherit $\mathrm{eq}$, they also inherit the newly added $\mathrm{repr}$: $e_{1}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{false}$ and $e_{2}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{true}$.
+These produce structurally distinct overlay trees (one converges to an abstraction selecting its first argument, the other to an abstraction selecting its second), so $e_{1}\not\cong_{1}e_{2}$.
+Since $e_{1}\cong_{0}e_{2}$ but $e_{1}\not\cong_{1}e_{2}$, we have ${\cong_{0}}\neq{(\cong_{1}|_{\mathscr{L}_{0}})}$.
+By Theorem 3.14(i) of Felleisen (Felleisen, 1991), $\mathscr{L}_{0}$ cannot macro-express the overlay facilities of $\mathscr{L}_{1}$.
+∎
+Corollary A.4.
+Overlay-calculus is strictly more expressive than the $\lambda$-calculus in the sense of Felleisen (Felleisen, 1991): the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
+Appendix B Multi-Path Self-Reference in Other Systems
+This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-path self-reference pattern that overlay-calculus handles naturally.
+Scala 3
 Consider two objects that independently extend an outer class, each providing its own copy of an inner trait:
@@
 Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer self-reference to a different object.
-Their intersection requires outer to simultaneously return Object1 and Object2—but DOT’s self variable is single-valued, making this intersection unrealizable.
+Their intersection requires outer to simultaneously return Object1 and Object2, but DOT’s self variable is single-valued, making this intersection unrealizable.
 The equivalent overlay-calculus definition is:
@@
 This is well-defined in overlay-calculus.
-Inside $\mathrm{MyInner}$, the reference $\mathrm{MyOuter}.\mathbf{this}$ has de Bruijn index $n=1$ (one scope level up from $\mathrm{MyInner}$ to $\mathrm{MyOuter}$).
+The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$.
 When $\mathrm{MyObjectA}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{MyObjectA})$ for overlay paths matching $\mathrm{MyInner}$’s definition site.
-It finds two inheritance-site paths—one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$—and returns both.
+It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both.
 Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{MyObjectA}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken.
 Since inheritance is idempotent, the two routes contribute the same properties and no disambiguation is needed.
-Appendix D Multi-Path Self-Reference in the NixOS Module System
-This appendix demonstrates that the NixOS module system rejects the multi-path self-reference pattern with a static error.
-The translation uses the module system’s own abstractions: deferredModule for traits (unevaluated module values imported into other fixpoints) and submoduleWith for objects (evaluated in their own fixpoints).
+NixOS Module System
+The NixOS module system rejects the same pattern with a static error.
+The translation uses the module system’s own abstractions: deferredModule for traits, which are unevaluated module values imported into other fixpoints, and submoduleWith for objects, which are evaluated in their own fixpoints.
 let
@@
 The rejection occurs because Object1.MyInner and Object2.MyInner each carry the full staticModules of the MyInner deferred module, including the declaration options.outer.
-When MyObjectA imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error—it cannot express that both declarations originate from the same trait and should be unified.
-The rejection is analogous to Scala’s “conflicting base types” error (Appendix C): both systems assume that each option (or type member) has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently.
+When MyObjectA imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error; it cannot express that both declarations originate from the same trait and should be unified.
+The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently.
 In overlay-calculus, the same pattern is well-defined because $\mathtt{overlays}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
-Appendix E Well-Definedness of the Semantic Functions
+Appendix C Well-Definedness of the Semantic Functions
 This appendix proves that the mutually recursive equations (1)–(6) have a unique least fixed point, and therefore the semantics of Section 3 is well-defined.
-E.1. The Product Lattice
+C.1. The Product Lattice
 Fix an AST with its primitive functions $\mathtt{defines}$ and $\mathtt{inherits}$.
@@
 The product $D=D_{\mathtt{properties}}\times D_{\mathtt{supers}}\times D_{\mathtt{overlays}}\times D_{\mathtt{bases}}\times D_{\mathtt{resolve}}\times D_{\mathtt{this}}$ is a complete lattice under the componentwise ordering, with bottom element $\bot=(\lambda x.\,\varnothing,\ldots,\lambda x.\,\varnothing)$.
-E.2. The Operator
+C.2. The Operator
 Equations (1)–(6) define a function $F:D\to D$ that takes a tuple of candidate interpretations $(\mathtt{properties},\mathtt{supers},\mathtt{overlays},\mathtt{bases},\mathtt{resolve},\mathtt{this})$ and produces a new tuple by evaluating the right-hand side of each equation.
@@
 The result of $\mathtt{this}(S,p_{\mathrm{def}},n)$ for all valid $(S,p_{\mathrm{def}},n)$ is therefore well-defined as a function of $\mathtt{supers}$, and it remains to show that it is monotone in $\mathtt{supers}$.
-E.3. Monotonicity
-Proposition E.1.
+C.3. Monotonicity
+Proposition C.1.
 The operator $F$ is monotone: if $\vec{d}\sqsubseteq\vec{d}^{\prime}$ in $D$, then $F(\vec{d})\sqsubseteq F(\vec{d}^{\prime})$.
@@
 For $p\neq()$:
-$\mathtt{overlays}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|(\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
+$\mathtt{overlays}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|{}\\ (\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
 The function $\mathtt{supers}$ appears as the set being iterated over.
@@
 Equation (4): $\mathtt{bases}$
-$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{overlay}}\in\mathtt{overlays}(p),\;(n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{overlay}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
+$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{overlay}}\in\mathtt{overlays}(p),\\ (n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{overlay}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
 Both $\mathtt{overlays}$ and $\mathtt{resolve}$ appear as sets being iterated over.
@@
 Positive in $\mathtt{supers}$.
-In summary, every mutually defined function appears on the right-hand side of every equation only inside the pattern $x\in f(\ldots)$ within a set comprehension—i.e., as the domain of an existential quantifier.
+In summary, every mutually defined function appears on the right-hand side of every equation only inside the pattern $x\in f(\ldots)$ within a set comprehension, i.e., as the domain of an existential quantifier.
 No complement, set difference, or negated membership test appears in any equation.
@@
 ∎
-E.4. Existence of the Least Fixed Point
-Theorem E.2.
+C.4. Existence of the Least Fixed Point
+Theorem C.2.
 The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overlays}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined as the least fixed point of $F$.
 Proof.
-By Proposition E.1, $F$ is monotone on the complete lattice $D$.
+By Proposition C.1, $F$ is monotone on the complete lattice $D$.
 By the Knaster–Tarski theorem (Tarski, 1955), $F$ has a least fixed point $\vec{d}^{*}=\bigsqcup_{n\geq 0}F^{n}(\bot)$, where $\bot$ assigns the empty set to every input.
@@
 A membership $\ell\in\mathtt{properties}(p)$ holds in $\vec{d}^{*}$ if and only if it appears at some finite stage $F^{k}(\bot)$, which corresponds to a finite derivation tree using the equations.
-When no finite derivation exists, the membership is absent from the least fixed point and the query is undefined—the analogue of divergence.
+When no finite derivation exists, the membership is absent from the least fixed point and the query is undefined, the analogue of divergence.
 ∎
-Appendix F Böhm Tree Correspondence: Proofs
+Appendix D Böhm Tree Correspondence: Proofs
 This appendix contains the full proofs of Adequacy (Theorem 4.2) and Full Abstraction (Theorem 4.3) stated in Section 4.1.
-F.1. Böhm Trees
+D.1. Böhm Trees
 We briefly recall the definition of Böhm trees (Barendregt, 1984).
@@
 The Böhm tree $\mathrm{BT}(M)$ of a $\lambda$-term $M$ is an infinite labeled tree defined by:
-$\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\text{if }M\to^{*}_{h}\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{cases}$
-Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Barendregt, 1984; Hyland and Ong, 2000).
-F.2. Path Encoding
+$\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\begin{aligned} &\text{if }M\to^{*}_{h}\\[-2.0pt] &\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{aligned}\end{cases}$
+Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993).
+D.2. Path Encoding
 We define a correspondence between positions in the Böhm tree and paths in the overlay tree.
@@
 The abstraction shape at the end signals that a weak head normal form has been reached.
-F.3. Single-Path Lemma
+D.3. Single-Path Lemma
 The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path overlay inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
-Lemma F.1 (Single Path).
+Lemma D.1 (Single Path).
 For any closed ANF $\lambda$-term $M$, at every invocation of $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)$ during evaluation of $\mathcal{T}(M)$, the frontier set $S$ contains exactly one path.
@@
 This is a reference, not a record.
-The $\mathtt{this}$ resolution walks up $n$ scope levels.
+The $\mathtt{this}$ function walks up $n$ steps from the enclosing scope.
 By the inductive hypothesis on the enclosing translations, each step of $\mathtt{this}$ encounters exactly one matching pair in $\mathtt{supers}$.
@@
 ∎
-F.4. Substitution Lemma
+D.4. Substitution Lemma
 The core mechanism of the translation is that inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ plays the role of substitution. We make this precise.
-Lemma F.2 (Substitution).
+Lemma D.2 (Substitution).
 Let $M$ be an ANF term in which $x$ may occur free, and let $V$ be a closed value.
@@
 Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference.
-In the inherited tree $C$, $\mathtt{this}$ resolves the reference by walking up $n$ scope levels to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ overlays the $\mathrm{argument}$ slot with $\mathcal{T}(V)$.
-Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma F.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree.
+In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ overlays the $\mathrm{argument}$ slot with $\mathcal{T}(V)$.
+Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma D.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree.
 Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
@@
 Inductive case: $M=\mathbf{let}\;z=V_{1}\;V_{2}$ $\mathbf{in}\;M^{\prime}$
-Then $\mathcal{T}(M)=\{z\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$.
+Then
+$\mathcal{T}(M)=\{z\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}.$
 The substitution distributes: $M[V/x]=\mathbf{let}\;z=V_{1}[V/x]\;V_{2}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$.
@@
 Analogous to the $\mathbf{let}$-binding case with $\mathrm{tail}$ in place of $z$. ∎
-F.5. Convergence Preservation
+D.5. Convergence Preservation
 The convergence criterion (Definition 4.1) follows the $\mathrm{result}$ chain from the root.
@@
 We make this precise.
-Lemma F.3 (Result-step).
+Lemma D.3 (Result-step).
 Let $(\lambda x.\,M)\;V$ be a tail call (a $\beta$-redex). Then for every path $p$ and label $\ell$:
@@
 The $\mathrm{result}$ label at the root is defined as $\mathrm{tail}.\mathrm{result}$, so $\text{root}\mathbin{\triangleright}\mathrm{result}$ resolves to the $\mathrm{result}$ path inside the inheritance $\mathrm{tail}=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$.
-This inheritance is exactly the inherited tree $C$ of Lemma F.2, and $\mathrm{tail}\mathbin{\triangleright}\mathrm{result}$ is $C\mathbin{\triangleright}\mathrm{result}$.
-By Lemma F.2, the properties under $C\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M[V/x])$.
+This inheritance is exactly the inherited tree $C$ of Lemma D.2, and $\mathrm{tail}\mathbin{\triangleright}\mathrm{result}$ is $C\mathbin{\triangleright}\mathrm{result}$.
+By Lemma D.2, the properties under $C\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M[V/x])$.
 Let-binding variant. An analogous result holds for $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$ where $V_{1}=\lambda y.\,M^{\prime\prime}$. The translation gives:
@@
 Inside $\mathcal{T}(M^{\prime})$, each reference to $x$ is $x.\mathrm{result}$, which resolves to the $\mathrm{result}$ path inside the $x$ subtree $\{\mathcal{T}(\lambda y.\,M^{\prime\prime}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$.
-This is the inherited tree $C$ of Lemma F.2, so $x.\mathrm{result}$ has the same properties as $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$.
+This is the inherited tree $C$ of Lemma D.2, so $x.\mathrm{result}$ has the same properties as $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$.
 In the $\lambda$-calculus, $M=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$, and the head reduct is $N=M^{\prime}[(V_{1}\;V_{2})/x]$.
@@
 ∎
-Theorem F.4 (Convergence preservation).
+Theorem D.4 (Convergence preservation).
 If $M\to_{h}N$ (head reduction), then $\mathcal{T}(M){\Downarrow}$ if and only if $\mathcal{T}(N){\Downarrow}$.
@@
 A head reduction step in ANF takes one of two forms: a tail call $(\lambda x.\,M^{\prime})\;V\to_{h}M^{\prime}[V/x]$, or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[(V_{1}\;V_{2})/x]$ (since $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$ in the $\lambda$-calculus).
-In both cases, Lemma F.3 (tail-call and let-binding variants respectively) shows that the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n}$ in $\mathcal{T}(M)$ coincide with the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n-1}$ in $\mathcal{T}(N)$ for all $n\geq 1$, where $N$ is the head reduct.
+In both cases, Lemma D.3 (tail-call and let-binding variants respectively) shows that the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n}$ in $\mathcal{T}(M)$ coincide with the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n-1}$ in $\mathcal{T}(N)$ for all $n\geq 1$, where $N$ is the head reduct.
 Hence the abstraction shape appears at depth $n$ in the pre-reduct if and only if it appears at depth n1n-1 in the post-reduct.
 ∎
-F.6. Adequacy
+D.6. Adequacy
 A closed ANF term at the top level is either an abstraction (a value) or a computation (a $\mathbf{let}$-binding or tail call).
@@
 Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$.
-By Lemma F.3 (tail-call or let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$.
+By Lemma D.3 (tail-call or let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$.
 By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
@@
 In both cases, $M$ is a $\beta$-redex (since $V_{1}=\lambda x.\,M^{\prime}$ or $V_{1}=\lambda y.\,M^{\prime\prime}$ respectively, the only closed ANF value).
-By Lemma F.3 (tail-call or let-binding variant respectively), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct.
+By Lemma D.3 (tail-call or let-binding variant respectively), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct.
 So $\mathcal{T}(N)$ converges at depth n1n-1.
@@
 ∎
-F.7. Full Abstraction
+D.7. Full Abstraction
 Adequacy relates a single term’s head normal form to its overlay-convergence.
 Full abstraction lifts this to an equivalence between terms: the translation $\mathcal{T}$ preserves and reflects observational equivalence.
-Definition F.5 (Overlay-contextual equivalence).
+Definition D.5 (Overlay-contextual equivalence).
 For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
@@
 ($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$.
-Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Hyland and Ong, 2000): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa).
+Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa).
 By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$.
@@
 ∎
-The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Hyland and Ong, 2000).
+The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993).
 The overlay tree semantics inherits full abstraction from the $\lambda$-calculus without needing to define or compare absolute paths inside overlay trees.
-Appendix G CPS-Agnostic Programs
-In the $\lambda$-calculus, a program is either in continuation-passing style (CPS) or direct style.
-Converting between these representations requires explicit transformation—the CPS conversion (Plotkin, 1975; Sabry and Felleisen, 1993).
-In Overlay-Calculus, the distinction does not exist: the same program text is CPS or direct-style depending on how dependencies are assembled.
-Consider a generic computation step:
-$\mathrm{Step}\mapsto\{\mathrm{input}\mapsto\{\},\;\mathrm{process}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\}$
-The empty $\mathrm{process}$ slot is an interpretation point.
-If filled externally by an orchestrator, it acts as a continuation parameter (CPS interpretation).
-If inherited locally with an implementation, it acts as a direct dependency (direct-style interpretation).
-The $\mathrm{Step}$ code remains unchanged.
-CPS interpretation via external orchestration
-When $\mathrm{process}$ is injected from outside, the computation passes its intermediate result to an external handler:
-$\{\mathrm{Step}.\mathbf{this},\;$
-$\quad\mathrm{input}\mapsto\{\mathrm{data}\mapsto\{\}\},\;$
-$\quad\mathrm{process}\mapsto\{$
-$\qquad\mathrm{result}\mapsto\{\mathrm{handled}\mapsto\{\}\}$
-$\quad\}$
-The $\mathrm{process}$ slot acts as a continuation: it receives the result and determines what happens next. This is CPS-style programming.
-Direct-style interpretation via local inheritance
-Alternatively, $\mathrm{process}$ can be filled by local inheritance:
-$\{\mathrm{Step}.\mathbf{this},\;$
-$\quad\mathrm{input}\mapsto\{\mathrm{data}\mapsto\{\}\},\;$
-$\quad\mathrm{process.result}\mapsto\{\mathrm{computed}\mapsto\{\}\}$
-Here $\mathrm{process}$ is implemented directly within the inheritance. This is direct-style programming. The $\mathrm{Step}$ code is identical in both cases—only the assembly strategy differs.
-Representation independence
-This property is a form of representation independence (Reynolds, 1983) applied to control flow: client code using $\mathrm{Step}$ need not know whether $\mathrm{process}$ will be filled continuation-style or directly.
-Library authors write code once; users choose the control flow style at assembly time.
-The $\lambda$-calculus lacks this flexibility.
-A function written in direct style must be explicitly transformed (CPS conversion) before it can be used continuation-style.
-Overlay-Calculus programs are inherently in A-normal form (Flanagan et al., 1993), where every intermediate result has a name.
-However, Overlay-Calculus ANF differs critically from ANF in functional programming.
-In functional language compilers, ANF conversion requires specifying a target monad—either a concrete monad or an abstract one (Moggi, 1991).
-Compiler-internal ANF transformations (the ANF form used in intermediate representations) typically target the identity monad, which cannot express CPS transformations.
-This is analogous to how imperative compilers use SSA (Static Single Assignment) form, which similarly lacks control flow flexibility.
-Overlay-Calculus ANF, by contrast, has built-in dependency injection: empty API slots can be filled by external orchestrators.
-This capability enables the same ANF program to serve both CPS and direct-style interpretations without monad specification or transformation.
-Multiple symmetric continuations
-When interpreted continuation-style, Overlay-Calculus naturally supports multiple exit points:
-$\mathrm{Choice}\mapsto\{\mathrm{on\_success}$ $\mapsto\{\},\;$
-$\mathrm{on\_failure}$ $\mapsto\{\},\;$
-$\mathrm{on\_timeout}$ $\mapsto\{\},\;$
-$\mathrm{result}$ $\mapsto\{\}\}$
-All continuation slots are symmetric elements of the same record.
-Traditional CPS in the $\lambda$-calculus distinguishes a single continuation parameter; additional continuations must be encoded (e.g., as pairs or disjoint sums).
-Overlay-Calculus treats all continuation slots uniformly, eliminating the need for specialized encodings.
-Abstract and concrete are syntactically identical
-From a category-theoretic perspective, abstract algebra and concrete implementation have the same syntactic structure in Overlay-Calculus:
-Abstract: $\mathrm{API}\mapsto\{\}$
-Concrete: $\mathrm{API}\mapsto\{\text{implementation}\}$
-This is the abstract/concrete correspondence: the same functor applies to both.
-In traditional languages, abstract interfaces (e.g., Java interfaces, Haskell type classes) have different syntax from concrete implementations.
-In Overlay-Calculus, they are syntactically identical—only the interpretation differs.
-An empty API is simultaneously an abstract requirement (when viewed as a dependency) and a concrete extension point (when viewed as an injection site).
-The Selector and Rebuilder combinators (Appendix A) demonstrate this flexibility: their callback slots ($\mathrm{on\_zero}$, $\mathrm{on\_odd}$, $\mathrm{on\_even}$) can be filled externally (CPS interpretation) or inherited locally (direct interpretation).
-Practical implications
-A library component with empty API slots can be used in multiple contexts without modification:
-  • In synchronous code, fill slots with direct implementations
-  • In asynchronous code, fill slots with continuation handlers
-  • In effect systems, compose with algebraic effect handlers (Plotkin and Power, 2003)
-  • In test code, fill slots with mock implementations
-No dual APIs are needed.
-The same component serves all use cases.
-This eliminates a common source of code duplication in traditional languages, where async and sync versions of the same functionality must be maintained separately.
-The implementation (Yang, 2025b) includes executable examples of all three patterns described above: escape continuations analogous to call/cc, success/failure continuations analogous to exception handling, and protected computations with multiple symmetric exit points.
-Each example includes test cases verifying that the correct continuation path is taken.
-CPS-agnosticism is not a designed feature of Overlay-Calculus; it is an emergent property of records and inheritance.
-The calculus defines no notion of continuation—the three syntax constructs and six semantic functions of Section 3 are the entire language.
-Appendix H Asymmetry in Problem Difficulty
-A skeptical reader might ask whether Overlay-Calculus is simply the $\lambda$-calculus with different syntax.
-The test of whether two computational models are “the same” is whether they make the same problems easy and the same problems hard.
-They do not:
-System Expression Problem Random-access memory
-$\lambda$-calculus Object algebras (d. S. Oliveira and Cook, 2012), Explicit trie encoding
-finally tagless (Carette et al., 2009)
-SKI combinators Same difficulty as $\lambda$ Same difficulty as $\lambda$
-Overlay-Calculus Direct inheritance Records are tries
-Overlay-Calculus requires no additional machinery for either problem class; the $\lambda$-calculus requires elaborate encodings for both.
-The semantic domain of Overlay-Calculus (Section 3) is entirely first-order: paths are label sequences, overlays are the record trees produced by evaluation, AST nodes are pure data, and the resolution functions ($\mathtt{overlays}$, $\mathtt{supers}$, $\mathtt{resolve}$, etc.) are ordinary set-valued functions over these structures.
-No function types appear in the object language or its semantics.
-The crucial question is not what notation is used, but what computational patterns are primitive in the language itself.
-In Overlay-Calculus, the primitive operations are record construction, qualified-this reference, and inheritance.
-Function application is derived from these primitives (Section 4), not primitive itself.
-The asymmetry of Section 4.1 makes this concrete: embedding the $\lambda$-calculus into Overlay-Calculus requires five rules; the reverse requires sophisticated encoding.
-Why tree-level inheritance dissolves the Expression Problem.
-Each operation is an independent overlay that extends a factory’s subtree.
-Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited overlays.
-Crucially, this applies to return types: if operation $A$ returns values from factory $F$, and an independently defined overlay $B$ adds a property to $F$, then inheriting from $A$ and $B$ causes the values returned by $A$ to carry that property automatically—without modifying either $A$ or $B$.
-In object algebras (d.
-S.
-Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
-Multi-path qualified this (equation 6) is essential to this dissolution—not merely to solving the Expression Problem (for which many solutions exist (d.
-S.
-Oliveira and Cook, 2012; Carette et al., 2009)), but to making it not arise in the first place.
-In the case study of Section 5, $\mathrm{evaluation}.\mathrm{Constant}$ and $\mathrm{display}.\mathrm{Constant}$ both inherit $\mathrm{expression}.\mathrm{Constant}$.
-When the two operations are inherited, the resulting $\mathrm{Constant}$ inherits the $\mathrm{expression}$ schema through two independent routes.
-The reference $\mathrm{Constant}.\mathbf{this}.\mathrm{magnitude}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties.
-In Scala, the same pattern—two independently defined inner classes extending the same outer class’s trait—triggers a “conflicting base types” rejection (Appendix C), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously.
-The Expression Problem can still be solved in Scala using object algebras or similar patterns, but it cannot be made to not arise: free inheritance of independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.

### v4

@@
-A Calculus of Overlays
+A Calculus of Inheritance
 Abstract.
-Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, overlay-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming.
+Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming.
 It trivially embeds the $\lambda$-calculus, although the entire semantics rests solely on naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance does not arise.
 This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence (Barendregt, 1984).
-Overlay-calculus is distilled from the Overlay language, a practical implementation in which we observed further emergent phenomena: the Expression Problem dissolves, programs are function color blind (Nystrom, 2015), ordinary arithmetic yields the relational semantics of logic programming, and self-reference resolves to multiple targets, making overlay-calculus strictly more expressive than the $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991).
+Inheritance-calculus is distilled from MIXINv2, a practical implementation in which we observed further emergent phenomena: the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); self-reference resolves to multiple targets; and programs are immune to the Expression Problem (Wadler, 1998).
+This makes inheritance-calculus strictly more expressive than the $\lambda$-calculus in both common sense and Felleisen’s sense (Felleisen, 1991).
 These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
@@
 Declarative and configuration languages are ubiquitous in modern software engineering.
-Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or overlay.
+Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or merging.
 Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively.
@@
 Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures, that is, initial algebras in the sense of universal algebra.
-We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures, in the spirit of F-coalgebras (Rutten, 2000), where semantics is determined by the observer and a finite prefix suffices for any finite observation.
+We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures, where semantics is determined by the observer and a finite prefix suffices for any finite observation.
 This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set.
 Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives.
-Through iterative refinement, beginning with an early version implemented in Nix, the system evolved into the Overlay language (Yang, 2025b), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
-In developing real programs in the Overlay language, including boolean logic and natural number arithmetic (Section 5) and a web application (Section 5), we found that the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) is structurally absent: because records are sets, inheritance is inherently commutative, idempotent, and associative (Section 3).
+The reduction produced MIXIN, an early implementation in Nix, and subsequently MIXINv2 (Yang, 2026), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
+In developing real programs in MIXINv2, including boolean logic and natural number arithmetic (Section 5) and a web application (Section Function color blindness: ), we found that the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) is structurally absent: because records are sets, inheritance is inherently commutative, idempotent, and associative (Section 3).
 Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system).
 Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction.
-We also found that boolean logic and arithmetic, implemented as separate overlay files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism; that Church-encoded values automatically form trie structures; and that operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$, which is the relational semantics of logic programming (Section 5); and that the same web application overlay file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015), equivalently CPS-agnostic behavior (Section 5).
-This paper distills these observations into overlay-calculus, a minimal computational model for declarative programming.
+We also found that boolean logic and arithmetic, implemented as separate MIXINv2 files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism; that Church-encoded values automatically form trie structures; and that operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$, which is the relational semantics of logic programming (van Emden and Kowalski, 1976) (Section 5); and that the same web application MIXINv2 file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015), equivalently CPS-agnostic behavior (Section Function color blindness: ).
+This paper distills these observations into inheritance-calculus, a minimal computational model for declarative programming.
 Function application translates directly into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument.
-Since the $\lambda$-calculus embeds into overlay-calculus, overlay-calculus is Turing complete.
-The translation reveals an asymmetry: embedding the $\lambda$-calculus into overlay-calculus requires five local, syntax-directed rules (Section 4), but the reverse is impossible—overlay enables open extension that changes operational equivalence on $\lambda$-terms, establishing a strict separation in Felleisen’s sense of expressiveness (Felleisen, 1991) (Section 4.2).
-Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, overlay-calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
-All constructions in this paper have been implemented and tested in the Overlay language, available as open source (Yang, 2025b), with a test suite covering every example.
-[^1] The Overlay language is statically typed and therefore stricter than the untyped overlay-calculus.
-All examples in this paper are ported from the Overlay language and conform to both semantics.
+Since the $\lambda$-calculus embeds into inheritance-calculus, inheritance-calculus is Turing complete.
+The translation reveals an asymmetry: embedding the $\lambda$-calculus into inheritance-calculus requires five local, syntax-directed rules (Section 4), but the reverse is not macro-expressible under the equivalence notions fixed in Section 4.2 and Appendix A— inheritance-calculus enables open extension that changes operational equivalence on $\lambda$-terms, establishing a strict separation in Felleisen’s sense of expressiveness (Felleisen, 1991) (Section 4.2).
+Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
+All constructions in this paper have been implemented and tested in MIXINv2, available as open source (Yang, 2026), with a test suite covering every example.
+[^1] MIXINv2 is statically typed and therefore stricter than the untyped inheritance-calculus.
+All examples in this paper are ported from MIXINv2 and conform to both semantics.
 Elements added to satisfy the type checker are harmless for the algorithms presented here and may aid readability.
@@
 Contributions
-  • We present overlay-calculus, a minimal computational model for declarative programming that contains only overlay constructs (record, definition, inheritance) and no functions or scalar types.
-  • Inheritance operates at the tree level: a record is a set of elements, both definitions and inheritance sources, so inheriting from two overlays that share a label recursively merges their subtrees (deep merge), rather than overriding at the method level as in mixin and trait calculi.
+  • We present inheritance-calculus, a minimal computational model for declarative programming that contains only three constructs (record, definition, inheritance) and no functions or scalar types.
+  • Inheritance operates at the tree level: a record is a set of elements, both definitions and inheritance sources, so inheriting from two mixins that share a label recursively merges their subtrees (deep merge), rather than overriding at the method level as in prior mixin and trait calculi.
 Since records are sets, inheritance is inherently commutative, idempotent, and associative.
@@
   • Self-referential records require no fixed-point combinator.
-The $\lambda$-calculus translation maps each application to a inheritance; the resulting inheritance tree is the call stack.
+The $\lambda$-calculus translation maps each application to an inheritance; the resulting inheritance tree is the call stack.
 The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree: a first-order set-theoretic query, not a fixed-point construction.
@@
 Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala and the NixOS module system, Appendix B).
-Since inheritance is idempotent, all objects contribute equally; no disambiguation is needed.
-  • The $\lambda$-calculus (in A-normal form) translates into overlay-calculus in five rules (Section 4).
-The translation is fully abstract with respect to Böhm tree equivalence (Section 4.1), yielding a semantics of the lazy $\lambda$-calculus that requires no domain theory.
-  • We observe further properties: extensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section 5).
-  • Open extension via overlay changes the operational equivalence of $\lambda$-terms, establishing via Felleisen’s Theorem 3.14 (Felleisen, 1991) that the $\lambda$-calculus cannot macro-express overlay (Section 4.2).
+Since inheritance is idempotent, all objects contribute equally; this semantics does not introduce a route-priority disambiguation rule.
+  • The $\lambda$-calculus (in A-normal form) translates into inheritance-calculus in five rules (Section 4).
+The translation is fully abstract with respect to Böhm tree equivalence under $\lambda$-definable contexts (Section 4.1), yielding a semantics of the lazy $\lambda$-calculus that requires no domain theory.
+  • We observe further properties: extensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section Function color blindness: ).
+  • Open extension via inheritance changes the operational equivalence of $\lambda$-terms, establishing via Felleisen’s Theorem 3.14 (Felleisen, 1991) that the $\lambda$-calculus cannot macro-express inheritance-calculus (Section 4.2).
 Section 5 demonstrates these properties through a case study of Nat arithmetic.
@@
 The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding.
-There is no variable binding in overlay-calculus.
-The Overlay language also provides scalar types through a foreign-function interface (FFI); overlay-calculus is what remains after removing the FFI.
+There is no variable binding in inheritance-calculus.
+MIXINv2 also provides scalar types through a foreign-function interface (FFI); inheritance-calculus is what remains after removing the FFI.
 None of the examples in this paper uses the FFI; scalars did not arise.
@@
 An inheritance is written as a reference that navigates the scope hierarchy upward and then projects properties downward. The upward navigation admits two equivalent notations:
-  • Named form: $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, analogous to Java’s Outer.this.field.
+Named form:
+$\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, analogous to Java’s Outer.this.field.
 Each $\{\ldots\}$ record has a label in its enclosing scope; $\ell_{\mathrm{up}}$ names the target enclosing scope.
-  • Indexed form: $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n\geq 0$ is a de Bruijn index (de Bruijn, 1972): $n=0$ refers to the reference’s enclosing scope, and each increment of $n$ moves one scope level further out.
+Indexed form:
+$\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n\geq 0$ is a de Bruijn index (de Bruijn, 1972): $n=0$ refers to the reference’s enclosing scope, and each increment of $n$ moves one scope level further out.
 Both forms retrieve the target scope, that is, the fully inherited record after all inheritance has been applied, and then project properties $\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ from it.
@@
 A new scope level is created at each $\{\ldots\}$ record literal.
-The scope contains all properties of that overlay, including those inherited via inheritance sources.
-For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single overlay whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$.
+The scope contains all properties of that mixin, including those inherited via inheritance sources.
+For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single mixin whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$.
 Inheritance sources do not create additional scope levels.
@@
 Since this uses $\mathtt{defines}$, the shorthand can only access properties defined in record literals, not those inherited through inheritance; one of the two qualified forms is required for inherited properties or to bypass variable shadowing.
-3. Overlay Trees
+3. Mixin Trees
 Given an AST, what are the properties visible at each position in the tree? This section answers the question by defining a collection of mutually recursive functions over paths.
-Unlike the $\lambda$-calculus, which requires reduction rules ($\beta$-reduction) to specify how terms compute, overlay-calculus has no reduction steps.
-There is no rewriting of terms, no evaluation order, and no notion of a “result” that a term reduces to.
+A mixin tree has a single abstraction, the mixin: every mixin can serve as field, method, class, and object at the same time in object-oriented terms.
+Unlike the $\lambda$-calculus, which requires $\beta$-reduction to specify how terms compute, inheritance-calculus has no reduction steps: no rewriting of terms, no evaluation order, and no notion of a result that a term reduces to.
 Instead, the semantics is purely observational: one queries whether a label $\ell$ is a property at a path $p$ in the inheritance tree.
-The mutually recursive functions defined below, namely $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overlays}$, $\mathtt{bases}$, $\mathtt{resolve}$, and $\mathtt{this}$, constitute the entire semantics.
-This is the observational perspective introduced in Section 1: like an F-coalgebra (Rutten, 2000), an overlay tree’s behavior is determined by observation (property presence), not by reduction.
-Most of these functions have direct analogues in object-oriented languages: $\mathtt{Path}$ identifies a position in the tree, $\mathtt{init}$ and $\mathtt{last}$ navigate paths, $\mathtt{defines}$ extracts the locally defined members from the AST, $\mathtt{inherits}$ extracts the inheritance declarations from the AST, $\mathtt{properties}$ computes all visible members at a path including both locally defined and inherited ones, $\mathtt{supers}$ computes the transitive inheritance closure, and $\mathtt{bases}$ collects the direct base classes.
-One function has no object-oriented analogue: $\mathtt{overlays}$ collects the identity-sharing paths that arise when inheritance introduces multiple definitions of the same label at the same scope level.
+The mutually recursive functions defined below, namely $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, and $\mathtt{this}$, constitute the entire semantics.
+Each function has an analogue in object-oriented languages: $\mathtt{Path}$ identifies a position in the tree, $\mathtt{init}$ and $\mathtt{last}$ navigate paths, $\mathtt{defines}$ extracts the locally defined members from the AST, $\mathtt{inherits}$ extracts the inheritance declarations from the AST, $\mathtt{properties}$ computes all visible members at a path including both locally defined and inherited ones, $\mathtt{supers}$ computes the transitive inheritance closure, $\mathtt{bases}$ collects the direct base classes, and $\mathtt{overrides}$ generalizes method override to the tree setting, collecting the identity-sharing paths that arise when inheritance introduces multiple definitions of the same label at the same scope level.
 Two functions, the reference resolution function $\mathtt{resolve}$ and the qualified this resolution function $\mathtt{this}$, have object-oriented analogues whose semantics differ significantly; they are discussed in detail below.
@@
 The properties of a path are its own properties together with those inherited from all supers:
-(1) $\mathtt{properties}(p)=\bigl\{\;\ell\;\big|\;(\_,\;p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\;\bigr\}$
+(1) $\mathtt{properties}(p)=\bigl\{\;\ell\;\big|\;(\_,\;p_{\mathrm{override}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{override}})\;\bigr\}$
 The remainder of this section defines $\mathtt{supers}$, the transitive inheritance closure, and its dependencies.
 Supers
-Intuitively, $\mathtt{supers}(p)$ collects every path that $p$ inherits from: the identity-sharing paths $\mathtt{overlays}$ of $p$ itself, plus the $\mathtt{overlays}$ of each direct base $\mathtt{bases}$ of $p$, and so on transitively.
+Intuitively, $\mathtt{supers}(p)$ collects every path that $p$ inherits from: the $\mathtt{overrides}$ of $p$ itself, plus the $\mathtt{overrides}$ of each direct base $\mathtt{bases}$ of $p$, and so on transitively.
 Each result is paired with the inheritance-site context $\mathtt{init}(p_{\mathrm{base}})$ through which it is reached.
 This provenance is needed by qualified this resolution, i.e.
-the $\mathtt{this}$ function defined below, to map a definition-site overlay back to the inheritance-site paths that incorporate it:
-(2) $\mathtt{supers}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{base}}),\;p_{\mathrm{overlay}})\;\big|\;p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{overlay}}\in\mathtt{overlays}(p_{\mathrm{base}})\;\bigr\}$
+the $\mathtt{this}$ function defined below, to map a definition-site mixin back to the inheritance-site paths that incorporate it:
+(2) $\mathtt{supers}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{base}}),\;p_{\mathrm{override}})\;\big|\;p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{override}}\in\mathtt{overrides}(p_{\mathrm{base}})\;\bigr\}$
 Here $\mathtt{bases}^{*}$ denotes the reflexive-transitive closure of $\mathtt{bases}$.
-The $\mathtt{supers}$ formula depends on two functions: the identity-sharing paths $\mathtt{overlays}$ and the one-hop reference targets $\mathtt{bases}$.
-We define $\mathtt{overlays}$ first.
-Overlays
-Intuitively, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overlays}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge.
-Concretely, the overlays of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
-(3) $\mathtt{overlays}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &(\_,\;p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
+The $\mathtt{supers}$ formula depends on two functions: $\mathtt{overrides}$ and the one-hop reference targets $\mathtt{bases}$.
+We define $\mathtt{overrides}$ first.
+Overrides
+Because inheritance-calculus has no separate notion of “method,” a mixin simultaneously plays the role of a class and the role of a method in object-oriented languages.
+When a mixin is viewed as a method, a child scope should be able to find it through the parent’s label even if the child does not redefine that label—this is method override.
+Concretely, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overrides}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge.
+The overrides of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
+(3) $\mathtt{overrides}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &(\_,\;p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
 It remains to define $\mathtt{bases}$, the other dependency of $\mathtt{supers}$.
@@
 Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references, analogous to the direct base classes in object-oriented languages.
-Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overlays}$ one step:
+Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overrides}$ one step:
 […]
@@
 Qualified this resolution
-$\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Each step finds, among the supers of every path in the frontier $S$, those whose overlay component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier.
+$\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Each step finds, among the supers of every path in the frontier $S$, those whose override component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier.
 After $n$ steps the frontier contains the answer:
-(6) $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
+(6) $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{override}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{override}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
 At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a non-negative integer, the recursion terminates.
@@
 Observational semantics without reduction
-The definitions above constitute the complete semantics of overlay-calculus.
+The definitions above constitute the complete semantics of inheritance-calculus.
 There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments.
 In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur.
-In overlay-calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything: it is the value.
+In inheritance-calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything: it is the value.
 The functions $\mathtt{properties}$, $\mathtt{supers}$, etc.
@@
 When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point, analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness.
-This observation model is inspired by F-coalgebras (Rutten, 2000), where identity is determined by observable behavior rather than by internal structure.
+Identity is determined by observable behavior rather than by internal structure.
 4. Embedding the $\lambda$-Calculus
-The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to overlay-calculus.
+The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to inheritance-calculus.
 In ANF, every intermediate result is bound to a name and each expression contains at most one application:
@@
 Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results, and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus.
-The ANF structure matches overlay-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
-Let $\mathcal{T}$ denote the translation function. Each $\lambda$-abstraction introduces one scope level in the translated overlay tree.
+The ANF structure matches inheritance-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
+Let $\mathcal{T}$ denote the translation function. Each $\lambda$-abstraction introduces one scope level in the translated mixin tree.
 $x$ ($\lambda$-bound, index $n$)   $\longrightarrow$ $\uparrow^{\!n}.\mathrm{argument}$
@@
 In both cases the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
-The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to overlay-calculus: every ANF term has an image, and the construction is compositional.
-Since the $\lambda$-calculus is Turing complete, so is overlay-calculus.
+The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to inheritance-calculus: every ANF term has an image, and the construction is compositional.
+Since the $\lambda$-calculus is Turing complete, so is inheritance-calculus.
 An elementary semantics of the $\lambda$-calculus
@@
 This induced semantics is more elementary than traditional approaches in the mathematical machinery it requires:
-  • No function spaces.
+No function spaces:
 Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices.
@@
 The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix C), which requires only set-theoretic monotonicity, which is strictly weaker than Scott continuity.
-  • No substitution.
+No substitution:
 Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi.
 Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree; no substitution occurs at any point.
-  • No reduction rules.
+No reduction rules:
 There are no $\beta$-reduction steps, no evaluation contexts, and no choice between call-by-name and call-by-value (Plotkin, 1975).
@@
 The key step is that function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function; no function spaces are required.
-Section 4.1 shows that the convergence behavior of the translated overlay tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate, meaning a term converges iff it has a head normal form, and fully abstract, meaning two terms are overlay-contextually equivalent iff they have the same Böhm tree.
-Combining these two facts, elementary machinery and full abstraction, yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics that lives entirely in naive set theory.
+Section 4.1 shows that the convergence behavior of the translated mixin tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate, meaning a term converges iff it has a head normal form, and fully abstract with respect to $\lambda$-definable contexts, meaning two $\lambda$-terms are equivalent under the translation iff they have the same Böhm tree.
+Combining these two facts, elementary machinery and full abstraction, yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics (w.r.t.
+$\lambda$-definable contexts) that lives entirely in naive set theory.
 The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem.
@@
 Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Abramsky and McCusker’s game semantics (Abramsky and McCusker, 1995) achieves full abstraction but at the cost of even greater mathematical complexity.
-The overlay tree semantics achieves full abstraction with strictly simpler foundations than either.
+The mixin tree semantics achieves full abstraction with strictly simpler foundations than either.
 4.1. Böhm Tree Correspondence
-The translation $\mathcal{T}$ induces a semantics on the $\lambda$-calculus via the overlay tree equations (1)–(6).
+The translation $\mathcal{T}$ induces a semantics on the $\lambda$-calculus via the mixin tree equations (1)–(6).
 We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
-An overlay tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$.
+A mixin tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$.
 We write $T{\Uparrow}$ if no such depth exists.
-Definition 4.1 (Overlay-convergence).
-An overlay tree $T$ converges, written $T{\Downarrow}$, if there exists $n\geq 0$ such that:
+Definition 4.1 (Inheritance-convergence).
+A mixin tree $T$ converges, written $T{\Downarrow}$, if there exists $n\geq 0$ such that:
 $\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{properties}\!\bigl(\,\underbrace{\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
@@
 A closed ANF $\lambda$-term $M$ has a head normal form if and only if $\mathcal{T}(M){\Downarrow}$.
-Theorem 4.3 (Full Abstraction).
-For closed $\lambda$-terms $M$ and $N$, define $\mathcal{T}(M)\approx\mathcal{T}(N)$ iff for every closing context $C[\cdot]$, $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$.
+Theorem 4.3 (Full abstraction w.r.t. $\lambda$-definable contexts).
+For closed $\lambda$-terms $M$ and $N$, define $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ iff for every closing $\lambda$-calculus context $C[\cdot]$, $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$.
 Then:
-$\mathcal{T}(M)\approx\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
-Together, Adequacy and Full Abstraction establish that the overlay tree semantics is a computation mechanism for Böhm trees (Barendregt, 1984), the canonical fully abstract model of the lazy $\lambda$-calculus, using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory.
+$\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
+Here $\approx_{\lambda}$ quantifies over $\lambda$-calculus contexts only; inheritance-calculus contexts are strictly more discriminating (Section 4.2).
+Together, Adequacy and Full Abstraction establish that the translation $\mathcal{T}$ is a fully faithful embedding of the lazy $\lambda$-calculus into inheritance-calculus with respect to Böhm tree equivalence (Barendregt, 1984), using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory.
 This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract, as it identifies terms that observational equivalence distinguishes (Plotkin, 1975).
 The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix D.
+The relationship between $\approx_{\lambda}$ and full inheritance-contextual equivalence is clarified by the expressive asymmetry of the next section: inheritance-calculus contexts can separate $\lambda$-terms that are Böhm-tree equivalent, so ${\approx_{\lambda}}\subsetneq{(\approx_{\mathrm{inh}}|_{\mathrm{im}(\mathcal{T})})}$.
+This gap is precisely the source of inheritance-calculus’s additional expressiveness (Theorem A.3).
 4.2. Expressive Asymmetry
-Two $\lambda$-terms that are Böhm-tree equivalent—and therefore indistinguishable in the $\lambda$-calculus—can be separated in overlay-calculus by overlaying new observable projections onto an existing definition via open extension.
-By Felleisen’s expressiveness criterion (Felleisen, 1991), overlay-calculus is strictly more expressive than the $\lambda$-calculus: the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
+Two $\lambda$-terms that are Böhm-tree equivalent—and therefore indistinguishable in the $\lambda$-calculus—can be separated in inheritance-calculus by inheriting new observable projections onto an existing definition via open extension.
+By Felleisen’s expressiveness criterion (Felleisen, 1991), inheritance-calculus is strictly more expressive than the $\lambda$-calculus: the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
 The formal definitions, constructions, and proofs are given in Appendix A; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
-The underlying mechanism is open extension: overlay definitions compose via open recursion, so existing definitions gain new observable properties without modification.
+The underlying mechanism is open extension: mixin definitions compose via open recursion, so existing definitions gain new observable properties without modification.
 In the $\lambda$-calculus, a function’s result is closed—there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
 5. Case Study: Nat Arithmetic
-We implemented boolean logic and natural number arithmetic, both unary and binary, as overlay files (Yang, 2025b).
+We implemented boolean logic and natural number arithmetic, both unary and binary, as MIXINv2 files (Yang, 2026).
 This section traces the unary Nat case in detail.
@@
 NatData
-Following the Church-encoding convention (Yang, 2025b), we define a factory containing the abstract product type and its two constructors.
+Following the Church-encoding convention (Yang, 2026), we define a factory containing the abstract product type and its two constructors.
 A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field.
@@
 NatPlus
-Addition is a separate overlay file that inherits $\mathrm{NatData}$ and extends each constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field.
+Addition is a separate MIXINv2 file that inherits $\mathrm{NatData}$ and extends each constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field.
 The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
@@
 Tree-level inheritance and return types
-Each operation is an independent overlay that extends a factory’s subtree.
-Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited overlays.
+Each operation is an independent mixin that extends a factory’s subtree.
+Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited mixins.
 Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory.
@@
 The result $\mathrm{sum}$ therefore collects all paths reachable through any combination of the two operands, which is the Cartesian product.
-The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: $\mathtt{overlays}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.
+The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: $\mathtt{overrides}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.
 6. Discussion
 6.1. Emergent Phenomena
-Several features of overlay-calculus are primitive yet have no direct counterpart in the $\lambda$-calculus. Encoding any one of them requires pervasive, non-local changes to program structure.
-  1. (1)
-    Self-referential records with lazy evaluation.
-An overlay’s properties can reference other properties and enclosing scopes through qualified this without explicit binding.
+Several features of inheritance-calculus are primitive yet have no direct counterpart in the $\lambda$-calculus. Encoding any one of them requires pervasive, non-local changes to program structure.
+Self-referential records with lazy evaluation:
+A mixin’s properties can reference other properties and enclosing scopes through qualified this without explicit binding.
 In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (Boudol, 2004) (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
-  2. (2)
-    Open recursion with symmetric inheritance.
-Inheritance merges definitions from independent sources, all sharing the same inherited overlay tree.
+Open recursion with symmetric inheritance:
+Inheritance merges definitions from independent sources, all sharing the same inherited mixin tree.
 Encoding this in the $\lambda$-calculus requires maintaining an extensible dictionary of methods that can be merged from multiple directions.
 Cook’s denotational semantics of inheritance (Cook, 1989) devoted an entire dissertation to formalizing this pattern in a functional setting.
-  3. (3)
-    Commutative, idempotent merge.
-Inheritance in overlay-calculus is symmetric: a record’s elements form a set, so reordering them produces the same observable result.
+Commutative, idempotent merge:
+The inheritance operation is symmetric: a record’s elements form a set, so reordering them produces the same observable result.
 In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery (Harper and Pierce, 1991; Cardelli, 1994).
-  4. (4)
-    Relational semantics.
-Every overlay-calculus operation is defined by recursion on overlay trees, not on individual values.
-When an overlay tree is viewed as a trie, the correspondence with relational algebra becomes precise: a path from the root to a leaf encodes a single value (e.g., $\mathrm{Successor}\to\mathrm{predecessor}\to\mathrm{Zero}$ encodes the natural number $1$); a trie union encodes a set of such values, i.e., a unary relation.
+Relational semantics:
+Every inheritance-calculus operation is defined by recursion on mixin trees, not on individual values.
+When a mixin tree is viewed as a trie, the correspondence with relational algebra becomes precise: a path from the root to a leaf encodes a single value (e.g., $\mathrm{Successor}\to\mathrm{predecessor}\to\mathrm{Zero}$ encodes the natural number $1$); a trie union encodes a set of such values, i.e., a unary relation.
 An operation that takes two trie-union operands traverses every combination of their paths through $\mathtt{bases}$ (equation 4), which is a cross-join; the transitive closure $\mathtt{bases}^{*}$ in $\mathtt{supers}$ (equation 2) computes the union of all reachable paths; and $\mathtt{properties}$ (equation 1) projects the observable labels.
@@
 In the Case Study (Section 5), the $\mathrm{Plus}$ and $\mathrm{Equal}$ operations are defined on single Nat values.
-When applied to set-valued overlays $\{1,2\}$ and $\{3,4\}$, the same definitions, without any modification, produce the set of all pairwise sums $\{4,5,6\}$.
+When applied to set-valued mixins $\{1,2\}$ and $\{3,4\}$, the same definitions, without any modification, produce the set of all pairwise sums $\{4,5,6\}$.
 This is relational semantics in action: each operation distributes over every combination of its inputs.
@@
 No code change is needed to move from one to the other.
-  5. (5)
-    Function color blindness.
-We implemented a web application as overlay files (Yang, 2025b), initially targeting Python’s synchronous sqlite3 and http.server modules.
-The business logic overlay file contained no async-specific code.
-Later, we replaced the FFI layer with one built on aiosqlite and starlette; the business logic overlay file did not change.
-This is an instance of function color blindness (Nystrom, 2015): async and sync functions are distinct types in the host language, yet the overlay file is agnostic to the distinction because its empty property slots are filled at assembly time, with coroutine handlers under asyncio, or with direct implementations under synchronous execution.
+Function color blindness:
+In a web application (Yang, 2026) initially targeting Python’s synchronous sqlite3 and http.server modules, the business logic MIXINv2 file contained no async-specific code.
+When the FFI layer was replaced with one built on aiosqlite and starlette, the business logic MIXINv2 file did not change.
+This is an instance of function color blindness (Nystrom, 2015): async and sync functions are distinct types in the host language, yet the MIXINv2 file is agnostic to the distinction because its empty property slots are filled at assembly time, with coroutine handlers under asyncio, or with direct implementations under synchronous execution.
 More generally, empty slots can be filled with effect handlers (Plotkin and Power, 2003; Pretnar, 2015) under a CPS interpretation, or with direct implementations under a direct-style interpretation, without explicit monad machinery.
-Each of these features is primitive in overlay-calculus.
+Immunity to nonextensibility:
+Programming languages typically provide a mix of extensible and nonextensible constructs.
+In such languages, extensibility is opt-in: frameworks must encode nonextensible constructs into extensible constructs using sophisticated mechanisms (Liang et al., 1995; Lämmel and Jones, 2003; Löh and Hinze, 2006; Swierstra, 2008; Carette et al., 2009; d.
+S.
+Oliveira and Cook, 2012; Kiselyov et al., 2013; Wu et al., 2014; Kiselyov and Ishii, 2015; Wang and d.
+S.
+Oliveira, 2016; Leijen, 2017; Poulsen and van der Rest, 2023), and code that uses nonextensible constructs cannot be extended unless rewritten in the framework.
+In contrast, every construct of inheritance-calculus (Section 2) is extensible via deep merge (equation 3).
+There is no nonextensible construct, so it is impossible to write a nonextensible program.
+In the Case Study (Section 5), new operations and data types are added without modifying existing files.
+Tree-level inheritance (Section 5) and multi-path qualified this (equation 6; Appendix B) account for this.
+The Expressive Asymmetry (Section 4.2; Theorem A.3) provides the formal counterpart: the mixin trees in its separating construction are mechanical translations of $\lambda$-terms via $\mathcal{T}$ (Section 4), yet a new observable projection is added to them without modification.
+This is the extensibility that the Expression Problem (Wadler, 1998) demands.
+[^3] As an untyped calculus, inheritance-calculus cannot violate the extensibility requirement of the Expression Problem but does not encounter its type-safety requirement.
+A typed language built on inheritance-calculus (Yang, 2026) inherits the immunity to nonextensibility and need only ensure that its type system does not introduce nonextensible constructs.
+Each of these features is primitive in inheritance-calculus.
 In the $\lambda$-calculus, encoding them requires pervasive changes to program structure: every scope must thread a mutable store for lazy allocation, every method invocation must consult a dynamically constructed dictionary for open recursion, and every composition must be wrapped in order-insensitive infrastructure for commutativity.
 These are global, cascading transformations, not local rewrites of individual terms.
-6.2. Classification of Semantic Variants
-The goals of declarative programming rule out several properties of the model: separate primitives for methods, classes, types, and objects; mutable reference cells; domain-theoretic approximation; and asymmetric override.
-Ruling these out leaves a minimal syntax of three constructs: records, definitions, and inheritance.
-Two degrees of freedom remain.
-These dimensions are independent, and each combination yields a coherent semantics.
-They arise in any system whose definitions form a tree of named scopes, where each node is a path and references are path-dependent, even in systems ruled out above, such as Scala, which has separate primitives for methods, classes, objects, and types.
-Existing designs occupy several of the resulting cells:
-No deep merge Deep merge
-Early binding CUE (van Lohuizen, 2019)
-Dynamic scope union file systems, Jsonnet (Cunningham, 2014)
-Lexical scope Scala (Odersky et al., 2004) NixOS modules (NixOS Contributors, 2024)
-Multi-target $\mathtt{this}$ overlay-calculus
-Early binding
+In common sense, inheritance-calculus is strictly more expressive than the $\lambda$-calculus.
+6.2. Semantic Alternatives
+The preceding item established that inheritance-calculus is immune to nonextensibility. We now ask whether an alternative design could share this property by surveying every candidate known to us.
+Degrees of freedom
+Every language whose programs are tree-structured ASTs with references can be characterized by two finite sets: a set $\mathcal{N}$ of node types, each determining how a subtree organizes its children, and a set $\mathcal{R}$ of reference types, each determining how one position in the tree refers to another.
+The syntax is fixed by the choice of $\mathcal{N}$ and $\mathcal{R}$; the semantics is determined by how each node type structures composition and how each reference type resolves its target.
+Language Node types $\mathcal{N}$ Reference types $\mathcal{R}$
+Java (OOP) class, interface, method, block, conditional, loop, throw, lambda, … variable, field access, method call, extends/implements, import, …
+Haskell (FP) $\lambda$, application, let, case, data, type class, instance, … variable, constructor, type class dispatch, import, …
+Scala (multi-paradigm) class, trait, object, method, block, conditional, match, while, throw, … variable, member access, extends, with, implicits, import, …
+$\lambda$-calculus abstraction, application variable
+Inheritance-calculus record inheritance
+Immunity to nonextensibility
+A path $p$ in a program $P$ is extensible if there exists a program $Q$ such that composing $P$ with $Q$ through a reference makes new properties observable at $p$ without modifying $P$.
+A language is immune to nonextensibility if every path in every well-formed program is extensible.
+We now classify the design space to explain why we chose the semantics of Section 3.
+Two axes remain free: the composition mechanism—how the single reference type incorporates definitions; and the reference resolution—how the single reference type finds its target.
+The composition mechanism
+Every composition mechanism determines how the definitions of one abstraction are incorporated into another.
+The following table classifies the known candidates by whether they satisfy commutativity (C), idempotence (I), and associativity (A), and by their extensibility model.
+Candidate C I A Extensibility Representative
+Deep merge $\checkmark$ $\checkmark$ $\checkmark$ universal this paper
+Shallow merge $\times$ $\checkmark$ $\checkmark$ top-level only JS spread
+Conflict rejection $\checkmark$ $\checkmark$ $\checkmark$ none on overlap Harper–Pierce (Harper and Pierce, 1991)
+Priority merge $\times$ $\times$ $\checkmark$ universal Jsonnet +, Nix //
+$\beta$-reduction $\times$ $\times$ $\times$ opt-in Java, Haskell, Scala, …
+Conditional breaks monotonicity
+Deep merge:
+Merging same-label definitions recurses into their subtrees ($\mathtt{overrides}$, equation 3). Every label is an extension point, whether or not the original author anticipated extension.
+Shallow merge:
+Nested labels cannot be independently extended.
+Conflict rejection:
+Same-label definitions are rejected as errors.
+All three identities hold—but vacuously: the identities are satisfied because the compositions that would test them are refused.
+Conflict rejection prevents extending existing labels (Harper and Pierce, 1991)—the composability that the Expression Problem (Wadler, 1998) requires.
+Priority merge:
+$A+B\neq B+A$: one source takes precedence (Cunningham, 2014; Dolstra et al., 2010), introducing an ordering dependency.
+$\beta$-reduction:
+Only the parameters placed by the author are extension points; the function body is closed.
+To make a new aspect extensible, the original function must be rewritten to accept an additional parameter.
+This applies to every language whose composition mechanism is function application—functional, object-oriented, and multi-paradigm alike—because method calls, virtual dispatch, and type-class resolution all reduce to $\beta$-reduction.
+The Expression Problem (Wadler, 1998) is hard in all of them for the same reason: extensibility is opt-in.
+The frameworks that simulate universal extensibility (visitor pattern, object algebras (d.
+S.
+Oliveira and Cook, 2012), finally tagless interpreters (Carette et al., 2009)) are the machinery needed to work around this limitation.
+Conditional incorporation:
+Removing or conditionally including definitions breaks monotonicity, so the semantic equations (Appendix C) may have no fixed point (Tarski, 1955).
+Among the candidates above, deep merge is the only one we found that is commutative, idempotent, associative, and universally extensible.
+The other candidates each introduce a vulnerability to nonextensibility or break monotonicity.
+The reference resolution
+Once deep merge over records is fixed as the composition mechanism, four of the six semantic equations are determined by the syntax: $\mathtt{properties}$ (1) reads the labels of a record, $\mathtt{bases}$ (4) reads its inheritance sources, $\mathtt{supers}$ (2) takes the transitive closure, and $\mathtt{overrides}$ (3) implements the merge.
+The only remaining degree of freedom is how references resolve: the $\mathtt{resolve}$ (5) and $\mathtt{this}$ (6) equations.
+We examine the three known candidates.
+Early binding:
 With early binding, references resolve at definition time to a fixed path, independently of how the record is later inherited.
-CUE (van Lohuizen, 2019) exemplifies this: field references are statically resolved, and a field in a nested record cannot refer back to properties of its enclosing context contributed by later composition.
-Early binding combined with deep merge is coherent and practically useful, but cross-overlay cooperation requires explicit wiring rather than inheritance: a definition cannot observe properties contributed by overlays composed after it.
-Dynamic scope
-With dynamic scope, references resolve against the context at the point of use rather than the definition site.
-Early Lisp implementations used dynamic scope (McCarthy, 1978), where variable references are resolved against the run-time call stack.
+CUE (van Lohuizen, 2019) exemplifies this: field references are statically resolved, and a nested field cannot refer to properties of its enclosing context contributed by later composition.
+This prevents the mixin from serving as a call stack for the $\lambda$-calculus: $\beta$-reduction requires the callee to observe the argument supplied at the call site, which is a later composition.
+With early binding, references are frozen before this composition occurs.
+In the equations, $\mathtt{resolve}$ (5) calls $\mathtt{this}$ (6), which walks the inherited tree structure; late binding is what allows $\mathtt{this}$ to see properties contributed by later-composed mixins.
+Without a Turing-complete mixin, functions must be reintroduced as the composition mechanism, re-exposing the opt-in extensibility problem above.
+Dynamic scope:
+With dynamic scope, references resolve at the point of use rather than the definition site.
+Jsonnet (Cunningham, 2014) follows this approach: self always refers to the final merged object, not the object at the definition site.
 Union file systems exhibit the same property at the file-system level: a symbolic link’s relative path resolves against the mount context at dereference time, not against the layer that defined the link.
-Jsonnet (Cunningham, 2014) follows this approach: self always refers to the final merged object, not the object at the definition site.
-When the Scala multi-path example (Appendix B) is translated to Jsonnet, it does not error, but outer resolves to the merged MyObjectA rather than to Object1 or Object2 separately, silently discarding the distinction between the two inheritance routes.
-Dynamic scope is flexible but loses the ability to track which inheritance route a reference came from.
-Deep merge
-Without the second case of equation (3), $\mathtt{overlays}(p)$ would always return $\{p\}$, a single element.
-When two independently authored overlays both define a label with the same name at the same scope level, only the definition at the literal path $p$ would contribute to $\mathtt{supers}$ and hence to $\mathtt{properties}$; the other definitions would be silently invisible.
-This is the situation in most composition systems: pytest permits at most one fixture per name per scope; trait systems (Schärli et al., 2003) flag same-name methods as conflicts.
-Under any of these rules, two independently authored overlays that both extend an entity with the same name, which is the situation that arises naturally when solving the Expression Problem (Wadler, 1998), cannot be composed.
-The non-trivial case of equation (3) removes this limitation: it collects all paths that share the same label and merges their definitions rather than shadowing each other.
-This is the record-level counterpart of method override in object-oriented languages, and the mechanism by which the calculus admits extensibility in both dimensions of the Expression Problem (Wadler, 1998) (Section 5).
-Multi-target qualified this
-The qualified this reference $\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Scheme (Sussman and Steele Jr., 1975) introduced lexical scope, where a de Bruijn index resolves in a single step by indexing into the statically determined environment; this is predictable but rigid.
-Scala (Odersky et al., 2004) and the NixOS module system (NixOS Contributors, 2024) follow the lexical approach: each this reference is bound to a single statically known path.
-When inheritance introduces multiple inheritance routes to the same scope, Scala rejects the ambiguity at compile time (Appendix B) and the NixOS module system rejects the duplicate option declaration (Appendix B); neither can express that all routes are equally valid.
-Overlay-calculus combines the predictability of lexical scope with the flexibility of dynamic scope.
-The reference itself is lexically determined at the definition site as a de Bruijn index, so what $\mathtt{this}$ looks up is predictable.
-But because inheritance interleaves scopes from different overlays, the result depends on the inherited tree structure: $\mathtt{this}$ walks upward one scope level at a time, consulting $\mathtt{supers}$ at each level, and naturally tracks a set of inheritance-site paths rather than a single one.
-All routes contribute equally to the inherited result, and since inheritance is idempotent, duplicated contributions are harmless.
-The only observable is property presence, making multi-path resolution well-defined without disambiguation.
-Our first three implementations of the Overlay language used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972).
+Dynamic scope breaks the $\lambda$-calculus embedding (Section 4).
+The Substitution Lemma (Lemma D.2) relies on the base case $M=y$ ($y\neq x$): the de Bruijn index $m$ of $y$ walks up from the definition-site path $\mathtt{init}(p_{\mathrm{def}})$ to a scope level different from $x$’s binder, so the substitution $\mathrm{argument}\mapsto\mathcal{T}(V)$ at $x$’s scope has no effect.
+With dynamic scope this separation fails.
+[^4] Counterexample: $(\lambda x.\,\lambda y.\,x)\;V_{1}\;V_{2}$.
+The reference to $x$ inside $\lambda y.\,x$ translates to $\uparrow^{\!1}.\mathrm{argument}$ (de Bruijn index $1$).
+With late binding, $\mathtt{this}$ starts from the definition-site scope (the inner $\lambda$) and walks up one level to the outer $\lambda$, reaching the slot $\mathrm{argument}\mapsto\mathcal{T}(V_{1})$.
+With dynamic scope, the reference resolves from the use-site context, where the innermost enclosing scope is the application $\{C_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$; walking up one level reaches this scope’s $\mathrm{argument}$ slot, returning $\mathcal{T}(V_{2})$ instead of $\mathcal{T}(V_{1})$.
+This is the classical variable-capture failure of dynamic scoping.
+As with early binding, the mixin cannot serve as a complete replacement for functions.
+Single-target $\mathtt{this}$:
+Single-target $\mathtt{this}$ is a promising candidate: the Single-Path Lemma (D.1) shows that for translated $\lambda$-terms, the frontier set $S$ in $\mathtt{this}$ (6) always contains exactly one path, so the $\lambda$-calculus embedding does not require multi-target resolution.
+However, single-target $\mathtt{this}$ is vulnerable to nonextensibility for a different reason.
+When multiple inheritance routes reach the same scope, single-target resolution must reject the situation or select one route.
+The NixOS module system (NixOS Contributors, 2024) raises a duplicate-declaration error (Appendix B).
+This rejects nested extensions where two independently authored modules define the same option—a composition that arises when solving the Expression Problem (Wadler, 1998) (Section 5).
+Our first three implementations of MIXINv2 used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972).
 All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope.
-The difficulty was structural, not incidental: a closure captures a single environment, so a fixed-point combinator solves for a single self-reference; a stack-based environment is a linear chain where each scope level has one parent.
+The difficulty was structural: a closure captures a single environment, so a fixed-point combinator solves for a single self-reference; a stack-based environment is a linear chain where each scope level has one parent.
 Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix B).
 Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope.
-If the equations were restricted to single-target resolution, as in DOT’s single-valued self variable (Amin et al., 2016), they would need to reject structurally legitimate cases, adding complexity rather than removing it.
-Consequences of the alternative choices
-Choosing differently in any dimension requires introducing machinery that is absent from the equations of Section 3.
-With early binding, the equations must freeze reference targets at definition time, so a definition cannot observe properties contributed by later-composed overlays.
-With dynamic scope, the equations must resolve references against a runtime context rather than the definition-site structure, losing the ability to distinguish inheritance routes.
-Restricting $\mathtt{overlays}(p)$ to $\{p\}$ means same-label definitions from different inheritance sources no longer share identity, requiring either a method resolution order (MRO) (Barrett et al., 1996) to linearize them or a shadowing rule to select one and discard the rest, which is why systems without deep merge rarely support path-dependent multiple inheritance in practice.
-Restricting $\mathtt{this}$ to a single path requires a mechanism to select among inheritance routes (Scala rejects the ambiguity at compile time; the NixOS module system raises a duplicate-declaration error), and loses contributions from all non-selected routes.
-These mechanisms, including early-binding freezing, runtime context threading, MRO, shadowing, route selection, and error reporting, are more artificial than their absence.
-The equations of Section 3 contain none of them.
+The following table summarizes the vulnerabilities of all alternatives at both levels.
+Candidate Vulnerability Representative
+composition Shallow merge nested extensions not composable JS spread
+Conflict rejection same-label extensions rejected Harper–Pierce (Harper and Pierce, 1991)
+Priority merge ordering dependency Jsonnet + (Cunningham, 2014), Nix // (Dolstra et al., 2010)
+$\beta$-reduction opt-in extensibility Java, Haskell, Scala, …
+Conditional breaks monotonicity
+resolution Early binding mixin cannot replace function CUE (van Lohuizen, 2019)
+Dynamic scope mixin cannot replace function Jsonnet (Cunningham, 2014), union FS
+Single-target $\mathtt{this}$ nested extensions rejected NixOS modules (NixOS Contributors, 2024)
+No alternative we examined at either level is immune to nonextensibility.
+The semantics of inheritance-calculus was not chosen from a menu.
+We arrived at it by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-path compositions that the Expression Problem demands (three successive implementations confirmed this—see above).
+Deep merge with multi-target late-binding $\mathtt{this}$ is what remained.
 7. Related Work
@@
 None of these reductions exhibits emergent phenomena absent from the model it reduces.
+Semantic descriptions of the $\lambda$-calculus
+The semantics of the $\lambda$-calculus can be defined in several fundamentally different ways.
+Operational semantics takes $\beta$-reduction $(\lambda x.\,M)\,N\to M[N/x]$ as the primitive notion of computation (Church, 1936, 1941; Barendregt, 1984).
+This requires capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi (Abadi et al., 1991), and a choice of evaluation strategy: call-by-name versus call-by-value (Plotkin, 1975) or the lazy strategy of Abramsky (Abramsky, 1990).
+Denotational semantics (Scott, 1970, 1976) interprets terms in a mathematical domain, but giving meaning to $\beta$-reduction requires a reflexive domain $D\cong[D\to D]$, which has no set-theoretic solution by Cantor’s theorem; Scott’s $D_{\infty}$ resolves this by constructing $D$ as an inverse limit of continuous lattices (Scott, 1972).
+The resulting semantics is adequate but not fully abstract (Plotkin, 1977; Milner, 1977): it identifies terms that observational equivalence distinguishes (Plotkin, 1975).
+Game semantics (Abramsky et al., 2000; Hyland and Ong, 2000; Nickau, 1994) achieves full abstraction for PCF and for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995) at the cost of considerably more intricate mathematical machinery (arenas, strategies, innocence conditions).
+In each of these approaches, the function space $D\to D]$ is the central obstacle: operational semantics needs it implicitly (via substitution), denotational semantics constructs it explicitly, and game semantics refines it via strategies.
+None of these approaches yields a semantics of the lazy $\lambda$-calculus that simultaneously lives in naive set theory and is fully abstract: Scott’s $D_{\infty}$ is adequate but not fully abstract, and game semantics is fully abstract but requires machinery (arenas, strategies, innocence conditions) far beyond set theory.
+None of them eliminates the function space from the semantic domain; none avoids capture-avoiding substitution; and none dispenses with reduction rules or evaluation strategies.
+Set-theoretic models of the $\lambda$-calculus
+Several families of $\lambda$-calculus models avoid the inverse-limit construction of Scott’s $D_{\infty}$ by working in set-theoretic rather than domain-theoretic settings.
+Graph models (Engeler, 1981; Plotkin, 1993) take a web $A$ equipped with an injection $\iota:\mathcal{P}_{\mathrm{fin}}(A)\times A\hookrightarrow A$ and set $D=\mathcal{P}(A)$.
+Application is $f\cdot x=\{\,b\mid\exists\,S\subseteq x,\;\iota(S,b)\in f\,\}$: the function space $D\to D]$ is not constructed explicitly but encoded inside $D$ via $\iota$.
+The construction uses only powerset lattices, no Scott topology or continuity conditions.
+Graph models are sensible (they equate all unsolvable terms) but not fully abstract: Bucciarelli and Salibra (Bucciarelli and Salibra, 2008) proved that the greatest sensible graph theory equals the Böhm tree theory $\mathcal{B}$, which is strictly coarser than observational equivalence for the lazy $\lambda$-calculus.
+Filter models (Barendregt et al., 1983; Coppo and Dezani-Ciancaglini, 1978) interpret a term as the filter (upward-closed, intersection-closed set) of all intersection types derivable for it.
+The collection of all such filters forms an $\omega$-algebraic lattice.
+Barendregt, Coppo, and Dezani-Ciancaglini proved that type assignment is sound and complete with respect to the filter $\lambda$-model.
+Dezani-Ciancaglini et al.
+(Dezani-Ciancaglini et al., 1998) showed that the $\lambda$-theory of the BCD filter model coincides exactly with Böhm tree equality: two terms receive the same set of types if and only if they have the same Böhm tree.
+Filter models are more discriminating than graph models but are still not fully abstract for the lazy $\lambda$-calculus, and require the additional machinery of intersection type theory and algebraic lattices.
+Relational models (Ehrhard, 2005; Bucciarelli and Ehrhard, 2001) arise from the relational semantics of linear logic: the exponential modality !!
+is interpreted via finite multisets $\mathcal{M}_{\mathrm{fin}}(A)$, giving a co-Kleisli category whose morphisms are relations from $\mathcal{M}_{\mathrm{fin}}(A)$ to $B$.
+The multiset structure corresponds to non-idempotent intersection types (de Carvalho, 2018), which carry quantitative information (bounds on reduction length) that idempotent systems lack.
+Ehrhard (Ehrhard, 2005) showed that in the finiteness-space refinement of this model, standard fixed-point operators are not finitary, and a standard reflexive object for the pure $\lambda$-calculus cannot exist in this category.
+For the lazy $\lambda$-calculus, full abstraction has been achieved only through game semantics (Abramsky and Ong, 1993; Abramsky and McCusker, 1995); none of the three set-theoretic families attains it.
+All three families are semantic models of the $\lambda$-calculus: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure.
+None of them is a self-contained calculus with its own primitives into which the $\lambda$-calculus embeds.
+Moreover, all three identify terms more coarsely than observational equivalence—graph and filter models coincide with Böhm tree equality, while relational models cannot even host a standard reflexive object (Ehrhard, 2005)—and none provides a fully abstract semantics for the lazy $\lambda$-calculus without resorting to game-theoretic machinery.
 Denotational semantics of inheritance
@@
 Whether these three roles could be unified under a single construct remains an open question in the DOT line of work.
-DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure overlay inheritance (see Section 8).
+DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure mixin inheritance (see Section 8).
 Configuration languages
@@
 Module systems and deep merge
-The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same overlay semantics that distinguishes tree-level inheritance from mixin and trait composition.
+The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same deep-merge semantics that distinguishes tree-level inheritance from method-level composition in prior mixin and trait calculi.
 When two modules define the same nested path, their subtrees are merged rather than overridden.
@@
 When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix B).
-The module system cannot express that both routes contribute equally to the same option, that is, the multi-path situation that recursive merging naturally gives rise to, which motivated the coalgebraic semantics of Section 3.
+The module system cannot express that both routes contribute equally to the same option, that is, the multi-path situation that recursive merging naturally gives rise to, which motivated the observational semantics of Section 3.
 Modular software architectures
@@
 The symmetric, associative nature of inheritance (Section 3) explains why such systems work: components can be inherited in any order without changing behavior.
-These systems implement overlay-calculus patterns ad hoc.
+These systems implement inheritance-calculus patterns ad hoc.
 Union file systems
 Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view.
-Their semantics resembles overlay-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers.
-We implemented overlay-calculus directly as a union file system (Yang, 2025a), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives.
+Their semantics resembles inheritance-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers.
+We implemented inheritance-calculus directly as a union file system (Yang, 2025), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives.
 In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance: a file-system-as-compiler in which the directory tree is both the source program and the compilation target.
 8. Future Work
-The Overlay language (Yang, 2025b) is an executable implementation of overlay-calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the inheritance, and a foreign-function interface (FFI) that introduces scalar values from the host language.
+MIXINv2 (Yang, 2026) is an executable implementation of inheritance-calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the mixin tree, and a foreign-function interface (FFI) that introduces scalar values from the host language.
 The future work described below concerns the distance between this implementation and a fully practical language, spanning both theoretical foundations and library-level encodings.
-Type system
-The Overlay language’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result.
+8.1. Type System
+MIXINv2’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result.
 Formalizing the soundness of these checks, that is, proving that well-typed programs do not produce dangling references at runtime, is one direction of future work.
-Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how overlay-calculus relates to the $\lambda$-calculus: a Typed overlay-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with overlay inheritance.
+Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how inheritance-calculus relates to the $\lambda$-calculus: a Typed inheritance-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with mixin inheritance.
 A formal type system would be amenable to mechanization in proof assistants (Coq, Agda) for verifying type safety, progress, and preservation properties.
 Beyond soundness of the existing checks, additional type system features would strengthen the language.
-A key example is totality checking: verifying that a inheritance provides implementations for all required slots, not merely that references resolve.
-In untyped overlay-calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary: it defines a structural slot but imposes no constraint on what must fill it.
+A key example is totality checking: verifying that an inheritance provides implementations for all required slots, not merely that references resolve.
+In untyped inheritance-calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary: it defines a structural slot but imposes no constraint on what must fill it.
 Totality checking would enforce such declarations statically, rejecting inheritances that leave required slots unfilled.
-Standard library
+8.2. Standard Library
 The remaining items concern the standard library: they require no extensions to the calculus or type system, but involve nontrivial encodings within the existing framework.
-Overlay-calculus is open by default: inheritance can freely merge any two overlays, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$).
+Inheritance-calculus is open by default: inheritance can freely merge any two mixins, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$).
 This is not a defect; it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 5).
@@
 9. Conclusion
-The goals of declarative programming forced the semantics of overlay-calculus upon us.
-The emergent phenomena exhibit an expressive asymmetry between overlay-calculus and the $\lambda$-calculus.
+Immunity to nonextensibility guided the design of inheritance-calculus.
+The emergent phenomena exhibit an expressive asymmetry between inheritance-calculus and the $\lambda$-calculus.
 Inheritance is not a mechanism for organizing functions; functions are syntactic sugar for inheritance.
@@
 Theorem A.2 (Forward macro-expressibility).
-The translation $\mathcal{T}$ (Section 4) is a macro-expressible embedding of the $\lambda$-calculus into overlay-calculus.
+The translation $\mathcal{T}$ (Section 4) is a macro-expressible embedding of the $\lambda$-calculus into inheritance-calculus.
 Proof.
-Each of the five rules of $\mathcal{T}$ replaces one $\lambda$-calculus construct with an overlay-calculus expression that depends only on the construct’s immediate subexpressions.
+Each of the five rules of $\mathcal{T}$ replaces one $\lambda$-calculus construct with an inheritance-calculus expression that depends only on the construct’s immediate subexpressions.
 No rule inspects the surrounding context.
 ∎
-Theorem A.3 (Non-expressibility of overlay).
-There is no macro-expressible translation from overlay-calculus to the $\lambda$-calculus.
+Theorem A.3 (Non-expressibility of inheritance-calculus).
+There is no macro-expressible translation from inheritance-calculus to the $\lambda$-calculus.
 Proof.
-Let $\mathscr{L}_{1}$ be the overlay-calculus and let $\mathscr{L}_{0}$ be the sublanguage consisting of overlay trees in the image of the translation $\mathcal{T}$ (Section 4), so that $\mathscr{L}_{0}\subset\mathscr{L}_{1}$.
+Let $\mathscr{L}_{1}$ be the inheritance-calculus and let $\mathscr{L}_{0}$ be the sublanguage consisting of mixin trees in the image of the translation $\mathcal{T}$ (Section 4), so that $\mathscr{L}_{0}\subset\mathscr{L}_{1}$.
 Define the operational equivalence $\cong_{0}$ on $\mathscr{L}_{0}$ via Böhm trees: $\mathcal{T}(M)\cong_{0}\mathcal{T}(N)$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$.
-Define $\cong_{1}$ on $\mathscr{L}_{1}$ as overlay-contextual equivalence: $S\cong_{1}S^{\prime}$ iff for every overlay context $C[\cdot]$, $C[S]{\Downarrow}\Leftrightarrow C[S^{\prime}]{\Downarrow}$.
+Define $\cong_{1}$ on $\mathscr{L}_{1}$ as inheritance-contextual equivalence: $S\cong_{1}S^{\prime}$ iff for every inheritance-calculus context $C[\cdot]$, $C[S]{\Downarrow}\Leftrightarrow C[S^{\prime}]{\Downarrow}$.
 By the Böhm Tree Correspondence (Theorems 4.2 and 4.3), $\mathscr{L}_{1}$ conservatively extends $\mathscr{L}_{0}$: for closed $\lambda$-terms $M$, $\mathcal{T}(M){\Downarrow}$ if and only if $M$ has a head normal form.
@@
 Here $\uparrow^{\!1}.\mathrm{argument}$ in $x_{3}$ reaches the outer $\lambda a$ scope (one $\lambda$ up from $\lambda b$), and $\uparrow^{\!0}.\mathrm{argument}$ reaches $\lambda b$ itself.
-In overlay-calculus, applying a function record to an argument is: inherit the function, override its $\mathrm{argument}$ slot, and project $\mathrm{result}$.
-Define overlay trees $e_{1}$ and $e_{2}$ that apply $\mathrm{eq}$ to its two arguments via this mechanism:
+In inheritance-calculus, applying a function record to an argument is: inherit the function, override its $\mathrm{argument}$ slot, and project $\mathrm{result}$.
+Define mixin trees $e_{1}$ and $e_{2}$ that apply $\mathrm{eq}$ to its two arguments via this mechanism:
 $e_{1}\;=\;\{$ $\mathrm{eqFalse}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{false}\},$
@@
 […]
-$C={}$ $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument},\;\mathrm{secondOperand}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}\}\}$
+$C={}$ $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{$
+$\qquad\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument},\;\mathrm{secondOperand}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}\}\}$
 The de Bruijn indices count scope levels from $\mathrm{repr}$: $\uparrow^{\!0}$ is $\mathrm{repr}$ itself, $\uparrow^{\!1}$ is the $\lambda b$ scope, and $\uparrow^{\!2}$ is the $\lambda a$ scope.
@@
 Since $e_{1}$ and $e_{2}$ inherit $\mathrm{eq}$, they also inherit the newly added $\mathrm{repr}$: $e_{1}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{false}$ and $e_{2}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{true}$.
-These produce structurally distinct overlay trees (one converges to an abstraction selecting its first argument, the other to an abstraction selecting its second), so $e_{1}\not\cong_{1}e_{2}$.
+These produce structurally distinct mixin trees (one converges to an abstraction selecting its first argument, the other to an abstraction selecting its second), so $e_{1}\not\cong_{1}e_{2}$.
 Since $e_{1}\cong_{0}e_{2}$ but $e_{1}\not\cong_{1}e_{2}$, we have ${\cong_{0}}\neq{(\cong_{1}|_{\mathscr{L}_{0}})}$.
-By Theorem 3.14(i) of Felleisen (Felleisen, 1991), $\mathscr{L}_{0}$ cannot macro-express the overlay facilities of $\mathscr{L}_{1}$.
+By Theorem 3.14(i) of Felleisen (Felleisen, 1991), $\mathscr{L}_{0}$ cannot macro-express the inheritance facilities of $\mathscr{L}_{1}$.
 ∎
 Corollary A.4.
-Overlay-calculus is strictly more expressive than the $\lambda$-calculus in the sense of Felleisen (Felleisen, 1991): the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
+Inheritance-calculus is strictly more expressive than the $\lambda$-calculus in the sense of Felleisen (Felleisen, 1991): the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
 Appendix B Multi-Path Self-Reference in Other Systems
-This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-path self-reference pattern that overlay-calculus handles naturally.
+This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-path self-reference pattern that inheritance-calculus handles naturally.
 Scala 3
@@
 object Object2 extends MyOuter
-object MyObjectA extends Object1.MyInner
+object HasMultipleOuters extends Object1.MyInner
 with Object2.MyInner
-Scala 3 rejects MyObjectA with the error:
+Scala 3 rejects HasMultipleOuters with the error:
 trait MyInner is extended twice
-object MyObjectA cannot be instantiated since
+object HasMultipleOuters cannot be instantiated since
 it has conflicting base types
@@
 Their intersection requires outer to simultaneously return Object1 and Object2, but DOT’s self variable is single-valued, making this intersection unrealizable.
-The equivalent overlay-calculus definition is:
+The equivalent inheritance-calculus definition is:
 $\{\quad\mathrm{MyOuter}$ $\mapsto\{\mathrm{MyInner}\mapsto\{\mathrm{outer}\mapsto\mathrm{MyOuter}.\mathbf{this}\}\},$
@@
 $\mathrm{Object2}$ $\mapsto\{\mathrm{MyOuter}\},$
-$\mathrm{MyObjectA}$ $\mapsto\{\mathrm{Object1}.\mathrm{MyInner},\;\mathrm{Object2}.\mathrm{MyInner}\}\quad\}$
-This is well-defined in overlay-calculus.
+$\mathrm{HasMultipleOuters}$ $\mapsto\{\mathrm{Object1}.\mathrm{MyInner},\;\mathrm{Object2}.\mathrm{MyInner}\}\quad\}$
+This is well-defined in inheritance-calculus.
 The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$.
-When $\mathrm{MyObjectA}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{MyObjectA})$ for overlay paths matching $\mathrm{MyInner}$’s definition site.
+When $\mathrm{HasMultipleOuters}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{HasMultipleOuters})$ for override paths matching $\mathrm{MyInner}$’s definition site.
 It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both.
-Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{MyObjectA}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken.
-Since inheritance is idempotent, the two routes contribute the same properties and no disambiguation is needed.
+Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{HasMultipleOuters}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken.
+Since inheritance is idempotent, the two routes contribute the same properties in this semantics, so no route-priority disambiguation rule is introduced.
 NixOS Module System
@@
 modules = [ toplevel.config.MyOuter ];
-        # object MyObjectA extends
+# object HasMultipleOuters extends
 # Object1.MyInner with Object2.MyInner
-        options.MyObjectA = lib.mkOption {
+options.HasMultipleOuters = lib.mkOption {
 default = { };
@@
 in
-  builtins.attrNames result.config.MyObjectA.outer
+builtins.attrNames result.config.HasMultipleOuters.outer
 The NixOS module system rejects this with:
-error: The option ‘MyObjectA.outer’
+error: The option HasMultipleOuters.outer
 in ‘<unknown-file>’
@@
 The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently.
-In overlay-calculus, the same pattern is well-defined because $\mathtt{overlays}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
+In inheritance-calculus, the same pattern is well-defined because $\mathtt{overrides}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
 Appendix C Well-Definedness of the Semantic Functions
@@
 $D_{\mathtt{properties}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{L})$ (sets of labels)
-$D_{\mathtt{supers}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P}\times\mathcal{P})$ (sets of site–overlay pairs)
-$D_{\mathtt{overlays}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
+$D_{\mathtt{supers}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P}\times\mathcal{P})$ (sets of site–override pairs)
+$D_{\mathtt{overrides}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
 $D_{\mathtt{bases}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
@@
 Each $D_{i}$ is a complete lattice under the pointwise subset ordering: $f\sqsubseteq g$ iff $f(x)\subseteq g(x)$ for all $x$.
-The product $D=D_{\mathtt{properties}}\times D_{\mathtt{supers}}\times D_{\mathtt{overlays}}\times D_{\mathtt{bases}}\times D_{\mathtt{resolve}}\times D_{\mathtt{this}}$ is a complete lattice under the componentwise ordering, with bottom element $\bot=(\lambda x.\,\varnothing,\ldots,\lambda x.\,\varnothing)$.
+The product $D=D_{\mathtt{properties}}\times D_{\mathtt{supers}}\times D_{\mathtt{overrides}}\times D_{\mathtt{bases}}\times D_{\mathtt{resolve}}\times D_{\mathtt{this}}$ is a complete lattice under the componentwise ordering, with bottom element $\bot=(\lambda x.\,\varnothing,\ldots,\lambda x.\,\varnothing)$.
 C.2. The Operator
-Equations (1)–(6) define a function $F:D\to D$ that takes a tuple of candidate interpretations $(\mathtt{properties},\mathtt{supers},\mathtt{overlays},\mathtt{bases},\mathtt{resolve},\mathtt{this})$ and produces a new tuple by evaluating the right-hand side of each equation.
+Equations (1)–(6) define a function $F:D\to D$ that takes a tuple of candidate interpretations $(\mathtt{properties},\mathtt{supers},\mathtt{overrides},\mathtt{bases},\mathtt{resolve},\mathtt{this})$ and produces a new tuple by evaluating the right-hand side of each equation.
 The transitive closure $\mathtt{bases}^{*}$ appearing in equation (2) is derived from $\mathtt{bases}$: for any fixed $\mathtt{bases}$, the set $\mathtt{bases}^{*}(p)$ is defined as the least fixed point of the monotone operator $X\mapsto\{p\}\cup\bigcup_{q\in X}\mathtt{bases}(q)$ on $\mathcal{P}(\mathcal{P})$.
@@
 Equation (1): $\mathtt{properties}$
-$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\,\bigr\}.$
+$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{override}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{override}})\,\bigr\}.$
 The function $\mathtt{supers}$ appears as the set being iterated over ($\in\mathtt{supers}(p)$).
@@
 Equation (2): $\mathtt{supers}$
-$\mathtt{supers}(p)=\bigl\{\,(\mathtt{init}(p_{\mathrm{base}}),p_{\mathrm{overlay}})\bigm|p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{overlay}}\in\mathtt{overlays}(p_{\mathrm{base}})\,\bigr\}$.
-Both $\mathtt{bases}^{*}$ and $\mathtt{overlays}$ appear as sets being iterated over.
+$\mathtt{supers}(p)=\bigl\{\,(\mathtt{init}(p_{\mathrm{base}}),p_{\mathrm{override}})\bigm|p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{override}}\in\mathtt{overrides}(p_{\mathrm{base}})\,\bigr\}$.
+Both $\mathtt{bases}^{*}$ and $\mathtt{overrides}$ appear as sets being iterated over.
 As noted above, $\mathtt{bases}^{*}$ is monotone in $\mathtt{bases}$.
-Enlarging either $\mathtt{bases}$ or $\mathtt{overlays}$ can only add elements to the result.
-Positive in $\mathtt{bases}$ and $\mathtt{overlays}$.
-Equation (3): $\mathtt{overlays}$
+Enlarging either $\mathtt{bases}$ or $\mathtt{overrides}$ can only add elements to the result.
+Positive in $\mathtt{bases}$ and $\mathtt{overrides}$.
+Equation (3): $\mathtt{overrides}$
 For $p\neq()$:
-$\mathtt{overlays}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|{}\\ (\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
+$\mathtt{overrides}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|{}\\ (\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
 The function $\mathtt{supers}$ appears as the set being iterated over.
@@
 Equation (4): $\mathtt{bases}$
-$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{overlay}}\in\mathtt{overlays}(p),\\ (n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{overlay}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
-Both $\mathtt{overlays}$ and $\mathtt{resolve}$ appear as sets being iterated over.
+$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{override}}\in\mathtt{overrides}(p),\\ (n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{override}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
+Both $\mathtt{overrides}$ and $\mathtt{resolve}$ appear as sets being iterated over.
 The function $\mathtt{inherits}$ is an AST primitive.
-Enlarging either $\mathtt{overlays}$ or $\mathtt{resolve}$ can only add targets to the result.
-Positive in $\mathtt{overlays}$ and $\mathtt{resolve}$.
+Enlarging either $\mathtt{overrides}$ or $\mathtt{resolve}$ can only add targets to the result.
+Positive in $\mathtt{overrides}$ and $\mathtt{resolve}$.
 Equation (5): $\mathtt{resolve}$
@@
 Equation (6): $\mathtt{this}$
-For $n>0$: $\mathtt{this}(S,p_{\mathrm{def}},n)=\mathtt{this}(S^{\prime},\mathtt{init}(p_{\mathrm{def}}),n-1)$ where $S^{\prime}=\bigl\{\,p_{\mathrm{site}}\bigm|p_{\mathrm{current}}\in S,\;(p_{\mathrm{site}},p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\,\bigr\}$.
+For $n>0$: $\mathtt{this}(S,p_{\mathrm{def}},n)=\mathtt{this}(S^{\prime},\mathtt{init}(p_{\mathrm{def}}),n-1)$ where $S^{\prime}=\bigl\{\,p_{\mathrm{site}}\bigm|p_{\mathrm{current}}\in S,\;(p_{\mathrm{site}},p_{\mathrm{override}})\in\mathtt{supers}(p_{\mathrm{current}}),\;p_{\mathrm{override}}=p_{\mathrm{def}}\,\bigr\}$.
 The function $\mathtt{supers}$ appears as the set being iterated over.
-The filtering condition $p_{\mathrm{overlay}}=p_{\mathrm{def}}$ is an equality test on path data, not involving any mutually defined function.
+The filtering condition $p_{\mathrm{override}}=p_{\mathrm{def}}$ is an equality test on path data, not involving any mutually defined function.
 Enlarging $\mathtt{supers}$ can only enlarge $S^{\prime}$, and by induction on $n$ (the base case $n=0$ returns $S$ unchanged), the recursive call is monotone in $S^{\prime}$.
@@
 Theorem C.2.
-The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overlays}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined as the least fixed point of $F$.
+The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined as the least fixed point of $F$.
 Proof.
@@
 D.2. Path Encoding
-We define a correspondence between positions in the Böhm tree and paths in the overlay tree.
-When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting overlay tree has a specific shape:
+We define a correspondence between positions in the Böhm tree and paths in the mixin tree.
+When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting mixin tree has a specific shape:
   • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$ (the abstraction shape).
@@
 In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body (peeling off one abstraction layer) and entering the $i$-th argument $M_{i}$ of the head variable.
-In the overlay tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
-Rather than comparing absolute paths across different overlay trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection.
-Recall the definition of overlay-convergence (Definition 4.1).
+In the mixin tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
+Rather than comparing absolute paths across different mixin trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection.
+Recall the definition of inheritance-convergence (Definition 4.1).
 Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation.
@@
 D.3. Single-Path Lemma
-The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path overlay inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
+The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path mixin inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
 Lemma D.1 (Single Path).
@@
 This is a record literal with $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$ and no inheritance sources, so $\mathtt{inherits}=\varnothing$.
-At the root path $p$, we have $\mathtt{overlays}(p)=\{p\}$ (a single path) since there are no inheritance sources to introduce additional branches.
+At the root path $p$, we have $\mathtt{overrides}(p)=\{p\}$ (a single path) since there are no inheritance sources to introduce additional branches.
 Inside $\mathcal{T}(M)$, any reference to $x$ is $\uparrow^{\!n}.\mathrm{argument}$ where $n$ is the de Bruijn index pointing to this scope level.
-Since $\mathcal{T}(\lambda x.\,M)$ introduces exactly one scope level, there is exactly one overlay at that level, hence $\mathtt{this}$ finds exactly one matching pair $(p_{\mathrm{site}},\;p_{\mathrm{overlay}})$ in $\mathtt{supers}$.
+Since $\mathcal{T}(\lambda x.\,M)$ introduces exactly one scope level, there is exactly one mixin at that level, hence $\mathtt{this}$ finds exactly one matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ in $\mathtt{supers}$.
 Case $\mathcal{T}(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
-The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level ($\mathtt{inherits}=\varnothing$), so $\mathtt{overlays}(\text{root})=\{\text{root}\}$.
+The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level ($\mathtt{inherits}=\varnothing$), so $\mathtt{overrides}(\text{root})=\{\text{root}\}$.
 Inside the $x$ subtree, the inheritance $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$ has exactly one inheritance source ($\mathcal{T}(V_{1})$), creating exactly one $\mathtt{bases}$ entry.
@@
 Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference.
-In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ overlays the $\mathrm{argument}$ slot with $\mathcal{T}(V)$.
+In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ inherits into the $\mathrm{argument}$ slot, providing $\mathcal{T}(V)$.
 Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma D.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree.
@@
 $\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)\;\Longleftrightarrow\;\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}(M[V/x])$
-That is, following one $\mathrm{result}$ projection in the tail call’s overlay tree yields the same properties as the root of the reduct’s overlay tree.
+That is, following one $\mathrm{result}$ projection in the tail call’s mixin tree yields the same properties as the root of the reduct’s mixin tree.
 Proof.
@@
 D.7. Full Abstraction
-Adequacy relates a single term’s head normal form to its overlay-convergence.
+Adequacy relates a single term’s head normal form to its inheritance-convergence.
 Full abstraction lifts this to an equivalence between terms: the translation $\mathcal{T}$ preserves and reflects observational equivalence.
-Definition D.5 (Overlay-contextual equivalence).
-For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
+Definition D.5 ($\lambda$-contextual equivalence under $\mathcal{T}$).
+For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
 $\mathcal{T}(C[M]){\Downarrow}\;\Longleftrightarrow\;\mathcal{T}(C[N]){\Downarrow}$
-This definition observes only convergence (a single bit), but quantifying over all contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior.
+The subscript $\lambda$ emphasizes that the quantification ranges over $\lambda$-calculus contexts, not over all inheritance-calculus contexts.
+This definition observes only convergence (a single bit), but quantifying over all $\lambda$-contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior within the $\lambda$-definable fragment.
 Proof of Theorem 4.3.
-The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the overlay tree $\mathcal{T}(C[M])$ is determined by the overlay tree $\mathcal{T}(M)$ and the translation of the context.
+The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the mixin tree $\mathcal{T}(C[M])$ is determined by the mixin tree $\mathcal{T}(M)$ and the translation of the context.
 This means $\mathcal{T}$ preserves the context structure needed for the argument.
@@
 By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$.
-Hence $\mathcal{T}(M)\approx\mathcal{T}(N)$.
+Hence $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$.
 ($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$.
@@
 By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$.
-Hence $\mathcal{T}(M)\not\approx\mathcal{T}(N)$.
+Hence $\mathcal{T}(M)\not\approx_{\lambda}\mathcal{T}(N)$.
 ∎
 The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993).
-The overlay tree semantics inherits full abstraction from the $\lambda$-calculus without needing to define or compare absolute paths inside overlay trees.
+The mixin tree semantics inherits full abstraction from the $\lambda$-calculus without needing to define or compare absolute paths inside mixin trees.

### v6

@@
 Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming.
-It trivially embeds the $\lambda$-calculus, although the entire semantics rests solely on naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance does not arise.
-This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence (Barendregt, 1984).
-Inheritance-calculus is distilled from MIXINv2, a practical implementation in which we observed further emergent phenomena: the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); self-reference resolves to multiple targets; and programs are immune to the Expression Problem (Wadler, 1998).
+By unifying classes, methods, objects, and properties under a single record abstraction, the calculus models inheritance simply as set union.
+Consequently, composition is inherently commutative, idempotent, and associative, structurally eliminating the multiple-inheritance linearization problem.
+Its semantics is first-order, denotational, and fixpoint-computable.
+These three properties extend to the lazy $\lambda$-calculus, since Böhm tree equivalence (Barendregt, 1984) is fully abstract for a sublanguage of inheritance-calculus, except that recursive $\lambda$-programs that diverge under $\beta$-reduction may converge under $\mathrm{lfp}(T_{P})$.
+Inheritance-calculus is distilled from MIXINv2, a practical implementation in which the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); self-reference resolves to multiple targets; and programs are immune to the Expression Problem (Wadler, 1998).
 This makes inheritance-calculus strictly more expressive than the $\lambda$-calculus in both common sense and Felleisen’s sense (Felleisen, 1991).
-These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
-declarative programming, inheritance, denotational semantics, self-referential records, expression problem, $\lambda$-calculus, Böhm trees, configuration languages
+declarative programming, inheritance, fixpoint semantics, self-referential records, expression problem, first-order semantics, $\lambda$-calculus, Böhm trees, configuration languages
 This work was conducted independently prior to the author’s employment at Figure AI.
-$c$s: Theory of computation Denotational semanticsccs: Theory of computation Object oriented constructsccs: Theory of computation Functional constructsccs: Theory of computation Program schemesccs: Software and its engineering Formal language definitionsccs: Software and its engineering Extensible languagesccs: Software and its engineering Object oriented languages
+$c$s: Theory of computation Program semanticsccs: Theory of computation Object oriented constructsccs: Theory of computation Functional constructsccs: Theory of computation Program schemesccs: Software and its engineering Formal language definitionsccs: Software and its engineering Extensible languagesccs: Software and its engineering Object oriented languages
 1. Introduction
@@
 Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or merging.
-Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively.
+Among these, the NixOS module system stands out: its recursive attribute set merging with fixpoint semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively.
 Yet no computational theory explains why this mechanism is so expressive.
@@
 The NixOS module system already suggests that such a model exists: its inheritance-based composition over recursive records, without explicit functions, is expressive enough to configure entire operating systems.
-Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures, that is, initial algebras in the sense of universal algebra.
-We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures, where semantics is determined by the observer and a finite prefix suffices for any finite observation.
-This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set.
+Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures—initial algebras in the sense of universal algebra.
+We challenge this assumption.
+Configuration values are better understood as lazily observable, possibly infinite structures, where semantics is determined by the observer and a finite prefix suffices for any finite observation.
+The nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set.
 Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives.
 The reduction produced MIXIN, an early implementation in Nix, and subsequently MIXINv2 (Yang, 2026), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
-In developing real programs in MIXINv2, including boolean logic and natural number arithmetic (Section 5) and a web application (Section Function color blindness: ), we found that the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) is structurally absent: because records are sets, inheritance is inherently commutative, idempotent, and associative (Section 3).
-Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system).
-Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction.
-We also found that boolean logic and arithmetic, implemented as separate MIXINv2 files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism; that Church-encoded values automatically form trie structures; and that operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$, which is the relational semantics of logic programming (van Emden and Kowalski, 1976) (Section 5); and that the same web application MIXINv2 file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015), equivalently CPS-agnostic behavior (Section Function color blindness: ).
+In developing real programs in MIXINv2, including boolean logic, natural number arithmetic, and a web application, the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) did not arise.
+Inheritance is inherently commutative, idempotent, and associative (Section 3).
+Self-reference naturally resolves to multiple objects rather than one.
+Prior systems such as Scala and the NixOS module system assume single-object resolution and reject the multi-path situation.
+Self-referential records require no fixed-point combinator in the object language.
+The mixin tree is the call stack, and qualified this resolution is a first-order set-theoretic query over paths.
+We also found that boolean logic and arithmetic, implemented as separate MIXINv2 files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism.
+Church-encoded values automatically form trie structures, and operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$—the relational semantics of logic programming (van Emden and Kowalski, 1976) (Section 5).
+Furthermore, the same web application MIXINv2 file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015) (Section Function color blindness: ).
 This paper distills these observations into inheritance-calculus, a minimal computational model for declarative programming.
@@
 Since the $\lambda$-calculus embeds into inheritance-calculus, inheritance-calculus is Turing complete.
-The translation reveals an asymmetry: embedding the $\lambda$-calculus into inheritance-calculus requires five local, syntax-directed rules (Section 4), but the reverse is not macro-expressible under the equivalence notions fixed in Section 4.2 and Appendix A— inheritance-calculus enables open extension that changes operational equivalence on $\lambda$-terms, establishing a strict separation in Felleisen’s sense of expressiveness (Felleisen, 1991) (Section 4.2).
+The translation reveals an asymmetry: embedding the $\lambda$-calculus into inheritance-calculus requires five local, syntax-directed rules (Section 4), but the lazy $\lambda$-calculus cannot macro-express the inheritance constructors.
+Inheritance-calculus enables open extension that changes operational equivalence on $\lambda$-terms, establishing that inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991) (Section 4.2).
 Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
 All constructions in this paper have been implemented and tested in MIXINv2, available as open source (Yang, 2026), with a test suite covering every example.
-[^1] MIXINv2 is statically typed and therefore stricter than the untyped inheritance-calculus.
+[^1] MIXINv2 separates symbol resolution (which resolves lexical references and qualified-this references) from runtime mixin evaluation, but symbol resolution is performed just-in-time per symbol rather than ahead of time, so it does not provide static type checking in the traditional sense.
+Nevertheless, MIXINv2’s resolution phase is stricter than the untyped inheritance-calculus.
 All examples in this paper are ported from MIXINv2 and conform to both semantics.
-Elements added to satisfy the type checker are harmless for the algorithms presented here and may aid readability.
+Elements added to satisfy the resolver are harmless for the algorithms presented here and may aid readability.
 All tests are mechanically verified against expected output snapshots.
 Contributions
-  • We present inheritance-calculus, a minimal computational model for declarative programming that contains only three constructs (record, definition, inheritance) and no functions or scalar types.
-  • Inheritance operates at the tree level: a record is a set of elements, both definitions and inheritance sources, so inheriting from two mixins that share a label recursively merges their subtrees (deep merge), rather than overriding at the method level as in prior mixin and trait calculi.
-Since records are sets, inheritance is inherently commutative, idempotent, and associative.
-The linearization problem inherent in prior mixin and trait calculi does not arise.
-  • The semantics is purely observational: there are no reduction rules, no evaluation order, and no notion of a result that a term reduces to.
-The observer drives computation by querying paths in a lazily constructed tree.
-  • Self-referential records require no fixed-point combinator.
-The $\lambda$-calculus translation maps each application to an inheritance; the resulting inheritance tree is the call stack.
-The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree: a first-order set-theoretic query, not a fixed-point construction.
-  • Self-reference resolves to multiple objects, not one.
-Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala and the NixOS module system, Appendix B).
-Since inheritance is idempotent, all objects contribute equally; this semantics does not introduce a route-priority disambiguation rule.
-  • The $\lambda$-calculus (in A-normal form) translates into inheritance-calculus in five rules (Section 4).
-The translation is fully abstract with respect to Böhm tree equivalence under $\lambda$-definable contexts (Section 4.1), yielding a semantics of the lazy $\lambda$-calculus that requires no domain theory.
-  • We observe further properties: extensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section Function color blindness: ).
-  • Open extension via inheritance changes the operational equivalence of $\lambda$-terms, establishing via Felleisen’s Theorem 3.14 (Felleisen, 1991) that the $\lambda$-calculus cannot macro-express inheritance-calculus (Section 4.2).
+  • We present inheritance-calculus, a minimal computational model for declarative programming that contains three constructs (record, definition, inheritance).
+  • The mixin is a single abstraction that unifies class, method, object, and property.
+Method-level overriding is subsumed by deep merge of mixin trees; inheritance is commutative, idempotent, and associative, and the linearization problem in OOP does not arise.
+  • The semantics is observational: the observer drives computation by querying paths in a lazily constructed tree.
+The six semantic equations are first-order [^2] First-order in three senses simultaneously: the semantic domain contains only paths (finite sequences of labels) and sets of paths, with no function spaces, closures, or strategies; the six equations are first-order functions whose inputs and outputs are paths and sets of paths, never functions; and the judgments are first-order inductive definitions (Aczel, 1977) over paths and labels, where quantifiers range over paths and labels, never over functions or predicates., denotational, and fixpoint-computable.
+  • The mixin tree is the call stack. Self-reference resolves to multiple objects. Since inheritance is idempotent, all objects contribute equally.
+  • The $\lambda$-calculus macro-expresses into inheritance-calculus in five local rules (Section 4).
+  • The translation gives the lazy $\lambda$-calculus a first-order semantics (Section 4.1): the prior higher-order semantic frameworks are shown to be faithful to a sublanguage of inheritance-calculus’s first-order equations.
+Appendix E reformulates this semantics directly on ANF.
+As a corollary, some $\lambda$-terms that diverge in the $\lambda$-calculus converge under $\mathrm{lfp}(T_{P})$ after translation (Section 4; Remark B.5).
+  • Inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991) (Section 4.2).
+  • We observe further properties: immunity to nonextensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section Function color blindness: ).
 Section 5 demonstrates these properties through a case study of Nat arithmetic.
 2. Syntax
-Let $e$ denote an expression. Let $\ell$ denote a label (property name). Let $k$ denote a non-negative integer (path length).
+Let $e$ denote an expression. Let $\ell$ denote a label (property name). Let $k$ denote a nonnegative integer (path length).
 $e\quad::=\quad$ $\{c_{1},\;\ldots,\;c_{m}\}$ $\text{(}m\geq 0\text{)}$
@@
 Because expressions are sets, element order is irrelevant and duplicates have no effect; inheritance is inherently commutative, idempotent, and associative.
-The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding.
-There is no variable binding in inheritance-calculus.
+The $\mapsto$ in $\ell\mapsto e$ defines a property.
 MIXINv2 also provides scalar types through a foreign-function interface (FFI); inheritance-calculus is what remains after removing the FFI.
@@
 3. Mixin Trees
-Given an AST, what are the properties visible at each position in the tree? This section answers the question by defining a collection of mutually recursive functions over paths.
-A mixin tree has a single abstraction, the mixin: every mixin can serve as field, method, class, and object at the same time in object-oriented terms.
-Unlike the $\lambda$-calculus, which requires $\beta$-reduction to specify how terms compute, inheritance-calculus has no reduction steps: no rewriting of terms, no evaluation order, and no notion of a result that a term reduces to.
-Instead, the semantics is purely observational: one queries whether a label $\ell$ is a property at a path $p$ in the inheritance tree.
-The mutually recursive functions defined below, namely $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, and $\mathtt{this}$, constitute the entire semantics.
-Each function has an analogue in object-oriented languages: $\mathtt{Path}$ identifies a position in the tree, $\mathtt{init}$ and $\mathtt{last}$ navigate paths, $\mathtt{defines}$ extracts the locally defined members from the AST, $\mathtt{inherits}$ extracts the inheritance declarations from the AST, $\mathtt{properties}$ computes all visible members at a path including both locally defined and inherited ones, $\mathtt{supers}$ computes the transitive inheritance closure, $\mathtt{bases}$ collects the direct base classes, and $\mathtt{overrides}$ generalizes method override to the tree setting, collecting the identity-sharing paths that arise when inheritance introduces multiple definitions of the same label at the same scope level.
-Two functions, the reference resolution function $\mathtt{resolve}$ and the qualified this resolution function $\mathtt{this}$, have object-oriented analogues whose semantics differ significantly; they are discussed in detail below.
-All functions are pure and may be cached (memoised) without changing semantics.
+In object-oriented programming, a mixin (Bracha and Cook, 1990) is a fragment of behavior that can be composed onto a class via inheritance.
+Inheritance-calculus has a single abstraction, the deep-mergeable mixin, [^3] Where no ambiguity arises, we write simply “mixin.” The distinction from the Bracha–Cook mixin (Bracha and Cook, 1990) is detailed in Section 7.
+which can simultaneously serve as field, method, class, and object, because there are no opaque methods: every value is a transparent record, and inheritance always performs a deep merge of subtrees.
+We call the resulting structure a mixin tree.
+Its semantics is purely observational: given an inheritance-calculus program, one queries whether a label $\ell$ is a property at a given position $p$.
+This section defines the semantics via six mutually recursive first-order functions on paths.
+The six functions map object-oriented concepts to set-theoretic operations: $\mathtt{properties}$ and $\mathtt{supers}$ generalize member lookup and inheritance chains, $\mathtt{overrides}$ and $\mathtt{bases}$ capture method override and base classes, and $\mathtt{resolve}$ and $\mathtt{this}$ handle name resolution and self-reference.
+3.1. Definitions
 Path
 A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ell_{2},\ldots,\ell_{n})$ that identifies a position in the tree.
-[^2] Since paths are referentially transparent, an implementation may intern paths so that structural equality reduces to pointer equality.
 We write $p$ for a path.
@@
 Given a path $p$ and a label $\ell$, $p\mathbin{\triangleright}\ell$ is the sequence $p$ extended with $\ell$.
-For a non-root path, the parent path $\mathtt{init}(p)$ is $p$ with its last element removed, and the final label $\mathtt{last}(p)$ is the last element of $p$.
+For a nonroot path, the parent path $\mathtt{init}(p)$ is $p$ with its last element removed, and the final label $\mathtt{last}(p)$ is the last element of $p$.
 For convenience, we sometimes write $\ell$ for $\mathtt{last}(p)$ when the path is clear from context.
@@
 A reference contributes $(n,\;\ell_{\mathrm{down},*})$ to $\mathtt{inherits}(p)$.
-Since records are sets, the order of elements is irrelevant and duplicate elements have no effect.
 Both $\mathtt{defines}$ and $\mathtt{inherits}$ are pure data; they are not functions of runtime state.
@@
 Overrides
-Because inheritance-calculus has no separate notion of “method,” a mixin simultaneously plays the role of a class and the role of a method in object-oriented languages.
-When a mixin is viewed as a method, a child scope should be able to find it through the parent’s label even if the child does not redefine that label—this is method override.
+Since a mixin can play the role of a method, a child scope should be able to find it through the parent’s label even if the child does not redefine that label; this is method override.
 Concretely, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overrides}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge.
@@
 (6) $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{override}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{override}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
-At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a non-negative integer, the recursion terminates.
+At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one.
+Since $n$ is a nonnegative integer, the recursion terminates.
+[^4] If no matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ exists at some step, the frontier becomes the empty set, and $\mathtt{resolve}$ returns $\varnothing$: the reference has no target, and no properties are inherited.
+This is the most self-consistent semantics for the untyped calculus.
+MIXINv2 (Yang, 2026) is stricter: its resolver rejects references whose frontier becomes empty, reporting a resolution error at compile time.
 This completes the chain of definitions needed to compute $\mathtt{properties}(p)$.
-Observational semantics without reduction
-The definitions above constitute the complete semantics of inheritance-calculus.
-There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments.
-In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur.
-In inheritance-calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything: it is the value.
-The functions $\mathtt{properties}$, $\mathtt{supers}$, etc.
-determine what is observable at each path, and that is all.
-Computation arises not from rewriting terms but from querying an inductively deeper path in a lazily constructed tree: the observer drives the computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand.
-Formally, the mutually recursive equations (1)–(6) define set membership inductively: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these equations.
-The corresponding operator on the product lattice of path sets is monotone, and its least fixed point exists by the Knaster–Tarski theorem (Tarski, 1955).
-The detailed well-definedness argument, which identifies the lattice, verifies positivity of each equation, and handles the transitive closure $\mathtt{bases}^{*}$, is given in Appendix C.
-When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point, analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness.
-Identity is determined by observable behavior rather than by internal structure.
+3.2. Recursive Evaluation
+The six equations (1)–(6) define $\mathtt{properties}(p)$ by mutual recursion: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these functions.
+The functions are the interpreter—the observer drives the computation by choosing which path to inspect, and each equation unfolds on demand.
+Because the functions are pure and order-independent, the reflexive-transitive closure $\mathtt{bases}^{*}$ in (2) may be explored breadth-first or depth-first without affecting the result; paths may be interned so that structural equality reduces to pointer equality; and the six functions form a memoizable dynamic-programming recurrence keyed by paths.
+Operationally, evaluation uses tabled evaluation (Tamaki and Sato, 1986; Chen and Warren, 1996): each query is memoised on first encounter, and re-entrant calls (cycles in the dependency graph) return the current accumulator rather than diverging.
+Denotationally, the six equations define a monotone immediate consequence operator $T_{P}$ (van Emden and Kowalski, 1976) whose right-hand sides use only $\cup$, $\in$, equality, and positive set comprehension, with no negation, set difference, or complement.
+By the Knaster–Tarski theorem (Tarski, 1955), $T_{P}$ has a unique least fixed point $\mathrm{lfp}(T_{P})$, which serves as the denotational semantics.
+Appendix A proves that tabled evaluation never returns a wrong answer and analyses termination.
 4. Embedding the $\lambda$-Calculus
 The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to inheritance-calculus.
-In ANF, every intermediate result is bound to a name and each expression contains at most one application:
+In ANF, every intermediate result is bound to a name and each expression contains at most one application.
+We extend the standard ANF grammar with a value binding $\mathbf{let}\;x=V\;\mathbf{in}\;M$, which gives a name to a value that is not an application result: [^5] Standard ANF allows values inline as operands; the value-binding production names them so that every subexpression of the translated program is accessible via a lexical reference (Appendix C).
 […]
-Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results, and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus.
+Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results (and, optionally, naming values), and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus.
 The ANF structure matches inheritance-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
 Let $\mathcal{T}$ denote the translation function. Each $\lambda$-abstraction introduces one scope level in the translated mixin tree.
-$x$ ($\lambda$-bound, index $n$)   $\longrightarrow$ $\uparrow^{\!n}.\mathrm{argument}$
+$x$ ($\lambda$-bound, index $n^{\prime}$)    $\longrightarrow$ $\uparrow^{\!n^{\prime}}.\mathrm{argument}$
 $x$ (let-bound)    $\longrightarrow$ $x.\mathrm{result}$
 $\lambda x.\,M$    $\longrightarrow$ $\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
-$\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$   $\longrightarrow$ $\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
-$V_{1}\;V_{2}$ (tail call)   $\longrightarrow$ $\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
+$\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$    $\longrightarrow$ $\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},$
+   $\longrightarrow$ $\mathrm{result}\mapsto\mathcal{T}(M)\}$
+$\mathbf{let}\;x=V\;\mathbf{in}\;M$    $\longrightarrow$ $\{x\mapsto\{\mathrm{result}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
+$V_{1}\;V_{2}$ (tail call)    $\longrightarrow$ $\{\mathrm{tailCall}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},$
+   $\longrightarrow$ $\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\}$
 […]
 An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$; we call this the abstraction shape.
-In the $\mathbf{let}$-binding rule, the name $x$ binds to the inherited record $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$, and $\mathcal{T}(M)$ may reference $x.\mathrm{result}$ to obtain the value of the application.
-In the tail-call rule, the fresh label $\mathrm{tail}$ serves the same purpose, and $\mathrm{result}$ projects the answer.
-In both cases the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
-The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to inheritance-calculus: every ANF term has an image, and the construction is compositional.
+In the application $\mathbf{let}$-binding rule, the name $x$ binds to the inherited record $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$, and $\mathcal{T}(M)$ may reference $x.\mathrm{result}$ to obtain the value of the application.
+In the value $\mathbf{let}$-binding rule, $x$ wraps $\mathcal{T}(V)$ under a $\mathrm{result}$ child so that the uniform accessor $x.\mathrm{result}$ yields $\mathcal{T}(V)$, the translated value itself.
+This extra wrapper is semantically inert but proof-relevant: it lets the same $\mathrm{result}$-following argument used for applications extend uniformly to value $\mathbf{let}$-bindings in the adequacy proof below.
+In the tail-call rule, the fresh label $\mathrm{tailCall}$ serves the same purpose, and $\mathrm{result}$ projects the answer.
+In the application and tail-call cases, the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
+The six rules above are a complete translation from the $\lambda$-calculus (via ANF) to inheritance-calculus: every ANF term has an image, and the construction is compositional.
 Since the $\lambda$-calculus is Turing complete, so is inheritance-calculus.
-An elementary semantics of the $\lambda$-calculus
-Composing the translation $\mathcal{T}$ with the semantic functions of Section 3 induces a semantics on the $\lambda$-calculus itself.
-Given a $\lambda$-term $M$, one translates it to $\mathcal{T}(M)$ and then determines convergence behavior by querying $\ell\in\mathtt{properties}(p)$ via equations (1)–(6), following the $\mathrm{result}$ chain until the abstraction shape appears or the fixed point yields the empty set.
-This induced semantics is more elementary than traditional approaches in the mathematical machinery it requires:
-No function spaces:
-Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices.
-Here, the semantic domain is a product of powerset lattices, with no function spaces, no continuity conditions, and no inverse limits.
-The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix C), which requires only set-theoretic monotonicity, which is strictly weaker than Scott continuity.
-No substitution:
-Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi.
-Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree; no substitution occurs at any point.
-No reduction rules:
-There are no $\beta$-reduction steps, no evaluation contexts, and no choice between call-by-name and call-by-value (Plotkin, 1975).
-The semantics is observational: one queries path membership, and the mutually recursive equations unfold on demand.
-The key step is that function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function; no function spaces are required.
-Section 4.1 shows that the convergence behavior of the translated mixin tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate, meaning a term converges iff it has a head normal form, and fully abstract with respect to $\lambda$-definable contexts, meaning two $\lambda$-terms are equivalent under the translation iff they have the same Böhm tree.
-Combining these two facts, elementary machinery and full abstraction, yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics (w.r.t.
-$\lambda$-definable contexts) that lives entirely in naive set theory.
-The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem.
-No continuous lattices, no inverse limits, no Scott topology, no game-theoretic machinery.
-Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Abramsky and McCusker’s game semantics (Abramsky and McCusker, 1995) achieves full abstraction but at the cost of even greater mathematical complexity.
-The mixin tree semantics achieves full abstraction with strictly simpler foundations than either.
 4.1. Böhm Tree Correspondence
-The translation $\mathcal{T}$ induces a semantics on the $\lambda$-calculus via the mixin tree equations (1)–(6).
+The translation $\mathcal{T}$ induces an elementary semantics on the $\lambda$-calculus via the mixin tree equations (1)–(6).
 We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
@@
 $\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{properties}\!\bigl(\,\underbrace{\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
-in the least fixed point of Section 3.
+in the semantics of Section 3.
 Theorem 4.2 (Adequacy).
 A closed ANF $\lambda$-term $M$ has a head normal form if and only if $\mathcal{T}(M){\Downarrow}$.
-Theorem 4.3 (Full abstraction w.r.t. $\lambda$-definable contexts).
+Theorem 4.3 (First-order semantics of the lazy $\lambda$-calculus).
 For closed $\lambda$-terms $M$ and $N$, define $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ iff for every closing $\lambda$-calculus context $C[\cdot]$, $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$.
@@
 Here $\approx_{\lambda}$ quantifies over $\lambda$-calculus contexts only; inheritance-calculus contexts are strictly more discriminating (Section 4.2).
-Together, Adequacy and Full Abstraction establish that the translation $\mathcal{T}$ is a fully faithful embedding of the lazy $\lambda$-calculus into inheritance-calculus with respect to Böhm tree equivalence (Barendregt, 1984), using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory.
-This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract, as it identifies terms that observational equivalence distinguishes (Plotkin, 1975).
-The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix D.
-The relationship between $\approx_{\lambda}$ and full inheritance-contextual equivalence is clarified by the expressive asymmetry of the next section: inheritance-calculus contexts can separate $\lambda$-terms that are Böhm-tree equivalent, so ${\approx_{\lambda}}\subsetneq{(\approx_{\mathrm{inh}}|_{\mathrm{im}(\mathcal{T})})}$.
-This gap is precisely the source of inheritance-calculus’s additional expressiveness (Theorem A.3).
+The theorem establishes that the lazy $\lambda$-calculus has a first-order semantics: the translation $\mathcal{T}$ embeds the $\lambda$-calculus into a sublanguage of inheritance-calculus, using only recursive evaluation of set-comprehension equations over paths and sets of paths.
+The prior higher-order semantic frameworks of (Scott, 1972; Wadsworth, 1976; Abramsky et al., 2000; Hyland and Ong, 2000; Barendregt, 1984) are thereby shown to be faithful to this first-order semantics: they agree with the equations on Böhm tree equivalence (Barendregt, 1984).
+[^7] Composing $\mathcal{T}$ with the equations of Section 3 yields a first-order semantics of the $\lambda$-calculus in which function application is record inheritance and convergence is determined by path queries.
+Appendix E reformulates this semantics directly on ANF, without passing through inheritance-calculus.
+The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix B.
+The relationship between $\approx_{\lambda}$ and full inheritance-contextual equivalence is clarified by the expressive asymmetry of the next section: contexts in the language universe $\mathscr{U}$ (Definition C.3) that mix $\lambda$-calculus and inheritance-calculus constructors can separate $\lambda$-terms that are Böhm-tree equivalent, so ${\approx_{\lambda}}\subsetneq{(\approx_{\mathscr{U}}|_{L_{0}})}$ where $L_{0}$ is the lazy $\lambda$-calculus.
+This gap is precisely the source of inheritance-calculus’s additional expressiveness (Theorem C.7).
 4.2. Expressive Asymmetry
-Two $\lambda$-terms that are Böhm-tree equivalent—and therefore indistinguishable in the $\lambda$-calculus—can be separated in inheritance-calculus by inheriting new observable projections onto an existing definition via open extension.
-By Felleisen’s expressiveness criterion (Felleisen, 1991), inheritance-calculus is strictly more expressive than the $\lambda$-calculus: the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
-The formal definitions, constructions, and proofs are given in Appendix A; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
+Two $\lambda$-terms that are Böhm-tree equivalent, and therefore indistinguishable in the $\lambda$-calculus, can be separated by a context that mixes $\lambda$-calculus and inheritance-calculus constructors, inheriting new observable projections onto an existing definition via open extension.
+By Felleisen’s expressiveness criterion (Felleisen, 1991), inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus (Theorem C.8): the $\lambda$-calculus can be macro-expressed in inheritance-calculus via ANF conversion and $\mathcal{T}$ (Lemma C.5, Theorem C.6), but the lazy $\lambda$-calculus cannot macro-express inheritance (Theorem C.7).
+The formal definitions, constructions, and proofs are given in Appendix C; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
 The underlying mechanism is open extension: mixin definitions compose via open recursion, so existing definitions gain new observable properties without modification.
-In the $\lambda$-calculus, a function’s result is closed—there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
+In the $\lambda$-calculus, a function’s result is closed: there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
 5. Case Study: Nat Arithmetic
@@
 This section traces the unary Nat case in detail.
-The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names are constructors and operations; lowerCamelCase names are fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action.
+The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names serve as data constructors and operations; lowerCamelCase names serve as fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action.
 Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
 NatData
-Following the Church-encoding convention (Yang, 2026), we define a factory containing the abstract product type and its two constructors.
+Following the Church-encoding convention (Yang, 2026), [^8] The data representation (Zero/Successor) is a Peano encoding, an initial algebra of the functor $F(X)=1+X$.
+The visitor pattern (NatVisitor, below) adds the Church-encoding behavior: a fold (catamorphism) over this algebra.
+These two concerns are defined in separate files and composed via inheritance, which is itself an instance of the Expression Problem.
+we define a factory containing the abstract product type and two labels that serve as data constructors.
 A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field.
@@
 NatPlus
-Addition is a separate MIXINv2 file that inherits $\mathrm{NatData}$ and extends each constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field.
+Addition is a separate MIXINv2 file that inherits $\mathrm{NatData}$ and extends each data constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field.
 The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
@@
 To test that $2+3=5$, we needed equality on Nat values.
-Equality requires case analysis on the constructor; we use the Visitor pattern.
-$\mathrm{NatVisitor}$ adds a $\mathrm{Visitor}$ operation to each constructor, with a $\mathrm{Visit}$ field that delegates to the appropriate $\mathrm{VisitZero}$ or $\mathrm{VisitSuccessor}$ branch:
+Equality requires case analysis on the data constructor; we use the visitor pattern.
+$\mathrm{NatVisitor}$ adds an $\mathrm{Acceptance}$ operation to each data constructor, containing a $\mathrm{VisitorMap}$ with the appropriate $\mathrm{ZeroVisitor}$ or $\mathrm{SuccessorVisitor}$ branch, and an $\mathrm{Accepted}$ field that delegates to the matching branch:
 $\mathrm{NatVisitor}\mapsto\{\;$ $\mathrm{NatData},$
 $\mathrm{NatFactory}\mapsto\{$
-$\quad\mathrm{Product}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{Visit}\mapsto\{\}\}\},$
-$\quad\mathrm{Zero}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitZero}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitZero}\}\}\},$
-$\quad\mathrm{Successor}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitSuccessor}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitSuccessor}\}\}\}\}\}$
+$\quad\mathrm{Product}\mapsto\{\mathrm{Acceptance}\mapsto\{\mathrm{Accepted}\mapsto\{\}\}\},$
+$\quad\mathrm{Zero}\mapsto\{\mathrm{Acceptance}\mapsto\{\mathrm{VisitorMap}\mapsto\{\mathrm{ZeroVisitor}\mapsto\{\}\},$
+$\qquad\mathrm{Accepted}\mapsto\{\mathrm{VisitorMap}.\mathrm{ZeroVisitor}\}\}\},$
+$\quad\mathrm{Successor}\mapsto\{\mathrm{Acceptance}\mapsto\{\mathrm{VisitorMap}\mapsto\{\mathrm{SuccessorVisitor}\mapsto\{\}\},$
+$\qquad\mathrm{Accepted}\mapsto\{\mathrm{VisitorMap}.\mathrm{SuccessorVisitor}\}\}\}\}\}$
 $\mathrm{BooleanData}$ defines the output sort:
@@
 $\mathrm{Boolean}\mapsto\{\mathrm{BooleanFactory},\;\mathrm{BooleanFactory}.\mathrm{Product}\}\}$
-$\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each constructor.
+$\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each data constructor.
 $\mathrm{Zero}$ is equal only to another $\mathrm{Zero}$; $\mathrm{Successor}$ is equal only to another $\mathrm{Successor}$ with an equal predecessor.
-Case analysis on $\mathrm{other}$ is done via its $\mathrm{Visitor}$:
+Case analysis on $\mathrm{other}$ is done via its $\mathrm{Acceptance}$:
 $\mathrm{NatEquality}\mapsto\{\;$ $\mathrm{NatVisitor},\;\mathrm{BooleanData},$
@@
 $\qquad\mathrm{Zero}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product}\},$
-$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
-$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{True}\}\},$
-$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
-$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
-$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\},$
+$\hskip 18.49988pt\mathrm{OtherAcceptance}\mapsto\{\mathrm{other}.\mathrm{Acceptance},$
+$\hskip 18.49988pt\;\mathrm{VisitorMap}\mapsto\{\mathrm{ZeroVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{True}\}\},$
+$\hskip 18.49988pt\;\qquad\mathrm{SuccessorVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\}\},$
+$\hskip 18.49988pt\;\mathrm{Accepted}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
+$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherAcceptance}.\mathrm{Accepted}.\mathrm{equal}\}\}\},$
 $\qquad\mathrm{Successor}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product},\;\mathrm{predecessor}\mapsto\{\mathrm{Product}\}\},$
@@
 $\hskip 18.49988pt\;\mathrm{other}\mapsto\{\mathrm{Equal}.\mathbf{this}.\mathrm{other}.\mathrm{predecessor}\}\},$
-$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
-$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
-$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{RecursiveEquality}.\mathrm{equal}\}\},$
-$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
-$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\}\}\}$
+$\hskip 18.49988pt\mathrm{OtherAcceptance}\mapsto\{\mathrm{other}.\mathrm{Acceptance},$
+$\hskip 18.49988pt\;\mathrm{VisitorMap}\mapsto\{\mathrm{ZeroVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
+$\hskip 18.49988pt\;\qquad\mathrm{SuccessorVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{RecursiveEquality}.\mathrm{equal}\}\}\},$
+$\hskip 18.49988pt\;\mathrm{Accepted}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
+$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherAcceptance}.\mathrm{Accepted}.\mathrm{equal}\}\}\}\}\}$
 Composing $\mathrm{NatPlus}$ with $\mathrm{NatEquality}$ by inheriting both, with no modifications to either file, gives the resulting Nat values both $\mathrm{Plus}$ and $\mathrm{Equal}$ automatically.
@@
 Each operation is an independent mixin that extends a factory’s subtree.
-Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited mixins.
+Inheritance recursively merges subtrees that share a label, so every label in the factory acquires all operations from all inherited mixins.
 Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory.
@@
 The reference $\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties.
-In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix B), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously.
+In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix D), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously.
 The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.
 Church-encoded Nats are tries
-A Church-encoded Nat is a record whose structure mirrors the constructor path used to build it: $\mathrm{Zero}$ is a flat record; $\mathrm{S}(\mathrm{Zero})$ is a record with a $\mathrm{predecessor}$ field pointing to a $\mathrm{Zero}$ record; and so on.
+A Church-encoded Nat is a record whose structure mirrors the data-constructor path used to build it: $\mathrm{Zero}$ is a flat record; $\mathrm{S}(\mathrm{Zero})$ is a record with a $\mathrm{predecessor}$ field pointing to a $\mathrm{Zero}$ record; and so on.
 This is exactly the shape of a trie.
@@
 6.1. Emergent Phenomena
-Several features of inheritance-calculus are primitive yet have no direct counterpart in the $\lambda$-calculus. Encoding any one of them requires pervasive, non-local changes to program structure.
-Self-referential records with lazy evaluation:
-A mixin’s properties can reference other properties and enclosing scopes through qualified this without explicit binding.
-In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (Boudol, 2004) (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
-Open recursion with symmetric inheritance:
-Inheritance merges definitions from independent sources, all sharing the same inherited mixin tree.
-Encoding this in the $\lambda$-calculus requires maintaining an extensible dictionary of methods that can be merged from multiple directions.
-Cook’s denotational semantics of inheritance (Cook, 1989) devoted an entire dissertation to formalizing this pattern in a functional setting.
-Commutative, idempotent merge:
-The inheritance operation is symmetric: a record’s elements form a set, so reordering them produces the same observable result.
-In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery (Harper and Pierce, 1991; Cardelli, 1994).
+In developing real programs in MIXINv2, we encountered three phenomena that we did not design for.
+They can be seen in the source code (Yang, 2026) and the Case Study (Section 5).
+Below we explain why they arise from the three primitives.
 Relational semantics:
-Every inheritance-calculus operation is defined by recursion on mixin trees, not on individual values.
-When a mixin tree is viewed as a trie, the correspondence with relational algebra becomes precise: a path from the root to a leaf encodes a single value (e.g., $\mathrm{Successor}\to\mathrm{predecessor}\to\mathrm{Zero}$ encodes the natural number $1$); a trie union encodes a set of such values, i.e., a unary relation.
-An operation that takes two trie-union operands traverses every combination of their paths through $\mathtt{bases}$ (equation 4), which is a cross-join; the transitive closure $\mathtt{bases}^{*}$ in $\mathtt{supers}$ (equation 2) computes the union of all reachable paths; and $\mathtt{properties}$ (equation 1) projects the observable labels.
-The carrier is a trie rather than a flat relation, and observation is coinductive—the observer drives computation by querying progressively deeper paths—rather than Datalog’s bottom-up materialization, but the relational structure of the computation is the same.
-In the Case Study (Section 5), the $\mathrm{Plus}$ and $\mathrm{Equal}$ operations are defined on single Nat values.
-When applied to set-valued mixins $\{1,2\}$ and $\{3,4\}$, the same definitions, without any modification, produce the set of all pairwise sums $\{4,5,6\}$.
-This is relational semantics in action: each operation distributes over every combination of its inputs.
-The behavior is native: the semantic equations of Section 3 resolve every property to a set of paths, so single-valued and set-valued computation are the same mechanism.
-No code change is needed to move from one to the other.
-Function color blindness:
-In a web application (Yang, 2026) initially targeting Python’s synchronous sqlite3 and http.server modules, the business logic MIXINv2 file contained no async-specific code.
-When the FFI layer was replaced with one built on aiosqlite and starlette, the business logic MIXINv2 file did not change.
-This is an instance of function color blindness (Nystrom, 2015): async and sync functions are distinct types in the host language, yet the MIXINv2 file is agnostic to the distinction because its empty property slots are filled at assembly time, with coroutine handlers under asyncio, or with direct implementations under synchronous execution.
-More generally, empty slots can be filled with effect handlers (Plotkin and Power, 2003; Pretnar, 2015) under a CPS interpretation, or with direct implementations under a direct-style interpretation, without explicit monad machinery.
+Single-valued computation and set-valued computation are the same mechanism.
+The semantic equations of Section 3 resolve every property to a set of paths, so operations defined on single values automatically distribute over all combinations of set-valued inputs (Section 5).
+When a mixin tree is viewed as a trie, the correspondence with relational algebra becomes precise.
+[^9] Inheritance-calculus to relational algebra: path from root to leaf == tuple; trie union == relation (set of tuples); $\mathtt{bases}$ (equation 4) == cross-join; $\mathtt{supers}$ (equation 2) == union; $\mathtt{properties}$ (equation 1) == projection.
+The carrier is a trie rather than a flat relation, and observation is coinductive: the observer drives computation by querying progressively deeper paths.
+This contrasts with Datalog’s bottom-up materialization, but the relational structure of the computation is the same.
+Appendix F makes this correspondence precise with a systematic encoding of Datalog into inheritance-calculus, demonstrating that variable joins, constant matches, multi-body rules, and recursive rules are all expressible using only the three primitives of Section 2.
 Immunity to nonextensibility:
-Programming languages typically provide a mix of extensible and nonextensible constructs.
-In such languages, extensibility is opt-in: frameworks must encode nonextensible constructs into extensible constructs using sophisticated mechanisms (Liang et al., 1995; Lämmel and Jones, 2003; Löh and Hinze, 2006; Swierstra, 2008; Carette et al., 2009; d.
+It is structurally impossible to write a nonextensible program.
+Every construct of inheritance-calculus (Section 2) is extensible via deep merge (equation 3).
+The Case Study (Section 5) and the Expressive Asymmetry (Section 4.2; Theorem C.7) demonstrate and formalize this: new operations and data types are added, and new observable projections are introduced.
+This is the extensibility that the Expression Problem (Wadler, 1998) demands.
+[^10] As an untyped calculus, inheritance-calculus cannot violate the extensibility requirement of the Expression Problem but does not encounter its type-safety requirement.
+A typed language built on inheritance-calculus (Yang, 2026) inherits the immunity to nonextensibility and need only ensure that its type system does not introduce nonextensible constructs.
+Programming languages typically provide a mix of extensible and nonextensible constructs, where extensibility is opt-in: frameworks must encode nonextensible constructs into extensible constructs using sophisticated mechanisms (Liang et al., 1995; Lämmel and Jones, 2003; Löh and Hinze, 2006; Swierstra, 2008; Carette et al., 2009; d.
 S.
@@
 Oliveira, 2016; Leijen, 2017; Poulsen and van der Rest, 2023), and code that uses nonextensible constructs cannot be extended unless rewritten in the framework.
-In contrast, every construct of inheritance-calculus (Section 2) is extensible via deep merge (equation 3).
-There is no nonextensible construct, so it is impossible to write a nonextensible program.
-In the Case Study (Section 5), new operations and data types are added without modifying existing files.
-Tree-level inheritance (Section 5) and multi-path qualified this (equation 6; Appendix B) account for this.
-The Expressive Asymmetry (Section 4.2; Theorem A.3) provides the formal counterpart: the mixin trees in its separating construction are mechanical translations of $\lambda$-terms via $\mathcal{T}$ (Section 4), yet a new observable projection is added to them without modification.
-This is the extensibility that the Expression Problem (Wadler, 1998) demands.
-[^3] As an untyped calculus, inheritance-calculus cannot violate the extensibility requirement of the Expression Problem but does not encounter its type-safety requirement.
-A typed language built on inheritance-calculus (Yang, 2026) inherits the immunity to nonextensibility and need only ensure that its type system does not introduce nonextensible constructs.
-Each of these features is primitive in inheritance-calculus.
-In the $\lambda$-calculus, encoding them requires pervasive changes to program structure: every scope must thread a mutable store for lazy allocation, every method invocation must consult a dynamically constructed dictionary for open recursion, and every composition must be wrapped in order-insensitive infrastructure for commutativity.
-These are global, cascading transformations, not local rewrites of individual terms.
-In common sense, inheritance-calculus is strictly more expressive than the $\lambda$-calculus.
+Section 6.2 analyzes why deep merge is the only composition mechanism that achieves this.
+Function color blindness:
+Business logic in MIXINv2 declares abstract property slots; different assemblies fill them with synchronous or asynchronous implementations via inheritance.
+Because the slot’s color varies at assembly time without affecting the code that uses it, the same MIXINv2 file runs under both synchronous and asynchronous runtimes (Yang, 2026), an instance of function color blindness (Nystrom, 2015).
+The mechanism is dependency injection via inheritance: business logic declares empty slots, and the assembly inherits an FFI module that fills them, so the same business logic accepts implementations of any color.
+[^11] Function color (Nystrom, 2015) arises when effectful functions have signatures $a\to f\,b$ (Kleisli arrows), requiring monadic bind to compose and making the effect $f$ visible at every call site.
+Because inheritance-calculus composes program structure rather than executing effects, this type distinction does not arise at the calculus level.
+The color is determined at assembly time by the choice of inherited implementation, and effect execution is delegated to the host language.
+Unstructured control flow, such as capturing and resuming continuations as in call/cc, cannot be abstracted this way, as it requires a global CPS transformation that inheritance-calculus does not express.
 6.2. Semantic Alternatives
@@
 Degrees of freedom
-Every language whose programs are tree-structured ASTs with references can be characterized by two finite sets: a set $\mathcal{N}$ of node types, each determining how a subtree organizes its children, and a set $\mathcal{R}$ of reference types, each determining how one position in the tree refers to another.
+The analysis below assumes that the AST is a named tree:every edge from parent to child carries a label, so that every node is uniquely identified by the path from the root.
+For configuration languages this is naturally the case: every key in a JSON, YAML, or Nix attribute set is a label, and nesting produces paths.
+[^12] Nix is a functional language with first-class functions; its attribute sets alone do not form a named-tree DSL.
+The NixOS module system (NixOS Contributors, 2024), however, restricts the user-facing interface to nested attribute sets composed by recursive merging, and it is this DSL-level structure that naturally forms a named tree.
+For mainstream programming languages the AST is not directly a named tree, since anonymous expression positions and implicit scopes lack labels.
+After ANF transformation (Flanagan et al., 1993), however, every intermediate result is bound to a name, and the remaining anonymous positions [^13] For example, the body of a $\mathbf{let}$.
+can be assigned synthetic fresh labels such as $\mathrm{result}$ and $\mathrm{tailCall}$ in Section 4, yielding a named tree.
+Under this premise, every language whose programs are such named trees with references can be characterized by two finite sets: a set $\mathcal{N}$ of node types, each determining how a subtree organizes its children, and a set $\mathcal{R}$ of reference types, each determining how one position in the tree refers to another.
 The syntax is fixed by the choice of $\mathcal{N}$ and $\mathcal{R}$; the semantics is determined by how each node type structures composition and how each reference type resolves its target.
@@
 We now classify the design space to explain why we chose the semantics of Section 3.
-Two axes remain free: the composition mechanism—how the single reference type incorporates definitions; and the reference resolution—how the single reference type finds its target.
+Two axes remain free: the composition mechanism, which determines how the single reference type incorporates definitions; and the reference resolution, which determines how the single reference type finds its target.
 The composition mechanism
@@
 $\beta$-reduction $\times$ $\times$ $\times$ opt-in Java, Haskell, Scala, …
-Conditional breaks monotonicity
+Conditional causes divergence
 Deep merge:
@@
 Same-label definitions are rejected as errors.
-All three identities hold—but vacuously: the identities are satisfied because the compositions that would test them are refused.
-Conflict rejection prevents extending existing labels (Harper and Pierce, 1991)—the composability that the Expression Problem (Wadler, 1998) requires.
+All three identities hold, but vacuously: the identities are satisfied because the compositions that would test them are refused.
+Conflict rejection prevents extending existing labels (Harper and Pierce, 1991), precisely the composability that the Expression Problem (Wadler, 1998) requires.
 Priority merge:
-$A+B\neq B+A$: one source takes precedence (Cunningham, 2014; Dolstra et al., 2010), introducing an ordering dependency.
+One source takes precedence (Cunningham, 2014; Dolstra et al., 2010), so $A+B\neq B+A$, introducing an ordering dependency.
 $\beta$-reduction:
@@
 To make a new aspect extensible, the original function must be rewritten to accept an additional parameter.
-This applies to every language whose composition mechanism is function application—functional, object-oriented, and multi-paradigm alike—because method calls, virtual dispatch, and type-class resolution all reduce to $\beta$-reduction.
+This applies to every language whose composition mechanism is function application, whether functional, object-oriented, or multi-paradigm, because method calls, virtual dispatch, and type-class resolution all reduce to $\beta$-reduction.
 The Expression Problem (Wadler, 1998) is hard in all of them for the same reason: extensibility is opt-in.
-The frameworks that simulate universal extensibility (visitor pattern, object algebras (d.
+The frameworks that simulate immunity to nonextensibility (visitor pattern, object algebras (d.
 S.
@@
 Conditional incorporation:
-Removing or conditionally including definitions breaks monotonicity, so the semantic equations (Appendix C) may have no fixed point (Tarski, 1955).
-Among the candidates above, deep merge is the only one we found that is commutative, idempotent, associative, and universally extensible.
-The other candidates each introduce a vulnerability to nonextensibility or break monotonicity.
+Removing or conditionally including definitions can introduce cycles in the dependency graph, causing the recursive evaluation (Appendix A) to diverge.
+Among the candidates above, deep merge is the only one we found that is commutative, idempotent, associative, and immune to nonextensibility.
+The other candidates each introduce a vulnerability to nonextensibility or cause divergence.
 The reference resolution
@@
 Dynamic scope breaks the $\lambda$-calculus embedding (Section 4).
-The Substitution Lemma (Lemma D.2) relies on the base case $M=y$ ($y\neq x$): the de Bruijn index $m$ of $y$ walks up from the definition-site path $\mathtt{init}(p_{\mathrm{def}})$ to a scope level different from $x$’s binder, so the substitution $\mathrm{argument}\mapsto\mathcal{T}(V)$ at $x$’s scope has no effect.
+The Substitution Lemma (Lemma B.2) relies on the base case $M=y$ ($y\neq x$): the de Bruijn index $m$ of $y$ walks up from the definition-site path $\mathtt{init}(p_{\mathrm{def}})$ to a scope level different from $x$’s binder, so the substitution $\mathrm{argument}\mapsto\mathcal{T}(V)$ at $x$’s scope has no effect.
 With dynamic scope this separation fails.
@@
 Single-target $\mathtt{this}$:
-Single-target $\mathtt{this}$ is a promising candidate: the Single-Path Lemma (D.1) shows that for translated $\lambda$-terms, the frontier set $S$ in $\mathtt{this}$ (6) always contains exactly one path, so the $\lambda$-calculus embedding does not require multi-target resolution.
+Single-target $\mathtt{this}$ is a promising candidate: the Single-Path Lemma (B.1) shows that for translated $\lambda$-terms, the frontier set $S$ in $\mathtt{this}$ (6) always contains exactly one path, so the $\lambda$-calculus embedding does not require multi-target resolution.
 However, single-target $\mathtt{this}$ is vulnerable to nonextensibility for a different reason.
 When multiple inheritance routes reach the same scope, single-target resolution must reject the situation or select one route.
-The NixOS module system (NixOS Contributors, 2024) raises a duplicate-declaration error (Appendix B).
-This rejects nested extensions where two independently authored modules define the same option—a composition that arises when solving the Expression Problem (Wadler, 1998) (Section 5).
+The NixOS module system (NixOS Contributors, 2024) raises a duplicate-declaration error (Appendix D).
+This rejects nested extensions where two independently authored modules define the same option, a composition that arises when solving the Expression Problem (Wadler, 1998) (Section 5).
 Our first three implementations of MIXINv2 used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972).
 All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope.
-The difficulty was structural: a closure captures a single environment, so a fixed-point combinator solves for a single self-reference; a stack-based environment is a linear chain where each scope level has one parent.
-Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix B).
+The difficulty was structural.
+A closure captures a single environment, so a fixed-point combinator solves for a single self-reference.
+A stack-based environment is a linear chain where each scope level has one parent.
+Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix D).
 Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope.
@@
 $\beta$-reduction opt-in extensibility Java, Haskell, Scala, …
-Conditional breaks monotonicity
+Conditional causes divergence
 resolution Early binding mixin cannot replace function CUE (van Lohuizen, 2019)
@@
 The semantics of inheritance-calculus was not chosen from a menu.
-We arrived at it by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-path compositions that the Expression Problem demands (three successive implementations confirmed this—see above).
+We arrived at it by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-path compositions that the Expression Problem demands.
+[^15] Three successive implementations confirmed this; see the discussion above.
 Deep merge with multi-target late-binding $\mathtt{this}$ is what remained.
 7. Related Work
+7.1. Computational Models and $\lambda$-Calculus Semantics
 Minimal Turing-complete models
@@
 The bracket abstraction that compiles $\lambda$-terms to SKI is mechanical but incurs exponential blow-up in term size; no new computational patterns emerge beyond those already present in the $\lambda$-calculus.
+Semantically, the two share the same models: Meyer (Meyer, 1982) showed that combinatory algebras satisfying five additional conditions are exactly lambda algebras.
+No independent semantic domain for SKI has ever been exhibited from which $\lambda$-calculus semantics could be derived.
+The direction has always been the reverse: Scott’s $D_{\infty}$ was built for the $\lambda$-calculus, and SKI was interpreted inside it afterwards.
 Dolan (Dolan, 2013) showed that the x86 mov instruction alone is Turing complete by exploiting addressing modes as implicit arithmetic on mutable memory; the resulting programs are correct but provide no abstraction mechanisms beyond those of the RAM Machine.
@@
 Operational semantics takes $\beta$-reduction $(\lambda x.\,M)\,N\to M[N/x]$ as the primitive notion of computation (Church, 1936, 1941; Barendregt, 1984).
-This requires capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi (Abadi et al., 1991), and a choice of evaluation strategy: call-by-name versus call-by-value (Plotkin, 1975) or the lazy strategy of Abramsky (Abramsky, 1990).
-Denotational semantics (Scott, 1970, 1976) interprets terms in a mathematical domain, but giving meaning to $\beta$-reduction requires a reflexive domain $D\cong[D\to D]$, which has no set-theoretic solution by Cantor’s theorem; Scott’s $D_{\infty}$ resolves this by constructing $D$ as an inverse limit of continuous lattices (Scott, 1972).
+This requires capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi (Abadi et al., 1991).
+It also requires a choice of evaluation strategy: call-by-name versus call-by-value (Plotkin, 1975) or the lazy strategy of Abramsky (Abramsky, 1990).
+Denotational semantics (Scott, 1970, 1976) interprets terms in a mathematical domain, but giving meaning to $\beta$-reduction requires a reflexive domain $D\cong[D\to D]$, which has no set-theoretic solution by Cantor’s theorem.
+Scott’s $D_{\infty}$ resolves this by constructing $D$ as an inverse limit of continuous lattices (Scott, 1972).
 The resulting semantics is adequate but not fully abstract (Plotkin, 1977; Milner, 1977): it identifies terms that observational equivalence distinguishes (Plotkin, 1975).
@@
 In each of these approaches, the function space $D\to D]$ is the central obstacle: operational semantics needs it implicitly (via substitution), denotational semantics constructs it explicitly, and game semantics refines it via strategies.
-None of these approaches yields a semantics of the lazy $\lambda$-calculus that simultaneously lives in naive set theory and is fully abstract: Scott’s $D_{\infty}$ is adequate but not fully abstract, and game semantics is fully abstract but requires machinery (arenas, strategies, innocence conditions) far beyond set theory.
-None of them eliminates the function space from the semantic domain; none avoids capture-avoiding substitution; and none dispenses with reduction rules or evaluation strategies.
+None of these approaches is first-order; [^16] Function spaces in denotational semantics, strategies in game semantics, and substitution in operational semantics.
+none eliminates the function space from the semantic domain; and none unifies denotational and operational semantics into a single first-order definition.
+All three also inherit $\beta$-reduction’s divergence on recursive programs: a naïve transitive closure on a cyclic graph, for instance, diverges in every framework.
 Set-theoretic models of the $\lambda$-calculus
 Several families of $\lambda$-calculus models avoid the inverse-limit construction of Scott’s $D_{\infty}$ by working in set-theoretic rather than domain-theoretic settings.
-Graph models (Engeler, 1981; Plotkin, 1993) take a web $A$ equipped with an injection $\iota:\mathcal{P}_{\mathrm{fin}}(A)\times A\hookrightarrow A$ and set $D=\mathcal{P}(A)$.
-Application is $f\cdot x=\{\,b\mid\exists\,S\subseteq x,\;\iota(S,b)\in f\,\}$: the function space $D\to D]$ is not constructed explicitly but encoded inside $D$ via $\iota$.
-The construction uses only powerset lattices, no Scott topology or continuity conditions.
-Graph models are sensible (they equate all unsolvable terms) but not fully abstract: Bucciarelli and Salibra (Bucciarelli and Salibra, 2008) proved that the greatest sensible graph theory equals the Böhm tree theory $\mathcal{B}$, which is strictly coarser than observational equivalence for the lazy $\lambda$-calculus.
-Filter models (Barendregt et al., 1983; Coppo and Dezani-Ciancaglini, 1978) interpret a term as the filter (upward-closed, intersection-closed set) of all intersection types derivable for it.
-The collection of all such filters forms an $\omega$-algebraic lattice.
-Barendregt, Coppo, and Dezani-Ciancaglini proved that type assignment is sound and complete with respect to the filter $\lambda$-model.
-Dezani-Ciancaglini et al.
-(Dezani-Ciancaglini et al., 1998) showed that the $\lambda$-theory of the BCD filter model coincides exactly with Böhm tree equality: two terms receive the same set of types if and only if they have the same Böhm tree.
-Filter models are more discriminating than graph models but are still not fully abstract for the lazy $\lambda$-calculus, and require the additional machinery of intersection type theory and algebraic lattices.
-Relational models (Ehrhard, 2005; Bucciarelli and Ehrhard, 2001) arise from the relational semantics of linear logic: the exponential modality !!
-is interpreted via finite multisets $\mathcal{M}_{\mathrm{fin}}(A)$, giving a co-Kleisli category whose morphisms are relations from $\mathcal{M}_{\mathrm{fin}}(A)$ to $B$.
-The multiset structure corresponds to non-idempotent intersection types (de Carvalho, 2018), which carry quantitative information (bounds on reduction length) that idempotent systems lack.
-Ehrhard (Ehrhard, 2005) showed that in the finiteness-space refinement of this model, standard fixed-point operators are not finitary, and a standard reflexive object for the pure $\lambda$-calculus cannot exist in this category.
-For the lazy $\lambda$-calculus, full abstraction has been achieved only through game semantics (Abramsky and Ong, 1993; Abramsky and McCusker, 1995); none of the three set-theoretic families attains it.
-All three families are semantic models of the $\lambda$-calculus: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure.
-None of them is a self-contained calculus with its own primitives into which the $\lambda$-calculus embeds.
-Moreover, all three identify terms more coarsely than observational equivalence—graph and filter models coincide with Böhm tree equality, while relational models cannot even host a standard reflexive object (Ehrhard, 2005)—and none provides a fully abstract semantics for the lazy $\lambda$-calculus without resorting to game-theoretic machinery.
+These models eliminate Scott topology but not the function space itself: they remain higher-order, encoding $D\to D]$ inside $D$ via injections, intersection types, or multiset relations.
+Graph models (Engeler, 1981; Plotkin, 1993) encode functions via a web injection on powerset lattices; Bucciarelli and Salibra (Bucciarelli and Salibra, 2008) proved that the greatest sensible graph theory equals the Böhm tree theory $\mathcal{B}$.
+Filter models (Barendregt et al., 1983; Coppo and Dezani-Ciancaglini, 1978) interpret a term as the set of intersection types derivable for it; Dezani-Ciancaglini et al.
+(Dezani-Ciancaglini et al., 1998) showed that the BCD filter model’s $\lambda$-theory coincides exactly with Böhm tree equality.
+Relational models (Ehrhard, 2005; Bucciarelli and Ehrhard, 2001) arise from the relational semantics of linear logic via finite multisets; Ehrhard (Ehrhard, 2005) showed that a standard reflexive object for the pure $\lambda$-calculus cannot exist in this category.
+All three families are semantic models: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure.
+None is a self-contained calculus with its own primitives into which the $\lambda$-calculus embeds.
+None is first-order, since the function space is encoded, not eliminated.
+And none provides a fully abstract denotational semantics of the lazy $\lambda$-calculus without higher-order machinery.
+Recursive equations and fixpoint semantics
+Van Emden and Kowalski (van Emden and Kowalski, 1976) showed that the semantics of predicate logic as a programming language admits three equivalent characterizations: operational (SLD resolution), model-theoretic (least Herbrand model), and fixed-point (least fixed point of the immediate consequence operator $T_{P}$ on the powerset of ground atoms).
+Aczel (Aczel, 1977) gave a systematic treatment of inductive definitions via monotone operators on powerset lattices, providing the logical foundations that Datalog-style semantics rests on.
+Leroy and Grall (Leroy and Grall, 2006) treated divergence in big-step semantics via coinductive definitions; in inheritance-calculus, divergence already emerges as non-termination of recursive evaluation, so no separate coinductive formulation is needed.
+None of these prior frameworks has been applied to a first-order domain of paths to obtain a fully abstract semantics of the lazy $\lambda$-calculus without function spaces, substitution, or reduction rules.
+In particular, Van Emden and Kowalski’s $T_{P}$ computes $\mathrm{lfp}$ for recursive Datalog programs, but has not been used to give recursive $\lambda$-programs a fixpoint semantics in which they converge instead of diverging.
+7.2. Inheritance, Objects, and Records
 Denotational semantics of inheritance
-Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records and inheritance as composition of generators, which are functions from self to complete object, and wrappers, which are functions that modify generators, solved by a fixed-point construction.
-Cook’s key insight, that inheritance is a general mechanism applicable to any form of recursive definition and not only object-oriented methods, is one of the starting points of the present work.
-The model has four primitives—records, functions, generators, and wrappers—and requires a fixed-point combinator; composition is asymmetric, as a wrapper modifies a generator but not vice versa.
-Mixin-based inheritance
-Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses, that is, functions from a superclass parameter to a subclass, unifying the inheritance mechanisms of Smalltalk, Beta, and CLOS under a single model.
-Because mixin application is function composition, it is neither commutative nor idempotent: applying the same mixin twice may produce different results, and applying two mixins in different orders may yield different linearizations.
-The C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order; whether linearization is inherent to inheritance or an artifact of non-commutativity had not been established.
-Mixin composition operates at the method level: applying a mixin replaces or extends individual methods in a flat namespace.
-There is no recursive merging of nested structure: composing two mixins that define the same method name results in an override, not deep merge.
-Finally, mixins presuppose the $\lambda$-calculus as their computational foundation: method bodies are functions, and mixin application is function composition.
-Mixins organize and compose functions but cannot replace them; without the underlying $\lambda$-calculus, there is nothing to compose.
-The mixin mechanism itself is not Turing complete.
-Traits
+Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records.
+Inheritance is composition of generators (functions from self to complete object) and wrappers (functions that modify generators), solved by a fixed-point construction.
+Cook’s key insight is that inheritance is a general mechanism applicable to any form of recursive definition, not only object-oriented methods.
+This insight is one of the starting points of the present work.
+The model has four primitives (records, functions, generators, and wrappers) and requires a fixed-point combinator.
+Composition is asymmetric: a wrapper modifies a generator but not vice versa.
+The fixed-point construction relies on Scott continuity in a domain of functions.
+The resulting denotations are not directly executable, so a separate operational semantics is needed to define method dispatch and object construction at runtime.
+Mixins and traits
+Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses—functions from a superclass parameter to a subclass.
+Because mixin application is function composition, it is neither commutative nor idempotent; the C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order.
 Schärli et al.
-(Schärli et al., 2003) introduced traits as composable units of behavior.
-Trait composition is symmetric and commutative: the sum of two traits is order-independent, which is a departure from mixin ordering.
-However, when two traits provide methods with the same name, a conflict arises and must be explicitly resolved by the programmer through exclusion or aliasing.
-Ducasse et al.
-(Ducasse et al., 2006) extended the model with a flattening property: the semantics of a class is independent of whether its methods originate from traits or are defined directly.
-Like mixins, trait composition is flat: it operates on individual method names, not on nested structures, so same-name definitions are conflicts rather than recursive merges.
-And like mixins, traits presuppose the $\lambda$-calculus: trait methods are functions, and the computational power resides in those functions, not in the trait mechanism itself.
-Family polymorphism and virtual classes
-Ernst (Ernst, 2001) introduced family polymorphism, where families of mutually dependent classes can be refined together in subclasses.
-Ernst, Ostermann, and Cook (Ernst et al., 2006) formalized this in the virtual class calculus, where classes are class-valued attributes of objects, accessed via path expressions of the form this.out.C.
-The path this.out navigates to an enclosing scope, analogous to qualified this ($\ell_{\mathrm{up}}.\mathbf{this}$) in this paper.
-A notable restriction is that each such path resolves to a single class in a single enclosing object.
-When multiple inheritance routes lead to the same scope, a common situation under inheritance, single-valued resolution must choose one route and discard the others, losing information.
-The framework provides no mechanism for set-valued resolution.
-Object calculi
-Abadi and Cardelli (Abadi and Cardelli, 1996) developed a comprehensive theory of objects, treating objects rather than classes as the primitive notion.
-Their calculi include method update and self-referential objects, but methods are fundamentally functions that take self as a parameter, and object extension is asymmetric since method override replaces the previous definition.
-Boudol (Boudol, 2004) studied the recursive record semantics of objects in a call-by-value setting, showing that self-referential records require an unsafe fixed-point operator.
-Whether self-reference can be handled without a fixed-point operator in the object language, for instance through coalgebraic observation rather than reduction, was not explored in these calculi.
-Record calculi
-Harper and Pierce (Harper and Pierce, 1991) presented a record calculus based on symmetric concatenation with a type system that tracks the absence of fields.
-When two records define the same label, concatenation is undefined and raises a type error; the calculus leaves same-label composition undefined.
-Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) gave a type system for records and variants in ML.
-These calculi operate on flat records with no nesting and treat records as data, that is, initial algebras: a record is a finite map from labels to values, fully determined at construction time.
-Nested records, self-reference, and lazy observation, which are the properties needed for records to serve as a computational substrate, are outside their scope.
-DOT calculus
+(Schärli et al., 2003) introduced traits, whose composition is commutative but rejects same-name conflicts (Ducasse et al., 2006).
+Both mechanisms operate at the method level: composing two definitions of the same name results in an override or a conflict, not a recursive merge of nested structure.
+Both presuppose the $\lambda$-calculus: method bodies are functions, and the composition mechanism itself is not Turing complete.
+Neither mixins nor traits are therefore deep-mergeable.
+Object and record calculi
+Abadi and Cardelli (Abadi and Cardelli, 1996) treat objects as the primitive notion, but methods are functions that take self as a parameter and object extension is asymmetric.
+Boudol (Boudol, 2004) showed that self-referential records require an unsafe fixed-point operator whose well-definedness depends on the evaluation strategy.
+Harper and Pierce (Harper and Pierce, 1991) gave a record calculus with symmetric concatenation that rejects same-label composition; Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) introduced row polymorphism for records and variants in ML.
+All of these calculi operate on flat records: there is no recursive merging of nested structure, no lazy observation, and self-reference requires an explicit fixed-point combinator.
+Family polymorphism and DOT
+Ernst (Ernst, 2001) introduced family polymorphism; Ernst, Ostermann, and Cook (Ernst et al., 2006) formalized it in the virtual class calculus, where classes are accessed via path expressions this.out.C.
 Amin et al.
-(Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala’s type system.
-DOT paths navigate through objects to reach nested type members, providing a rich dependent type structure.
-DOT’s self variable in the object constructor $\nu(x:T)\,d$ binds $x$ to a single object; when multiple inheritance routes lead to the same enclosing scope, the resulting intersection type is unrealizable and Scala rejects it at compile time (Appendix B).
-DOT stratifies its syntax into terms and types, with types further divided into abstract type members bounded by $L..U$ and concrete definitions; each stratum requires its own subtyping rules.
-Whether these three roles could be unified under a single construct remains an open question in the DOT line of work.
-DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure mixin inheritance (see Section 8).
+(Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala.
+In both frameworks, each path resolves to a single class in a single enclosing object.
+When multiple inheritance routes lead to the same scope, single-valued resolution must choose one route and discard the others; DOT’s self variable $\nu(x:T)\,d$ binds $x$ to a single object, and Scala rejects the resulting intersection type at compile time (Appendix D).
+Neither framework provides a mechanism for set-valued resolution.
+7.3. Configuration Languages and Module Systems
 Configuration languages
@@
 The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance.
-When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix B).
+When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix D).
 The module system cannot express that both routes contribute equally to the same option, that is, the multi-path situation that recursive merging naturally gives rise to, which motivated the observational semantics of Section 3.
+7.4. Our Companion Work
 Modular software architectures
@@
 The symmetric, associative nature of inheritance (Section 3) explains why such systems work: components can be inherited in any order without changing behavior.
-These systems implement inheritance-calculus patterns ad hoc.
+MIXINv2 (Yang, 2026) is itself a dependency injection framework: each scope declares its dependencies as named slots, and composition resolves them by name across scope boundaries, making the inheritance-calculus patterns explicit rather than ad hoc.
 Union file systems
 Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view.
-Their semantics resembles inheritance-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers.
+Their semantics resembles inheritance-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, noncommutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers.
 We implemented inheritance-calculus directly as a union file system (Yang, 2025), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives.
@@
 The remaining items concern the standard library: they require no extensions to the calculus or type system, but involve nontrivial encodings within the existing framework.
-Inheritance-calculus is open by default: inheritance can freely merge any two mixins, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$).
-This is not a defect; it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 5).
-However, many operations (e.g., equality testing) assume that values are linear, inhabiting exactly one constructor.
-The observer pattern (Section 5) does not enforce this: applied to a multi-constructor value, all callbacks fire and their results are inherited.
+Inheritance-calculus is open by default: inheritance can freely merge any two mixins, and a value may simultaneously inhabit multiple data constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$).
+This is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 5).
+However, many operations (e.g., equality testing) assume that values are linear, inhabiting exactly one data constructor.
+The observer pattern (Section 5) does not enforce this: applied to a multi-data-constructor value, all callbacks fire and their results are inherited.
 Schema validation can be implemented at the library level using existing primitives.
-An observer can test whether a particular constructor is present, returning a Boolean; a second Boolean dispatch then branches on the result.
-By chaining such tests, a factory can validate that a value matches exactly one constructor and carries the required fields.
+An observer can test whether a particular data constructor is present, returning a Boolean; a second Boolean dispatch then branches on the result.
+By chaining such tests, a factory can validate that a value matches exactly one data constructor and carries the required fields.
 This does not restrict inheritance itself, which remains unconstrained, but provides a way to detect invariant violations before they propagate.
 The same technique, an observer returning Boolean followed by Boolean dispatch, yields closed pattern matching.
-The observer pattern of Section 5 is inherently open: new constructor cases can be added through inheritance.
-Closed matching, where exactly one branch is taken, can be encoded as a visitor chain: a linked list of if-then-else nodes, each testing one constructor and falling through on mismatch.
+The observer pattern of Section 5 is inherently open: new data-constructor cases can be added through inheritance.
+Closed matching, where exactly one branch is taken, can be encoded as a visitor chain: a linked list of if-then-else nodes, each testing one data constructor and falling through on mismatch.
 This is analogous to GHC.Generics’ $\mathrm{(:+:)}$ sum representation (Magalhães et al., 2010), where each link corresponds to one summand.
@@
 The emergent phenomena exhibit an expressive asymmetry between inheritance-calculus and the $\lambda$-calculus.
-Inheritance is not a mechanism for organizing functions; functions are syntactic sugar for inheritance.
-Appendix A Expressive Asymmetry: Proofs
-Definition A.1 (Macro-expressibility, after Felleisen (Felleisen, 1991)).
+Functions are syntactic sugar for inheritance.
+Appendix A Well-Definedness of the Semantic Functions
+The six semantic equations (1)–(6) are first-order: their inputs and outputs are paths and sets of paths, with no function spaces, closures, or higher-order constructs.
+This section establishes well-definedness by showing that the six equations define a monotone operator whose least fixed point is computed soundly by tabled evaluation (Tamaki and Sato, 1986; Chen and Warren, 1996).
+Definition A.1 (Dependency graph).
+Fix an AST with its primitive functions $\mathtt{defines}$ and $\mathtt{inherits}$.
+The dependency graph $G$ has as vertices all semantic-function applications ($\mathtt{properties}(p)$, $\mathtt{supers}(p)$, $\mathtt{overrides}(p)$, $\mathtt{bases}(p)$, $\mathtt{resolve}(\ldots)$, $\mathtt{this}(\ldots)$) that arise from the AST.
+There is a directed edge from vertex $u$ to vertex $v$ whenever computing $u$ via equations (1)–(6) requires the value of $v$.
+Definition A.2 (Immediate consequence operator $T_{P}$).
+Let $\mathcal{Q}$ be the set of all semantic-function queries (vertices of $G$).
+For each query $q\in\mathcal{Q}$, let $\mathcal{V}_{q}$ be the set of answer values of the appropriate shape for $q$ (paths, sets of paths, or sets of pairs of paths, depending on the equation defining $q$).
+A query–answer map is an element of the pointwise product $\prod_{q\in\mathcal{Q}}\mathcal{P}(\mathcal{V}_{q})$; equivalently, it assigns to each query $q$ a set $M(q)\subseteq\mathcal{V}_{q}$ of candidate answers.
+The immediate consequence operator (van Emden and Kowalski, 1976) $T_{P}$ maps a query–answer map $M$ to a new query–answer map $T_{P}(M)$ by evaluating, for each query $q$, the right-hand side of the equation that defines $q$, interpreting every sub-query $q^{\prime}$ by the current approximation $M(q^{\prime})$.
+Lemma A.3 (Monotonicity).
+The operator $T_{P}$ is monotone on the complete lattice $\Bigl(\prod_{q\in\mathcal{Q}}\mathcal{P}(\mathcal{V}_{q}),\;\subseteq\Bigr)$.
+Proof.
+Each equation’s right-hand side is built from $\cup$, $\in$, equality tests, and positive set comprehension $\{\,x\mid x\in S,\ldots\,\}$.
+None of the six equations uses negation, set difference, or complement.
+Therefore, if $M\subseteq M^{\prime}$ (pointwise), then $T_{P}(M)\subseteq T_{P}(M^{\prime})$.
+∎
+Theorem A.4 (Existence and uniqueness of $\mathrm{lfp}(T_{P})$).
+The operator $T_{P}$ has a unique least fixed point $\mathrm{lfp}(T_{P})$.
+Proof.
+Each powerset lattice $(\mathcal{P}(\mathcal{V}_{q}),\subseteq)$ is complete, so the pointwise product lattice $\Bigl(\prod_{q\in\mathcal{Q}}\mathcal{P}(\mathcal{V}_{q}),\subseteq\Bigr)$ is also complete.
+By the Knaster–Tarski theorem (Tarski, 1955), every monotone function on a complete lattice has a unique least fixed point.
+∎
+Theorem A.5 (Well-definedness).
+The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined: their denotational semantics is $\mathrm{lfp}(T_{P})$, the unique least fixed point of the immediate consequence operator.
+Theorem A.6 (Tabling soundness).
+Tabled evaluation of equations (1)–(6) never returns a wrong answer, even when the dependency graph $G$ is infinite. Specifically:
+  1. (1)
+    Soundness (no false positives).
+Every fact returned by tabling was produced by evaluating some equation’s right-hand side on its inputs.
+By induction on the evaluation trace, every returned fact belongs to $\mathrm{lfp}(T_{P})$.
+  2. (2)
+    Monotone accumulation. The tabling accumulator only grows between iterations (via set union); no fact is ever removed.
+  3. (3)
+    Correct fixpoint detection.
+If no sub-query was re-entered during evaluation, all sub-queries were fully computed with no approximation, and the result is exact.
+If re-entry occurred but the accumulator is stable (no new facts were added), the accumulator equals $\mathrm{lfp}(T_{P})$ restricted to the reachable queries, because it is a pre-fixed point reached by ascending from $\bot$.
+  4. (4)
+    Sound failure.
+When tabling cannot determine the answer, whether because the iteration budget is exhausted or because of an infinite non-cyclic chain, it raises an error rather than returning a wrong answer.
+Proof.
+For (1), each returned fact is the output of some equation’s right-hand side applied to previously returned facts; by induction, all inputs are in $\mathrm{lfp}(T_{P})$, so the output is too (by monotonicity and the fixed-point property).
+For (2), the implementation accumulates via $A\leftarrow A\cup T_{P}(A)$, which is monotone in $A$.
+For (3), when no reentry occurs, the evaluation is ordinary memoised recursion on a finite acyclic reachable subgraph, so every sub-query is computed exactly and no approximation is involved.
+When reentry occurs, let $R\subseteq\mathcal{Q}$ be the set of queries reachable from the initial query, and let $T_{P}^{R}$ denote the restriction of $T_{P}$ to maps on $R$.
+If the accumulator is stable, i.e.
+$A=A\cup T_{P}^{R}(A)$, then $T_{P}^{R}(A)\subseteq A$, so $A$ is a pre-fixed point of $T_{P}^{R}$.
+By (1), every fact ever added to the accumulator was produced by some right-hand side evaluation and therefore belongs to $\mathrm{lfp}(T_{P}^{R})$. Hence $A\subseteq\mathrm{lfp}(T_{P}^{R})$.
+Conversely, by the Knaster–Tarski characterization of the least fixed point as the least pre-fixed point, $\mathrm{lfp}(T_{P}^{R})$ is contained in every pre-fixed point of $T_{P}^{R}$, in particular in $A$.
+Thus $\mathrm{lfp}(T_{P}^{R})\subseteq A$.
+Therefore $A=\mathrm{lfp}(T_{P}^{R})$, i.e. the stable accumulator equals the exact denotation on the reachable queries.
+For (4), budget exhaustion and stack overflow both produce exceptions, never silent wrong answers. ∎
+Termination
+Whether evaluation terminates depends on the program:
+  • Acyclic, finite reachable subgraph. One pass suffices; evaluation terminates.
+  • Cyclic, finite per-query domain. Multiple passes converge by the ascending chain condition on a finite lattice (Bancilhon and Ramakrishnan, 1986). This is the Datalog case (Section F.8).
+  • Infinite reachable subgraph. Evaluation may diverge; this is inherent to Turing-completeness and not a defect of the evaluation strategy.
+Appendix B Böhm Tree Correspondence: Proofs
+This appendix contains the full proofs of Adequacy (Theorem 4.2) and the first-order semantics theorem (Theorem 4.3) stated in Section 4.1.
+B.1. Böhm Trees
+We briefly recall the definition of Böhm trees (Barendregt, 1984).
+A head reduction $M\to_{h}M^{\prime}$ contracts the outermost $\beta$-redex only: if $M$ has the form $(\lambda x.\,e)\;v\;M_{1}\cdots M_{k}$, then $M\to_{h}e[v/x]\;M_{1}\cdots M_{k}$.
+A term $M$ is in head normal form (HNF) if it has no head redex, i.e., $M=\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ where $y$ is a variable.
+The Böhm tree $\mathrm{BT}(M)$ of a $\lambda$-term $M$ is an infinite labeled tree defined by:
+$\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\begin{aligned} &\text{if }M\to^{*}_{h}\\[-2.0pt] &\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{aligned}\end{cases}$
+Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Abramsky and Ong, 1993).
+B.2. Path Encoding
+We define a correspondence between positions in the Böhm tree and paths in the mixin tree.
+When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting mixin tree has a specific shape:
+  • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$ (the abstraction shape).
+  • A $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ translates to a record with own properties $\{x,\mathrm{result}\}$, where $x$ holds the encapsulated application $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$.
+  • A tail call $V_{1}\;V_{2}$ translates to a record with own properties $\{\mathrm{tailCall},\mathrm{result}\}$, where $\mathrm{tailCall}$ holds the encapsulated application and $\mathrm{result}$ projects $\mathrm{tailCall}.\mathrm{result}$.
+  • A variable $x$ with de Bruijn index $n$ translates to an indexed reference $\uparrow^{\!n}.\mathrm{argument}$.
+Each Böhm tree position is a sequence of navigation steps from the root.
+In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body (peeling off one abstraction layer) and entering the $i$-th argument $M_{i}$ of the head variable.
+In the mixin tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
+Rather than comparing absolute paths across different mixin trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection.
+Recall the definition of inheritance-convergence (Definition 4.1).
+Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation.
+The abstraction shape at the end signals that a weak head normal form has been reached.
+B.3. Single-Path Lemma
+The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path mixin inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
+Lemma B.1 (Single Path).
+For any closed ANF $\lambda$-term $M$, at every invocation of $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)$ during evaluation of $\mathcal{T}(M)$, the frontier set $S$ contains exactly one path.
+Proof.
+By case analysis on the translation rules of $\mathcal{T}$.
+Since $\mathcal{T}$ is compositional, the structural invariant established for each rule propagates to all inherited subtrees: every subtree appearing in the mixin tree is itself the image of some sub-term under $\mathcal{T}$, so the invariant holds throughout the runtime composition.
+Case $\mathcal{T}(\lambda x.\,M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
+This is a record literal with $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$ and no inheritance sources, so $\mathtt{inherits}=\varnothing$.
+At the root path $p$, we have $\mathtt{overrides}(p)=\{p\}$ (a single path) since there are no inheritance sources to introduce additional branches.
+Inside $\mathcal{T}(M)$, any reference to $x$ is $\uparrow^{\!n}.\mathrm{argument}$ where $n$ is the de Bruijn index pointing to this scope level.
+Since $\mathcal{T}(\lambda x.\,M)$ introduces exactly one scope level, there is exactly one mixin at that level, hence $\mathtt{this}$ finds exactly one matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ in $\mathtt{supers}$.
+Case $\mathcal{T}(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
+The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level ($\mathtt{inherits}=\varnothing$), so $\mathtt{overrides}(\text{root})=\{\text{root}\}$.
+Inside the $x$ subtree, the inheritance $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$ has exactly one inheritance source ($\mathcal{T}(V_{1})$), creating exactly one $\mathtt{bases}$ entry.
+By induction on $V_{1}$, the single-path property holds inside $\mathcal{T}(V_{1})$.
+Inside $\mathcal{T}(M)$, the induction hypothesis applies directly.
+Case $\mathcal{T}(\mathbf{let}\;x=V\;\mathbf{in}\;M)=\{x\mapsto\{\mathrm{result}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
+The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level, so $\mathtt{overrides}(\text{root})=\{\text{root}\}$.
+The $x$ subtree is itself a record literal with own property $\{\mathrm{result}\}$ and no inheritance sources.
+Thus neither entering the bound value nor entering the body introduces branching in $\mathtt{supers}$.
+By induction, the single-path property holds inside both $\mathcal{T}(V)$ and $\mathcal{T}(M)$.
+Case $\mathcal{T}(V_{1}\;V_{2})=\{\mathrm{tailCall}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\}$
+Identical to the $\mathbf{let}$-binding case: the outer record has own properties $\{\mathrm{tailCall},\mathrm{result}\}$ with no inheritance sources at the outer level, and the $\mathrm{tailCall}$ subtree contains exactly one inheritance source.
+Case $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$
+This is a reference, not a record.
+The $\mathtt{this}$ function walks up $n$ steps from the enclosing scope.
+By the compositionality argument above, every scope level encountered during the walk, including inherited subtrees reached through reference resolution, is the image of some sub-term under $\mathcal{T}$, so each step of $\mathtt{this}$ encounters exactly one matching pair in $\mathtt{supers}$.
+The key structural invariant is the following.
+For every subtree generated by $\mathcal{T}$, its root has either no inheritance source or exactly one inheritance source.
+Moreover, the only roots with an inheritance source are the application wrappers introduced by the application $\mathbf{let}$-binding and tail-call rules, and in both cases that source is unique and encapsulated under a distinguished child ($x$ or $\mathrm{tailCall}$), never at the enclosing scope level.
+Therefore every scope level traversed by $\mathtt{this}$ contributes at most one matching pair in $\mathtt{supers}$, so the frontier set remains a singleton throughout evaluation.
+∎
+B.4. Substitution Lemma
+The core mechanism of the translation is that inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ plays the role of substitution. We make this precise.
+Lemma B.2 (Substitution).
+Let $M$ be an ANF term in which $x$ may occur free, and let $V$ be a closed value.
+Define the inherited tree $C=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$.
+Then for every path $p$ under $C\mathbin{\triangleright}\mathrm{result}$ and every label $\ell$:
+$\ell\in\mathtt{properties}(p)\text{ in }C\mathbin{\triangleright}\mathrm{result}\quad\Longleftrightarrow\quad\ell\in\mathtt{properties}(p)\text{ in }\mathcal{T}(M[V/x])$
+where $M[V/x]$ is the usual capture-avoiding substitution.
+Proof.
+By structural induction on $M$.
+Base case: $M=x$
+Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference.
+In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ inherits into the $\mathrm{argument}$ slot, providing $\mathcal{T}(V)$.
+Since $C$ is the $\mathrm{tailCall}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma B.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree.
+Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
+Base case: $M=y$ where $y\neq x$ and $y$ is $\lambda$-bound
+Then $\mathcal{T}(y)=\uparrow^{\!m}.\mathrm{argument}$ where $m$ is the de Bruijn index of $y$.
+This walks up to the scope that binds $y$, which is different from the scope that binds $x$, so the inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ at $x$’s binding scope has no effect.
+Since $M[V/x]=y$, the properties are identical.
+Base case: $M=y$ where $y\neq x$ and $y$ is let-bound
+Then $\mathcal{T}(y)=y.\mathrm{result}$.
+This is a lexical reference to the sibling property $y$ in the enclosing record, so the inherited tree $C$ does not alter its resolution.
+Since $M[V/x]=y$, the properties are identical.
+Inductive case: $M=\lambda y.\,M^{\prime}$
+Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$.
+The subtree at $C\mathbin{\triangleright}\mathrm{result}$ is another abstraction-shaped record.
+References to $x$ inside $\mathcal{T}(M^{\prime})$ resolve through $\mathtt{this}$ in exactly the same way (with one additional scope level from the inner $\lambda$), and by induction on $M^{\prime}$, the properties under $C\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M^{\prime}[V/x])$.
+Since $(\lambda y.\,M^{\prime})[V/x]=\lambda y.\,(M^{\prime}[V/x])$ (assuming $y$ is fresh), the result follows.
+Inductive case: $M=\mathbf{let}\;z=V_{1}\;V_{2}$ $\mathbf{in}\;M^{\prime}$
+Then
+$\mathcal{T}(M)=\{z\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}.$
+The substitution distributes: $M[V/x]=\mathbf{let}\;z=V_{1}[V/x]\;V_{2}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$.
+By induction on $V_{1}$, $V_{2}$, and $M^{\prime}$, the properties under each subtree ($z$ and $\mathrm{result}$) coincide, since references to $x$ inside each subtree resolve to $\mathcal{T}(V)$ via the same $\mathtt{this}$ mechanism.
+Inductive case: $M=\mathbf{let}\;z=V^{\prime}\;\mathbf{in}\;M^{\prime}$
+Then
+$\mathcal{T}(M)=\{z\mapsto\{\mathrm{result}\mapsto\mathcal{T}(V^{\prime})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}.$
+The substitution distributes: $M[V/x]=\mathbf{let}\;z=V^{\prime}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$.
+By induction on $V^{\prime}$ and $M^{\prime}$, the properties under both subtrees ($z$ and $\mathrm{result}$) coincide.
+In particular, occurrences of $z$ in $M^{\prime}$ are translated as the lexical reference $z.\mathrm{result}$ on both sides, so the wrapper introduced by the value $\mathbf{let}$-binding is preserved by the substitution.
+Inductive case: $M=V_{1}\;V_{2}$ (tail call)
+Analogous to the $\mathbf{let}$-binding case with $\mathrm{tailCall}$ in place of $z$. ∎
+B.5. Convergence Preservation
+The convergence criterion (Definition 4.1) follows the $\mathrm{result}$ chain from the root.
+Each step in this chain corresponds to one head reduction step in the $\lambda$-calculus.
+We make this precise.
+Lemma B.3 (Result-step).
+Let $(\lambda x.\,M)\;V$ be a tail call (a $\beta$-redex). Then for every path $p$ and label $\ell$:
+$\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)\;\Longleftrightarrow\;\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}(M[V/x])$
+That is, following one $\mathrm{result}$ projection in the tail call’s mixin tree yields the same properties as the root of the reduct’s mixin tree.
+Proof.
+The translation gives:
+$\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)=\{\mathrm{tailCall}\mapsto\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\}$
+The $\mathrm{result}$ label at the root is defined as $\mathrm{tailCall}.\mathrm{result}$, so $\text{root}\mathbin{\triangleright}\mathrm{result}$ resolves to the $\mathrm{result}$ path inside the inheritance $\mathrm{tailCall}=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$.
+This inheritance is exactly the inherited tree $C$ of Lemma B.2, and $\mathrm{tailCall}\mathbin{\triangleright}\mathrm{result}$ is $C\mathbin{\triangleright}\mathrm{result}$.
+By Lemma B.2, the properties under $C\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M[V/x])$.
+Let-binding variant. An analogous result holds for $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$ where $V_{1}=\lambda y.\,M^{\prime\prime}$. The translation gives:
+$\mathcal{T}\bigl(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\bigr)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$
+Since $\mathrm{result}\mapsto\mathcal{T}(M^{\prime})$ is a property definition, the subtree at $\text{root}\mathbin{\triangleright}\mathrm{result}$ is $\mathcal{T}(M^{\prime})$.
+Inside $\mathcal{T}(M^{\prime})$, each reference to $x$ is $x.\mathrm{result}$, which resolves to the $\mathrm{result}$ path inside the $x$ subtree $\{\mathcal{T}(\lambda y.\,M^{\prime\prime}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$.
+This is the inherited tree $C$ of Lemma B.2, so $x.\mathrm{result}$ has the same properties as $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$.
+In the $\lambda$-calculus, $M=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$, and the head reduct is $N=M^{\prime}[(V_{1}\;V_{2})/x]$.
+In $\mathcal{T}(N)$, each occurrence of $x$ is replaced by $(V_{1}\;V_{2})$, whose result, by the tail-call variant above, also yields $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$.
+Therefore the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$ for all $p$.
+Value let-binding variant. An analogous result holds for $\mathbf{let}\;x=V\;\mathbf{in}\;M^{\prime}$. The translation gives:
+$\mathcal{T}\bigl(\mathbf{let}\;x=V\;\mathbf{in}\;M^{\prime}\bigr)=\{x\mapsto\{\mathrm{result}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}.$
+Since $\mathrm{result}\mapsto\mathcal{T}(M^{\prime})$ is a property definition, the subtree at $\text{root}\mathbin{\triangleright}\mathrm{result}$ is $\mathcal{T}(M^{\prime})$.
+Inside $\mathcal{T}(M^{\prime})$, each occurrence of $x$ is translated as the lexical reference $x.\mathrm{result}$, which resolves to the subtree $\mathcal{T}(V)$.
+This is exactly the translation of the reduct $M^{\prime}[V/x]$.
+Therefore the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(\mathbf{let}\;x=V\;\mathbf{in}\;M^{\prime})$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(M^{\prime}[V/x])$ for all $p$.
+∎
+Theorem B.4 (Convergence preservation).
+If $M\to_{h}N$ (head reduction), then $\mathcal{T}(M){\Downarrow}$ if and only if $\mathcal{T}(N){\Downarrow}$.
+Proof.
+A head reduction step in ANF takes one of three forms: a tail call $(\lambda x.\,M^{\prime})\;V\to_{h}M^{\prime}[V/x]$, an application $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[(V_{1}\;V_{2})/x]$, or a value $\mathbf{let}$-binding $\mathbf{let}\;x=V\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[V/x]$.
+The second and third forms are simply the ANF presentation of one $\beta$-step in the underlying $\lambda$-term.
+In all three cases, Lemma B.3 (tail-call, application let-binding, and value let-binding variants) shows that the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n}$ in $\mathcal{T}(M)$ coincide with the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n-1}$ in $\mathcal{T}(N)$ for all $n\geq 1$, where $N$ is the head reduct.
+Hence the abstraction shape appears at depth $n$ in the pre-reduct if and only if it appears at depth n1n-1 in the post-reduct.
+∎
+Extension to $\mathrm{lfp}(T_{P})$ semantics
+The proof above works directly with the six equations (1)–(6), not with any particular evaluation strategy.
+We distinguish two consequences of switching from single-pass evaluation to multi-pass tabled evaluation (computing $\mathrm{lfp}(T_{P})$ via iterated application of $T_{P}$; see Appendix A).
+Head-reduction preservation is maintained.:
+The Result-Step Lemma (Lemma B.3) establishes a structural isomorphism between the equation systems of $\mathcal{T}(M)$ and $\mathcal{T}(N)$, shifted by one $\mathrm{result}$ step.
+This isomorphism preserves monotonicity, so $\mathrm{lfp}(T_{P})$ yields the same answer on both sides (modulo the depth shift): if $\mathcal{T}(M)$ converges at depth $n$ under $\mathrm{lfp}$, then $\mathcal{T}(N)$ converges at depth n1n-1, and vice versa.
+For $\lambda$-calculus–translated programs, the Single-Path Lemma (Lemma B.1) ensures that no cycles arise in the dependency graph (each scope level has exactly one inheritance source), so single-pass and multi-pass evaluation produce identical results in this fragment.
+Agreement with the $\lambda$-calculus is not maintained in general.:
+For programs beyond the $\lambda$-calculus fragment that introduce cyclic dependencies, $\mathrm{lfp}(T_{P})$ may converge where the $\lambda$-calculus diverges.
+This means $\mathrm{lfp}(T_{P})$ evaluation is strictly more defined than $\beta$-reduction on the translated image: $M{\Downarrow}_{\lambda}$ implies $\mathcal{T}(M){\Downarrow}_{\mathrm{lfp}}$, but not conversely.
+As a minimal illustration, the $\lambda$-term $\mathbf{let}\;x=x\;\mathbf{in}\;x$ diverges under $\beta$-reduction (unfolding $x$ produces $x$ again, ad infinitum). Translation $\mathcal{T}$ gives
+$\{x\mapsto\{\mathrm{result}\mapsto x.\mathrm{result}\},\;\mathrm{result}\mapsto x.\mathrm{result}\}$
+where $x.\mathrm{result}$ inherits from itself, creating a cyclic dependency.
+Under single-pass evaluation ($\mathit{max\_fixpoint\_iterations}=0$), the cycle raises $\bot$, matching the $\lambda$-calculus divergence.
+Under $\mathrm{lfp}(T_{P})$, fixpoint iteration detects the cycle and converges: $\mathrm{properties}(\mathrm{root})=\{x,\mathrm{result}\}$ and $\mathrm{properties}(\mathrm{root}\triangleright\mathrm{result})=\varnothing$, reflecting that the value is undefined (no abstraction shape).
+Appendix F gives a more motivated example (Datalog transitive closure on cyclic graphs).
+B.6. Adequacy
+A closed ANF term at the top level is either an abstraction (a value) or a computation (an application $\mathbf{let}$-binding, a value $\mathbf{let}$-binding, or a tail call).
+An abstraction translates to a record whose root immediately has the abstraction shape; a computation’s root has the form $\{x,\mathrm{result}\}$ or $\{\mathrm{tailCall},\mathrm{result}\}$, and one must follow the $\mathrm{result}$ chain to find the eventual value.
+Proof of Theorem 4.2.
+($\Rightarrow$) Suppose $M\to^{*}_{h}\lambda x.\,M^{\prime}$ in $k$ head reduction steps. We show $\mathcal{T}(M)$ converges at depth $\leq k$ by induction on $k$.
+Base case ($k=0$): $M$ is already an abstraction $\lambda x.\,M^{\prime}$.
+Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$, whose root has $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$.
+These labels are directly in $\mathtt{defines}$ of the root, so the recursive evaluation returns them without further recursion, and $\mathcal{T}(M)$ converges at depth $n=0$.
+Inductive step ($k\geq 1$): $M$ is not an abstraction, so $M$ is either a tail call, an application $\mathbf{let}$-binding, or a value $\mathbf{let}$-binding.
+Each form corresponds to one head $\beta$-step in the underlying $\lambda$-term.
+Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$.
+Note that $N$ is closed: head reduction substitutes a closed value into a term whose only free variable is the substitution target, preserving closedness.
+By Lemma B.3 (tail-call, application let-binding, or value let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$.
+By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
+($\Leftarrow$) Suppose $\mathcal{T}(M){\Downarrow}$ at depth $n$. We show $M$ has a head normal form by induction on $n$.
+Base case ($n=0$): the root of $\mathcal{T}(M)$ has abstraction shape $\{\mathrm{argument},\mathrm{result}\}$.
+Only the abstraction rule $\mathcal{T}(\lambda x.\,M^{\prime})$ produces a root with own property $\mathrm{argument}$.
+(For tail calls and $\mathbf{let}$-bindings, the root has $\mathtt{defines}=\{\mathrm{tailCall},\mathrm{result}\}$ or $\{x,\mathrm{result}\}$, and $\mathrm{argument}$ cannot appear via $\mathtt{supers}$ because $\mathtt{inherits}=\varnothing$ at the root, giving $\mathtt{bases}(\text{root})=\varnothing$.) Therefore $M$ is an abstraction, hence already in head normal form.
+Inductive step ($n\geq 1$): the abstraction shape appears at depth $n$ but not at depth 0.
+By the base case argument, $M$ is not an abstraction, so at the head position of this closed top-level ANF term, $M$ is either a tail call, an application $\mathbf{let}$-binding, or a value $\mathbf{let}$-binding.
+In each case, the head form is a $\beta$-redex in the underlying $\lambda$-term: the function position is occupied by a closed ANF value, and the only closed ANF value that can appear at head position is a $\lambda$-abstraction.
+By Lemma B.3 (tail-call, application let-binding, or value let-binding variant), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct.
+So $\mathcal{T}(N)$ converges at depth n1n-1.
+Since $M$ is closed and head reduction preserves closedness (substituting a closed value for the only free variable), $N$ is a closed ANF term.
+By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$.
+∎
+B.7. First-Order Semantics of the Lazy $\lambda$-Calculus
+Adequacy relates a single term’s head normal form to its inheritance-convergence.
+The following theorem lifts this to an equivalence between terms, establishing that the translation $\mathcal{T}$ gives the lazy $\lambda$-calculus a first-order semantics that preserves and reflects Böhm tree equivalence.
+Definition B.5 ($\lambda$-contextual equivalence under $\mathcal{T}$).
+For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
+$\mathcal{T}(C[M]){\Downarrow}\;\Longleftrightarrow\;\mathcal{T}(C[N]){\Downarrow}$
+The subscript $\lambda$ emphasizes that the quantification ranges over $\lambda$-calculus contexts, not over all inheritance-calculus contexts.
+This definition observes only convergence (a single bit), but quantifying over all $\lambda$-contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior within the $\lambda$-definable fragment.
+Proof of Theorem 4.3.
+The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the mixin tree $\mathcal{T}(C[M])$ is determined by the mixin tree $\mathcal{T}(M)$ and the translation of the context.
+This means $\mathcal{T}$ preserves the context structure needed for the argument.
+($\Leftarrow$) Suppose $\mathrm{BT}(M)=\mathrm{BT}(N)$.
+Böhm tree equivalence is a congruence (Barendregt, 1984), so for every context $C[\cdot]$, $\mathrm{BT}(C[M])=\mathrm{BT}(C[N])$.
+In particular, $C[M]$ has a head normal form iff $C[N]$ does.
+By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$.
+Hence $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$.
+($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$.
+Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa).
+By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$.
+Hence $\mathcal{T}(M)\not\approx_{\lambda}\mathcal{T}(N)$.
+∎
+The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993).
+The prior higher-order frameworks are thereby shown to be faithful to the first-order semantics, without needing to define or compare absolute paths inside mixin trees.
+Appendix C Expressive Asymmetry: Proofs
+Definition C.1 (Macro-expressibility, after Felleisen (Felleisen, 1991)).
 A construct $C$ of language $\mathscr{L}_{1}$ is macro-expressible in language $\mathscr{L}_{0}$ if there exists a translation $\mathcal{E}$ such that for every program $P$ containing occurrences of $C$, the translation $\mathcal{E}(P)$ is obtained by replacing each occurrence of $C$ with an $\mathscr{L}_{0}$ expression that depends only on $C$’s subexpressions, leaving all other constructs of $P$ unchanged.
-Theorem A.2 (Forward macro-expressibility).
+Lemma C.2 (Composition of macro-expressible translations).
+If $\mathcal{E}_{1}$ macro-expresses the constructs of $\mathscr{L}_{1}$ in $\mathscr{L}_{0}$ and $\mathcal{E}_{2}$ macro-expresses the constructs of $\mathscr{L}_{2}$ in $\mathscr{L}_{1}$, then the composition $\mathcal{E}_{1}\circ\mathcal{E}_{2}$ macro-expresses the constructs of $\mathscr{L}_{2}$ in $\mathscr{L}_{0}$.
+Proof.
+Each construct $F$ of $\mathscr{L}_{2}$ has a fixed syntactic abstraction $A_{F}^{(2)}$ over $\mathscr{L}_{1}$.
+Each $\mathscr{L}_{1}$-construct $G$ appearing in $A_{F}^{(2)}$ has a fixed syntactic abstraction $A_{G}^{(1)}$ over $\mathscr{L}_{0}$.
+Substituting $A_{G}^{(1)}$ for every $G$ in $A_{F}^{(2)}$ yields a fixed syntactic abstraction over $\mathscr{L}_{0}$ for $F$, satisfying E4.
+∎
+Definition C.3 (Language universe).
+Let $\mathscr{U}$ be the language whose constructors are the union of the constructors of inheritance-calculus, the constructors of the lazy $\lambda$-calculus ($\lambda$-abstraction, application, variable), and $\mathbf{let}$-binding.
+The semantics of $\mathscr{U}$ is given by translating every $\mathscr{U}$-program to inheritance-calculus in three steps:
+  1. (1)
+    Convert every maximal $\lambda$-calculus subexpression (i.e. every subtree whose root and all internal nodes use only $\lambda$-calculus constructors) to A-normal form (Flanagan et al., 1993).
+  2. (2)
+    Apply the translation $\mathcal{T}$ (Section 4) to every ANF $\lambda$-calculus subexpression.
+  3. (3)
+    Evaluate the resulting inheritance-calculus program by the rules of Section 3.
+When a $\mathscr{U}$-expression interleaves $\lambda$-calculus and inheritance-calculus constructors (e.g.
+a $\mathbf{let}$-binding whose bound value is an inheritance-calculus expression), the translation is applied compositionally from leaves to root: each $\lambda$-calculus construct is replaced by its inheritance-calculus equivalent via $\mathcal{T}$, treating inheritance-calculus subexpressions as opaque values.
+Since each rule of $\mathcal{T}$ (Section 4) is a fixed syntactic abstraction depending only on the translated subexpressions, this bottom-up application is well-defined and yields a pure inheritance-calculus program for step (3).
+A $\mathscr{U}$-program $P$ converges, $\mathrm{eval}_{\mathscr{U}}(P)$, iff its translation converges.
+The lazy $\lambda$-calculus and inheritance-calculus are conservative restrictions of $\mathscr{U}$: each is obtained by removing the other’s constructors.
+For a pure inheritance-calculus program, steps (1) and (2) are vacuous, so $\mathrm{eval}_{\mathscr{U}}$ agrees with the inheritance-calculus semantics.
+For a pure $\lambda$-calculus program, the composition of steps (1)–(3) agrees with lazy evaluation by Adequacy (Theorem 4.2).
+The three-step translation is effective (ANF conversion and $\mathcal{T}$ are syntactic transformations) and the recursive evaluation of inheritance-calculus (Section 3) is a semi-decision procedure, so $\mathrm{eval}_{\mathscr{U}}$ is recursively enumerable, as required by Felleisen’s framework (Felleisen, 1991).
+Lemma C.4 ($\mathbf{let}$-binding is macro-expressible in the lazy $\lambda$-calculus).
+The translation $\mathbf{let}\;x=e_{1}\;\mathbf{in}\;e_{2}\;\mapsto\;(\lambda x.\,e_{2})\;e_{1}$ is a macro-expressible elimination of $\mathbf{let}$ in the lazy $\lambda$-calculus.
+Proof.
+The translation replaces each $\mathbf{let}$-binding with a $\lambda$-expression that depends only on the two subexpressions $e_{1}$ and $e_{2}$.
+This is Proposition 3.7 of Felleisen (Felleisen, 1991).
+∎
+Lemma C.5 (Equal expressiveness of the lazy $\lambda$-calculus and its ANF fragment).
+In $\mathscr{U}$, the lazy $\lambda$-calculus and the ANF lazy $\lambda$-calculus (Section 4) are equally macro-expressive.
+Proof.
+Every ANF term is a $\lambda$-term, so the embedding of the ANF fragment into the full $\lambda$-calculus is the identity, which is trivially macro-expressible.
+For the reverse direction, the ANF transformation (Flanagan et al., 1993) consists of two steps, each macro-expressible:
+  1. (1)
+    Application flattening.
+Replace each nested application $(e_{1}\;e_{2})\;e_{3}$ with $\mathbf{let}\;x=e_{1}\;e_{2}\;\mathbf{in}\;x\;e_{3}$.
+The syntactic abstraction is $A(\cdot_{1},\cdot_{2},\cdot_{3})=\mathbf{let}\;x=\cdot_{1}\;\cdot_{2}\;\mathbf{in}\;x\;\cdot_{3}$, depending only on the three subexpressions.
+  2. (2)
+    Value lifting.
+Replace each free value $V$ used in an application position with $\mathbf{let}\;x=V\;\mathbf{in}\;x$.
+The syntactic abstraction is $A(\cdot_{1})=\mathbf{let}\;x=\cdot_{1}\;\mathbf{in}\;x$.
+By Lemma C.4, each resulting $\mathbf{let}$-binding is itself macro-expressible in the $\lambda$-calculus.
+By Lemma C.2, composing these macro-expressible translations yields a macro-expressible translation from the full $\lambda$-calculus to its ANF fragment.
+∎
+Theorem C.6 (Forward macro-expressibility).
 The translation $\mathcal{T}$ (Section 4) is a macro-expressible embedding of the $\lambda$-calculus into inheritance-calculus.
 Proof.
-Each of the five rules of $\mathcal{T}$ replaces one $\lambda$-calculus construct with an inheritance-calculus expression that depends only on the construct’s immediate subexpressions.
-No rule inspects the surrounding context.
+We verify condition E4 of Felleisen (Felleisen, 1991, Definition 3.11): for each $\lambda$-calculus construct $F$ of arity $a$, there must exist a fixed $a$-ary syntactic abstraction $A_{F}(\cdot_{1},\ldots,\cdot_{a})$ over inheritance-calculus such that $\mathcal{T}(F(e_{1},\ldots,e_{a}))=A_{F}(\mathcal{T}(e_{1}),\ldots,\mathcal{T}(e_{a}))$.
+We check each rule of $\mathcal{T}$ (Section 4):
+  1. (1)
+    $\lambda$-bound variable ($x$ with de Bruijn index $n^{\prime}$).
+This is a nullary construct (no subexpressions).
+The syntactic abstraction is the constant $A=\uparrow^{\!n^{\prime}}.\mathrm{argument}$.
+  2. (2)
+    Let-bound variable ($x$). Nullary. $A=x.\mathrm{result}$.
+  3. (3)
+    $\lambda$-abstraction ($\lambda x.\,M$). Unary in the body $M$. $A(\cdot_{1})=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\cdot_{1}\}$.
+  4. (4)
+    Application let-binding ($\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$).
+Ternary in $V_{1}$, $V_{2}$, and $M$.
+$A(\cdot_{1},\cdot_{2},\cdot_{3})=\{x\mapsto\{\cdot_{1},\;\mathrm{argument}\mapsto\cdot_{2}\},\;\mathrm{result}\mapsto\cdot_{3}\}$.
+  5. (5)
+    Value let-binding ($\mathbf{let}\;x=V\;\mathbf{in}\;M$). Binary in $V$ and $M$. $A(\cdot_{1},\cdot_{2})=\{x\mapsto\{\mathrm{result}\mapsto\cdot_{1}\},\;\mathrm{result}\mapsto\cdot_{2}\}$.
+  6. (6)
+    Tail call ($V_{1}\;V_{2}$). Binary.
+    $A(\cdot_{1},\cdot_{2})=\{\mathrm{tailCall}\mapsto\{\cdot_{1},\;\mathrm{argument}\mapsto\cdot_{2}\},\;\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\}$.
+In each case the syntactic abstraction $A_{F}$ is a fixed inheritance-calculus context whose holes are filled by the recursive translations of the subexpressions, satisfying E4. ∎
+Theorem C.7 (Nonexpressibility of inheritance).
+The lazy $\lambda$-calculus cannot macro-express the inheritance constructors of $\mathscr{U}$ (Definition C.3).
+Proof.
+We apply Theorem 3.14(i) of Felleisen (Felleisen, 1991) to the pair $L_{0}=\text{lazy $\lambda$-calculus}$, $L_{1}=\mathscr{U}$.
+By Definition C.3, $\mathscr{U}=L_{0}+\{$mixin-tree constructors}\} is a conservative extension of $L_{0}$ (Definition 3.2 of Felleisen (Felleisen, 1991)).
+We verify the four conditions:
+  1. (i)
+    $L_{0}\subseteq L_{1}$: The syntax of $\mathscr{U}$ includes all $L_{0}$-phrases.
+  2. (ii)
+    No new $L_{0}$-phrases: The mixin-tree constructors do not introduce new phrases in $L_{0}$.
+  3. (iii)
+    Disjoint constructors: The mixin-tree constructors are disjoint from the $L_{0}$-constructors.
+  4. (iv)
+    Semantics preserved: For pure $\lambda$-terms, steps (1) and (2) of Definition C.3 are vacuous and step (3) yields lazy evaluation by Adequacy (Theorem 4.2), so $\mathrm{eval}_{\mathscr{U}}$ agrees with $\mathrm{eval}_{L_{0}}$ on $L_{0}$-phrases.
+The theorem requires us to exhibit two $L_{0}$-terms $M$ and $N$ with $M\cong_{0}N$ but $M\not\cong_{1}N$, where $\cong_{0}$ is operational equivalence in $L_{0}$ and $\cong_{1}$ is operational equivalence in $\mathscr{U}$.
+Operational equivalence in $L_{0}$
+By Böhm tree equivalence (Abramsky, 1990) (Theorem 4.3), for closed $\lambda$-terms $M$ and $N$: $M\cong_{0}N$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$.
+Separating pair
+Define Church booleans and Church boolean equality: $\mathrm{true}=\lambda t.\,\lambda f.\,t$, $\mathrm{false}=\lambda t.\,\lambda f.\,f$, and $\mathrm{eq}=\lambda a.\,\lambda b.\,(a\;b)\;(b\;\mathrm{false}\;\mathrm{true})$.
+Let
+$M=\mathrm{eq}\;\mathrm{false}\;\mathrm{false}\qquad N=\mathrm{eq}\;\mathrm{true}\;\mathrm{true}.$
+Both reduce to Church $\mathrm{true}$ in the lazy $\lambda$-calculus, so $\mathrm{BT}(M)=\mathrm{BT}(N)$ and hence $M\cong_{0}N$.
+Distinguishing $\mathscr{U}$-context
+We now construct a $\mathscr{U}$-context that separates $M$ and $N$.
+$M$ and $N$ are syntactically closed $\lambda$-terms; by Definition C.3, their semantics in $\mathscr{U}$ is given by translating them to inheritance-calculus via ANF conversion followed by $\mathcal{T}$.
+Under the extended ANF convention (Section 4), every free value used in an application is first bound to a name:
+$\mathrm{ANF}(M)={}$ $\mathbf{let}\;\mathrm{eq}=\lambda a.\,\lambda b.\,\mathbf{let}\;x_{1}=a\;b\;\mathbf{in}\;$
+$\qquad\mathbf{let}\;f_{1}=\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;x_{2}=b\;f_{1}\;\mathbf{in}\;$
+$\qquad\mathbf{let}\;t_{1}=\mathrm{true}\;\mathbf{in}\;\mathbf{let}\;x_{3}=x_{2}\;t_{1}\;\mathbf{in}\;x_{1}\;x_{3}$
+$\mathbf{in}\;\;\mathbf{let}\;f_{1}=\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;\mathrm{eqFalse}=\mathrm{eq}\;f_{1}$
+$\quad\;\mathbf{in}\;\mathbf{let}\;f_{2}=\mathrm{false}\;\mathbf{in}\;\mathrm{eqFalse}\;f_{2}$
+Applying $\mathcal{T}$:
+$\mathcal{T}(\mathrm{ANF}(M))\;=\;\{$ $\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\mathcal{T}(\lambda a.\,\lambda b.\,\cdots)\},$
+$f_{1}\mapsto\{\mathrm{result}\mapsto\mathcal{T}(\mathrm{false})\},$
+$\mathrm{eqFalse}\mapsto\{\mathrm{eq}.\mathrm{result},\;\mathrm{argument}\mapsto f_{1}.\mathrm{result}\},$
+$f_{2}\mapsto\{\mathrm{result}\mapsto\mathcal{T}(\mathrm{false})\},$
+$\mathrm{tailCall}\mapsto\{\mathrm{eqFalse}.\mathrm{result},\;\mathrm{argument}\mapsto f_{2}.\mathrm{result}\},$
+$\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\;\}$
+and symmetrically for $N$ with $\mathrm{true}$ in place of $\mathrm{false}$.
+Note that $\mathrm{eq}$ is a named child of $\mathcal{T}(\mathrm{ANF}(M))$, bound via a value let-binding.
+Inside the translation of $\mathrm{eq}$, the inner $\lambda b$ scope contains the let-bindings for $x_{1}$, $f_{1}$, $x_{2}$, $t_{1}$, $x_{3}$, and a tail call.
+In particular, $\uparrow^{\!1}.\mathrm{argument}$ reaches the outer $\lambda a$ scope and $\uparrow^{\!0}.\mathrm{argument}$ reaches $\lambda b$ itself.
+[^17] These indices use the condensed convention, counting only $\lambda$-scopes.
+As noted in the translation footnote (Section 4), each let-binding and tail call introduces an additional scope level in inheritance-calculus; the formal index is correspondingly larger, but the target scope is the same.
+The $\mathscr{U}$-context uses inheritance to add a $\mathrm{repr}$ record inside the $\lambda b$ scope of $\mathrm{eq}$ that projects the first operand. Define the context $C[\cdot]$ as:
+$C[\cdot]={}$ $\mathbf{let}\;\mathrm{extended}=\{[\cdot],\;\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{$
+$\quad\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument}\}\}\}\}\}$
+$\mathbf{in}\;\mathrm{extended}.\mathrm{tailCall}.\mathrm{repr}.\mathrm{firstOperand}$
+Here $\cdot$] is the hole.
+Note the two levels of $\mathrm{result}$ nesting in the mixin: the outer $\mathrm{result}$ enters the value let-binding wrapper for $\mathrm{eq}$, and the inner $\mathrm{result}$ enters the $\lambda a$ scope.
+The second $\mathrm{result}$ then contains the $\lambda b$ scope, so $\mathrm{repr}$ is placed at the $\lambda b$ scope level.
+The mixin $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\cdots\}\}\}\}$ is an inheritance-calculus expression: it uses only the mixin-tree constructors of $\mathscr{U}$.
+The outer $\mathbf{let}$-binding and projection are $\lambda$-calculus constructs (Lemma C.4).
+Thus $C[\cdot]$ is a valid $\mathscr{U}$-context.
+The de Bruijn index $\uparrow^{\!2}$ in $\mathrm{repr}$ counts scope levels (in the condensed convention): $\uparrow^{\!0}$ is $\mathrm{repr}$ itself, $\uparrow^{\!1}$ is the $\lambda b$ scope, and $\uparrow^{\!2}$ is the $\lambda a$ scope.
+So $\mathrm{firstOperand}$ projects $\lambda a$’s argument.
+Since $\mathrm{eqFalse}$ (respectively $\mathrm{eqTrue}$) inherits $\mathrm{eq}.\mathrm{result}$, and $\mathrm{tailCall}$ inherits $\mathrm{eqFalse}.\mathrm{result}$ (the $\lambda b$ scope), the $\mathrm{repr}$ record placed by the mixin at the $\lambda b$ scope level propagates through the inheritance chain to $\mathrm{tailCall}$.
+In $C[M]$: the $\mathrm{eqFalse}$ binding supplies $\mathrm{argument}\mapsto f_{1}.\mathrm{result}$ (i.e.
+$\mathrm{false}$) to the $\lambda a$ scope, so $\mathrm{firstOperand}$ evaluates to $\mathrm{false}=\lambda t.\,\lambda f.\,f$.
+In $C[N]$: the $\mathrm{eqTrue}$ binding supplies $\mathrm{argument}\mapsto\mathrm{true}$, so $\mathrm{firstOperand}$ evaluates to $\mathrm{true}=\lambda t.\,\lambda f.\,t$.
+Define $D[\cdot]=(\lambda x.\,x\;\Omega\;I)\;C[\cdot]$.
+Then $\mathrm{eval}_{\mathscr{U}}$ holds for $D[M]$ (since $\mathrm{false}\;\Omega\;I\to I$, which converges) but not for $D[N]$ (since $\mathrm{true}\;\Omega\;I\to\Omega$, which diverges).
+Therefore $M\not\cong_{1}N$.
+Conclusion
+We have $M\cong_{0}N$ but $M\not\cong_{1}N$, so ${\cong_{0}}\neq{(\cong_{1}|_{L_{0}})}$.
+By Theorem 3.14(i) of Felleisen (Felleisen, 1991), the lazy $\lambda$-calculus cannot macro-express the inheritance constructors of $\mathscr{U}$.
 ∎
-Theorem A.3 (Non-expressibility of inheritance-calculus).
-There is no macro-expressible translation from inheritance-calculus to the $\lambda$-calculus.
+Theorem C.8 (Expressive asymmetry).
+In the language universe $\mathscr{U}$ (Definition C.3), inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus: it can macro-express every $\mathscr{U}$-construct that the $\lambda$-calculus can, but not conversely (Definition 3.17 of Felleisen (Felleisen, 1991)).
 Proof.
-Let $\mathscr{L}_{1}$ be the inheritance-calculus and let $\mathscr{L}_{0}$ be the sublanguage consisting of mixin trees in the image of the translation $\mathcal{T}$ (Section 4), so that $\mathscr{L}_{0}\subset\mathscr{L}_{1}$.
-Define the operational equivalence $\cong_{0}$ on $\mathscr{L}_{0}$ via Böhm trees: $\mathcal{T}(M)\cong_{0}\mathcal{T}(N)$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$.
-Define $\cong_{1}$ on $\mathscr{L}_{1}$ as inheritance-contextual equivalence: $S\cong_{1}S^{\prime}$ iff for every inheritance-calculus context $C[\cdot]$, $C[S]{\Downarrow}\Leftrightarrow C[S^{\prime}]{\Downarrow}$.
-By the Böhm Tree Correspondence (Theorems 4.2 and 4.3), $\mathscr{L}_{1}$ conservatively extends $\mathscr{L}_{0}$: for closed $\lambda$-terms $M$, $\mathcal{T}(M){\Downarrow}$ if and only if $M$ has a head normal form.
-Define Church booleans via the translation $\mathcal{T}$ (Section 4): $\mathrm{true}=\lambda t.\,\lambda f.\,t$ and $\mathrm{false}=\lambda t.\,\lambda f.\,f$. In de Bruijn notation:
-$\mathrm{true}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}$
-$\mathrm{false}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!0}.\mathrm{argument}\}\}$
-Define Church boolean equality $\mathrm{eq}=\lambda a.\,\lambda b.\,(a\;b)\;(b\;\mathrm{false}\;\mathrm{true})$. In ANF:
-$\mathrm{eq}=\lambda a.\,\lambda b.\,\mathbf{let}\;x_{1}=b\;\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;x_{2}=x_{1}\;\mathrm{true}\;\mathbf{in}\;\mathbf{let}\;x_{3}=a\;b\;\mathbf{in}\;x_{3}\;x_{2}$
-Applying $\mathcal{T}$ (Section 4):
-$\mathrm{eq}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},$
-$\quad x_{1}\mapsto\{\uparrow^{\!0}.\mathrm{argument},\;\mathrm{argument}\mapsto\mathrm{false}\},$
-$\quad x_{2}\mapsto\{x_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
-$\quad x_{3}\mapsto\{\uparrow^{\!1}.\mathrm{argument},\;\mathrm{argument}\mapsto\uparrow^{\!0}.\mathrm{argument}\},$
-$\quad\_\mathrm{tail}\mapsto\{x_{3}.\mathrm{result},\;\mathrm{argument}\mapsto x_{2}.\mathrm{result}\},$
-$\quad\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}\}$
-Here $\uparrow^{\!1}.\mathrm{argument}$ in $x_{3}$ reaches the outer $\lambda a$ scope (one $\lambda$ up from $\lambda b$), and $\uparrow^{\!0}.\mathrm{argument}$ reaches $\lambda b$ itself.
-In inheritance-calculus, applying a function record to an argument is: inherit the function, override its $\mathrm{argument}$ slot, and project $\mathrm{result}$.
-Define mixin trees $e_{1}$ and $e_{2}$ that apply $\mathrm{eq}$ to its two arguments via this mechanism:
-$e_{1}\;=\;\{$ $\mathrm{eqFalse}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{false}\},$
-$\mathrm{applied}\mapsto\{\mathrm{eqFalse}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{false}\},$
-$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
-$e_{2}\;=\;\{$ $\mathrm{eqTrue}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{true}\},$
-$\mathrm{applied}\mapsto\{\mathrm{eqTrue}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
-$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
-That is, $e_{1}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{false}\;\mathrm{false}))$ and $e_{2}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{true}\;\mathrm{true}))$, where $\mathrm{ANF}$ converts a $\lambda$-term to A-normal form by naming every intermediate application (Section 4).
-Since both $\mathrm{ANF}$ and $\mathcal{T}$ are macro-expressible, their composition is too, so $e_{1},e_{2}\in\mathscr{L}_{0}$.
-Both $e_{1}.\mathrm{result}$ and $e_{2}.\mathrm{result}$ evaluate to Church $\mathrm{true}$.
-By Full Abstraction (Theorem 4.3), $\mathrm{BT}(e_{1})=\mathrm{BT}(e_{2})$, so $e_{1}\cong_{0}e_{2}$.
-[…]
-$C={}$ $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{$
-$\qquad\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument},\;\mathrm{secondOperand}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}\}\}$
-The de Bruijn indices count scope levels from $\mathrm{repr}$: $\uparrow^{\!0}$ is $\mathrm{repr}$ itself, $\uparrow^{\!1}$ is the $\lambda b$ scope, and $\uparrow^{\!2}$ is the $\lambda a$ scope.
-Thus $\mathrm{firstOperand}$ projects $\lambda a$’s argument and $\mathrm{secondOperand}$ projects $\lambda b$’s argument.
-In the composite scope $\{S,\,C\}$, the definition of $\mathrm{eq}$ gains $\mathrm{repr}$ as a new sibling of $x_{1},x_{2},x_{3}$ inside the $\lambda b$ scope.
-By open recursion, the de Bruijn references $\uparrow^{\!2}.\mathrm{argument}$ and $\uparrow^{\!1}.\mathrm{argument}$ resolve through the same scope chain as the existing let-bindings.
-Since $e_{1}$ and $e_{2}$ inherit $\mathrm{eq}$, they also inherit the newly added $\mathrm{repr}$: $e_{1}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{false}$ and $e_{2}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{true}$.
-These produce structurally distinct mixin trees (one converges to an abstraction selecting its first argument, the other to an abstraction selecting its second), so $e_{1}\not\cong_{1}e_{2}$.
-Since $e_{1}\cong_{0}e_{2}$ but $e_{1}\not\cong_{1}e_{2}$, we have ${\cong_{0}}\neq{(\cong_{1}|_{\mathscr{L}_{0}})}$.
-By Theorem 3.14(i) of Felleisen (Felleisen, 1991), $\mathscr{L}_{0}$ cannot macro-express the inheritance facilities of $\mathscr{L}_{1}$.
+We must show: (a) inheritance-calculus can macro-express every $\mathscr{U}$-construct that the lazy $\lambda$-calculus can macro-express, and (b) the converse fails.
+Part (b): the lazy $\lambda$-calculus $\not\geq$ inheritance-calculus
+The $\lambda$-calculus contains its own constructors ($\lambda$, application, variable) and can macro-express $\mathbf{let}$-binding (Lemma C.4), but cannot macro-express the inheritance constructors of $\mathscr{U}$ (Theorem C.7).
+Inheritance-calculus contains those constructors.
+Therefore the lazy $\lambda$-calculus is not at least as expressive as inheritance-calculus with respect to $\mathscr{U}$.
+Part (a): inheritance-calculus $\geq$ the lazy $\lambda$-calculus
+We show that inheritance-calculus can macro-express every $\mathscr{U}$-construct that the $\lambda$-calculus contains or can macro-express.
+The $\lambda$-calculus contains $\lambda$-abstraction, application, and variable; it can macro-express $\mathbf{let}$-binding (Lemma C.4).
+We must show that inheritance-calculus can macro-express all four of these constructs.
+The translation proceeds in two stages:
+  1. (1)
+    ANF conversion.
+The A-normal form transformation (Flanagan et al., 1993) replaces nested applications with $\mathbf{let}$-bindings; it is macro-expressible in the $\lambda$-calculus by Lemma C.5.
+In the reverse direction, $\mathbf{let}$-bindings are macro-expressible in the $\lambda$-calculus by Proposition 3.7 of Felleisen (Felleisen, 1991) (restated as Lemma C.4).
+Thus ANF conversion is a macro-expressible equivalence between the $\lambda$-calculus and its ANF fragment.
+  2. (2)
+    The translation $\mathcal{T}$.
+Each ANF $\lambda$-calculus construct maps to a fixed syntactic abstraction over inheritance-calculus (Theorem C.6), satisfying condition E4 of Felleisen (Felleisen, 1991, Definition 3.11).
+By Lemma C.2, composing the two macro-expressible translations gives a macro-expressible translation from the $\lambda$-calculus (with $\mathbf{let}$) into inheritance-calculus.
+Since inheritance-calculus trivially contains its own constructors, it can macro-express every $\mathscr{U}$-construct that the $\lambda$-calculus contains or can macro-express.
 ∎
-Corollary A.4.
-Inheritance-calculus is strictly more expressive than the $\lambda$-calculus in the sense of Felleisen (Felleisen, 1991): the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
-Appendix B Multi-Path Self-Reference in Other Systems
+Appendix D Multi-Path Self-Reference in Other Systems
 This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-path self-reference pattern that inheritance-calculus handles naturally.
@@
 Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{HasMultipleOuters}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken.
-Since inheritance is idempotent, the two routes contribute the same properties in this semantics, so no route-priority disambiguation rule is introduced.
+Because attribute resolution aggregates results via set union (which is intrinsically idempotent), identical property definitions from the two routes merge without conflict.
+Thus, the calculus eliminates the need for any artificial route-priority or linearization rules.
 NixOS Module System
@@
 In inheritance-calculus, the same pattern is well-defined because $\mathtt{overrides}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
-Appendix C Well-Definedness of the Semantic Functions
-This appendix proves that the mutually recursive equations (1)–(6) have a unique least fixed point, and therefore the semantics of Section 3 is well-defined.
-C.1. The Product Lattice
-Fix an AST with its primitive functions $\mathtt{defines}$ and $\mathtt{inherits}$.
-Let $\mathcal{L}$ denote the set of all labels and $\mathcal{P}$ the set of all paths (finite sequences of labels).
-Define the following domains, one for each semantic function:
-$D_{\mathtt{properties}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{L})$ (sets of labels)
-$D_{\mathtt{supers}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P}\times\mathcal{P})$ (sets of site–override pairs)
-$D_{\mathtt{overrides}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
-$D_{\mathtt{bases}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
-$D_{\mathtt{resolve}}$ $=(\mathcal{P}\times\mathcal{P}\times\mathbb{N}\times\mathcal{L}^{*})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
-$D_{\mathtt{this}}$ $=(\mathcal{P}(\mathcal{P})\times\mathcal{P}\times\mathbb{N})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
-Each $D_{i}$ is a complete lattice under the pointwise subset ordering: $f\sqsubseteq g$ iff $f(x)\subseteq g(x)$ for all $x$.
-The product $D=D_{\mathtt{properties}}\times D_{\mathtt{supers}}\times D_{\mathtt{overrides}}\times D_{\mathtt{bases}}\times D_{\mathtt{resolve}}\times D_{\mathtt{this}}$ is a complete lattice under the componentwise ordering, with bottom element $\bot=(\lambda x.\,\varnothing,\ldots,\lambda x.\,\varnothing)$.
-C.2. The Operator
-Equations (1)–(6) define a function $F:D\to D$ that takes a tuple of candidate interpretations $(\mathtt{properties},\mathtt{supers},\mathtt{overrides},\mathtt{bases},\mathtt{resolve},\mathtt{this})$ and produces a new tuple by evaluating the right-hand side of each equation.
-The transitive closure $\mathtt{bases}^{*}$ appearing in equation (2) is derived from $\mathtt{bases}$: for any fixed $\mathtt{bases}$, the set $\mathtt{bases}^{*}(p)$ is defined as the least fixed point of the monotone operator $X\mapsto\{p\}\cup\bigcup_{q\in X}\mathtt{bases}(q)$ on $\mathcal{P}(\mathcal{P})$.
-If $\mathtt{bases}\sqsubseteq\mathtt{bases}^{\prime}$ (pointwise), then $\mathtt{bases}^{*}\sqsubseteq\mathtt{bases}^{\prime*}$ (pointwise), since enlarging the one-step relation can only enlarge the transitive closure.
-The $\mathtt{this}$ function involves a recursion on the de Bruijn index $n$.
-Since $n$ decreases by one at each step and is a non-negative integer, this recursion terminates for any fixed interpretation of $\mathtt{supers}$.
-The result of $\mathtt{this}(S,p_{\mathrm{def}},n)$ for all valid $(S,p_{\mathrm{def}},n)$ is therefore well-defined as a function of $\mathtt{supers}$, and it remains to show that it is monotone in $\mathtt{supers}$.
-C.3. Monotonicity
-Proposition C.1.
-The operator $F$ is monotone: if $\vec{d}\sqsubseteq\vec{d}^{\prime}$ in $D$, then $F(\vec{d})\sqsubseteq F(\vec{d}^{\prime})$.
-Proof.
-It suffices to show that each component of $F$ is monotone in every argument, i.e., that enlarging any of the input interpretations (in the $\subseteq$ ordering) can only enlarge the output.
-We verify this by inspecting the right-hand side of each equation and confirming that every occurrence of a mutually defined function appears in a positive position—as the domain of an existential quantifier in a set comprehension (Aczel, 1977).
-Equation (1): $\mathtt{properties}$
-$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{override}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{override}})\,\bigr\}.$
-The function $\mathtt{supers}$ appears as the set being iterated over ($\in\mathtt{supers}(p)$).
-The predicate $\ell\in\mathtt{defines}(\cdot)$ involves only the AST primitive $\mathtt{defines}$, not any mutually defined function.
-Enlarging $\mathtt{supers}(p)$ can only add pairs, hence can only add labels to the result.
-Positive in $\mathtt{supers}$.
-Equation (2): $\mathtt{supers}$
-$\mathtt{supers}(p)=\bigl\{\,(\mathtt{init}(p_{\mathrm{base}}),p_{\mathrm{override}})\bigm|p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{override}}\in\mathtt{overrides}(p_{\mathrm{base}})\,\bigr\}$.
-Both $\mathtt{bases}^{*}$ and $\mathtt{overrides}$ appear as sets being iterated over.
-As noted above, $\mathtt{bases}^{*}$ is monotone in $\mathtt{bases}$.
-Enlarging either $\mathtt{bases}$ or $\mathtt{overrides}$ can only add elements to the result.
-Positive in $\mathtt{bases}$ and $\mathtt{overrides}$.
-Equation (3): $\mathtt{overrides}$
-For $p\neq()$:
-$\mathtt{overrides}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|{}\\ (\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
-The function $\mathtt{supers}$ appears as the set being iterated over.
-The filtering condition $\mathtt{last}(p)\in\mathtt{defines}(\cdot)$ involves only the AST primitive.
-Enlarging $\mathtt{supers}$ can only add branches, hence can only add paths to the result.
-Positive in $\mathtt{supers}$.
-Equation (4): $\mathtt{bases}$
-$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{override}}\in\mathtt{overrides}(p),\\ (n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{override}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
-Both $\mathtt{overrides}$ and $\mathtt{resolve}$ appear as sets being iterated over.
-The function $\mathtt{inherits}$ is an AST primitive.
-Enlarging either $\mathtt{overrides}$ or $\mathtt{resolve}$ can only add targets to the result.
-Positive in $\mathtt{overrides}$ and $\mathtt{resolve}$.
-Equation (5): $\mathtt{resolve}$
-$\mathtt{resolve}(\ldots)=\bigl\{\,p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\bigm|p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\mathtt{init}(p_{\mathrm{def}}),n)\,\bigr\}.$
-The function $\mathtt{this}$ appears as the set being iterated over. Enlarging $\mathtt{this}$ can only add paths to the result. Positive in $\mathtt{this}$.
-Equation (6): $\mathtt{this}$
-For $n>0$: $\mathtt{this}(S,p_{\mathrm{def}},n)=\mathtt{this}(S^{\prime},\mathtt{init}(p_{\mathrm{def}}),n-1)$ where $S^{\prime}=\bigl\{\,p_{\mathrm{site}}\bigm|p_{\mathrm{current}}\in S,\;(p_{\mathrm{site}},p_{\mathrm{override}})\in\mathtt{supers}(p_{\mathrm{current}}),\;p_{\mathrm{override}}=p_{\mathrm{def}}\,\bigr\}$.
-The function $\mathtt{supers}$ appears as the set being iterated over.
-The filtering condition $p_{\mathrm{override}}=p_{\mathrm{def}}$ is an equality test on path data, not involving any mutually defined function.
-Enlarging $\mathtt{supers}$ can only enlarge $S^{\prime}$, and by induction on $n$ (the base case $n=0$ returns $S$ unchanged), the recursive call is monotone in $S^{\prime}$.
-Positive in $\mathtt{supers}$.
-In summary, every mutually defined function appears on the right-hand side of every equation only inside the pattern $x\in f(\ldots)$ within a set comprehension, i.e., as the domain of an existential quantifier.
-No complement, set difference, or negated membership test appears in any equation.
-Therefore $F$ is a monotone operator on the complete lattice $D$.
-∎
-C.4. Existence of the Least Fixed Point
-Theorem C.2.
-The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined as the least fixed point of $F$.
-Proof.
-By Proposition C.1, $F$ is monotone on the complete lattice $D$.
-By the Knaster–Tarski theorem (Tarski, 1955), $F$ has a least fixed point $\vec{d}^{*}=\bigsqcup_{n\geq 0}F^{n}(\bot)$, where $\bot$ assigns the empty set to every input.
-The components of $\vec{d}^{*}$ are the semantic functions of Section 3.
-Concretely, the least fixed point is the union of the finite approximation chain: $\bot\sqsubseteq F(\bot)\sqsubseteq F^{2}(\bot)\sqsubseteq\cdots$.
-A membership $\ell\in\mathtt{properties}(p)$ holds in $\vec{d}^{*}$ if and only if it appears at some finite stage $F^{k}(\bot)$, which corresponds to a finite derivation tree using the equations.
-When no finite derivation exists, the membership is absent from the least fixed point and the query is undefined, the analogue of divergence.
-∎
-Appendix D Böhm Tree Correspondence: Proofs
-This appendix contains the full proofs of Adequacy (Theorem 4.2) and Full Abstraction (Theorem 4.3) stated in Section 4.1.
-D.1. Böhm Trees
-We briefly recall the definition of Böhm trees (Barendregt, 1984).
-A head reduction $M\to_{h}M^{\prime}$ contracts the outermost $\beta$-redex only: if $M$ has the form $(\lambda x.\,e)\;v\;M_{1}\cdots M_{k}$, then $M\to_{h}e[v/x]\;M_{1}\cdots M_{k}$.
-A term $M$ is in head normal form (HNF) if it has no head redex, i.e., $M=\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ where $y$ is a variable.
-The Böhm tree $\mathrm{BT}(M)$ of a $\lambda$-term $M$ is an infinite labeled tree defined by:
-$\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\begin{aligned} &\text{if }M\to^{*}_{h}\\[-2.0pt] &\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{aligned}\end{cases}$
-Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993).
-D.2. Path Encoding
-We define a correspondence between positions in the Böhm tree and paths in the mixin tree.
-When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting mixin tree has a specific shape:
-  • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$ (the abstraction shape).
-  • A $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ translates to a record with own properties $\{x,\mathrm{result}\}$, where $x$ holds the encapsulated application $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$.
-  • A tail call $V_{1}\;V_{2}$ translates to a record with own properties $\{\_\mathrm{tail},\mathrm{result}\}$, where $\mathrm{tail}$ holds the encapsulated application and $\mathrm{result}$ projects $\mathrm{tail}.\mathrm{result}$.
-  • A variable $x$ with de Bruijn index $n$ translates to an indexed reference $\uparrow^{\!n}.\mathrm{argument}$.
-Each Böhm tree position is a sequence of navigation steps from the root.
-In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body (peeling off one abstraction layer) and entering the $i$-th argument $M_{i}$ of the head variable.
-In the mixin tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
-Rather than comparing absolute paths across different mixin trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection.
-Recall the definition of inheritance-convergence (Definition 4.1).
-Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation.
-The abstraction shape at the end signals that a weak head normal form has been reached.
-D.3. Single-Path Lemma
-The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path mixin inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
-Lemma D.1 (Single Path).
-For any closed ANF $\lambda$-term $M$, at every invocation of $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)$ during evaluation of $\mathcal{T}(M)$, the frontier set $S$ contains exactly one path.
-Proof.
-By induction on the structure of the ANF term.
-Case $\mathcal{T}(\lambda x.\,M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
-This is a record literal with $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$ and no inheritance sources, so $\mathtt{inherits}=\varnothing$.
-At the root path $p$, we have $\mathtt{overrides}(p)=\{p\}$ (a single path) since there are no inheritance sources to introduce additional branches.
-Inside $\mathcal{T}(M)$, any reference to $x$ is $\uparrow^{\!n}.\mathrm{argument}$ where $n$ is the de Bruijn index pointing to this scope level.
-Since $\mathcal{T}(\lambda x.\,M)$ introduces exactly one scope level, there is exactly one mixin at that level, hence $\mathtt{this}$ finds exactly one matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ in $\mathtt{supers}$.
-Case $\mathcal{T}(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
-The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level ($\mathtt{inherits}=\varnothing$), so $\mathtt{overrides}(\text{root})=\{\text{root}\}$.
-Inside the $x$ subtree, the inheritance $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$ has exactly one inheritance source ($\mathcal{T}(V_{1})$), creating exactly one $\mathtt{bases}$ entry.
-By induction on $V_{1}$, the single-path property holds inside $\mathcal{T}(V_{1})$.
-Inside $\mathcal{T}(M)$, the induction hypothesis applies directly.
-Case $\mathcal{T}(V_{1}\;V_{2})=\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
-Identical to the $\mathbf{let}$-binding case: the outer record has own properties $\{\_\mathrm{tail},\mathrm{result}\}$ with no inheritance sources at the outer level, and the $\mathrm{tail}$ subtree contains exactly one inheritance source.
-Case $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$
-This is a reference, not a record.
-The $\mathtt{this}$ function walks up $n$ steps from the enclosing scope.
-By the inductive hypothesis on the enclosing translations, each step of $\mathtt{this}$ encounters exactly one matching pair in $\mathtt{supers}$.
-The key structural invariant is that the translation $\mathcal{T}$ never produces two independent inheritance sources at the same scope level that both define the same label.
-In every case, application inheritances are encapsulated inside a named property ($x$ or $\mathrm{tail}$), and each such encapsulated inheritance has exactly one inheritance source.
-This prevents diamond inheritance and ensures single-path resolution throughout.
-∎
-D.4. Substitution Lemma
-The core mechanism of the translation is that inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ plays the role of substitution. We make this precise.
-Lemma D.2 (Substitution).
-Let $M$ be an ANF term in which $x$ may occur free, and let $V$ be a closed value.
-Define the inherited tree $C=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$.
-Then for every path $p$ under $C\mathbin{\triangleright}\mathrm{result}$ and every label $\ell$:
-$\ell\in\mathtt{properties}(p)\text{ in }C\mathbin{\triangleright}\mathrm{result}\quad\Longleftrightarrow\quad\ell\in\mathtt{properties}(p)\text{ in }\mathcal{T}(M[V/x])$
-where $M[V/x]$ is the usual capture-avoiding substitution.
-Proof.
-By structural induction on $M$.
-Base case: $M=x$
-Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference.
-In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ inherits into the $\mathrm{argument}$ slot, providing $\mathcal{T}(V)$.
-Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma D.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree.
-Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
-Base case: $M=y$ where $y\neq x$
-Then $\uparrow^{\!m}.\mathrm{argument}$ (where $m$ is the de Bruijn index of $y$, $m\neq n$) walks up to a different scope level, so the inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ at the binding scope of $x$ has no effect.
-Since $M[V/x]=y$, the properties are identical.
-Inductive case: $M=\lambda y.\,M^{\prime}$
-Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$.
-The subtree at $C\mathbin{\triangleright}\mathrm{result}$ is another abstraction-shaped record.
-References to $x$ inside $\mathcal{T}(M^{\prime})$ resolve through $\mathtt{this}$ in exactly the same way (with one additional scope level from the inner $\lambda$), and by induction on $M^{\prime}$, the properties under $C\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M^{\prime}[V/x])$.
-Since $(\lambda y.\,M^{\prime})[V/x]=\lambda y.\,(M^{\prime}[V/x])$ (assuming $y$ is fresh), the result follows.
-Inductive case: $M=\mathbf{let}\;z=V_{1}\;V_{2}$ $\mathbf{in}\;M^{\prime}$
-Then
-$\mathcal{T}(M)=\{z\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}.$
-The substitution distributes: $M[V/x]=\mathbf{let}\;z=V_{1}[V/x]\;V_{2}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$.
-By induction on $V_{1}$, $V_{2}$, and $M^{\prime}$, the properties under each subtree ($z$ and $\mathrm{result}$) coincide, since references to $x$ inside each subtree resolve to $\mathcal{T}(V)$ via the same $\mathtt{this}$ mechanism.
-Inductive case: $M=V_{1}\;V_{2}$ (tail call)
-Analogous to the $\mathbf{let}$-binding case with $\mathrm{tail}$ in place of $z$. ∎
-D.5. Convergence Preservation
-The convergence criterion (Definition 4.1) follows the $\mathrm{result}$ chain from the root.
-Each step in this chain corresponds to one head reduction step in the $\lambda$-calculus.
-We make this precise.
-Lemma D.3 (Result-step).
-Let $(\lambda x.\,M)\;V$ be a tail call (a $\beta$-redex). Then for every path $p$ and label $\ell$:
-$\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)\;\Longleftrightarrow\;\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}(M[V/x])$
-That is, following one $\mathrm{result}$ projection in the tail call’s mixin tree yields the same properties as the root of the reduct’s mixin tree.
-Proof.
-The translation gives:
-$\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)=\{\_\mathrm{tail}\mapsto\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
-The $\mathrm{result}$ label at the root is defined as $\mathrm{tail}.\mathrm{result}$, so $\text{root}\mathbin{\triangleright}\mathrm{result}$ resolves to the $\mathrm{result}$ path inside the inheritance $\mathrm{tail}=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$.
-This inheritance is exactly the inherited tree $C$ of Lemma D.2, and $\mathrm{tail}\mathbin{\triangleright}\mathrm{result}$ is $C\mathbin{\triangleright}\mathrm{result}$.
-By Lemma D.2, the properties under $C\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M[V/x])$.
-Let-binding variant. An analogous result holds for $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$ where $V_{1}=\lambda y.\,M^{\prime\prime}$. The translation gives:
-$\mathcal{T}\bigl(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\bigr)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$
-Since $\mathrm{result}\mapsto\mathcal{T}(M^{\prime})$ is a property definition, the subtree at $\text{root}\mathbin{\triangleright}\mathrm{result}$ is $\mathcal{T}(M^{\prime})$.
-Inside $\mathcal{T}(M^{\prime})$, each reference to $x$ is $x.\mathrm{result}$, which resolves to the $\mathrm{result}$ path inside the $x$ subtree $\{\mathcal{T}(\lambda y.\,M^{\prime\prime}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$.
-This is the inherited tree $C$ of Lemma D.2, so $x.\mathrm{result}$ has the same properties as $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$.
-In the $\lambda$-calculus, $M=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$, and the head reduct is $N=M^{\prime}[(V_{1}\;V_{2})/x]$.
-In $\mathcal{T}(N)$, each occurrence of $x$ is replaced by $(V_{1}\;V_{2})$, whose result, by the tail-call variant above, also yields $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$.
-Therefore the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$ for all $p$.
-∎
-Theorem D.4 (Convergence preservation).
-If $M\to_{h}N$ (head reduction), then $\mathcal{T}(M){\Downarrow}$ if and only if $\mathcal{T}(N){\Downarrow}$.
-Proof.
-A head reduction step in ANF takes one of two forms: a tail call $(\lambda x.\,M^{\prime})\;V\to_{h}M^{\prime}[V/x]$, or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[(V_{1}\;V_{2})/x]$ (since $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$ in the $\lambda$-calculus).
-In both cases, Lemma D.3 (tail-call and let-binding variants respectively) shows that the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n}$ in $\mathcal{T}(M)$ coincide with the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n-1}$ in $\mathcal{T}(N)$ for all $n\geq 1$, where $N$ is the head reduct.
-Hence the abstraction shape appears at depth $n$ in the pre-reduct if and only if it appears at depth n1n-1 in the post-reduct.
-∎
-D.6. Adequacy
-A closed ANF term at the top level is either an abstraction (a value) or a computation (a $\mathbf{let}$-binding or tail call).
-An abstraction translates to a record whose root immediately has the abstraction shape; a computation’s root has the form $\{x,\mathrm{result}\}$ or $\{\_\mathrm{tail},\mathrm{result}\}$, and one must follow the $\mathrm{result}$ chain to find the eventual value.
-Proof of Theorem 4.2.
-($\Rightarrow$) Suppose $M\to^{*}_{h}\lambda x.\,M^{\prime}$ in $k$ head reduction steps. We show $\mathcal{T}(M)$ converges at depth $\leq k$ by induction on $k$.
-Base case ($k=0$): $M$ is already an abstraction $\lambda x.\,M^{\prime}$.
-Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$, whose root has $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$.
-These labels appear at stage $F^{1}(\bot)$ of the approximation chain, so $\mathcal{T}(M)$ converges at depth $n=0$.
-Inductive step ($k\geq 1$): $M$ is not an abstraction, so $M$ is either a tail call or a $\mathbf{let}$-binding, both of which are $\beta$-redexes.
-Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$.
-By Lemma D.3 (tail-call or let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$.
-By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
-($\Leftarrow$) Suppose $\mathcal{T}(M){\Downarrow}$ at depth $n$. We show $M$ has a head normal form by induction on $n$.
-Base case ($n=0$): the root of $\mathcal{T}(M)$ has abstraction shape $\{\mathrm{argument},\mathrm{result}\}$.
-Only the abstraction rule $\mathcal{T}(\lambda x.\,M^{\prime})$ produces a root with own property $\mathrm{argument}$.
-(For tail calls and $\mathbf{let}$-bindings, the root has $\mathtt{defines}=\{\_\mathrm{tail},\mathrm{result}\}$ or $\{x,\mathrm{result}\}$, and $\mathrm{argument}$ cannot appear via $\mathtt{supers}$ because $\mathtt{inherits}=\varnothing$ at the root, giving $\mathtt{bases}(\text{root})=\varnothing$.) Therefore $M$ is an abstraction, hence already in head normal form.
-Inductive step ($n\geq 1$): the abstraction shape appears at depth $n$ but not at depth 0.
-By the base case argument, $M$ is not an abstraction, so $M$ is either a tail call $V_{1}\;V_{2}$ or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$.
-In both cases, $M$ is a $\beta$-redex (since $V_{1}=\lambda x.\,M^{\prime}$ or $V_{1}=\lambda y.\,M^{\prime\prime}$ respectively, the only closed ANF value).
-By Lemma D.3 (tail-call or let-binding variant respectively), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct.
-So $\mathcal{T}(N)$ converges at depth n1n-1.
-By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$.
-∎
-D.7. Full Abstraction
-Adequacy relates a single term’s head normal form to its inheritance-convergence.
-Full abstraction lifts this to an equivalence between terms: the translation $\mathcal{T}$ preserves and reflects observational equivalence.
-Definition D.5 ($\lambda$-contextual equivalence under $\mathcal{T}$).
-For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
-$\mathcal{T}(C[M]){\Downarrow}\;\Longleftrightarrow\;\mathcal{T}(C[N]){\Downarrow}$
-The subscript $\lambda$ emphasizes that the quantification ranges over $\lambda$-calculus contexts, not over all inheritance-calculus contexts.
-This definition observes only convergence (a single bit), but quantifying over all $\lambda$-contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior within the $\lambda$-definable fragment.
-Proof of Theorem 4.3.
-The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the mixin tree $\mathcal{T}(C[M])$ is determined by the mixin tree $\mathcal{T}(M)$ and the translation of the context.
-This means $\mathcal{T}$ preserves the context structure needed for the argument.
-($\Leftarrow$) Suppose $\mathrm{BT}(M)=\mathrm{BT}(N)$.
-Böhm tree equivalence is a congruence (Barendregt, 1984), so for every context $C[\cdot]$, $\mathrm{BT}(C[M])=\mathrm{BT}(C[N])$.
-In particular, $C[M]$ has a head normal form iff $C[N]$ does.
-By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$.
-Hence $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$.
-($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$.
-Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa).
-By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$.
-Hence $\mathcal{T}(M)\not\approx_{\lambda}\mathcal{T}(N)$.
-∎
-The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993).
-The mixin tree semantics inherits full abstraction from the $\lambda$-calculus without needing to define or compare absolute paths inside mixin trees.
+Appendix E First-Order Semantics of the ANF Lazy $\lambda$-Calculus
+This appendix presents a first-order semantics of the lazy $\lambda$-calculus in A-normal form (ANF).
+The semantics is given by mutually recursive set-comprehension equations whose well-definedness follows from the same termination argument as in Appendix A.
+The central idea is that function application is modeled as tree grafting: applying a function record to an argument splices the function’s subtree into the call site, and variable references navigate the resulting tree structure.
+ANF is essential.
+Traditional $\lambda$-calculus syntax allows anonymous subexpressions such as (fx)(f\;x) in (fx)y(f\;x)\;y that have no label and cannot form paths.
+ANF restores path-addressability by binding every intermediate result to a name, so that every node in the evaluation tree has a unique path address.
+Syntax
+Let $M$ denote a computation and $V$ a value.
+[…]
+The grammar includes both application let-bindings and value let-bindings (the latter bind a name to a value without application).
+Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results; ANF is neither a restriction nor an extension of the $\lambda$-calculus.
+Variables use de Bruijn indices: $\uparrow^{\!n}$ counts the number of enclosing scope levels (record braces) between the reference and its binder, following the same convention as inheritance-calculus.
+In particular, a $\lambda$-bound variable that crosses $k$ intervening $\mathbf{let}$-binding scopes and $n$ intervening $\lambda$-abstraction scopes is written $\uparrow^{\!n+k}$, not $\uparrow^{\!n}$.
+Let-bound variables are referenced by their binding name $x$.
+Evaluation tree
+An evaluation tree has a single abstraction: every node can serve as function, argument, binding, and result at the same time.
+A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ldots,\ell_{k})$ identifying a position in the mixin tree.
+We write $p$ for a path, $()$ for the root, $p\mathbin{\triangleright}\ell$ for extension, $\mathtt{init}(p)$ for the parent path, and $\mathtt{last}(p)$ for the final label.
+AST
+An ANF term is parsed into an AST that provides two functions at each path $p$:
+  • $\mathtt{children}(p)$: the set of labels locally defined at $p$.
+  • $\mathtt{refs}(p)$: the set of reference pairs $(n,\;\ell_{*})$, where $n$ is a de Bruijn index and $\ell_{*}$ is a list of downward projections.
+Each ANF construct populates these as follows.
+An abstraction $\lambda x.\,M$ at path $p$:
+$\mathtt{children}(p)$ $=\{\mathrm{argument},\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
+The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$; the subtree at $p\mathbin{\triangleright}\mathrm{argument}$ is empty ($\mathtt{children}=\varnothing$, $\mathtt{refs}=\varnothing$).
+A let-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ at path $p$:
+$\mathtt{children}(p)$ $=\{x,\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
+The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$.
+The subtree at $p\mathbin{\triangleright}x$ represents the application $V_{1}\;V_{2}$ and is determined by the form of $V_{1}$:
+  • If $V_{1}$ is a variable (either $\uparrow^{\!m}$ or a let-bound name), then
+    $\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{argument}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\{(n_{1},\;\ell_{*})\},$
+    where the reference $(n_{1},\;\ell_{*})$ encodes the translation of $V_{1}$ (see below).
+  • If $V_{1}=\lambda y.\,M^{\prime}$, the $\lambda$-literal is inlined:
+    $\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{argument},\;\mathrm{result}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\varnothing,$
+    and the subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{result}$ is determined by $M^{\prime}$.
+In both cases the subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{argument}$ is determined by $V_{2}$.
+A value let-binding $\mathbf{let}\;x=V\;\mathbf{in}\;M$ at path $p$:
+$\mathtt{children}(p)$ $=\{x,\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
+The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$. The subtree at $p\mathbin{\triangleright}x$ wraps the value in a $\mathrm{result}$ scope:
+$\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{result}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\varnothing.$
+The subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{result}$ is determined by $V$.
+The accessor $x.\mathrm{result}$ thus yields the translated value, matching the convention for application let-bindings.
+A tail call $V_{1}\;V_{2}$ at path $p$:
+$\mathtt{children}(p)$ $=\{\mathrm{tailCall},\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
+The subtree at $p\mathbin{\triangleright}\mathrm{tailCall}$ is analogous to the application node $p\mathbin{\triangleright}x$ above: if $V_{1}$ is a variable, $\mathtt{children}=\{\mathrm{argument}\}$ with $\mathtt{refs}=\{(n_{1},\;\ell_{*})\}$; if $V_{1}$ is a $\lambda$-literal, $\mathtt{children}=\{\mathrm{argument},\;\mathrm{result}\}$ with $\mathtt{refs}=\varnothing$ and the $\lambda$-body inlined.
+The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is a forwarding reference:
+$\mathtt{children}(p\mathbin{\triangleright}\mathrm{result})$ $=\varnothing,$ $\mathtt{refs}(p\mathbin{\triangleright}\mathrm{result})$ $=\{(0,\;(\mathrm{tailCall},\;\mathrm{result}))\}.$
+Here $(0,\;(\mathrm{tailCall},\;\mathrm{result}))$ navigates to the enclosing scope ($p$ itself, since $n=0$ from $p\mathbin{\triangleright}\mathrm{result}$’s perspective), then projects $\mathrm{tailCall}.\mathrm{result}$.
+A $\lambda$-bound variable $\uparrow^{\!n}$ used as an inheritance source at an application node $p$ contributes $(n,\;(\mathrm{argument}))$ to $\mathtt{refs}(p)$: navigate up $n$ scope levels, then project $\mathrm{argument}$.
+The same applies when $\uparrow^{\!n}$ occurs as a bare value in tail position or as an argument value, contributing the reference to $\mathtt{refs}$ at the corresponding path.
+A let-bound variable $x$ at path $p$ is resolved lexically: find the nearest ancestor $p^{\prime}$ of $p$ such that $x\in\mathtt{children}(p^{\prime})$, let $n=|p|-|p^{\prime}|-1$, and contribute $(n,\;(x,\;\mathrm{result}))$ to $\mathtt{refs}(p)$, projecting $x.\mathrm{result}$ from the enclosing scope.
+Note that since the de Bruijn index $n$ counts all scope levels (not just $\lambda$-abstractions), it accounts for intervening $\mathbf{let}$-binding and tail-call scopes.
+For example, in $\lambda.\,\lambda.\,\mathbf{let}\;r=\uparrow^{\!2}\;\uparrow^{\!0}\;\mathbf{in}\;r$, the outer-$\lambda$ variable has de Bruijn index $2$ (crossing the $\mathbf{let}$-scope and the inner-$\lambda$ scope), not $1$.
+Both $\mathtt{children}$ and $\mathtt{refs}$ are pure data extracted from the AST; they are not functions of runtime state.
+Labels
+The labels observable at a path are those locally defined at any graft of any transitive callee:
+(7) $\mathtt{labels}(p)=\bigl\{\;\ell\;\big|\;p_{\mathrm{step}}\in\mathtt{callee}^{*}(p),\;p_{\mathrm{graft}}\in\mathtt{grafts}(p_{\mathrm{step}}),\;\ell\in\mathtt{children}(p_{\mathrm{graft}})\;\bigr\}$
+Grafts
+When a function is applied, its subtree is grafted at the call site.
+A child label $\ell$ of the callee also becomes a child of the caller; the child at the callee’s $\ell$ position and the child at the caller’s $\ell$ position (if any) share the same identity.
+The grafts of $p$ collect all such identity-sharing positions:
+(8) $\mathtt{grafts}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{callee}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &p_{\mathrm{callee}}\in\mathtt{callee}^{*}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{children}(p_{\mathrm{callee}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
+Here $\mathtt{callee}^{*}$ denotes the reflexive-transitive closure of $\mathtt{callee}$, defined next.
+Callee
+The callee of a path $p$ is the set of paths that $p$ is grafted from. Concretely, $\mathtt{callee}$ resolves every reference in $p$’s grafts one step:
+[…]
+where $\ell_{*}=(\ell_{1},\ldots,\ell_{k})$.
+The $\mathtt{init}(p_{\mathrm{graft}})$ passes the reference’s enclosing scope to $\mathtt{scope}$, matching the convention of reference resolution (5), which calls $\mathtt{this}(\{p_{\mathrm{site}}\},\;\mathtt{init}(p_{\mathrm{def}}),\;n)$.
+Scope resolution
+$\mathtt{scope}$ navigates upward through the evaluation tree to find the binding site of a de Bruijn reference.
+Each ANF construct ($\lambda$-abstraction, $\mathbf{let}$-binding, or tail call) introduces one scope level; crossing a scope boundary means finding the call site that grafted the current subtree.
+$\mathtt{scope}$ takes three arguments: an evaluation-site path $p_{\mathrm{site}}$ (where the grafted subtree lives), a definition-site path $p_{\mathrm{def}}$ (the position in the original AST), and a de Bruijn index $n$.
+(10) $\mathtt{scope}(p_{\mathrm{site}},\;p_{\mathrm{def}},\;n)=\begin{cases}p_{\mathrm{site}}&\text{if }n=0\\[6.0pt] \mathtt{scope}\!\left(p_{\mathrm{caller}},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
+where $p_{\mathrm{caller}}$ is the unique path satisfying $(p_{\mathrm{caller}},\;p_{\mathrm{def}})\in\mathtt{callee}^{\mathrm{ctx}}(p_{\mathrm{site}})$, and $\mathtt{callee}^{\mathrm{ctx}}$ pairs each element of $\mathtt{callee}^{*}(p_{\mathrm{site}})$ with its provenance:
+(11) $\mathtt{callee}^{\mathrm{ctx}}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{step}}),\;p_{\mathrm{graft}})\;\big|\;p_{\mathrm{step}}\in\mathtt{callee}^{*}(p),\;p_{\mathrm{graft}}\in\mathtt{grafts}(p_{\mathrm{step}})\;\bigr\}$
+At each step, $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a nonnegative integer, the recursion terminates.
+This completes the chain of definitions needed to compute $\mathtt{labels}(p)$.
+Observational fixpoint semantics
+The definitions above constitute the complete semantics.
+Computation arises from querying progressively deeper paths in a lazily constructed tree: the observer drives computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand.
+Formally, the mutually recursive equations (7)–(11) define set membership recursively: $\ell\in\mathtt{labels}(p)$ holds if and only if it can be established by a finite chain of applications of these equations.
+Well-definedness follows from the same termination argument as for the mixin tree equations (Appendix A): if the recursive evaluation terminates, the reachable dependency subgraph is finite and acyclic, and the equations have a unique solution.
+When no finite derivation exists for a given query, the recursive evaluation does not terminate, analogous to divergence.
+Convergence
+An ANF term $M$ converges, written $M{\Downarrow}$, if there exists $n\geq 0$ such that
+$\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{labels}\!\bigl(\,\underbrace{\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
+in the semantics defined by the equations above.
+This corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: $M$ converges iff it has a head normal form (Theorem 4.2), and two terms are equivalent under this semantics iff they have the same Böhm tree (Theorem 4.3).
+The proofs, given in Appendix B, apply verbatim: each equation above is the specialization of the corresponding mixin tree equation (Section 3) to terms in the image of the translation $\mathcal{T}$ (Section 4), where the Single-Path Lemma (Lemma B.1) ensures that the set-valued $\mathtt{this}$ function (equation 6) degenerates to the single-valued $\mathtt{scope}$ (equation 10).
+Appendix F Encoding Datalog in Inheritance-Calculus
+Section 6.1 observed that inheritance-calculus natively produces the relational semantics of logic programming.
+This appendix makes that observation precise by giving a systematic encoding of Datalog into inheritance-calculus.
+The encoding uses no extensions to the calculus; it is expressible in the three primitives of Section 2.
+All examples in this appendix have executable MIXINv2 implementations (Yang, 2026).
+There are three ingredients to keep separate.
+First, relations are stored as tries, so tuple lookup is indexed by prefixes rather than by scanning a flat set of facts.
+Second, the control skeleton is written in the visitor-pattern style already used in Section 5; from a functional-programming viewpoint, this is the same role played by Scott-encoded case analysis.
+Third, inheritance union and tabled evaluation give the resulting program its relational, least-fixed-point semantics.
+The appendix proceeds in that order: data representation first, then control flow, then recursion.
+F.1. Relations as Tries
+A binary relation over a finite domain $D=\{a,b,c,\ldots\}$ is represented as a trie: a mixin tree whose paths encode tuples.
+The trie is built from three base scopes: $\mathrm{Relation}$ (abstract base), $\mathrm{Cons}$ (non-empty node), and $\mathrm{Nil}$ (terminator).
+Each constant $d\in D$ inherits from $\mathrm{Cons}$ and carries a $\mathrm{TailMap}$ that provides per-constant sub-tries:
+$\mathrm{Relation}\mapsto\{\},\quad\mathrm{Cons}\mapsto\{\mathrm{Relation},\;\mathrm{Tail}\mapsto\{\mathrm{Relation}\}\},\quad\mathrm{Nil}\mapsto\{\mathrm{Relation}\}$
+$d\mapsto\{\mathrm{Cons},\;\mathrm{Tail}\mapsto\{\},\;\mathrm{TailMap}\mapsto\{d\mapsto\{\mathrm{Tail}\}\}\}$
+A fact $R(d_{1},d_{2})$ is encoded as a trie path of depth 2, with a $\mathrm{Nil}$ terminator. The per-constant sub-tries are accessed via $\mathrm{TailMap}.d$:
+$R(a,b)\;\leadsto\quad R\mapsto\{a,\;\mathrm{TailMap}\mapsto\{a\mapsto\{b,\;\mathrm{TailMap}\mapsto\{b\mapsto\{\mathrm{Nil}\}\}\}\}\}$
+Multiple facts for the same relation are accumulated by inheritance (trie union). Adding $R(b,c)$:
+$R\mapsto\{a,\;b,\;\mathrm{TailMap}\mapsto\{$ $a\mapsto\{b,\;\mathrm{TailMap}\mapsto\{b\mapsto\{\mathrm{Nil}\}\}\},$
+$b\mapsto\{c,\;\mathrm{TailMap}\mapsto\{c\mapsto\{\mathrm{Nil}\}\}\}\}\}$
+The point of the trie representation is not merely that it is a tree-shaped encoding.
+It also provides the indexing structure used by the subsequent joins: once a constant $d$ has been selected, the encoding moves directly to $\mathrm{TailMap}.d$ and continues from the matching sub-trie.
+This is why the appendix speaks about tries rather than about an arbitrary encoding of finite sets of tuples.
+F.2. Infrastructure
+Each constant $d$ carries four groups of boilerplate definitions used by the encoding.
+These play a role analogous to derived type-class instances in Haskell: each constant must provide a fixed set of properties that the encoding’s generic patterns rely on.
+Inheritance-calculus has no built-in derivation mechanism, so the boilerplate is written out manually; however, the amount of boilerplate per constant is $O(1)$ and does not grow with the domain size $|D|$, so adding a new constant does not require modifying any existing scope, preserving the expression-problem property.
+Acceptance:
+The visitor-pattern infrastructure from Section 5.
+Operationally, this is the object-oriented spelling of a case split on the current constant.
+Each constant $d$ defines a $\mathrm{VisitorMap}$ containing a branch keyed by $d$, a $\mathrm{Visitor}$ with a $\mathrm{Visited}$ slot for per-constant results, and an $\mathrm{Accepted}$ property that delegates to the matching branch’s $\mathrm{Visited}$ value.
+This is the dispatch mechanism: given a scope that may contain multiple constants, $\mathrm{Acceptance}$ selects the branch corresponding to each constant present and collects the per-constant $\mathrm{Visited}$ results.
+WithVisitorTail:
+Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisitorTail}$.
+This provides the right-hand (visitor) side of a join: when two relations’ $\mathrm{VisitorMap}$ entries are composed, $\mathrm{VisitorTail}$ carries the joined relation’s row data for the matched constant.
+WithVisiteeTail:
+Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisiteeTail}$.
+This provides the left-hand (visitee) side: the relation being iterated.
+Together, $\mathrm{WithVisitorTail}$ and $\mathrm{WithVisiteeTail}$ enable value-level joins where only constants present in both relations produce output.
+Replacement:
+Given a $\mathrm{NewTail}$ value, produces a $\mathrm{Replaced}$ scope that inherits the label $d$ but substitutes $\mathrm{TailMap}.d$ with $\mathrm{NewTail}$.
+This is the output construction mechanism: it builds output tuples by substituting new sub-tries into constant wrappers.
+Crucially, $\mathrm{Replacement}$ is defined on the per-constant $\mathrm{TailMap}$ entry (not on the merged scope), ensuring that output construction operates on individual constants rather than their union.
+As a running example, we use the executable two-hop encoding $\mathrm{twoHop}(X,Z)\mathrel{{:}\!{-}}\mathrm{edge}(X,Y),\;\mathrm{edge}(Y,Z)$ from RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026).
+Its top-level scopes are named after the stages of the encoding: _XAcceptance iterates the first column, _YAcceptance performs the shared-variable join, _NilAcceptance0 confirms the end of the first fact, _ZAcceptance iterates the output column, and _NilAcceptance1 triggers the final Replacement chain.
+The schematic encodings below should be read as abstractions of that executable MIXINv2 file.
+If one temporarily ignores relational accumulation and reads each dispatch as an ordinary case split, the control flow is: choose $X$ from the first relation, choose a matching $Y$, confirm that the first tuple is complete, choose $Z$ from the second relation’s matching row, confirm that the second tuple is complete, and emit $(X,Z)$.
+The rest of the appendix explains how that familiar single-valued control skeleton is realised by visitor dispatch over tries and then lifted to relational semantics by inheritance union.
+F.3. Encoding Rules
+A Datalog rule $H(\bar{X})\mathrel{{:}\!{-}}B_{1}(\bar{Y}_{1}),\ldots,B_{n}(\bar{Y}_{n})$ is encoded by nesting three mechanisms inside that control skeleton:
+  1. (1)
+    Column iteration.
+For each column position, an $\mathrm{Acceptance}$ dispatch iterates over the constants present.
+The $\mathrm{VisitorMap}$ is populated from $\mathrm{WithVisiteeTail}$, so each matched constant’s $\mathrm{Visitor}$ receives $\mathrm{VisiteeTail}$: the per-constant sub-trie at that position.
+The result of processing each constant is placed in $\mathrm{Visited}$, and $\mathrm{Accepted}$ collects all per-constant results.
+(Details in Sections F.4, F.5, and F.6.)
+  2. (2)
+    Join.
+For each shared variable between body atoms, the inner $\mathrm{Acceptance}$’s $\mathrm{VisitorMap}$ inherits from both $\mathrm{WithVisitorTail}$ (from the joined relation) and $\mathrm{WithVisiteeTail}$ (from the iterated relation).
+A constant $d$ produces a $\mathrm{Visitor}$ entry only if $d$ appears in both relations, achieving a value-level join.
+The $\mathrm{Visitor}$ then receives both $\mathrm{VisiteeTail}$ and $\mathrm{VisitorTail}$, providing access to both relations’ data for the matched constant.
+  3. (3)
+    Output construction.
+Inside the innermost $\mathrm{Nil}$ confirmation (verifying that each body atom’s tuple is complete), chain $\mathrm{Replacement}$ operations from the innermost output column outward, each setting $\mathrm{NewTail}$ to the previous $\mathrm{Replaced}$ value, terminated by $\mathrm{Nil}$.
+The final $\mathrm{Replaced}$ value is assigned to $\mathrm{Visited}$.
+F.4. Variable Join
+Consider the simplest join: $\mathrm{composed}(X,Z)\mathrel{{:}\!{-}}\alpha(X,Y),\;\beta(Y,Z)$.
+The executable self-join instance of this pattern is RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026), whose scope names are shown in parentheses below.
+This is the right place to see the division of labour from the section introduction: visitor dispatch provides the control flow, the trie provides direct access to the matching tails, and $\mathtt{supers}$ accumulates the outputs produced by each successful match.
+The shared variable $Y$ requires matching constants between $\alpha$’s second column and $\beta$’s first column.
+The encoding nests five $\mathrm{Acceptance}$ dispatches: iterate $X$ from $\alpha$, join on $Y$ between $\alpha$ and $\beta$, confirm $\mathrm{Nil}$ after $Y$ in $\alpha$, iterate $Z$ from $\beta$’s matched row, and confirm $\mathrm{Nil}$ after $Z$.
+We present each level as a named sub-scope.
+Level 1: iterate $X$ from $\alpha$
+(Executable scope: _XAcceptance.)
+$\mathrm{xAcc}\mapsto\{\alpha.\mathrm{Acceptance},\;\mathrm{VisitorMap}\mapsto\{\alpha.\mathrm{WithVisiteeTail}\},$
+$\qquad\mathrm{Visitor}\mapsto\{\_X\mapsto\{\mathrm{VisiteeTail}\},\;\ldots,\;\mathrm{Visited}\mapsto\{\mathrm{yAcc}.\mathrm{Accepted}\}\}\}$
+Level 2: join $\alpha$’s $Y$-column with $\beta$’s first column
+(Executable scope: _YAcceptance.)
+The $\mathrm{VisitorMap}$ inherits from both $\beta.\mathrm{WithVisitorTail}$ and $X.\mathrm{WithVisiteeTail}$, producing an entry for constant $d$ only if $d$ appears in both $\alpha$’s second column (via _X\_X) and $\beta$’s first column.
+$\mathrm{yAcc}\mapsto\{\_X.\mathrm{Acceptance},$
+$\qquad\mathrm{VisitorMap}\mapsto\{\beta.\mathrm{WithVisitorTail},\;\_X.\mathrm{WithVisiteeTail}\},$
+$\qquad\mathrm{Visitor}\mapsto\{\_Y_{0}\mapsto\{\mathrm{VisiteeTail}\},\;\mathrm{VisitorTail}\mapsto\{\beta.\mathrm{Tail}\},$
+$\hskip 18.49988pt\ldots,\;\mathrm{Visited}\mapsto\{\mathrm{nilAcc}_{0}.\mathrm{Accepted}\}\}\}$
+Level 3: confirm $\mathrm{Nil}$ after $Y$ in $\alpha$
+(Executable scope: _NilAcceptance0.)
+This checks that $\alpha(X,Y)$ is a complete two-column fact (its trie path terminates at $\mathrm{Nil}$ after $Y$).
+Inside the $\mathrm{Nil}$ branch, the encoding proceeds to iterate $\beta$’s second column via $\mathrm{VisitorTail}$:
+$\mathrm{nilAcc}_{0}\mapsto\{\_Y_{0}.\mathrm{Acceptance},\;\mathrm{VisitorMap}\mapsto\{\mathrm{Nil}\mapsto\{\ldots,\;\mathrm{Visited}\mapsto\{\mathrm{zAcc}.\mathrm{Accepted}\}\}\}\}$
+Level 4: iterate $Z$ from $\beta$’s matched row
+(Executable scope: _ZAcceptance.)
+$\mathrm{zAcc}\mapsto\{\mathrm{VisitorTail}.\mathrm{Acceptance},\;\mathrm{VisitorMap}\mapsto\{\mathrm{VisitorTail}.\mathrm{WithVisiteeTail}\},$
+$\qquad\mathrm{Visitor}\mapsto\{\_Z\mapsto\{\mathrm{VisiteeTail}\},\;\ldots,\;\mathrm{Visited}\mapsto\{\mathrm{nilAcc}_{1}.\mathrm{Accepted}\}\}\}$
+Level 5: confirm $\mathrm{Nil}$ after $Z$ and construct output
+(Executable scope: _NilAcceptance1, with _ZReplacement and _XReplacement.)
+$\mathrm{nilAcc}_{1}\mapsto\{\mathrm{VisiteeTail}.\mathrm{Acceptance},\;\mathrm{VisitorMap}\mapsto\{\mathrm{Nil}\mapsto\{$
+$\qquad\mathrm{zR}\mapsto\{\_Z.\mathrm{Replacement},\;\mathrm{NewTail}\mapsto\{\mathrm{Nil}\}\},$
+$\qquad\mathrm{xR}\mapsto\{\_X.\mathrm{Replacement},\;\mathrm{NewTail}\mapsto\{\mathrm{zR}.\mathrm{Replaced}\}\},$
+$\qquad\mathrm{Visited}\mapsto\{\mathrm{xR}.\mathrm{Replaced}\}\}\}\}$
+Final result
+$\mathrm{composed}\mapsto\{\mathrm{xAcc}.\mathrm{Accepted}\}$
+Semantics
+At Level 2, the combined $\mathrm{VisitorMap}$ ensures that a constant $d$ produces a $\mathrm{Visitor}$ only when $d$ appears in both $\alpha$’s second column and $\beta$’s first column.
+The $\mathrm{Visitor}$ for each matched constant $d$ independently receives $\mathrm{VisiteeTail}$ ($\alpha$’s data for $d$) and $\mathrm{VisitorTail}$ ($\beta$’s row for $d$).
+Output construction (Level 5) operates inside this per-constant $\mathrm{Visitor}$, so $\mathrm{Replacement}$ acts on the individual constant’s data, not on the merged scope.
+Each matched constant $d$ produces its own output tuples in $\mathrm{Visited}$, and these are collected by $\mathrm{Accepted}$ via trie union ($\mathtt{supers}$).
+This yields standard Datalog tuple-level semantics: only constants that actually match between both relations contribute to the output.
+For the concrete instance $\mathrm{edge}=\{(a,b),\;(b,c),\;(c,a)\}$, as encoded in RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026), Level 1 selects $a$, Level 2 matches $b$, Level 3 confirms that $\mathrm{edge}(a,b)$ ends in $\mathrm{Nil}$, Level 4 iterates $\mathrm{edge}.\mathrm{TailMap}.b$ and selects $c$, and Level 5 confirms $\mathrm{Nil}$ again before the two $\mathrm{Replacement}$ operations reconstruct $(a,c)$.
+F.5. Constant Match and Navigation
+Constants in Datalog atoms use two mechanisms, neither of which is a join.
+They are simpler than variable joins because no value from a second relation needs to be matched against the currently chosen constant:
+Type existence match
+An atom $R(X,b)$ with a constant in the second position requires checking that constant $b$ exists in $R$’s second column at the matched row.
+This is encoded as an $\mathrm{Acceptance}$ with a hand-crafted $\mathrm{VisitorMap}$ containing only the $b$ branch (instead of composing from $\mathrm{WithVisitorTail}$, which would create branches for all constants):
+$\mathrm{VisitorMap}\mapsto\{b\mapsto\{\ldots\}\}$
+If $b$ exists in the dispatching scope, the $b$ branch fires; otherwise, $\mathrm{Accepted}$ receives no matching branch.
+Direct navigation
+An atom $R(b,Z)$ with a constant in the first position navigates directly to $b$’s sub-trie:
+$\mathrm{edgeB}_{Z}\mapsto\{R.\mathrm{TailMap}.b\}$
+This bypasses the visitor mechanism entirely, accessing the per-constant $\mathrm{TailMap}$ entry to restrict the scope to only $b$’s data.
+Example
+$\mathrm{viaB}(X,Z)\mathrel{{:}\!{-}}\mathrm{edge}(X,b),\;\mathrm{edge}(b,Z)$ combines both mechanisms.
+The executable MIXINv2 encoding is RelationalConstant.mixin.yaml (Yang, 2026).
+The encoding iterates $\mathrm{edge}$ for $X$ (Level 1), then checks for constant $b$ in $X$’s row with a hand-crafted $\mathrm{VisitorMap}$ (Level 2), navigates $\mathrm{edge}.\mathrm{TailMap}.b$ for $Z$ (Level 3), confirms $\mathrm{Nil}$ (Level 4), and constructs $(X,Z)$ via $\mathrm{Replacement}$.
+F.6. Multi-Body Rules
+Rules with $n>2$ body atoms chain additional join levels.
+Each shared variable introduces one combined $\mathrm{WithVisitorTail}$$\mathrm{WithVisiteeTail}$ join inside the previous level’s $\mathrm{Visitor}$, and the output construction is placed inside the innermost $\mathrm{Nil}$ confirmation.
+So the binary-rule encoding is the whole story structurally: larger bodies only add more nested copies of the same pattern.
+Three-body rule
+$\mathrm{chain}(X,W)\mathrel{{:}\!{-}}\alpha(X,Y),\;\beta(Y,Z),\;\gamma(Z,W)$ requires two joins: on $Y$ (between $\alpha$ and $\beta$) and on $Z$ (between $\beta$ and $\gamma$).
+The structure is: iterate $X$ from $\alpha$, join on $Y$ (getting both $\alpha$’s and $\beta$’s data), confirm $\mathrm{Nil}$ after $Y$ in $\alpha$, then inside the $\mathrm{Nil}$ branch join $\beta$’s $Z$-column with $\gamma$’s first column, iterate $W$ from $\gamma$’s matched row, confirm $\mathrm{Nil}$, and construct $(X,W)$.
+The $Z$-join is nested inside the $Y$-join’s $\mathrm{Nil}$ confirmation, ensuring that $\beta$’s data for the matched $Y$ constant (available via $\mathrm{VisitorTail}$) is used to iterate $\beta$’s second column.
+Mixed constants and variables
+$\mathrm{constJoin}(X,W)\mathrel{{:}\!{-}}\mathrm{edge}(X,Y),\;\mathrm{edge}(Y,b),\;\mathrm{edge}(b,W)$ combines a variable join on $Y$, a constant match on $b$ (hand-crafted $\mathrm{VisitorMap}$ with only the $b$ entry), and direct navigation via $\mathrm{edge}.\mathrm{TailMap}.b$ for $W$.
+The corresponding executable MIXINv2 file is RelationalConstantJoin.mixin.yaml (Yang, 2026).
+F.7. Multi-Variable Join
+When a rule shares multiple variables between two body atoms, each shared variable requires its own $\mathrm{Acceptance}$ check.
+The checks are chained: the outer $\mathrm{Acceptance}$ joins on one column pair using combined $\mathrm{WithVisitorTail}$$\mathrm{WithVisiteeTail}$, and its $\mathrm{Visitor}$ contains an inner $\mathrm{Acceptance}$ joining on another column pair.
+Example
+$\mathrm{multiJoin}(X,Y)\mathrel{{:}\!{-}}r(X,Y),\;s(X,Y)$ shares both $X$ and $Y$.
+The executable MIXINv2 encoding is RelationalMultiVariableJoin.mixin.yaml (Yang, 2026).
+The outer $\mathrm{Acceptance}$ joins $r$’s and $s$’s first columns (variable $X$) using combined $\mathrm{WithVisitorTail}$$\mathrm{WithVisiteeTail}$.
+Inside the outer $\mathrm{Visitor}$, an inner $\mathrm{Acceptance}$ joins $r$’s and $s$’s second columns (variable $Y$) similarly.
+Both joins must match for the output to be produced, and the per-constant $\mathrm{Visited}$ pattern ensures that only constants present in both relations contribute.
+F.8. Recursive Rules
+Recursive Datalog rules, such as transitive closure $\mathrm{path}(X,Y)\mathrel{{:}\!{-}}\mathrm{edge}(X,Z),\;\mathrm{path}(Z,Y)$, are encoded identically to non-recursive rules: the head relation $\mathrm{path}$ appears both as a defined scope and as a reference in the body (providing $\mathrm{WithVisitorTail}$ for the $Z$-join).
+This should not be read as ordinary self-reference of a single-valued lazy program.
+Operationally, tabled evaluation (Section 3, Appendix A) computes $\mathrm{lfp}(T_{P})$: each query into $\mathrm{path}$ triggers further evaluation through the recursive body, and the trie union ($\mathtt{supers}$) accumulates all reachable tuples.
+This corresponds exactly to Naive evaluation (Bancilhon and Ramakrishnan, 1986) of the $T_{P}$ operator (van Emden and Kowalski, 1976): each iteration applies every rule to the current fact set and adds newly derived facts.
+Because Datalog-encoded relations have finitely many constants per query, the per-query answer domain is finite, and convergence is guaranteed by the ascending chain condition on a finite lattice.
+Semi-Naive evaluation (Bancilhon and Ramakrishnan, 1986), which restricts each iteration to rules that use at least one newly derived fact, is a natural optimisation left for future work.
+F.9. Semantic Correspondence with Standard Datalog
+The encoding produces standard Datalog tuple-level semantics.
+Putting the previous subsections together: tries provide indexed tuple storage, visitor dispatch supplies the control skeleton, and per-constant $\mathrm{Replacement}$ keeps output construction local to each witness.
+The key mechanism is the per-constant $\mathrm{Visited}$ pattern: output construction ($\mathrm{Replacement}$) occurs inside each matched constant’s $\mathrm{Visitor}$ scope, operating on the individual constant’s data rather than on the merged scope.
+The $\mathrm{Accepted}$ field then collects all per-constant $\mathrm{Visited}$ results via trie union ($\mathtt{supers}$).
+For example, given $\mathrm{edge}=\{(a,b),\;(b,c),\;(c,a)\}$ (a 3-cycle), the two-hop self-join $\mathrm{twoHop}(X,Z)\mathrel{{:}\!{-}}\mathrm{edge}(X,Y),\;\mathrm{edge}(Y,Z)$ produces exactly $3$ pairs: $\{(a,c),\;(b,a),\;(c,b)\}$, matching standard Datalog.
+This is the behavior exhibited by RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026).
+When the $Y$-join matches constant $b$ (present in both $\mathrm{edge}$’s second column from $(a,b)$ and $\mathrm{edge}$’s first column), the $\mathrm{Visitor}$ for $b$ receives $\mathrm{VisitorTail}=\mathrm{edge}.\mathrm{TailMap}.b$ (containing only $c$), and $\mathrm{Replacement}$ constructs a single output tuple (a,c)(a,c)—not all of $\{a,b,c\}$.
+This value-level precision arises because each constant’s $\mathrm{Replacement}$ operates on $\mathrm{TailMap}.d$ (the per-constant sub-trie) rather than on the merged trie.
+The same point is visible in the repeated-visitor self-join regression example (Yang, 2026), where repeated visitor dispatch prevents witnesses from different constants from being merged spuriously.
+Although $\mathtt{bases}$ (equation 4) still computes cross-joins at the primitive level, the encoding routes computation through per-constant $\mathrm{TailMap}$ entries before any merging occurs, achieving the same effect as standard Datalog’s tuple-level evaluation.
+Acknowledgements.
+Claude Code, GitHub Copilot, Gemini Code Assist, and Codex were utilized to generate sections of this Work, including text, tables, graphs, code, formulas, citations, etc.

### v7

@@
 Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming.
-By unifying classes, methods, objects, and properties under a single record abstraction, the calculus models inheritance simply as set union.
+By unifying modules, classes, objects, methods, fields, and locals under a single record abstraction, the calculus models inheritance simply as set union.
 Consequently, composition is inherently commutative, idempotent, and associative, structurally eliminating the multiple-inheritance linearization problem.
-Its semantics is first-order, denotational, and fixpoint-computable.
-These three properties extend to the lazy $\lambda$-calculus, since Böhm tree equivalence (Barendregt, 1984) is fully abstract for a sublanguage of inheritance-calculus, except that recursive $\lambda$-programs that diverge under $\beta$-reduction may converge under $\mathrm{lfp}(T_{P})$.
-Inheritance-calculus is distilled from MIXINv2, a practical implementation in which the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); self-reference resolves to multiple targets; and programs are immune to the Expression Problem (Wadler, 1998).
+Its semantics is first-order (van Emden and Kowalski, 1976; Reynolds, 1972; Aczel, 1977), denotational, and computable by tabling (Tamaki and Sato, 1986), even for cyclic inheritance hierarchies.
+These three properties extend to the $\lambda$-calculus, since Böhm tree equivalence (Barendregt, 1984) is fully abstract for the first-iteration approximation of a sublanguage of inheritance-calculus.
+As a corollary, this establishes a convergence hierarchy $\text{eager}\subsetneq\text{lazy}\subsetneq\text{fixpoint}$ among $\lambda$-calculi sharing the same $\lambda$-syntax.
+Inheritance-calculus is distilled from MIXINv2, a practical implementation in which the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); $\mathtt{this}$ resolves to multiple targets; and programs are immune to nonextensibility in the sense of the Expression Problem (Wadler, 1998).
 This makes inheritance-calculus strictly more expressive than the $\lambda$-calculus in both common sense and Felleisen’s sense (Felleisen, 1991).
@@
 Declarative and configuration languages are ubiquitous in modern software engineering.
-Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or merging.
-Among these, the NixOS module system stands out: its recursive attribute set merging with fixpoint semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively.
-Yet no computational theory explains why this mechanism is so expressive.
+Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, ), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or merging.
+Among these, the NixOS module system stands out: its recursive attribute set merging with fixpoint semantics and deferred modules (NixOS Contributors, ) allows independently authored modules to extend the system configuration without modifying existing modules; in practice, the ecosystem has not been troubled by the multi-dimensional extension difficulty analogous to the Expression Problem (Wadler, 1998).
+[^1] As a rough proxy for ecosystem scale: according to Repology (Repology Contributors, 2025), nixpkgs contains over 113,000 packages, nearly three times the count of Debian or Ubuntu, with a comparable number of contributors.
+Not every package exercises the module system’s extensibility, but the same caveat applies to other distributions.
+The mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively.
+Yet no computational theory captures the extensibility that this mechanism exhibits.
 The $\lambda$-calculus serves as the foundational computational model for functional programming.
@@
 Configuration languages are inherently declarative: their values are immutable and they have no first-class functions.
+Note that a function body is closed: its internal structure cannot be extended without modifying the definition, whereas the configuration languages above all compose by extending existing structure from outside.
 At first glance, Turing completeness appears incompatible with these constraints.
@@
 RAM Machine $\checkmark$ $\times$ $\checkmark$
+Table 1. Classical computational models and the declarative gap.
 The Turing Machine and RAM Machine require mutable state; the $\lambda$-calculus requires first-class functions.
-A computational model that is Turing complete, immutable, and function-free would sit in the empty cell of this table.
+A computational model that is Turing complete, immutable, and function-free would sit in the empty cell of Table 1.
 The NixOS module system already suggests that such a model exists: its inheritance-based composition over recursive records, without explicit functions, is expressive enough to configure entire operating systems.
-Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures—initial algebras in the sense of universal algebra.
-We challenge this assumption.
-Configuration values are better understood as lazily observable, possibly infinite structures, where semantics is determined by the observer and a finite prefix suffices for any finite observation.
-The nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set.
+Most configuration languages operate on finite, well-founded structures (initial algebras in the sense of universal algebra).
+Nix is an exception: although designed specifically as a configuration language for NixOS (Dolstra, 2006), it is a complete programming language whose values are lazily observable and possibly infinite.
+The nixpkgs package collection (over 100,000 packages) behaves as a lazily observed structure, allowing any single package to be evaluated without materializing the entire set.
 Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives.
-The reduction produced MIXIN, an early implementation in Nix, and subsequently MIXINv2 (Yang, 2026), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
-In developing real programs in MIXINv2, including boolean logic, natural number arithmetic, and a web application, the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) did not arise.
-Inheritance is inherently commutative, idempotent, and associative (Section 3).
-Self-reference naturally resolves to multiple objects rather than one.
-Prior systems such as Scala and the NixOS module system assume single-object resolution and reject the multi-path situation.
-Self-referential records require no fixed-point combinator in the object language.
-The mixin tree is the call stack, and qualified this resolution is a first-order set-theoretic query over paths.
-We also found that boolean logic and arithmetic, implemented as separate MIXINv2 files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism.
-Church-encoded values automatically form trie structures, and operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$—the relational semantics of logic programming (van Emden and Kowalski, 1976) (Section 5).
-Furthermore, the same web application MIXINv2 file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015) (Section Function color blindness: ).
+The reduction produced MIXIN, an early implementation in Nix, and subsequently MIXINv2 (Yang, 2026), now implemented in Python: a declarative programming language with only three constructs (record literals, definitions, and inheritance) and no functions or let-bindings.
+[^2] The executable MIXINv2 implementation (Yang, 2026) contains every example in this paper.
+In developing real programs in MIXINv2, including boolean logic, natural number arithmetic, and a web application, several phenomena arose from off-label use of the adopted design patterns.
+Section 3 shows how the three constructs give rise to a single abstraction that simultaneously serves as module, class, object, method, field, and local binding.
+The linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) did not arise: inheritance is inherently commutative, idempotent, and associative, and $\mathtt{this}$ naturally resolves to multiple targets rather than one, where prior systems such as Scala and the NixOS module system reject the multi-target situation.
+Section 5 traces how boolean logic and arithmetic, implemented as separate files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism, and how ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976).
+Section 6.1 discusses these and further emergent phenomena, including function color blindness (Nystrom, 2015).
 This paper distills these observations into inheritance-calculus, a minimal computational model for declarative programming.
-Function application translates directly into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument.
-Since the $\lambda$-calculus embeds into inheritance-calculus, inheritance-calculus is Turing complete.
-The translation reveals an asymmetry: embedding the $\lambda$-calculus into inheritance-calculus requires five local, syntax-directed rules (Section 4), but the lazy $\lambda$-calculus cannot macro-express the inheritance constructors.
-Inheritance-calculus enables open extension that changes operational equivalence on $\lambda$-terms, establishing that inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991) (Section 4.2).
-Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
-All constructions in this paper have been implemented and tested in MIXINv2, available as open source (Yang, 2026), with a test suite covering every example.
-[^1] MIXINv2 separates symbol resolution (which resolves lexical references and qualified-this references) from runtime mixin evaluation, but symbol resolution is performed just-in-time per symbol rather than ahead of time, so it does not provide static type checking in the traditional sense.
-Nevertheless, MIXINv2’s resolution phase is stricter than the untyped inheritance-calculus.
-All examples in this paper are ported from MIXINv2 and conform to both semantics.
-Elements added to satisfy the resolver are harmless for the algorithms presented here and may aid readability.
-All tests are mechanically verified against expected output snapshots.
-Contributions
-  • We present inheritance-calculus, a minimal computational model for declarative programming that contains three constructs (record, definition, inheritance).
-  • The mixin is a single abstraction that unifies class, method, object, and property.
-Method-level overriding is subsumed by deep merge of mixin trees; inheritance is commutative, idempotent, and associative, and the linearization problem in OOP does not arise.
-  • The semantics is observational: the observer drives computation by querying paths in a lazily constructed tree.
-The six semantic equations are first-order [^2] First-order in three senses simultaneously: the semantic domain contains only paths (finite sequences of labels) and sets of paths, with no function spaces, closures, or strategies; the six equations are first-order functions whose inputs and outputs are paths and sets of paths, never functions; and the judgments are first-order inductive definitions (Aczel, 1977) over paths and labels, where quantifiers range over paths and labels, never over functions or predicates., denotational, and fixpoint-computable.
-  • The mixin tree is the call stack. Self-reference resolves to multiple objects. Since inheritance is idempotent, all objects contribute equally.
-  • The $\lambda$-calculus macro-expresses into inheritance-calculus in five local rules (Section 4).
-  • The translation gives the lazy $\lambda$-calculus a first-order semantics (Section 4.1): the prior higher-order semantic frameworks are shown to be faithful to a sublanguage of inheritance-calculus’s first-order equations.
-Appendix E reformulates this semantics directly on ANF.
-As a corollary, some $\lambda$-terms that diverge in the $\lambda$-calculus converge under $\mathrm{lfp}(T_{P})$ after translation (Section 4; Remark B.5).
-  • Inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991) (Section 4.2).
-  • We observe further properties: immunity to nonextensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section Function color blindness: ).
-Section 5 demonstrates these properties through a case study of Nat arithmetic.
+Section 3 defines the semantics via six mutually recursive first-order functions on paths and sets of paths, with no function spaces, closures, or higher-order constructs; the functions are computable by tabling (Tamaki and Sato, 1986).
+Section 4 shows that these functions extend to the $\lambda$-calculus, embedding it.
+Remark B.5 derives a convergence hierarchy among eager, lazy, and fixpoint $\lambda$-calculi sharing the same syntax as a corollary.
+Section 4.2 establishes that the lazy $\lambda$-calculus cannot macro-express the inheritance constructors, so inheritance-calculus is strictly more expressive in Felleisen’s sense (Felleisen, 1991).
 2. Syntax
-Let $e$ denote an expression. Let $\ell$ denote a label (property name). Let $k$ denote a nonnegative integer (path length).
+Let $e$ denote an expression, $\ell$ a label serving as a property name, and $k$ a nonnegative integer serving as the length of a downward projection.
 $e\quad::=\quad$ $\{c_{1},\;\ldots,\;c_{m}\}$ $\text{(}m\geq 0\text{)}$
@@
 $c\quad::=\quad$ $\ell\mapsto e$ (definition)
-$\mid\quad$ $[\ell_{\mathrm{up}}.\mathbf{this}.\mid\uparrow^{\!n}.]\;\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ $\text{(inheritance, }k\geq 0\text{ with qualifier,}$
-$k\geq 1\text{ without)}$
+$\mid\quad$ $(\ell_{\mathrm{up}}.\mathbf{this}\mid\uparrow^{\!n}).\;\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ (inheritance: qualifier + downward path)
+$\mid\quad$ $\ell_{\mathrm{up}}.\mathbf{this}\mid\uparrow^{\!n}$ (inheritance: qualifier only)
+$\mid\quad$ $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ (inheritance: downward path only)
 An expression is a set of elements enclosed in braces.
@@
 Because expressions are sets, element order is irrelevant and duplicates have no effect; inheritance is inherently commutative, idempotent, and associative.
-The $\mapsto$ in $\ell\mapsto e$ defines a property.
+The $\mapsto$ in $\ell\mapsto e$ defines a property named $\ell$ with value $e$.
 MIXINv2 also provides scalar types through a foreign-function interface (FFI); inheritance-calculus is what remains after removing the FFI.
-None of the examples in this paper uses the FFI; scalars did not arise.
+None of the examples in this paper uses the FFI; no scalar values are needed in the examples.
 Inheritance
@@
 A new scope level is created at each $\{\ldots\}$ record literal.
+[^3] We use record for the syntactic construct, scope for the name-resolution context it creates, and mixin for its role as a composable unit; in inheritance-calculus these are three views of the same entity.
 The scope contains all properties of that mixin, including those inherited via inheritance sources.
@@
   • $\mathrm{Foo}.\mathbf{this}$ (equivalently, $\uparrow^{\!n}$ where $\mathrm{Foo}$ is $n$ steps above the enclosing scope) retrieves the record named $\mathrm{Foo}$.
-  • $\mathrm{Foo}.\mathbf{this}.\ell$ (equivalently, $\uparrow^{\!n}.\ell$) projects property $\ell$ from $\mathrm{Foo}$.
+  • $\mathrm{Foo}.\mathbf{this}.\ell$, or equivalently, $\uparrow^{\!n}.\ell$, projects property $\ell$ from $\mathrm{Foo}$.
   • When $k=0$, the reference inherits the entire enclosing scope, which may produce an infinitely deep tree.
@@
 Inheritance-calculus has a single abstraction, the deep-mergeable mixin, [^4] Where no ambiguity arises, we write simply “mixin.” The distinction from the Bracha–Cook mixin (Bracha and Cook, 1990) is detailed in Section 7.
-which can simultaneously serve as field, method, class, and object, because there are no opaque methods: every value is a transparent record, and inheritance always performs a deep merge of subtrees.
+which can simultaneously serve as module, class, object, method, field, and local binding, because there are no opaque methods: every value is a transparent record, and inheritance always performs a deep merge of subtrees.
 We call the resulting structure a mixin tree.
 Its semantics is purely observational: given an inheritance-calculus program, one queries whether a label $\ell$ is a property at a given position $p$.
-This section defines the semantics via six mutually recursive first-order functions on paths.
-The six functions map object-oriented concepts to set-theoretic operations: $\mathtt{properties}$ and $\mathtt{supers}$ generalize member lookup and inheritance chains, $\mathtt{overrides}$ and $\mathtt{bases}$ capture method override and base classes, and $\mathtt{resolve}$ and $\mathtt{this}$ handle name resolution and self-reference.
+This section defines the semantics via six mutually recursive first-order [^5] First-order in three senses simultaneously: the semantic domain contains only paths (finite sequences of labels) and sets of paths, with no function spaces, closures, or strategies (van Emden and Kowalski, 1976); the six equations are first-order functions whose inputs and outputs are paths and sets of paths, never functions (Reynolds, 1972); and the judgments are first-order inductive definitions (Aczel, 1977) over paths and labels, where quantifiers range over paths and labels, never over functions or predicates.
+functions on paths.
+The six functions map object-oriented concepts to set-theoretic operations: $\mathtt{properties}$ and $\mathtt{supers}$ generalize member lookup and inheritance chains, $\mathtt{overrides}$ and $\mathtt{bases}$ capture method override and base classes, and $\mathtt{resolve}$ and $\mathtt{this}$ handle name resolution and qualified $\mathtt{this}$ resolution.
 3.1. Definitions
@@
 An expression $e$ (Section 2) is parsed into an AST. The AST provides two primitive functions at each path $p$:
-  • $\mathtt{defines}(p)$: the set of labels that have subtrees at $p$.
+  • $\mathtt{defines}(p)$: the set of labels $\ell$ for which $p$ contains a definition $\ell\mapsto e$.
   • $\mathtt{inherits}(p)$: the set of reference pairs $(n,\;\ell_{\mathrm{down},*})$, where $n$ is the de Bruijn index and $\ell_{\mathrm{down},*}$ is the list of downward projections.
@@
 Each result is paired with the inheritance-site context $\mathtt{init}(p_{\mathrm{base}})$ through which it is reached.
-This provenance is needed by qualified this resolution, i.e.
-the $\mathtt{this}$ function defined below, to map a definition-site mixin back to the inheritance-site paths that incorporate it:
+This provenance is needed by qualified $\mathtt{this}$ resolution, that is, the $\mathtt{this}$ function defined below, to map a definition-site mixin back to the inheritance-site paths that incorporate it:
 (2) $\mathtt{supers}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{base}}),\;p_{\mathrm{override}})\;\big|\;p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{override}}\in\mathtt{overrides}(p_{\mathrm{base}})\;\bigr\}$
@@
 (5) $\mathtt{resolve}(p_{\mathrm{site}},\;p_{\mathrm{def}},\;n,\;\ell_{\mathrm{down},*})=\\ \bigl\{\;p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\ell_{\mathrm{down},k}\;\big|\;p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\;\mathtt{init}(p_{\mathrm{def}}),\;n)\;\bigr\}$
-When multiple routes exist due to multi-path inheritance, they may yield different target paths; all are collected. We now define $\mathtt{this}$.
-Qualified this resolution
+When multiple routes exist due to multi-target $\mathtt{this}$ resolution, they may yield different target paths; all are collected. We now define $\mathtt{this}$.
+Qualified $\mathtt{this}$ resolution
 $\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Each step finds, among the supers of every path in the frontier $S$, those whose override component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier.
@@
 Since $n$ is a nonnegative integer, the recursion terminates.
-[^4] If no matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ exists at some step, the frontier becomes the empty set, and $\mathtt{resolve}$ returns $\varnothing$: the reference has no target, and no properties are inherited.
+[^6] If no matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ exists at some step, the frontier becomes the empty set, and $\mathtt{resolve}$ returns $\varnothing$: the reference has no target; no properties are inherited.
 This is the most self-consistent semantics for the untyped calculus.
@@
 The six equations (1)–(6) define $\mathtt{properties}(p)$ by mutual recursion: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these functions.
-The functions are the interpreter—the observer drives the computation by choosing which path to inspect, and each equation unfolds on demand.
+The functions are the interpreter: the observer drives the computation by choosing which path to inspect, and each equation unfolds on demand.
 Because the functions are pure and order-independent, the reflexive-transitive closure $\mathtt{bases}^{*}$ in (2) may be explored breadth-first or depth-first without affecting the result; paths may be interned so that structural equality reduces to pointer equality; and the six functions form a memoizable dynamic-programming recurrence keyed by paths.
-Operationally, evaluation uses tabled evaluation (Tamaki and Sato, 1986; Chen and Warren, 1996): each query is memoised on first encounter, and re-entrant calls (cycles in the dependency graph) return the current accumulator rather than diverging.
+Operationally, evaluation uses tabled evaluation (Tamaki and Sato, 1986): each query is memoised on first encounter, and re-entrant calls, which correspond to cycles in the dependency graph, return the current accumulator rather than diverging.
 Denotationally, the six equations define a monotone immediate consequence operator $T_{P}$ (van Emden and Kowalski, 1976) whose right-hand sides use only $\cup$, $\in$, equality, and positive set comprehension, with no negation, set difference, or complement.
 By the Knaster–Tarski theorem (Tarski, 1955), $T_{P}$ has a unique least fixed point $\mathrm{lfp}(T_{P})$, which serves as the denotational semantics.
-Appendix A proves that tabled evaluation never returns a wrong answer and analyses termination.
+This denotational characterization formalizes the observational semantics introduced above: the observer-driven queries are computed by the same fixed-point equations.
+Whether evaluation terminates depends on the program.
+Recall that $\mathtt{supers}$ (equation 2) takes the reflexive-transitive closure of $\mathtt{bases}$ (equation 4); the resulting graph is the inheritance hierarchy.
+Since each inheritance source occupies a distinct node in the mixin tree, this hierarchy is directly readable from the tree structure.
+Programs whose inheritance hierarchy is acyclic terminate in one pass.
+When the hierarchy contains cycles, evaluation terminates when each query has a finite answer domain: the ascending chain condition on the finite lattice guarantees convergence (Bancilhon and Ramakrishnan, 1986), as in Datalog (Appendix F).
+When the hierarchy is infinite, evaluation may diverge; this is inherent to Turing completeness.
+Appendix A proves that tabled evaluation never returns a wrong answer and formalizes these termination conditions.
 4. Embedding the $\lambda$-Calculus
@@
 […]
-Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results (and, optionally, naming values), and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus.
+Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results.
+[^8] This can optionally include naming values as well; since every ANF term is also a $\lambda$-term, ANF is neither a restriction nor an extension of the $\lambda$-calculus.
 The ANF structure matches inheritance-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
@@
 In the value $\mathbf{let}$-binding rule, $x$ wraps $\mathcal{T}(V)$ under a $\mathrm{result}$ child so that the uniform accessor $x.\mathrm{result}$ yields $\mathcal{T}(V)$, the translated value itself.
-This extra wrapper is semantically inert but proof-relevant: it lets the same $\mathrm{result}$-following argument used for applications extend uniformly to value $\mathbf{let}$-bindings in the adequacy proof below.
+This extra wrapper is semantically inert but proof-relevant: it lets the same $\mathrm{result}$-following argument used for applications extend uniformly to value $\mathbf{let}$-bindings in the adequacy proof (Appendix B).
 In the tail-call rule, the fresh label $\mathrm{tailCall}$ serves the same purpose, and $\mathrm{result}$ projects the answer.
@@
 Since the $\lambda$-calculus is Turing complete, so is inheritance-calculus.
+[^10] MIXINv2 separates symbol resolution (which resolves lexical references and qualified $\mathtt{this}$ references) from runtime mixin evaluation, but symbol resolution is performed just-in-time per symbol rather than ahead of time, so it does not provide static type checking in the traditional sense.
+Nevertheless, MIXINv2’s resolution phase is stricter than the untyped inheritance-calculus.
+All examples in this paper are ported from MIXINv2 and conform to both semantics.
+Elements added to satisfy the resolver are harmless for the algorithms presented here and may aid readability.
 4.1. Böhm Tree Correspondence
-The translation $\mathcal{T}$ induces an elementary semantics on the $\lambda$-calculus via the mixin tree equations (1)–(6).
-We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
+The translation $\mathcal{T}$ embeds the $\lambda$-calculus into a sublanguage of inheritance-calculus.
+Intuitively, a mixin tree augments the AST with additional semantic structure; for translated $\lambda$-terms, the projection of this additional structure is isomorphic to the Böhm tree.
 A mixin tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$.
-We write $T{\Uparrow}$ if no such depth exists.
-Definition 4.1 (Inheritance-convergence).
-A mixin tree $T$ converges, written $T{\Downarrow}$, if there exists $n\geq 0$ such that:
-$\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{properties}\!\bigl(\,\underbrace{\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
-in the semantics of Section 3.
-Theorem 4.2 (Adequacy).
-A closed ANF $\lambda$-term $M$ has a head normal form if and only if $\mathcal{T}(M){\Downarrow}$.
-Theorem 4.3 (First-order semantics of the lazy $\lambda$-calculus).
-For closed $\lambda$-terms $M$ and $N$, define $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ iff for every closing $\lambda$-calculus context $C[\cdot]$, $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$.
-Then:
-$\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
-Here $\approx_{\lambda}$ quantifies over $\lambda$-calculus contexts only; inheritance-calculus contexts are strictly more discriminating (Section 4.2).
-The theorem establishes that the lazy $\lambda$-calculus has a first-order semantics: the translation $\mathcal{T}$ embeds the $\lambda$-calculus into a sublanguage of inheritance-calculus, using only recursive evaluation of set-comprehension equations over paths and sets of paths.
-The prior higher-order semantic frameworks of (Scott, 1972; Wadsworth, 1976; Abramsky et al., 2000; Hyland and Ong, 2000; Barendregt, 1984) are thereby shown to be faithful to this first-order semantics: they agree with the equations on Böhm tree equivalence (Barendregt, 1984).
-[^7] Composing $\mathcal{T}$ with the equations of Section 3 yields a first-order semantics of the $\lambda$-calculus in which function application is record inheritance and convergence is determined by path queries.
-Appendix E reformulates this semantics directly on ANF, without passing through inheritance-calculus.
-The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix B.
+The translation preserves and reflects convergence (adequacy).
+Write $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ when $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$ for every closing $\lambda$-calculus context $C[\cdot]$; then $\approx_{\lambda}$ coincides with Böhm tree equivalence (Barendregt, 1984).
+Formal statements and proofs are in Appendix B.
+[^11] Appendix E applies the same first-order technique directly on ANF to give the $\lambda$-calculus a self-contained fixpoint semantics, without passing through inheritance-calculus.
+Note that $\approx_{\lambda}$ quantifies over $\lambda$-calculus contexts only; inheritance-calculus contexts are strictly more discriminating (Section 4.2).
+Böhm tree equivalence is fully abstract for the first-iteration approximation $T_{P}\!\uparrow\!1$ of the equations.
+The full $\mathrm{lfp}(T_{P})$ is strictly more defined.
+For example, $\mathbf{let}\;x=x\;\mathbf{in}\;x$ diverges under $\beta$-reduction, but its translation $\{x\mapsto\{\mathrm{result}\mapsto x.\mathrm{result}\},\;\mathrm{result}\mapsto x.\mathrm{result}\}$ stabilizes under $\mathrm{lfp}(T_{P})$: [^12] This simple cycle could also be resolved by a fixed-point combinator such as $Y$.
+The $\mathrm{lfp}(T_{P})$ semantics is more general: it computes fixed points at the container level, e.g.
+finding a set $S$ such that $S=F(S)$.
+Recursive Datalog on cyclic graphs (Appendix F, Section F.8) requires exactly this, and neither $Y$ nor Haskell’s mfix (RecursiveDo) can express it.
+tabled evaluation detects the cycle and returns $\mathtt{properties}(\mathrm{root})=\{x,\mathrm{result}\}$ with no abstraction shape, reflecting an undefined value rather than diverging (Remark B.5).
+[^13] We call the sublanguage of inheritance-calculus corresponding to $\lambda$-terms, evaluated under $\mathrm{lfp}(T_{P})$ rather than $T_{P}\!\uparrow\!1$, the fixpoint $\lambda$-calculus; Remark B.5 elaborates.
 The relationship between $\approx_{\lambda}$ and full inheritance-contextual equivalence is clarified by the expressive asymmetry of the next section: contexts in the language universe $\mathscr{U}$ (Definition C.3) that mix $\lambda$-calculus and inheritance-calculus constructors can separate $\lambda$-terms that are Böhm-tree equivalent, so ${\approx_{\lambda}}\subsetneq{(\approx_{\mathscr{U}}|_{L_{0}})}$ where $L_{0}$ is the lazy $\lambda$-calculus.
@@
 The formal definitions, constructions, and proofs are given in Appendix C; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
-The underlying mechanism is open extension: mixin definitions compose via open recursion, so existing definitions gain new observable properties without modification.
+The underlying mechanism is open extension: mixin definitions compose via open extension, so existing definitions gain new observable properties without modification.
 In the $\lambda$-calculus, a function’s result is closed: there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
@@
 We implemented boolean logic and natural number arithmetic, both unary and binary, as MIXINv2 files (Yang, 2026).
+[^14] The examples in this section are rendered in inheritance-calculus notation; the executable MIXINv2 sources are available as supplementary material.
 This section traces the unary Nat case in detail.
-The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names serve as data constructors and operations; lowerCamelCase names serve as fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action.
+The implementation follows a declarative object-oriented style: each concern is one top-level mixin; UpperCamelCase names serve as data constructors and operations; lowerCamelCase names serve as fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action.
 Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
 NatData
-Following the Church-encoding convention (Yang, 2026), [^8] The data representation (Zero/Successor) is a Peano encoding, an initial algebra of the functor $F(X)=1+X$.
-The visitor pattern (NatVisitor, below) adds the Church-encoding behavior: a fold (catamorphism) over this algebra.
-These two concerns are defined in separate files and composed via inheritance, which is itself an instance of the Expression Problem.
-we define a factory containing the abstract product type and two labels that serve as data constructors.
+The data representation (Zero/Successor) is a Peano encoding, an initial algebra of the functor $F(X)=1+X$.
+The visitor pattern (NatVisitor, below) adds Church-encoding behavior: a fold (catamorphism) over this algebra.
+These two concerns are defined in separate files and composed via inheritance (Yang, 2026), which is itself an instance of the Expression Problem.
+Following this convention, we define a factory containing the abstract product type and two labels that serve as data constructors.
 A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field.
-The file name $\mathrm{NatData}$ is itself the top-level scope:
+The name $\mathrm{NatData}$ is the top-level scope:
 $\mathrm{NatData}\mapsto\{\;$ $\mathrm{NatFactory}\mapsto\{\mathrm{Product}\mapsto\{\},$
@@
 NatPlus
-Addition is a separate MIXINv2 file that inherits $\mathrm{NatData}$ and extends each data constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field.
+Addition is a separate mixin that inherits $\mathrm{NatData}$ and extends each data constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field.
 The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
@@
 $\qquad\mathrm{sum}\mapsto\{\mathrm{\_recursiveAddition}.\mathrm{sum}\}\}\}\}\}$
-NatEquality
-To test that $2+3=5$, we needed equality on Nat values.
+NatVisitor
 Equality requires case analysis on the data constructor; we use the visitor pattern.
@@
 $\mathrm{Boolean}\mapsto\{\mathrm{BooleanFactory},\;\mathrm{BooleanFactory}.\mathrm{Product}\}\}$
+NatEquality
+To test that $2+3=5$, we needed equality on Nat values.
+The Cartesian product test (page 5) will apply this same $\mathrm{Equal}$ to a trie union of multiple Nat values without modification.
 $\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each data constructor.
@@
 $\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherAcceptance}.\mathrm{Accepted}.\mathrm{equal}\}\}\}\}\}$
-Composing $\mathrm{NatPlus}$ with $\mathrm{NatEquality}$ by inheriting both, with no modifications to either file, gives the resulting Nat values both $\mathrm{Plus}$ and $\mathrm{Equal}$ automatically.
-The concrete numerals live in a shared file that inherits only $\mathrm{NatData}$:
+Composing $\mathrm{NatPlus}$ with $\mathrm{NatEquality}$ by inheriting both, with no modifications to either mixin, gives the resulting Nat values both $\mathrm{Plus}$ and $\mathrm{Equal}$ automatically.
+The concrete numerals live in a shared mixin that inherits only $\mathrm{NatData}$:
 $\mathrm{NatConstants}\mapsto\{\;$ $\mathrm{NatData},$
@@
 $\mathrm{Test2plus3}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Five}.\mathrm{Equal},\;\mathrm{other}\mapsto\{\mathrm{Addition}.\mathrm{sum}\}\}\}$
-Querying $\mathrm{Test}.\mathrm{Test2plus3}.\mathrm{equal}$ via the host-language API returns $\mathrm{True}$.
-No existing file was modified.
-$\mathrm{NatEquality}$, a new operation, and $\mathrm{Boolean}$, a new data type, each required only a new file: extensibility in both dimensions of the Expression Problem (Wadler, 1998).
+The path $\mathrm{Test}.\mathrm{Test2plus3}.\mathrm{equal}$ inherits $\mathrm{True}$.
+No existing mixin was modified.
+$\mathrm{NatEquality}$, a new operation, and $\mathrm{Boolean}$, a new data type, each required only a new mixin: extensibility in both dimensions of the Expression Problem (Wadler, 1998).
 Tree-level inheritance and return types
@@
 Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory.
-Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either file.
+Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either mixin.
 In object algebras (d.
@@
 Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
-Multi-path qualified this (equation 6) is essential.
-In the test file above, $\mathrm{NatPlus}.\mathrm{Successor}$ and $\mathrm{NatEquality}.\mathrm{Successor}$ both inherit $\mathrm{NatData}.\mathrm{Successor}$.
+Multi-target qualified $\mathtt{this}$ (equation 6) is essential.
+In the test mixin above, $\mathrm{NatPlus}.\mathrm{Successor}$ and $\mathrm{NatEquality}.\mathrm{Successor}$ both inherit $\mathrm{NatData}.\mathrm{Successor}$.
 When the two operations are composed, the resulting $\mathrm{Successor}$ inherits the $\mathrm{NatData}$ schema through two independent routes.
@@
 In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix D), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously.
-The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.
+The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-target $\mathtt{this}$ resolution to shared schemas, which DOT’s single-valued self variable cannot accommodate.
 Church-encoded Nats are tries
@@
 This is exactly the shape of a trie.
+[^15] The unary Nat trie can be viewed as a sparse linked list whose depth equals the number it represents: in a trie union, some $\mathrm{Successor}$ nodes along the list point to $\mathrm{Zero}$ (leaf nodes), so not every position in the list is occupied.
+The MIXINv2 source (Yang, 2026) includes a binary natural number (BinNat) arithmetic whose trie depth is logarithmic.
 A scope that simultaneously inherits two Nat values is a trie union.
@@
 $\mathrm{Result}.\mathrm{sum}$ evaluates to the trie $\{4,5,6\}$, the set of all pairwise sums.
-$\mathrm{Equal}$ then applies to every pair drawn from $\{4,5,6\}\times\{4,5,6\}$; querying $\mathrm{CartesianTest}.\mathrm{Check}.\mathrm{equal}$ via the host-language API returns $\{\mathrm{True},\mathrm{False}\}$.
+$\mathrm{Equal}$ then applies to every pair drawn from $\{4,5,6\}\times\{4,5,6\}$; the path $\mathrm{CartesianTest}.\mathrm{Check}.\mathrm{equal}$ inherits both $\mathrm{True}$ and $\mathrm{False}$.
 The Cartesian product arises directly from the semantic equations.
@@
 The result $\mathrm{sum}$ therefore collects all paths reachable through any combination of the two operands, which is the Cartesian product.
-The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: $\mathtt{overrides}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.
+The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: [^16] Appendix F gives the systematic encoding of Datalog into inheritance-calculus.
+$\mathtt{overrides}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.
 6. Discussion
 6.1. Emergent Phenomena
-In developing real programs in MIXINv2, we encountered three phenomena that we did not design for.
+In developing real programs in MIXINv2, three phenomena arose from off-label use of the adopted design patterns.
+[^17] MIXINv2 programs adopt two standard patterns: data types follow a factory pattern isomorphic to Scala’s sealed trait with case classes; pattern matching uses Scott encoding.
+Both are named after Gang-of-Four terminology (Gamma et al., 1994) (factory, visitor, acceptance), with verbs replaced by nouns since the calculus has no methods.
 They can be seen in the source code (Yang, 2026) and the Case Study (Section 5).
@@
 When a mixin tree is viewed as a trie, the correspondence with relational algebra becomes precise.
-[^9] Inheritance-calculus to relational algebra: path from root to leaf == tuple; trie union == relation (set of tuples); $\mathtt{bases}$ (equation 4) == cross-join; $\mathtt{supers}$ (equation 2) == union; $\mathtt{properties}$ (equation 1) == projection.
 The carrier is a trie rather than a flat relation, and observation is coinductive: the observer drives computation by querying progressively deeper paths.
@@
 6.2. Semantic Alternatives
-The preceding item established that inheritance-calculus is immune to nonextensibility. We now ask whether an alternative design could share this property by surveying every candidate known to us.
+The preceding item established that inheritance-calculus is immune to nonextensibility.
+We adopt this as a design goal and ask whether an alternative design could share this property by surveying every candidate known to us.
 Degrees of freedom
@@
 [^20] Nix is a functional language with first-class functions; its attribute sets alone do not form a named-tree DSL.
-The NixOS module system (NixOS Contributors, 2024), however, restricts the user-facing interface to nested attribute sets composed by recursive merging, and it is this DSL-level structure that naturally forms a named tree.
+The NixOS module system (NixOS Contributors, ), however, restricts the user-facing interface to nested attribute sets composed by recursive merging, and it is this DSL-level structure that naturally forms a named tree.
 For mainstream programming languages the AST is not directly a named tree, since anonymous expression positions and implicit scopes lack labels.
@@
 Haskell (FP) $\lambda$, application, let, case, data, type class, instance, … variable, constructor, type class dispatch, import, …
-Scala (multi-paradigm) class, trait, object, method, block, conditional, match, while, throw, … variable, member access, extends, with, implicits, import, …
+Scala [^22] Scala is a multi-paradigm language. class, trait, object, method, block, conditional, match, while, throw, … variable, member access, extends, with, implicits, import, …
 $\lambda$-calculus abstraction, application variable
@@
 To make a new aspect extensible, the original function must be rewritten to accept an additional parameter.
-This applies to every language whose composition mechanism is function application, whether functional, object-oriented, or multi-paradigm, because method calls, virtual dispatch, and type-class resolution all reduce to $\beta$-reduction.
-The Expression Problem (Wadler, 1998) is hard in all of them for the same reason: extensibility is opt-in.
+Any composition mechanism based on function application inherits this limitation, because $\beta$-reduction closes the function body.
+The Expression Problem (Wadler, 1998) is hard in such languages for the same reason: extensibility is opt-in.
 The frameworks that simulate immunity to nonextensibility (visitor pattern, object algebras (d.
@@
 Conditional incorporation:
-Removing or conditionally including definitions can introduce cycles in the dependency graph, causing the recursive evaluation (Appendix A) to diverge.
+Removing or conditionally including definitions based on the presence of other definitions creates a dependency cycle: whether a definition exists depends on whether another definition exists, which in turn depends on the first.
+The recursive evaluation (Appendix A) cannot reach a fixed point when addition and removal alternate, causing divergence.
+This differs from the divergence inherent to Turing completeness: Turing-complete programs diverge because computation is unbounded, whereas conditional incorporation diverges because the immediate consequence operator $T_{P}$ oscillates rather than ascending monotonically.
 Among the candidates above, deep merge is the only one we found that is commutative, idempotent, associative, and immune to nonextensibility.
@@
 When multiple inheritance routes reach the same scope, single-target resolution must reject the situation or select one route.
-The NixOS module system (NixOS Contributors, 2024) raises a duplicate-declaration error (Appendix D).
+The NixOS module system (NixOS Contributors, ) raises a duplicate-declaration error (Appendix D).
 This rejects nested extensions where two independently authored modules define the same option, a composition that arises when solving the Expression Problem (Wadler, 1998) (Section 5).
-Our first three implementations of MIXINv2 used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972).
+Our first three implementations of MIXINv2 used standard techniques for $\mathtt{this}$ resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972).
 All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope.
 The difficulty was structural.
-A closure captures a single environment, so a fixed-point combinator solves for a single self-reference.
+A closure captures a single environment, so a fixed-point combinator solves for a single $\mathtt{this}$ target.
 A stack-based environment is a linear chain where each scope level has one parent.
-Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix D).
+Both data structures embed a single-valued assumption incompatible with the multi-target situation that multiple inheritance naturally produces (Appendix D).
 Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope.
@@
 Dynamic scope mixin cannot replace function Jsonnet (Cunningham, 2014), union FS
-Single-target $\mathtt{this}$ nested extensions rejected NixOS modules (NixOS Contributors, 2024)
+Single-target $\mathtt{this}$ nested extensions rejected NixOS modules (NixOS Contributors, )
 No alternative we examined at either level is immune to nonextensibility.
-The semantics of inheritance-calculus was not chosen from a menu.
-We arrived at it by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-path compositions that the Expression Problem demands.
-[^15] Three successive implementations confirmed this; see the discussion above.
+We arrived at this semantics by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-target compositions that the Expression Problem demands.
 Deep merge with multi-target late-binding $\mathtt{this}$ is what remained.
@@
 Minimal Turing-complete models
-Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus reduces to three combinators $S$, $K$, II (II is redundant: $I=SKK$).
+Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus reduces to three combinators $S$, $K$, II  [^24] The $I$ combinator is redundant since $I=SKK$.
+.
 The bracket abstraction that compiles $\lambda$-terms to SKI is mechanical but incurs exponential blow-up in term size; no new computational patterns emerge beyond those already present in the $\lambda$-calculus.
@@
 The Turing Machine itself, while universal, lacks random access: reading a cell at distance $d$ requires $d$ sequential tape movements.
-Each of these models achieves Turing completeness by removing structure from an existing model: fewer abstractions, worse ergonomics, no new expressiveness.
-None of these reductions exhibits emergent phenomena absent from the model it reduces.
+Each of these models achieves Turing completeness within the semantic framework of an existing model, and no independent semantic domain has been exhibited for any of them.
 Semantic descriptions of the $\lambda$-calculus
@@
 Scott’s $D_{\infty}$ resolves this by constructing $D$ as an inverse limit of continuous lattices (Scott, 1972).
-The resulting semantics is adequate but not fully abstract (Plotkin, 1977; Milner, 1977): it identifies terms that observational equivalence distinguishes (Plotkin, 1975).
-Game semantics (Abramsky et al., 2000; Hyland and Ong, 2000; Nickau, 1994) achieves full abstraction for PCF and for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995) at the cost of considerably more intricate mathematical machinery (arenas, strategies, innocence conditions).
-In each of these approaches, the function space $D\to D]$ is the central obstacle: operational semantics needs it implicitly (via substitution), denotational semantics constructs it explicitly, and game semantics refines it via strategies.
-None of these approaches is first-order; [^16] Function spaces in denotational semantics, strategies in game semantics, and substitution in operational semantics.
-none eliminates the function space from the semantic domain; and none unifies denotational and operational semantics into a single first-order definition.
-All three also inherit $\beta$-reduction’s divergence on recursive programs: a naïve transitive closure on a cyclic graph, for instance, diverges in every framework.
+The resulting semantics is adequate but not fully abstract (Plotkin, 1977; Milner, 1977): it identifies terms that observational equivalence distinguishes.
+Game semantics (Abramsky et al., 2000; Hyland and Ong, 2000; Nickau, 1994) achieves full abstraction for PCF and for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995), requiring arenas, strategies, and innocence conditions.
+In each of these approaches, the function space $D\to D]$ is the central obstacle: operational semantics needs it implicitly via substitution.
+In contrast, denotational semantics constructs it explicitly, and game semantics refines it via strategies.
+Each of these approaches involves higher-order structure [^25] Function spaces in denotational semantics, strategies in game semantics, and substitution in operational semantics.
+and retains the function space in its semantic domain.
+The denotational and operational accounts remain separate formalisms in each case.
+All three inherit $\beta$-reduction’s divergence on recursive programs: a naïve transitive closure on a cyclic graph, for instance, diverges in every framework.
 Set-theoretic models of the $\lambda$-calculus
@@
 Relational models (Ehrhard, 2005; Bucciarelli and Ehrhard, 2001) arise from the relational semantics of linear logic via finite multisets; Ehrhard (Ehrhard, 2005) showed that a standard reflexive object for the pure $\lambda$-calculus cannot exist in this category.
-All three families are semantic models: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure.
-None is a self-contained calculus with its own primitives into which the $\lambda$-calculus embeds.
-None is first-order, since the function space is encoded, not eliminated.
-And none provides a fully abstract denotational semantics of the lazy $\lambda$-calculus without higher-order machinery.
+All three families are semantic models: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure, rather than defining an independent calculus with its own primitives.
+Each encodes the function space rather than eliminating it, and each requires higher-order machinery to obtain full abstraction for the lazy $\lambda$-calculus.
 Recursive equations and fixpoint semantics
-Van Emden and Kowalski (van Emden and Kowalski, 1976) showed that the semantics of predicate logic as a programming language admits three equivalent characterizations: operational (SLD resolution), model-theoretic (least Herbrand model), and fixed-point (least fixed point of the immediate consequence operator $T_{P}$ on the powerset of ground atoms).
+Van Emden and Kowalski (van Emden and Kowalski, 1976) showed that the semantics of predicate logic as a programming language admits three equivalent characterizations: operational [^26] Typically SLD resolution.
+, model-theoretic [^27] Based on least Herbrand models.
+, and fixed-point [^28] Namely, the least fixed point of the immediate consequence operator $T_{P}$ on the powerset of ground atoms.
+.
 Aczel (Aczel, 1977) gave a systematic treatment of inductive definitions via monotone operators on powerset lattices, providing the logical foundations that Datalog-style semantics rests on.
-Leroy and Grall (Leroy and Grall, 2006) treated divergence in big-step semantics via coinductive definitions; in inheritance-calculus, divergence already emerges as non-termination of recursive evaluation, so no separate coinductive formulation is needed.
-None of these prior frameworks has been applied to a first-order domain of paths to obtain a fully abstract semantics of the lazy $\lambda$-calculus without function spaces, substitution, or reduction rules.
-In particular, Van Emden and Kowalski’s $T_{P}$ computes $\mathrm{lfp}$ for recursive Datalog programs, but has not been used to give recursive $\lambda$-programs a fixpoint semantics in which they converge instead of diverging.
+Leroy and Grall (Leroy and Grall, 2006) treated divergence in big-step semantics via coinductive definitions, requiring a separate coinductive judgment to account for non-termination alongside the inductive evaluation judgment.
+Van Emden and Kowalski’s $T_{P}$ computes $\mathrm{lfp}$ for recursive Datalog programs over a first-order domain.
+Conventional $\lambda$-calculus semantics requires function spaces in the semantic domain, and recursive $\lambda$-terms ordinarily diverge under $\beta$-reduction; using $T_{P}$ over a first-order domain to give the $\lambda$-calculus a fixpoint semantics has not been explored.
 7.2. Inheritance, Objects, and Records
@@
 Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records.
-Inheritance is composition of generators (functions from self to complete object) and wrappers (functions that modify generators), solved by a fixed-point construction.
+Inheritance is composition of generators  [^29] These are functions from self to complete object.
+and wrappers [^30] Wrappers are functions that modify generators.
+, solved by a fixed-point construction.
 Cook’s key insight is that inheritance is a general mechanism applicable to any form of recursive definition, not only object-oriented methods.
@@
 Mixins and traits
-Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses—functions from a superclass parameter to a subclass.
+Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses, that is, functions from a superclass parameter to a subclass.
 Because mixin application is function composition, it is neither commutative nor idempotent; the C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order.
@@
 Both presuppose the $\lambda$-calculus: method bodies are functions, and the composition mechanism itself is not Turing complete.
-Neither mixins nor traits are therefore deep-mergeable.
+Neither mixins nor traits are deep-mergeable.
 Object and record calculi
@@
 Harper and Pierce (Harper and Pierce, 1991) gave a record calculus with symmetric concatenation that rejects same-label composition; Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) introduced row polymorphism for records and variants in ML.
-All of these calculi operate on flat records: there is no recursive merging of nested structure, no lazy observation, and self-reference requires an explicit fixed-point combinator.
+All of these calculi operate on flat records: there is no recursive merging of nested structure, no lazy observation, and $\mathtt{this}$ requires an explicit fixed-point combinator.
 Family polymorphism and DOT
@@
 Configuration languages
-CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification (&) is commutative, associative, and idempotent.
+CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification, denoted by &, is commutative, associative, and idempotent.
 CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification.
 CUE’s lattice includes scalar types with a conflict semantics where unifying incompatible scalars yields $\bot$; whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question.
-Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative since the right-hand side wins on scalar conflicts, and deep merging requires explicit opt-in via the +: syntax) on a per-field basis, so deep merge is available but not the default.
+Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative since the right-hand side wins on scalar conflicts, and deep merging requires explicit opt-in via the +: syntax on a per-field basis, so deep merge is available but not the default.
 Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
 Module systems and deep merge
-The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same deep-merge semantics that distinguishes tree-level inheritance from method-level composition in prior mixin and trait calculi.
+The NixOS module system (Dolstra et al., 2010; NixOS Contributors, ) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same deep-merge semantics that distinguishes tree-level inheritance from method-level composition in prior mixin and trait calculi.
 When two modules define the same nested path, their subtrees are merged rather than overridden.
-The NixOS module system incorporates deep merge, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top.
+The NixOS module system incorporates deep merge, recursive $\mathtt{this}$, deferred evaluation of modules, and a rich type-checking layer on top.
 The same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity.
-The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance.
+The one aspect that appears not to have been fully addressed is the semantics of $\mathtt{this}$ under multiple inheritance.
 When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix D).
-The module system cannot express that both routes contribute equally to the same option, that is, the multi-path situation that recursive merging naturally gives rise to, which motivated the observational semantics of Section 3.
+The module system cannot express that both routes contribute equally to the same option, that is, the multi-target situation that recursive merging naturally gives rise to, which motivated the observational semantics of Section 3.
 7.4. Our Companion Work
@@
 Union file systems
-Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view.
-Their semantics resembles inheritance-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, noncommutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers.
+Union file systems, such as UnionFS, OverlayFS, and AUFS, layer multiple directory hierarchies to present a unified view.
+Their semantics resembles inheritance-calculus in that later layers override earlier layers and files from all layers remain accessible, but their designs differ in specific ways: inheritance is asymmetric, conflict resolution for scalar files is noncommutative, and there is no late-binding or dynamic dispatch across layers.
 We implemented inheritance-calculus directly as a union file system (Yang, 2025), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives.
@@
 Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how inheritance-calculus relates to the $\lambda$-calculus: a Typed inheritance-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with mixin inheritance.
-A formal type system would be amenable to mechanization in proof assistants (Coq, Agda) for verifying type safety, progress, and preservation properties.
+A formal type system would be amenable to mechanization in proof assistants, such as Coq and Agda, for verifying type safety, progress, and preservation properties.
 Beyond soundness of the existing checks, additional type system features would strengthen the language.
@@
 This is the natural trie semantics of the calculus, and the basis for solving the Expression Problem (Section 5).
-However, many operations (e.g., equality testing) assume that values are linear, inhabiting exactly one data constructor.
+However, many operations, such as equality testing, assume that values are linear, inhabiting exactly one data constructor.
 The observer pattern (Section 5) does not enforce this: applied to a multi-data-constructor value, all callbacks fire and their results are inherited.
@@
 Immunity to nonextensibility guided the design of inheritance-calculus.
-The emergent phenomena exhibit an expressive asymmetry between inheritance-calculus and the $\lambda$-calculus.
-Functions are syntactic sugar for inheritance.
+Section 6.2 eliminated every alternative composition mechanism and reference resolution, leaving deep merge with multi-target late-binding qualified $\mathtt{this}$ as the unique survivor.
+The mixin served as the single abstraction unifying module, class, object, method, field, and local binding; deep merge of mixin trees made inheritance commutative, idempotent, and associative, so the linearization problem of multiple inheritance did not arise.
+The six first-order equations of Section 3, denotational and computable by tabling (Tamaki and Sato, 1986), suffice to define the semantics; the observer drives computation by querying paths in a lazily constructed tree.
+These equations are a minimal departure from standard object-oriented semantics: the only change is the $\mathtt{overrides}$ equation (3), which replaces method-level shadowing with deep merge of subtrees.
+Object-oriented programming is traditionally classified as imperative; removing mutable state and function bodies yields pure object-oriented programming,313131“Pure” carries a double meaning: not impure, since values are immutable and composition has no side effects; and not mixed, since composition relies entirely on inheritance, without functional abstractions such as $\lambda$ or $\beta$-reduction.
+a minimal computational model for declarative programming that fills the empty cell of Table 1.
+Since $\mathtt{overrides}$ merges subtrees rather than method bodies, functions are no longer primitive.
+The $\lambda$-calculus macro-expressed into inheritance-calculus in six translation rules in Section 4.
+The Böhm tree correspondence of Section 4.1 extended this first-order semantics to the $\lambda$-calculus: the prior higher-order semantic frameworks proved faithful to the first-iteration approximation of a sublanguage of inheritance-calculus’s first-order equations; Appendix E reformulated this semantics directly on ANF.
+A convergence hierarchy among the eager, lazy, and fixpoint $\lambda$-calculi sharing the same $\lambda$-syntax followed as a corollary, detailed in Remark B.5.
+Inheritance-calculus further gains three phenomena that the $\lambda$-calculus cannot express; indeed, inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991) (Section 4.2).
+Section 5 established immunity to nonextensibility in the sense of the Expression Problem (Wadler, 1998).
+The case study of Section 5 showed that ordinary arithmetic yields the relational semantics of logic programming: cyclic inheritance hierarchies converge by the same ascending chain condition that governs recursive Datalog, and Appendix F gave a systematic encoding.
+Section Function color blindness: exhibited function color blindness (Nystrom, 2015).
+Inheritance is not a mechanism for organizing functions. Functions are syntactic sugar for inheritance.
 Appendix A Well-Definedness of the Semantic Functions
 The six semantic equations (1)–(6) are first-order: their inputs and outputs are paths and sets of paths, with no function spaces, closures, or higher-order constructs.
-This section establishes well-definedness by showing that the six equations define a monotone operator whose least fixed point is computed soundly by tabled evaluation (Tamaki and Sato, 1986; Chen and Warren, 1996).
+This section establishes well-definedness by showing that the six equations define a monotone operator whose least fixed point is computed soundly by tabled evaluation (Tamaki and Sato, 1986).
 Definition A.1 (Dependency graph).
@@
   2. (2)
-    Monotone accumulation. The tabling accumulator only grows between iterations (via set union); no fact is ever removed.
+    Monotone accumulation. The tabling accumulator only grows between iterations via set union; no fact is ever removed.
   3. (3)
@@
 If no sub-query was re-entered during evaluation, all sub-queries were fully computed with no approximation, and the result is exact.
-If re-entry occurred but the accumulator is stable (no new facts were added), the accumulator equals $\mathrm{lfp}(T_{P})$ restricted to the reachable queries, because it is a pre-fixed point reached by ascending from $\bot$.
+If re-entry occurred but the accumulator is stable, meaning no new facts were added, the accumulator equals $\mathrm{lfp}(T_{P})$ restricted to the reachable queries, because it is a pre-fixed point reached by ascending from $\bot$.
   4. (4)
@@
 When reentry occurs, let $R\subseteq\mathcal{Q}$ be the set of queries reachable from the initial query, and let $T_{P}^{R}$ denote the restriction of $T_{P}$ to maps on $R$.
-If the accumulator is stable, i.e.
-$A=A\cup T_{P}^{R}(A)$, then $T_{P}^{R}(A)\subseteq A$, so $A$ is a pre-fixed point of $T_{P}^{R}$.
+If the accumulator is stable, that is, $A=A\cup T_{P}^{R}(A)$, then $T_{P}^{R}(A)\subseteq A$, so $A$ is a pre-fixed point of $T_{P}^{R}$.
 By (1), every fact ever added to the accumulator was produced by some right-hand side evaluation and therefore belongs to $\mathrm{lfp}(T_{P}^{R})$. Hence $A\subseteq\mathrm{lfp}(T_{P}^{R})$.
@@
 Thus $\mathrm{lfp}(T_{P}^{R})\subseteq A$.
-Therefore $A=\mathrm{lfp}(T_{P}^{R})$, i.e. the stable accumulator equals the exact denotation on the reachable queries.
+Therefore $A=\mathrm{lfp}(T_{P}^{R})$, that is, the stable accumulator equals the exact denotation on the reachable queries.
 For (4), budget exhaustion and stack overflow both produce exceptions, never silent wrong answers. ∎
@@
   • Cyclic, finite per-query domain. Multiple passes converge by the ascending chain condition on a finite lattice (Bancilhon and Ramakrishnan, 1986). This is the Datalog case (Section F.8).
-  • Infinite reachable subgraph. Evaluation may diverge; this is inherent to Turing-completeness and not a defect of the evaluation strategy.
+  • Infinite reachable subgraph. Evaluation may diverge; this is inherent to Turing completeness and not a defect of the evaluation strategy.
 Appendix B Böhm Tree Correspondence: Proofs
-This appendix contains the full proofs of Adequacy (Theorem 4.2) and the first-order semantics theorem (Theorem 4.3) stated in Section 4.1.
+This appendix proves two results about the sublanguage of inheritance-calculus corresponding to the $\lambda$-calculus (Section 4.1).
+First, the first-iteration approximation $T_{P}\!\uparrow\!1$ is fully abstract with respect to Böhm tree equivalence (Theorems B.7 and B.9), establishing that the first iteration coincides with the lazy $\lambda$-calculus.
+Second, the full $\mathrm{lfp}(T_{P})$ semantics, the fixpoint $\lambda$-calculus, is strictly more defined: $\lambda$-terms that diverge under $\beta$-reduction may converge (Remark B.5), establishing a convergence hierarchy among the eager, lazy, and fixpoint $\lambda$-calculi.
 B.1. Böhm Trees
@@
 A head reduction $M\to_{h}M^{\prime}$ contracts the outermost $\beta$-redex only: if $M$ has the form $(\lambda x.\,e)\;v\;M_{1}\cdots M_{k}$, then $M\to_{h}e[v/x]\;M_{1}\cdots M_{k}$.
-A term $M$ is in head normal form (HNF) if it has no head redex, i.e., $M=\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ where $y$ is a variable.
+A term $M$ is in head normal form (HNF) if it has no head redex, that is, $M=\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ where $y$ is a variable.
 The Böhm tree $\mathrm{BT}(M)$ of a $\lambda$-term $M$ is an infinite labeled tree defined by:
 $\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\begin{aligned} &\text{if }M\to^{*}_{h}\\[-2.0pt] &\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{aligned}\end{cases}$
-Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Abramsky and Ong, 1993).
+Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995).
 B.2. Path Encoding
@@
 When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting mixin tree has a specific shape:
-  • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$ (the abstraction shape).
+  • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$, which is referred to as the abstraction shape.
   • A $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ translates to a record with own properties $\{x,\mathrm{result}\}$, where $x$ holds the encapsulated application $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$.
@@
 Each Böhm tree position is a sequence of navigation steps from the root.
-In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body (peeling off one abstraction layer) and entering the $i$-th argument $M_{i}$ of the head variable.
+In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body by peeling off one abstraction layer, and entering the $i$-th argument $M_{i}$ of the head variable.
 In the mixin tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
-Rather than comparing absolute paths across different mixin trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection.
-Recall the definition of inheritance-convergence (Definition 4.1).
+Rather than comparing absolute paths across different mixin trees, which would be sensitive to internal structure; hence, we observe only the convergence behavior accessible via the $\mathrm{result}$ projection.
+Recall the definition of inheritance-convergence (Definition B.6).
 Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation.
-The abstraction shape at the end signals that a weak head normal form has been reached.
+The abstraction shape at the end signals that a head normal form has been reached.
 B.3. Single-Path Lemma
-The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path mixin inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
-Lemma B.1 (Single Path).
+The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-target $\mathtt{this}$ resolution.
+For translated $\lambda$-terms, we show it degenerates to single-target resolution.
+Lemma B.1 (Single-Path).
 For any closed ANF $\lambda$-term $M$, at every invocation of $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)$ during evaluation of $\mathcal{T}(M)$, the frontier set $S$ contains exactly one path.
@@
 Let $M$ be an ANF term in which $x$ may occur free, and let $V$ be a closed value.
+[^32] The proof uses the Single-Path Lemma (Lemma B.1), which requires the inherited tree to occur within a closed term.
+This holds whenever $\lambda x.\,M$ applied to $V$ is a subterm of a closed ANF program.
 Define the inherited tree $C=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$.
@@
 The substitution distributes: $M[V/x]=\mathbf{let}\;z=V^{\prime}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$.
-By induction on $V^{\prime}$ and $M^{\prime}$, the properties under both subtrees ($z$ and $\mathrm{result}$) coincide.
+By induction on $V^{\prime}$ and $M^{\prime}$, the properties under both subtrees, specifically $z$ and $\mathrm{result}$, coincide.
 In particular, occurrences of $z$ in $M^{\prime}$ are translated as the lexical reference $z.\mathrm{result}$ on both sides, so the wrapper introduced by the value $\mathbf{let}$-binding is preserved by the substitution.
@@
 B.5. Convergence Preservation
-The convergence criterion (Definition 4.1) follows the $\mathrm{result}$ chain from the root.
+The convergence criterion (Definition B.6) follows the $\mathrm{result}$ chain from the root.
 Each step in this chain corresponds to one head reduction step in the $\lambda$-calculus.
@@
 Lemma B.3 (Result-step).
-Let $(\lambda x.\,M)\;V$ be a tail call (a $\beta$-redex). Then for every path $p$ and label $\ell$:
+Let $(\lambda x.\,M)\;V$ be a tail call, which is a $\beta$-redex. Then for every path $p$ and label $\ell$:
 $\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)\;\Longleftrightarrow\;\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}(M[V/x])$
@@
 ∎
-Extension to $\mathrm{lfp}(T_{P})$ semantics
-The proof above works directly with the six equations (1)–(6), not with any particular evaluation strategy.
-We distinguish two consequences of switching from single-pass evaluation to multi-pass tabled evaluation (computing $\mathrm{lfp}(T_{P})$ via iterated application of $T_{P}$; see Appendix A).
-Head-reduction preservation is maintained.:
+Remark B.5 (Convergence hierarchy).
+Section 4.1 showed that the Böhm tree correspondence holds at the first iteration $T_{P}\!\uparrow\!1$ and that $\mathrm{lfp}(T_{P})$ is strictly more defined.
+Here we make the connection precise.
+Tabled evaluation (Tamaki and Sato, 1986) computes $\mathrm{lfp}(T_{P})$ by iterating: $T_{P}\!\uparrow\!0=\bot$, $T_{P}\!\uparrow\!1=T_{P}(\bot)$, $T_{P}\!\uparrow\!2=T_{P}(T_{P}(\bot))$, … until stabilisation (Theorem A.6).
+$T_{P}\!\uparrow\!1$ corresponds to $\beta$-reduction.:
+$T_{P}\!\uparrow\!1$ evaluates each equation once, with re-entrant calls returning $\bot$ (the value from $T_{P}\!\uparrow\!0$).
 The Result-Step Lemma (Lemma B.3) establishes a structural isomorphism between the equation systems of $\mathcal{T}(M)$ and $\mathcal{T}(N)$, shifted by one $\mathrm{result}$ step.
-This isomorphism preserves monotonicity, so $\mathrm{lfp}(T_{P})$ yields the same answer on both sides (modulo the depth shift): if $\mathcal{T}(M)$ converges at depth $n$ under $\mathrm{lfp}$, then $\mathcal{T}(N)$ converges at depth n1n-1, and vice versa.
-For $\lambda$-calculus–translated programs, the Single-Path Lemma (Lemma B.1) ensures that no cycles arise in the dependency graph (each scope level has exactly one inheritance source), so single-pass and multi-pass evaluation produce identical results in this fragment.
-Agreement with the $\lambda$-calculus is not maintained in general.:
-For programs beyond the $\lambda$-calculus fragment that introduce cyclic dependencies, $\mathrm{lfp}(T_{P})$ may converge where the $\lambda$-calculus diverges.
-This means $\mathrm{lfp}(T_{P})$ evaluation is strictly more defined than $\beta$-reduction on the translated image: $M{\Downarrow}_{\lambda}$ implies $\mathcal{T}(M){\Downarrow}_{\mathrm{lfp}}$, but not conversely.
-As a minimal illustration, the $\lambda$-term $\mathbf{let}\;x=x\;\mathbf{in}\;x$ diverges under $\beta$-reduction (unfolding $x$ produces $x$ again, ad infinitum). Translation $\mathcal{T}$ gives
-$\{x\mapsto\{\mathrm{result}\mapsto x.\mathrm{result}\},\;\mathrm{result}\mapsto x.\mathrm{result}\}$
-where $x.\mathrm{result}$ inherits from itself, creating a cyclic dependency.
-Under single-pass evaluation ($\mathit{max\_fixpoint\_iterations}=0$), the cycle raises $\bot$, matching the $\lambda$-calculus divergence.
-Under $\mathrm{lfp}(T_{P})$, fixpoint iteration detects the cycle and converges: $\mathrm{properties}(\mathrm{root})=\{x,\mathrm{result}\}$ and $\mathrm{properties}(\mathrm{root}\triangleright\mathrm{result})=\varnothing$, reflecting that the value is undefined (no abstraction shape).
-Appendix F gives a more motivated example (Datalog transitive closure on cyclic graphs).
+This isomorphism is purely structural and holds for any number of iterations.
+Theorems B.7 and B.9 are proved at this level.
+$\mathrm{lfp}(T_{P})$ defines the fixpoint $\lambda$-calculus.:
+We call the sublanguage with $\mathrm{lfp}(T_{P})$ semantics the fixpoint $\lambda$-calculus; it shares the same syntax as the lazy $\lambda$-calculus but has strictly greater computational power: $M{\Downarrow}_{\lambda}$ implies $\mathcal{T}(M){\Downarrow}_{\mathrm{lfp}}$, but not conversely.
 B.6. Adequacy
+Definition B.6 (Inheritance-convergence).
+A mixin tree $T$ converges, written $T{\Downarrow}$, if there exists $n\geq 0$ such that:
+$\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{properties}\!\bigl(\,\underbrace{\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
+in the semantics of Section 3.
+Theorem B.7 (Adequacy).
+A closed ANF $\lambda$-term $M$ has a head normal form if and only if $\mathcal{T}(M){\Downarrow}$.
 A closed ANF term at the top level is either an abstraction (a value) or a computation (an application $\mathbf{let}$-binding, a value $\mathbf{let}$-binding, or a tail call).
 An abstraction translates to a record whose root immediately has the abstraction shape; a computation’s root has the form $\{x,\mathrm{result}\}$ or $\{\mathrm{tailCall},\mathrm{result}\}$, and one must follow the $\mathrm{result}$ chain to find the eventual value.
-Proof of Theorem 4.2.
+Proof of Theorem B.7.
 ($\Rightarrow$) Suppose $M\to^{*}_{h}\lambda x.\,M^{\prime}$ in $k$ head reduction steps. We show $\mathcal{T}(M)$ converges at depth $\leq k$ by induction on $k$.
-Base case ($k=0$): $M$ is already an abstraction $\lambda x.\,M^{\prime}$.
+Base case ($k=0$)
+$M$ is already an abstraction $\lambda x.\,M^{\prime}$.
 Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$, whose root has $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$.
 These labels are directly in $\mathtt{defines}$ of the root, so the recursive evaluation returns them without further recursion, and $\mathcal{T}(M)$ converges at depth $n=0$.
-Inductive step ($k\geq 1$): $M$ is not an abstraction, so $M$ is either a tail call, an application $\mathbf{let}$-binding, or a value $\mathbf{let}$-binding.
+Inductive step ($k\geq 1$)
+$M$ is not an abstraction, so $M$ is either a tail call, an application $\mathbf{let}$-binding, or a value $\mathbf{let}$-binding.
 Each form corresponds to one head $\beta$-step in the underlying $\lambda$-term.
@@
 ($\Leftarrow$) Suppose $\mathcal{T}(M){\Downarrow}$ at depth $n$. We show $M$ has a head normal form by induction on $n$.
-Base case ($n=0$): the root of $\mathcal{T}(M)$ has abstraction shape $\{\mathrm{argument},\mathrm{result}\}$.
+Base case ($n=0$)
+The root of $\mathcal{T}(M)$ has abstraction shape $\{\mathrm{argument},\mathrm{result}\}$.
 Only the abstraction rule $\mathcal{T}(\lambda x.\,M^{\prime})$ produces a root with own property $\mathrm{argument}$.
 (For tail calls and $\mathbf{let}$-bindings, the root has $\mathtt{defines}=\{\mathrm{tailCall},\mathrm{result}\}$ or $\{x,\mathrm{result}\}$, and $\mathrm{argument}$ cannot appear via $\mathtt{supers}$ because $\mathtt{inherits}=\varnothing$ at the root, giving $\mathtt{bases}(\text{root})=\varnothing$.) Therefore $M$ is an abstraction, hence already in head normal form.
-Inductive step ($n\geq 1$): the abstraction shape appears at depth $n$ but not at depth 0.
+Inductive step ($n\geq 1$)
+The abstraction shape appears at depth $n$ but not at depth 0.
 By the base case argument, $M$ is not an abstraction, so at the head position of this closed top-level ANF term, $M$ is either a tail call, an application $\mathbf{let}$-binding, or a value $\mathbf{let}$-binding.
-In each case, the head form is a $\beta$-redex in the underlying $\lambda$-term: the function position is occupied by a closed ANF value, and the only closed ANF value that can appear at head position is a $\lambda$-abstraction.
+In each case, the ANF grammar requires the function position to be a value $V_{1}$ ($V::=x\mid\lambda x.\,M$).
+Since $M$ is closed, $V_{1}$ cannot be a free variable, so $V_{1}$ must be a $\lambda$-abstraction; the head form is therefore a $\beta$-redex.
 By Lemma B.3 (tail-call, application let-binding, or value let-binding variant), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct.
 So $\mathcal{T}(N)$ converges at depth n1n-1.
-Since $M$ is closed and head reduction preserves closedness (substituting a closed value for the only free variable), $N$ is a closed ANF term.
+Since $M$ is closed and head reduction preserves closedness , which substitutes a closed value for the only free variable, $N$ is a closed ANF term.
 By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$.
@@
 The subscript $\lambda$ emphasizes that the quantification ranges over $\lambda$-calculus contexts, not over all inheritance-calculus contexts.
-This definition observes only convergence (a single bit), but quantifying over all $\lambda$-contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior within the $\lambda$-definable fragment.
-Proof of Theorem 4.3.
+This definition observes only convergence [^33] Convergence provides exactly a single bit of information.
+, but quantifying over all $\lambda$-contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior within the $\lambda$-definable fragment.
+Theorem B.9 (First-order semantics of the lazy $\lambda$-calculus).
+For closed $\lambda$-terms $M$ and $N$:
+$\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
+Proof of Theorem B.9.
 The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the mixin tree $\mathcal{T}(C[M])$ is determined by the mixin tree $\mathcal{T}(M)$ and the translation of the context.
@@
 In particular, $C[M]$ has a head normal form iff $C[N]$ does.
-By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$.
+By Theorem B.7, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$.
 Hence $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$.
 ($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$.
-Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa).
-By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$.
+Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges, or vice versa.
+By Theorem B.7, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$.
 Hence $\mathcal{T}(M)\not\approx_{\lambda}\mathcal{T}(N)$.
 ∎
-The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993).
-The prior higher-order frameworks are thereby shown to be faithful to the first-order semantics, without needing to define or compare absolute paths inside mixin trees.
+The proof rests on two pillars: Adequacy (Theorem B.7), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995).
+Neither pillar requires defining or comparing absolute paths inside mixin trees.
 Appendix C Expressive Asymmetry: Proofs
@@
 Definition C.3 (Language universe).
-Let $\mathscr{U}$ be the language whose constructors are the union of the constructors of inheritance-calculus, the constructors of the lazy $\lambda$-calculus ($\lambda$-abstraction, application, variable), and $\mathbf{let}$-binding.
+Let $\mathscr{U}$ be the language whose constructors are the union of the constructors of inheritance-calculus, the constructors of the lazy $\lambda$-calculus , including $\lambda$-abstraction, application, and variable, and $\mathbf{let}$-binding.
 The semantics of $\mathscr{U}$ is given by translating every $\mathscr{U}$-program to inheritance-calculus in three steps:
   1. (1)
-    Convert every maximal $\lambda$-calculus subexpression (i.e. every subtree whose root and all internal nodes use only $\lambda$-calculus constructors) to A-normal form (Flanagan et al., 1993).
+    Convert every maximal $\lambda$-calculus subexpression (that is, every subtree whose root and all internal nodes use only $\lambda$-calculus constructors) to A-normal form (Flanagan et al., 1993).
   2. (2)
@@
 A $\mathscr{U}$-program $P$ converges, $\mathrm{eval}_{\mathscr{U}}(P)$, iff its translation converges.
+Remark C.4.
 The lazy $\lambda$-calculus and inheritance-calculus are conservative restrictions of $\mathscr{U}$: each is obtained by removing the other’s constructors.
 For a pure inheritance-calculus program, steps (1) and (2) are vacuous, so $\mathrm{eval}_{\mathscr{U}}$ agrees with the inheritance-calculus semantics.
-For a pure $\lambda$-calculus program, the composition of steps (1)–(3) agrees with lazy evaluation by Adequacy (Theorem 4.2).
+For a pure $\lambda$-calculus program, the composition of steps (1)–(3) agrees with lazy evaluation by Adequacy (Theorem B.7).
 The three-step translation is effective (ANF conversion and $\mathcal{T}$ are syntactic transformations) and the recursive evaluation of inheritance-calculus (Section 3) is a semi-decision procedure, so $\mathrm{eval}_{\mathscr{U}}$ is recursively enumerable, as required by Felleisen’s framework (Felleisen, 1991).
@@
     $\lambda$-bound variable ($x$ with de Bruijn index $n^{\prime}$).
-This is a nullary construct (no subexpressions).
+This is a nullary construct and therefore has no subexpressions.
 The syntactic abstraction is the constant $A=\uparrow^{\!n^{\prime}}.\mathrm{argument}$.
@@
   3. (3)
-    $\lambda$-abstraction ($\lambda x.\,M$). Unary in the body $M$. $A(\cdot_{1})=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\cdot_{1}\}$.
+    $\lambda$-abstraction, denoted $\lambda x.\,M$. Unary in the body $M$. $A(\cdot_{1})=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\cdot_{1}\}$.
   4. (4)
@@
   4. (iv)
-    Semantics preserved: For pure $\lambda$-terms, steps (1) and (2) of Definition C.3 are vacuous and step (3) yields lazy evaluation by Adequacy (Theorem 4.2), so $\mathrm{eval}_{\mathscr{U}}$ agrees with $\mathrm{eval}_{L_{0}}$ on $L_{0}$-phrases.
+    Semantics preserved: For pure $\lambda$-terms, steps (1) and (2) of Definition C.3 are vacuous and step (3) yields lazy evaluation by Adequacy (Theorem B.7), so $\mathrm{eval}_{\mathscr{U}}$ agrees with $\mathrm{eval}_{L_{0}}$ on $L_{0}$-phrases.
 The theorem requires us to exhibit two $L_{0}$-terms $M$ and $N$ with $M\cong_{0}N$ but $M\not\cong_{1}N$, where $\cong_{0}$ is operational equivalence in $L_{0}$ and $\cong_{1}$ is operational equivalence in $\mathscr{U}$.
 Operational equivalence in $L_{0}$
-By Böhm tree equivalence (Abramsky, 1990) (Theorem 4.3), for closed $\lambda$-terms $M$ and $N$: $M\cong_{0}N$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$.
+By Böhm tree equivalence (Abramsky and Ong, 1993; Abramsky and McCusker, 1995) (Theorem B.9), for closed $\lambda$-terms $M$ and $N$: $M\cong_{0}N$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$.
 Separating pair
@@
 Since $\mathrm{eqFalse}$ (respectively $\mathrm{eqTrue}$) inherits $\mathrm{eq}.\mathrm{result}$, and $\mathrm{tailCall}$ inherits $\mathrm{eqFalse}.\mathrm{result}$ (the $\lambda b$ scope), the $\mathrm{repr}$ record placed by the mixin at the $\lambda b$ scope level propagates through the inheritance chain to $\mathrm{tailCall}$.
-In $C[M]$: the $\mathrm{eqFalse}$ binding supplies $\mathrm{argument}\mapsto f_{1}.\mathrm{result}$ (i.e.
-$\mathrm{false}$) to the $\lambda a$ scope, so $\mathrm{firstOperand}$ evaluates to $\mathrm{false}=\lambda t.\,\lambda f.\,f$.
+In $C[M]$: the $\mathrm{eqFalse}$ binding supplies $\mathrm{argument}\mapsto f_{1}.\mathrm{result}$ (that is, $\mathrm{false}$) to the $\lambda a$ scope, so $\mathrm{firstOperand}$ evaluates to $\mathrm{false}=\lambda t.\,\lambda f.\,f$.
 In $C[N]$: the $\mathrm{eqTrue}$ binding supplies $\mathrm{argument}\mapsto\mathrm{true}$, so $\mathrm{firstOperand}$ evaluates to $\mathrm{true}=\lambda t.\,\lambda f.\,t$.
@@
 ∎
-Appendix D Multi-Path Self-Reference in Other Systems
-This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-path self-reference pattern that inheritance-calculus handles naturally.
+Appendix D Multi-Target $\mathtt{this}$ Resolution in Other Systems
+This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-target $\mathtt{this}$ resolution pattern that inheritance-calculus handles naturally.
 Scala 3
@@
 In DOT, an object is constructed as $\nu(x:T)\,d$, where the self variable $x$ binds to a single object.
-Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer self-reference to a different object.
+Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer $\mathtt{this}$ to a different object.
 Their intersection requires outer to simultaneously return Object1 and Object2, but DOT’s self variable is single-valued, making this intersection unrealizable.
@@
 Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{HasMultipleOuters}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken.
-Because attribute resolution aggregates results via set union (which is intrinsically idempotent), identical property definitions from the two routes merge without conflict.
+Because attribute resolution aggregates results via set union (which is an intrinsically idempotent operation), identical property definitions from the two routes merge without conflict.
 Thus, the calculus eliminates the need for any artificial route-priority or linearization rules.
@@
 The rejection occurs because Object1.MyInner and Object2.MyInner each carry the full staticModules of the MyInner deferred module, including the declaration options.outer.
-When MyObjectA imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error; it cannot express that both declarations originate from the same trait and should be unified.
-The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently.
+When HasMultipleOuters imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error; it cannot express that both declarations originate from the same trait and should be unified.
+The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-target situation where two inheritance routes introduce the same declaration independently.
 In inheritance-calculus, the same pattern is well-defined because $\mathtt{overrides}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
-Appendix E First-Order Semantics of the ANF Lazy $\lambda$-Calculus
-This appendix presents a first-order semantics of the lazy $\lambda$-calculus in A-normal form (ANF).
-The semantics is given by mutually recursive set-comprehension equations whose well-definedness follows from the same termination argument as in Appendix A.
-The central idea is that function application is modeled as tree grafting: applying a function record to an argument splices the function’s subtree into the call site, and variable references navigate the resulting tree structure.
-ANF is essential.
-Traditional $\lambda$-calculus syntax allows anonymous subexpressions such as (fx)(f\;x) in (fx)y(f\;x)\;y that have no label and cannot form paths.
-ANF restores path-addressability by binding every intermediate result to a name, so that every node in the evaluation tree has a unique path address.
+Appendix E First-Order Semantics of the ANF Fixpoint $\lambda$-Calculus
+This appendix reformulates the semantics of Section 3, specialized to $\lambda$-terms via the Single-Path Lemma (Lemma B.1), directly on ANF syntax.
+The first-iteration approximation gives the lazy $\lambda$-calculus semantics.
+The Böhm tree has exactly two node types—scope and application—with separate path types and separate first-order functions for each.
+All dynamic functions are single-valued: the Single-Path Lemma (Lemma B.1) guarantees that no multi-inheritance arises for translated $\lambda$-terms.
 Syntax
@@
 […]
-The grammar includes both application let-bindings and value let-bindings (the latter bind a name to a value without application).
-Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results; ANF is neither a restriction nor an extension of the $\lambda$-calculus.
-Variables use de Bruijn indices: $\uparrow^{\!n}$ counts the number of enclosing scope levels (record braces) between the reference and its binder, following the same convention as inheritance-calculus.
-In particular, a $\lambda$-bound variable that crosses $k$ intervening $\mathbf{let}$-binding scopes and $n$ intervening $\lambda$-abstraction scopes is written $\uparrow^{\!n+k}$, not $\uparrow^{\!n}$.
+The grammar includes application let-bindings ($\mathbf{let}\;x=V_{1}\;V_{2}$), value let-bindings ($\mathbf{let}\;x=V$), tail calls ($V_{1}\;V_{2}$), and tail values ($V$).
+Variables use de Bruijn indices: $\uparrow^{\!n}$ counts only $\lambda$-abstractions between the reference and its binder, following the standard convention.
 Let-bound variables are referenced by their binding name $x$.
-Evaluation tree
-An evaluation tree has a single abstraction: every node can serve as function, argument, binding, and result at the same time.
-A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ldots,\ell_{k})$ identifying a position in the mixin tree.
-We write $p$ for a path, $()$ for the root, $p\mathbin{\triangleright}\ell$ for extension, $\mathtt{init}(p)$ for the parent path, and $\mathtt{last}(p)$ for the final label.
-AST
-An ANF term is parsed into an AST that provides two functions at each path $p$:
-  • $\mathtt{children}(p)$: the set of labels locally defined at $p$.
-  • $\mathtt{refs}(p)$: the set of reference pairs $(n,\;\ell_{*})$, where $n$ is a de Bruijn index and $\ell_{*}$ is a list of downward projections.
-Each ANF construct populates these as follows.
-An abstraction $\lambda x.\,M$ at path $p$:
-$\mathtt{children}(p)$ $=\{\mathrm{argument},\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
-The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$; the subtree at $p\mathbin{\triangleright}\mathrm{argument}$ is empty ($\mathtt{children}=\varnothing$, $\mathtt{refs}=\varnothing$).
-A let-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ at path $p$:
-$\mathtt{children}(p)$ $=\{x,\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
-The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$.
-The subtree at $p\mathbin{\triangleright}x$ represents the application $V_{1}\;V_{2}$ and is determined by the form of $V_{1}$:
-  • If $V_{1}$ is a variable (either $\uparrow^{\!m}$ or a let-bound name), then
-    $\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{argument}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\{(n_{1},\;\ell_{*})\},$
-    where the reference $(n_{1},\;\ell_{*})$ encodes the translation of $V_{1}$ (see below).
-  • If $V_{1}=\lambda y.\,M^{\prime}$, the $\lambda$-literal is inlined:
-    $\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{argument},\;\mathrm{result}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\varnothing,$
-    and the subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{result}$ is determined by $M^{\prime}$.
-In both cases the subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{argument}$ is determined by $V_{2}$.
-A value let-binding $\mathbf{let}\;x=V\;\mathbf{in}\;M$ at path $p$:
-$\mathtt{children}(p)$ $=\{x,\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
-The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$. The subtree at $p\mathbin{\triangleright}x$ wraps the value in a $\mathrm{result}$ scope:
-$\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{result}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\varnothing.$
-The subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{result}$ is determined by $V$.
-The accessor $x.\mathrm{result}$ thus yields the translated value, matching the convention for application let-bindings.
-A tail call $V_{1}\;V_{2}$ at path $p$:
-$\mathtt{children}(p)$ $=\{\mathrm{tailCall},\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
-The subtree at $p\mathbin{\triangleright}\mathrm{tailCall}$ is analogous to the application node $p\mathbin{\triangleright}x$ above: if $V_{1}$ is a variable, $\mathtt{children}=\{\mathrm{argument}\}$ with $\mathtt{refs}=\{(n_{1},\;\ell_{*})\}$; if $V_{1}$ is a $\lambda$-literal, $\mathtt{children}=\{\mathrm{argument},\;\mathrm{result}\}$ with $\mathtt{refs}=\varnothing$ and the $\lambda$-body inlined.
-The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is a forwarding reference:
-$\mathtt{children}(p\mathbin{\triangleright}\mathrm{result})$ $=\varnothing,$ $\mathtt{refs}(p\mathbin{\triangleright}\mathrm{result})$ $=\{(0,\;(\mathrm{tailCall},\;\mathrm{result}))\}.$
-Here $(0,\;(\mathrm{tailCall},\;\mathrm{result}))$ navigates to the enclosing scope ($p$ itself, since $n=0$ from $p\mathbin{\triangleright}\mathrm{result}$’s perspective), then projects $\mathrm{tailCall}.\mathrm{result}$.
-A $\lambda$-bound variable $\uparrow^{\!n}$ used as an inheritance source at an application node $p$ contributes $(n,\;(\mathrm{argument}))$ to $\mathtt{refs}(p)$: navigate up $n$ scope levels, then project $\mathrm{argument}$.
-The same applies when $\uparrow^{\!n}$ occurs as a bare value in tail position or as an argument value, contributing the reference to $\mathtt{refs}$ at the corresponding path.
-A let-bound variable $x$ at path $p$ is resolved lexically: find the nearest ancestor $p^{\prime}$ of $p$ such that $x\in\mathtt{children}(p^{\prime})$, let $n=|p|-|p^{\prime}|-1$, and contribute $(n,\;(x,\;\mathrm{result}))$ to $\mathtt{refs}(p)$, projecting $x.\mathrm{result}$ from the enclosing scope.
-Note that since the de Bruijn index $n$ counts all scope levels (not just $\lambda$-abstractions), it accounts for intervening $\mathbf{let}$-binding and tail-call scopes.
-For example, in $\lambda.\,\lambda.\,\mathbf{let}\;r=\uparrow^{\!2}\;\uparrow^{\!0}\;\mathbf{in}\;r$, the outer-$\lambda$ variable has de Bruijn index $2$ (crossing the $\mathbf{let}$-scope and the inner-$\lambda$ scope), not $1$.
-Both $\mathtt{children}$ and $\mathtt{refs}$ are pure data extracted from the AST; they are not functions of runtime state.
-Labels
-The labels observable at a path are those locally defined at any graft of any transitive callee:
-(7) $\mathtt{labels}(p)=\bigl\{\;\ell\;\big|\;p_{\mathrm{step}}\in\mathtt{callee}^{*}(p),\;p_{\mathrm{graft}}\in\mathtt{grafts}(p_{\mathrm{step}}),\;\ell\in\mathtt{children}(p_{\mathrm{graft}})\;\bigr\}$
-Grafts
-When a function is applied, its subtree is grafted at the call site.
-A child label $\ell$ of the callee also becomes a child of the caller; the child at the callee’s $\ell$ position and the child at the caller’s $\ell$ position (if any) share the same identity.
-The grafts of $p$ collect all such identity-sharing positions:
-(8) $\mathtt{grafts}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{callee}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &p_{\mathrm{callee}}\in\mathtt{callee}^{*}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{children}(p_{\mathrm{callee}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
-Here $\mathtt{callee}^{*}$ denotes the reflexive-transitive closure of $\mathtt{callee}$, defined next.
-Callee
-The callee of a path $p$ is the set of paths that $p$ is grafted from. Concretely, $\mathtt{callee}$ resolves every reference in $p$’s grafts one step:
-[…]
-where $\ell_{*}=(\ell_{1},\ldots,\ell_{k})$.
-The $\mathtt{init}(p_{\mathrm{graft}})$ passes the reference’s enclosing scope to $\mathtt{scope}$, matching the convention of reference resolution (5), which calls $\mathtt{this}(\{p_{\mathrm{site}}\},\;\mathtt{init}(p_{\mathrm{def}}),\;n)$.
-Scope resolution
-$\mathtt{scope}$ navigates upward through the evaluation tree to find the binding site of a de Bruijn reference.
-Each ANF construct ($\lambda$-abstraction, $\mathbf{let}$-binding, or tail call) introduces one scope level; crossing a scope boundary means finding the call site that grafted the current subtree.
-$\mathtt{scope}$ takes three arguments: an evaluation-site path $p_{\mathrm{site}}$ (where the grafted subtree lives), a definition-site path $p_{\mathrm{def}}$ (the position in the original AST), and a de Bruijn index $n$.
-(10) $\mathtt{scope}(p_{\mathrm{site}},\;p_{\mathrm{def}},\;n)=\begin{cases}p_{\mathrm{site}}&\text{if }n=0\\[6.0pt] \mathtt{scope}\!\left(p_{\mathrm{caller}},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
-where $p_{\mathrm{caller}}$ is the unique path satisfying $(p_{\mathrm{caller}},\;p_{\mathrm{def}})\in\mathtt{callee}^{\mathrm{ctx}}(p_{\mathrm{site}})$, and $\mathtt{callee}^{\mathrm{ctx}}$ pairs each element of $\mathtt{callee}^{*}(p_{\mathrm{site}})$ with its provenance:
-(11) $\mathtt{callee}^{\mathrm{ctx}}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{step}}),\;p_{\mathrm{graft}})\;\big|\;p_{\mathrm{step}}\in\mathtt{callee}^{*}(p),\;p_{\mathrm{graft}}\in\mathtt{grafts}(p_{\mathrm{step}})\;\bigr\}$
-At each step, $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a nonnegative integer, the recursion terminates.
-This completes the chain of definitions needed to compute $\mathtt{labels}(p)$.
-Observational fixpoint semantics
-The definitions above constitute the complete semantics.
-Computation arises from querying progressively deeper paths in a lazily constructed tree: the observer drives computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand.
-Formally, the mutually recursive equations (7)–(11) define set membership recursively: $\ell\in\mathtt{labels}(p)$ holds if and only if it can be established by a finite chain of applications of these equations.
-Well-definedness follows from the same termination argument as for the mixin tree equations (Appendix A): if the recursive evaluation terminates, the reachable dependency subgraph is finite and acyclic, and the equations have a unique solution.
-When no finite derivation exists for a given query, the recursive evaluation does not terminate, analogous to divergence.
-Convergence
-An ANF term $M$ converges, written $M{\Downarrow}$, if there exists $n\geq 0$ such that
-$\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{labels}\!\bigl(\,\underbrace{\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
-in the semantics defined by the equations above.
-This corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: $M$ converges iff it has a head normal form (Theorem 4.2), and two terms are equivalent under this semantics iff they have the same Böhm tree (Theorem 4.3).
-The proofs, given in Appendix B, apply verbatim: each equation above is the specialization of the corresponding mixin tree equation (Section 3) to terms in the image of the translation $\mathcal{T}$ (Section 4), where the Single-Path Lemma (Lemma B.1) ensures that the set-valued $\mathtt{this}$ function (equation 6) degenerates to the single-valued $\mathtt{scope}$ (equation 10).
+Two node types
+Each position in the Böhm tree is either a scope node or an application node. These have distinct path types $\mathcal{S}$ and $\mathcal{A}$ and distinct first-order functions.
+A scope node (path type $\mathcal{S}$) represents a body: a chain of let-bindings ending in a tail expression.
+A scope may or may not be a $\lambda$-abstraction (recorded by $\mathrm{is\_abstraction}$).
+Its edges are the names of application let-bindings $\{x_{1},\ldots,x_{k}\}$; value let-bindings do not create edges (see AST below).
+Navigation from a scope to a binding: $s\mathbin{\triangleright}x_{i}$ gives an application path $a\in\mathcal{A}$.
+An application node (path type $\mathcal{A}$) represents a function application $V_{1}\;V_{2}$ bound by a $\mathbf{let}$.
+In the AST, $\mathtt{head\_ref}$ and $\mathtt{arg\_ref}$ are references (de Bruijn index plus projection, or an inline $\lambda$-literal); in the evaluated Böhm tree, $\mathtt{head}$ and $\mathtt{argument}$ lazily resolve these references to scope nodes.
+Navigation from an application to a scope: $\mathtt{head}(a)\in\mathcal{S}$ and $\mathtt{argument}(a)\in\mathcal{S}$.
+The path types alternate strictly:
+$\mathcal{S}\xrightarrow{\;\mathbin{\triangleright}\,x_{i}\;}\mathcal{A}\xrightarrow{\;\mathtt{head}/\mathtt{argument}\;}\mathcal{S}\xrightarrow{\;\mathbin{\triangleright}\,y_{j}\;}\mathcal{A}\;\cdots$
+There is no scope-to-scope or application-to-application direct transition.
 Appendix F Encoding Datalog in Inheritance-Calculus
@@
 Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisitorTail}$.
-This provides the right-hand (visitor) side of a join: when two relations’ $\mathrm{VisitorMap}$ entries are composed, $\mathrm{VisitorTail}$ carries the joined relation’s row data for the matched constant.
+This provides the right-hand side of a join [^36] In the context of the visitor pattern, this acts as the visitor.
+: when two relations’ $\mathrm{VisitorMap}$ entries are composed, $\mathrm{VisitorTail}$ carries the joined relation’s row data for the matched constant.
 WithVisiteeTail:
 Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisiteeTail}$.
-This provides the left-hand (visitee) side: the relation being iterated.
+This provides the left-hand side [^37] This acts as the visitee.
+: the relation being iterated.
 Together, $\mathrm{WithVisitorTail}$ and $\mathrm{WithVisiteeTail}$ enable value-level joins where only constants present in both relations produce output.
@@
 Recursive Datalog rules, such as transitive closure $\mathrm{path}(X,Y)\mathrel{{:}\!{-}}\mathrm{edge}(X,Z),\;\mathrm{path}(Z,Y)$, are encoded identically to non-recursive rules: the head relation $\mathrm{path}$ appears both as a defined scope and as a reference in the body (providing $\mathrm{WithVisitorTail}$ for the $Z$-join).
-This should not be read as ordinary self-reference of a single-valued lazy program.
+This should not be read as ordinary $\mathtt{this}$ of a single-valued lazy program.
 Operationally, tabled evaluation (Section 3, Appendix A) computes $\mathrm{lfp}(T_{P})$: each query into $\mathrm{path}$ triggers further evaluation through the recursive body, and the trie union ($\mathtt{supers}$) accumulates all reachable tuples.
@@
 Semi-Naive evaluation (Bancilhon and Ramakrishnan, 1986), which restricts each iteration to rules that use at least one newly derived fact, is a natural optimisation left for future work.
+The recursive rule as monadic pseudocode
+The recursive rule above roughly corresponds to the following pseudocode (Haskell-like syntax with RebindableSyntax and OverloadedLists semantics, where do-notation desugars to a Set-based bind and list literals denote sets):
+{-# LANGUAGE RebindableSyntax, OverloadedLists #-}
+s >>= f = unions [f x | x <- s]
+return x = [x]
+guard True = [()]
+guard False = []
+edge = [("a","b"), ("b","c"), ("c","a")]
+path = edge <> do
+(x, y0) <- edge
+(y1, z) <- path
+guard (y0 == y1)
+return (x, z)-}
+Each line of the pseudocode has a direct counterpart in the inheritance-calculus encoding (Section F.4):
+edge:
+corresponds to the base-fact trie;
+path = edge <> do …:
+corresponds to $\mathrm{path}$ inheriting $\mathrm{edge}$ (base facts) and the recursive body (derived facts via $\mathtt{supers}$);
+(x, y0) <- edge:
+corresponds to Level 1 ($\mathrm{xAcc}$), iterating $X$ from $\mathrm{edge}$;
+(y1, z) <- path:
+corresponds to Level 2 ($\mathrm{yAcc}$), joining on $Y$ with $\mathrm{path}$, and Level 4 ($\mathrm{zAcc}$), iterating $Z$;
+guard (y0 == y1):
+corresponds to the $\mathrm{WithVisitorTail}\cap\mathrm{WithVisiteeTail}$ intersection in Level 2: a constant produces a $\mathrm{Visitor}$ entry only if it appears in both relations;
+return (x, z):
+corresponds to the $\mathrm{Replacement}$ chain in Level 5, constructing the output tuple $(X,Z)$.
+The only difference is the evaluation strategy: Haskell evaluates path lazily (diverges), whereas inheritance-calculus evaluates it via tabled $T_{P}$ iteration (converges to $\mathrm{lfp}$).
+However, this monadic program does not converge because path is defined recursively.
+[^38] Enabling the RecursiveDo extension does not help: mfix ties a knot at the element level via laziness, whereas the transitive closure equation requires a fixed point at the container level, that is, finding a set $S$ such that $S=F(S)$.
+Inheritance-calculus solves this through tabled evaluation (Section 3), which computes the least fixed point of the $T_{P}$ operator by iterating over the lattice of mixin trees, precisely the container-level fixed point that the pseudocode above cannot express.
 F.9. Semantic Correspondence with Standard Datalog
@@
 This is the behavior exhibited by RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026).
-When the $Y$-join matches constant $b$ (present in both $\mathrm{edge}$’s second column from $(a,b)$ and $\mathrm{edge}$’s first column), the $\mathrm{Visitor}$ for $b$ receives $\mathrm{VisitorTail}=\mathrm{edge}.\mathrm{TailMap}.b$ (containing only $c$), and $\mathrm{Replacement}$ constructs a single output tuple (a,c)(a,c)—not all of $\{a,b,c\}$.
+When the $Y$-join matches constant $b$ (present in both $\mathrm{edge}$’s second column from $(a,b)$ and $\mathrm{edge}$’s first column), the $\mathrm{Visitor}$ for $b$ receives $\mathrm{VisitorTail}=\mathrm{edge}.\mathrm{TailMap}.b$ (containing only $c$), and $\mathrm{Replacement}$ constructs a single output tuple $(a,c)$, rather than all of $\{a,b,c\}$.
 This value-level precision arises because each constant’s $\mathrm{Replacement}$ operates on $\mathrm{TailMap}.d$ (the per-constant sub-trie) rather than on the merged trie.